### v2

# Approximating the coefficients of the Bessel functions

Abstract.
For the type A, BC, and D root systems, we determine equivalent conditions between the coefficients of an exponential holomorphic function and the asymptotic values taken by the Dunkl bilinear form when one of its entries is the function. We establish these conditions over the $|\theta N|\rightarrow\infty$ regime for the type A and D root systems and over the $|\theta_{0}N|\rightarrow\infty,\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ regime for the type BC root system. We also generalize existing equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type A root system and over the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ regime for the type BC root system and prove new equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type D root system. Furthermore, we determine the asymptotics of the coefficients of the Bessel functions over the regimes that we have mentioned.
1. Introduction
For positive integers $N\geq 2$, we study the asymptotics of the Bessel functions associated to the irreducible root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. We denote the Bessel function associated to a finite root system $\mathcal{R}$ and multiplicity function $\theta$ as $J^{\mathcal{R}(\theta)}_{a}(x)$, where $a,x\in\mathbb{C}^{N}$, provided that the function exists. It is an eigenfunction of the Dunkl operators associated to $\mathcal{R}$ and $\theta$, which are introduced in [dunkloperators]. Furthermore, the function is symmetric with respect to the reflection group generated by $\mathcal{R}$ in the variables $a$ and $x$. For the definitions of root systems, multiplicity functions, and the Bessel functions, see Subsection˜2.1.
For certain choices of $\theta$, the paper [dunklbessel] shows that $J^{\mathcal{R}(\theta)}_{a}(x)$ exists, is unique, and is holomorphic over $a,x\in\mathbb{C}^{N}$, see Theorem˜2.8. The paper also discusses a nonsymmetric eigenfunction, see Theorem˜2.5; the formula relating the two eigenfunctions is given in Theorem˜2.8.
In this paper, we analyze the coefficients of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is fixed and $\mathcal{R}$ is one of $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$. The multiplicity function $\theta$ varies with $N$ and should be viewed as a function of $N$, although we denote it as $\theta$ rather than $\theta(N)$ for brevity. We outline the notation we use to record the root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$ and the corresponding multiplicity functions. Note that $e_{i}\triangleq[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}$ for $i\in[N]$.
  • For $\theta\in\mathbb{C}$, we let $A^{N-1}(\theta)$ denote the root system $A^{N-1}$ with multiplicity function assigning $\theta$ to the roots $e_{i}-e_{j}$ for distinct $i,j\in[N]$.
  • For $\theta_{0},\theta_{1}\in\mathbb{C}$, we let $BC^{N}(\theta_{0},\theta_{1})$ denote the root system $B^{N}$ or $C^{N}$ with multiplicity function assigning $\theta_{0}$ to the roots $e_{i}-e_{j}$, $e_{j}-e_{i}$, $e_{i}+e_{j}$, and $e_{i}-e_{j}$ for $i,j\in[N]$ such that $i<j$ and $\theta_{1}$ to the roots that are scalar multiples of $e_{i}$ for $i\in[N]$.
  • For $\theta\in\mathbb{C}$, we let $D^{N}(\theta)$ denote the root system $D^{N}$ with multiplicity function assigning $\theta$ to the roots $e_{i}-e_{j}$, $e_{j}-e_{i}$, $e_{i}+e_{j}$, and $e_{i}-e_{j}$ for $i,j\in[N]$ such that $i<j$.
Furthermore, we define the Dunkl operators associated with these root systems:
[…]
for $i\in[N]$, where $s_{ij}$ switches the $i$th and $j$th entries of an element of $\mathbb{C}^{N}$ for distinct $i,j\in[N]$ and $\tau_{i}$ flips the sign of the $i$th entry of an element of $\mathbb{C}^{N}$ for $i\in[N]$.
To compute the coefficients of $J_{a}^{\mathcal{R}(\theta)}$, it is equivalent to compute the values of
$f(\mathcal{D}_{1}(\mathcal{R}(\theta)),\ldots,\mathcal{D}_{N}(\mathcal{R}(\theta)))g(x_{1},\ldots,x_{N})$
for all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are homogeneous, have the same degree, and symmetric with respect to the reflection group generated by $\mathcal{R}$; for the explanation of why this is the case, see Lemma˜4.19. We compute the asymptotics of these values in Theorem˜1.1.
1.1. Main result
For the notation regarding partitions and noncrossing partitions in the following result, see Subsections˜2.2 and 2.3, respectively. We briefly mention that for $N\geq 1$, $\Gamma_{N}$ consists of partitions with all parts of size at most $N$, $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}$, and $\Gamma_{\text{even}}$ and $\Gamma_{\text{even};\,N}$ are the elements of $\Gamma$ and $\Gamma_{N}$, respectively, with all parts of even size; for $\lambda\in\Gamma$, $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$. Additionally, $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$. Furthermore, for $k\geq 1$, $NC(k)$ denotes the set of of noncrossing partitions of $[k]$ and $NC^{\text{even}}(k)$ denotes the set of noncrossing partitions of $[k]$ with all even block sizes.
We state the main result of this paper, which is over the regime $|\theta N|\rightarrow\infty$ for the $A^{N-1}$ and $D^{N}$ root systems and the regime $|\theta_{0}N|\rightarrow\infty$, $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ for the $BC^{N}$ root system. Observe that the first regime includes the case where $\theta$ is a fixed element of $\mathbb{C}^{\times}$ and the second regime includes the case where $\theta_{0}$ is a fixed element of $\mathbb{C}^{\times}$ and $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$.
Theorem 1.1.
Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
(A): Suppose $F^{A}_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$ for $N\geq 1$. Then, the following are equivalent.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(A^{N-1}(\theta))^{\nu_{i}}F_{N}^{A}}{(\theta N)^{|\nu|}N^{\ell(\nu)}}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
(B): Suppose $F^{BC}_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]$ for $N\geq 1$. Consider the following statements.
  1. (c)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{\theta_{0}N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (d)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(BC^{N}(\theta_{0},\theta_{1}))^{\nu_{i}}F_{N}^{BC}}{(\theta_{0}N)^{|\nu|}N^{\ell(\nu)}}$
    $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
Then, (c) implies (d), and if $c\not=-1$, then (d) implies (c).
(C): Suppose
$F_{N}^{D}(x_{1},\ldots,x_{N})$ $=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$
$\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]$
for $N\geq 1$. Consider the following statements.
  1. (e)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
  2. (f)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
  3. (g)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}F_{N}^{D}}{N^{\ell(\nu)}(\theta N)^{|\nu|}}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
  4. (h)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}F_{N}^{D}}{N^{\ell(\nu)}(\theta N)^{|\nu|}\prod_{j=1}^{N}(1+2(j-1)\theta)}$
    $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
Then, (e) and (g) are equivalent.
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (f) and (h) are equivalent.
Proof.
See Corollaries˜6.11, 7.5 and 8.22 for the proofs of (A), (B), and (C), respectively. ∎
First, we note that we actually prove the generalizations where the $N\rightarrow\infty$ limits of $\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$, $\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}$, and $\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$ can be nonzero when $\ell(\lambda)>1$, see Theorems˜6.10, 7.4 and 8.21. We present this corollary because it is more applicable to the setting where the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are set as $J_{a(N)}^{A^{N-1}(\theta)}$, $J_{a(N)}^{BC^{N}(\theta_{0},\theta_{1})}$, and $J_{a(N)}^{D^{N}(\theta)}$ in some sufficiently small neighborhood of the origin where the Bessel functions are nonzero so that their logarithms are holomorphic over the neighborhood; $a(N)\in\mathbb{C}^{N}$ varies with $N$.
For example, in the case where $F_{N}^{A}\triangleq J_{a(N)}^{A^{N-1}(\theta)}$, condition (a) implies that for $\nu\in\Gamma$,
$\lim_{N\rightarrow\infty}\frac{\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\left(\frac{a(N)_{j}}{\theta N}\right)^{\nu_{i}}}{N^{\ell(\nu)}}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
Then, we can deduce the convergence of the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ in terms of moments. Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin. We prove a generalized version of this result in Corollary˜10.18.
Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]. In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$. The main contribution of this paper is the set of equivalence relations in the theorem where some conditions are based on scaling the coefficients of the logarithms of the input functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ by varying powers of $\theta N$. The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed is mentioned in the appendix of [matrix]. However, in [limits_prob_measures], the coefficients are only scaled by $(\theta N)^{-1}$ and it is similarly the case that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
As we mention in Section˜6, a method to prove part (A) of Theorem˜1.1 is to prove Theorem˜6.1 using the results of [limits_prob_measures]. However, this method is not applicable to parts (B) and (C), so we develop a new approach which can be used in these settings in Subsection˜6.1. This approach is applicable to the $\theta N\rightarrow c\in\mathbb{C}$ regime as well.
Interestingly, in part (C) of Theorem˜1.1, we also consider the asymptotics of coefficients of terms with degree increasing to infinity. See Subsection˜1.6, Section˜8, and Section˜11 for more discussion regarding this direction.
1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
Observe that in Theorem˜1.1, we consider the following regimes:
  1. 1.
    The $N\rightarrow\infty$ limit of $|\theta N|$ is $\infty$ for $A^{N-1}(\theta)$ and $D^{N}(\theta)$.
  2. 2.
    The $N\rightarrow\infty$ limit of $|\theta_{0}N|$ is $\infty$ and the $N\rightarrow\infty$ limit of $\frac{\theta_{1}}{\theta_{0}N}$ is $c\in\mathbb{C}$ for $BC^{N}(\theta_{0},\theta_{1})$.
We consider the following regimes in addition to those listed above:
  1. 3.
    The $N\rightarrow\infty$ limit of $\theta N$ is $c\in\mathbb{C}$ for $A^{N-1}(\theta)$ and $D^{N}(\theta)$.
  2. 4.
    The $N\rightarrow\infty$ limit of $\theta_{0}N$ is $c_{0}\in\mathbb{C}$ and the $N\rightarrow\infty$ limit of $\theta_{1}$ is $c_{1}\in\mathbb{C}$ for $BC^{N}(\theta_{0},\theta_{1})$.
Note that [matrix] considers regime 3 for the root system $A^{N-1}$ while [rectangularmatrix] considers regime 4. The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well. Theorems that we prove over regimes 3 and 4 are Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix].
1.3. Related works
Some related works that we have already mentioned are [matrix, limits_prob_measures, rectangularmatrix]. These works are based on analyzing the Bessel generating functions of exponentially decaying measures to compute the asymptotic moments of the measures. While we focus on a broader setting, we discuss the applications of this paper’s results to exponentially decaying measures in Subsection˜10.5. Another work that is based on analyzing Bessel generating functions to compute asymptotic moments is [gaussianfluctuation], which considers when $\theta=1$. Furthermore, the papers [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov] consider similar results by analyzing the coefficients of Jack and Schur generating functions.
The paper [limits_prob_measures] introduces a framework for analyzing the asymptotic moments when the coefficients of terms of the Taylor series of the type A Bessel generating functions with two or more variables can have nonzero $N\rightarrow\infty$ limits after scaling by $(\theta N)^{-1}$. In fact, we are able to apply the results of the paper to obtain a straightforward proof of part (A) of Theorem˜1.1, as we mentioned earlier. In addition, we present a new proof method that is applicable to other settings, such as the $BC^{N}$ and $D^{N}$ root systems and regimes 3 and 4 that we described in Subsection˜1.2.
1.4. Weak convergence to the free convolution
Note that the definitions of $\theta(\mathcal{R})$ and $H(\mathcal{R})$ are included in Subsection˜2.1.
First, we discuss an integral representation of the nonsymmetric eigenfunction for the Dunkl operators. This integral representation assumes that the multiplicity function is nonnegative.
Theorem 1.2 ([rosler_positivity]).
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a\in\mathbb{R}^{N}$. There exists a unique Borel probability measure $\mu_{a}$ whose support is contained in the convex hull of $H(\mathcal{R})a$ such that
$E_{a}^{\mathcal{R}}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
for all $x\in\mathbb{C}^{N}$. Furthermore, $\text{supp}(\mu_{a})\cap H(\mathcal{R})a$ is nonempty and the Borel probability measure
$\mu_{a}^{\text{sym}}\triangleq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}\mu_{ha}$
is invariant with respect to the action of $H(\mathcal{R})$ and satisfies
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)$
for all $x\in\mathbb{C}^{N}$.
Remark 1.3.
From [dunkloperators, dunklbessel], if $\theta\in\theta(\mathcal{R})$ such that $\theta(r)$ has nonnegative real part for all $r\in\mathcal{R}$, then $\theta\in\Theta(\mathcal{R})$, or equivalently, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. See Subsection˜2.1 for the definitions of these notions. The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8. In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.2.
We state a well-known conjecture related to an integral representation of the product of two Bessel functions; it can easily be extended to the product of an arbitrary number of Bessel functions. The conjecture also assumes that the multiplicity function is nonnegative.
Conjecture 1.4.
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$. There exists a nonnegative probability measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ over $\mathbb{R}^{N}$ such that
$J^{\mathcal{R}(\theta)}_{a_{1}}(x)J^{\mathcal{R}(\theta)}_{a_{2}}(x)=\int_{\mathbb{R}^{N}}J^{\mathcal{R}(\theta)}_{a}(x)d\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}(a)$
for all $x\in\mathbb{C}^{N}$.
The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$. By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in ˜1.4 is satisfied. It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
The paper [rosler_product] shows that ˜1.4 holds in the context of radially symmetric Bessel functions. Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in ˜1.4. For the proofs of the corollaries, see Subsection˜10.1.
Corollary 1.5.
Assume that ˜1.4 is true. Furthermore, assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$. Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$ as defined in [nica_speicher_2006]*Definition 12.1.
Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$. Let $\tilde{\mu}$ be the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$.
  1. (A)
    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
  2. (B)
    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\tilde{\mu}$ as $N\rightarrow\infty$.
Remark 1.6.
By “always”, we mean that the statement is true for any choices of the measures $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ and $\mu_{a(N),b(N)}^{D^{N}(\theta)}$, since there may be multiple choices.
For the type BC root system, we can deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
Corollary 1.7.
Assume that ˜1.4 is true. Furthermore, assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$. Let $\mu$ be the rectangular free convolution with $\lambda$ set as $\frac{1}{1+c}$ of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ as defined in [rectangular_free_convolution]*Proposition-Definition 2.1. The distribution $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
We also note that analogues of part (A) of Corollary˜1.5 and Corollary˜1.7 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix]. We can similarly establish the analogue of part (B) of Corollary˜1.5 in the $\theta N\rightarrow c\in\mathbb{C}$ regime; its statement is almost the same as that of the analogue of part (A) of the corollary.
1.5. Uniform convergence of the Bessel functions
For a fixed value of $\lambda\in\Gamma$, we can compute the asymptotics of the coefficients of $p_{\lambda}(x)$ in the Taylor expansions of $J_{a}^{A^{N-1}(\theta)}(x)$, $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(x)$, and $J_{a}^{D^{N}(\theta)}(x)$, which are homogeneous polynomials of degree $|\lambda|$ in $a\in\mathbb{C}^{N}$. These computations are included in Section˜9.
[…]
Furthermore, in the case where $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$ or $\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$, we can compute the asymptotic coefficients when $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N}$ exists for all $k\in\mathbb{N}$. This setting is more general than that of the Vershik-Kerov sequences although $\theta$ is not fixed. Additionally, we require that $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer.
When $|\theta N|\rightarrow\infty$ and $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N(\theta N)^{k}}$ exists for all $k\in\mathbb{N}$, we can no longer compute the asymptotic coefficients; for example, in part (A) of Theorem˜1.1, $c_{\lambda}(N)$ has order $(\theta N)^{\ell(\lambda)}$ for $\lambda\in\Gamma$. In this case, we can still approximate the coefficients, although they will not converge.
However, even if the coefficients of the sequence of Bessel function converge, we have not yet determined that the sequence of functions uniformly converges. In order to prove uniform convergence over compact subsets of an open and simply connected domain, we follow the argument of [rank_infinity] and first prove that the Bessel functions are uniformly bounded and then apply Montel’s theorem, see Subsections˜10.3 and 10.4.
To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.2. An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
1.6. Coefficients of terms with all odd degrees in the type D Bessel function
The type D Bessel function is a symmetric linear combination of terms with all even and all odd degrees. To analyze the coefficients of the terms with all odd degrees of these functions, we compute the asymptotics of
(1) $\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}\prod_{j=1}^{N}x_{j}\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$
for $\nu,\lambda\in\Gamma_{\text{even}}$. These quantities are mentioned in part (h) of Theorem˜1.1 and are interesting because they are no longer polynomials in $N$ and $\theta$; however, they can be expressed as the product of a polynomial and an expression involving gamma functions, see Theorem˜8.16. In fact, we can consider analogous results for $A^{N-1}$ and $BC^{N}$, although terms with all odd degrees are not particularly important for analyzing the Bessel functions for these root systems so the results are not used while computing their coefficients, see Theorems˜12.1 and 12.3.
Perhaps more interesting are the asymptotics of (1) in the regime $\theta N\rightarrow c\in\mathbb{C}$ for $A^{N-1}$ and $D^{N}$ and the regime $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ for $BC^{N}$. For the corresponding leading order terms for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, see Theorems 12.4, 12.5, and 8.23, respectively. Of course, we are most interested in the $D^{N}$ case since that is when summands which are multiples of $\prod_{j=1}^{N}x_{j}$ are present in the Bessel function. However, it is interesting that we can apply the methods that we develop for the $D^{N}$ case to the $A^{N-1}$ and $BC^{N}$ cases.
1.7. Applying a graded ring of operators to a graded vector field
In Sections 3, 4, and 5, we discuss the applications of a graded ring of operators to a graded vector field. The paper [dunkl_singular_poly] discusses the applications of operators to a graded vector field. We extend this notion by considering a graded ring of operators. The results that we obtain are relevant to Dunkl operators and in particular the Dunkl bilinear form, see Example˜5.19; for the definition of the Dunkl bilinear form, see Subsection˜2.1. We also study the notion of invertible operators, which is the focus of [dunkl_singular_poly]. See Section˜4 for the definition of invertibility. This framework and especially the content of Section˜4 are useful for analyzing the Bessel functions.
1.8. Paper organization
In Section˜2, we define the Dunkl operators and notation regarding partitions and noncrossing partitions. In Section˜3, we introduce the setting of applying a graded ring of operators to a graded vector field and in Section˜4, we define the notion of an invertible graded ring of operators. Afterwards, in Section˜5, we discuss a representation of invertible graded rings of operators as sequences of invertible matrices and connect the framework introduced in Section˜3 to the Dunkl operators. In Sections 6, 7, and 8, we discuss the leading order terms of the Dunkl bilinear form and prove Theorem˜1.1 for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, respectively. Following this, in Section˜9, we determine the asymptotics of the coefficients of the terms of the Bessel functions that are homogeneous with a fixed degree. In Section˜10, we discuss applications of the results of this paper and in Section˜11, we present combinatorial expressions for the Dunkl bilinear form. In Section˜12, we analyze a special case of the leading order terms of the Dunkl bilinear form for the $A^{N-1}$ and $BC^{N}$ root systems.
2. Basic definitions and notation
2.1. Dunkl operators
Suppose $N\geq 1$ and that $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system. For $\alpha\in\mathcal{R}$, we let $r_{\alpha}$ denote the reflection $r_{\alpha}:x\mapsto x-2\left\langle x,\alpha\right\rangle\left\lVert\alpha\right\rVert_{2}^{-2}\alpha$. Let $H(\mathcal{R})$ be the finite reflection group generated by $r_{\alpha}$ for $\alpha\in\mathcal{R}$. For a function $f$ over $\mathbb{C}^{N}$, we define the action of $h\in H(\mathcal{R})$ over $f$ by $hf(x)\triangleq f(hx)$.
Let $\mathcal{R}^{+}$ be a set of positive roots in $\mathcal{R}$. Furthermore, let $\theta(\mathcal{R})$ be the set of multiplicity functions $\theta:\mathcal{R}\rightarrow\mathbb{C}$ such that $\theta(\alpha_{1})=\theta(\alpha_{2})$ for all $\alpha_{1},\alpha_{2}\in\mathcal{R}$ such that $r_{\alpha_{1}}$ and $r_{\alpha_{2}}$ are conjugates in $H$. When we write $\mathcal{R}$, we assume that $\mathcal{R}$ is a finite root system and when we write $\mathcal{R}(\theta)$ to denote a root system and a multiplicity function, it is implicit that $\theta\in\theta(\mathcal{R})$. Unless stated otherwise, $N$ is a positive integer and $\mathcal{R}\subset\mathbb{R}^{N}$.
Next, we define the Dunkl operator introduced in [dunkloperators]. For $u\in\mathbb{R}^{N}$, define the operator $\mathcal{D}_{u}(\mathcal{R}(\theta))$ over the ring $\mathbb{C}[[x_{1},\ldots,x_{N}]]$ of complex formal power series with variables $x_{1},\ldots,x_{N}$ by $\mathcal{D}_{u}(\mathcal{R}(\theta)):f\mapsto\left\langle\nabla_{\theta}f,u\right\rangle$, where
$\nabla_{\theta}f(x)\triangleq\nabla f(x)+\sum_{\alpha\in\mathcal{R}^{+}}\theta(\alpha)\frac{f(x)-f(r_{\alpha}x)}{\left\langle x,\alpha\right\rangle}\alpha.$
The definition of the Dunkl operator does not depend on the choice of $\mathcal{R}^{+}$, see [dunklbound]*Remark 2.4. Additionally, it is well known that the Dunkl operators are commutative, which is stated in the following lemma.
Lemma 2.1 ([dunkloperators]).
For $u_{1},u_{2}\in\mathbb{R}^{N}$, $\mathcal{D}(\mathcal{R}(\theta))_{u_{1}}\mathcal{D}(\mathcal{R}(\theta))_{u_{2}}=\mathcal{D}(\mathcal{R}(\theta))_{u_{2}}\mathcal{D}(\mathcal{R}(\theta))_{u_{1}}$ .
For $1\leq i\leq N$, define $\mathcal{D}_{i}(\mathcal{R}(\theta))\triangleq\mathcal{D}_{[\mathbf{1}\{i=j\}]_{1\leq j\leq N}^{T}}(\mathcal{R}(\theta))$. Furthermore, for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, we define $\mathcal{D}(\mathcal{R}(\theta))(f)$ to be the operator $f(\mathcal{D}_{1},\ldots,\mathcal{D}_{N})$; note that this operator is well defined by Lemma˜2.1. The following lemma is also well known.
Lemma 2.2 ([dunkloperators]).
Suppose $h\in H(\mathcal{R})$. Then, for all $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $h\mathcal{D}(\mathcal{R}(\theta))(f)h^{-1}=\mathcal{D}(\mathcal{R}(\theta))(hf)$.
By the previous lemma, we have that for all $h\in H(\mathcal{R})$, $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, and $g\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$, $h\mathcal{D}(\mathcal{R}(\theta))(f)g=\mathcal{D}(\mathcal{R}(\theta))(hf)hg$. We use this result later in the paper, for example to prove that a function exhibits symmetries after applications of Dunkl operators.
Furthermore, we have that $\mathcal{D}$ defines a bilinear form which is introduced in [dunkl_integralkernel]. For $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$, the Dunkl bilinear form is defined as
$f,g]_{\mathcal{R}(\theta)}\triangleq[1]\mathcal{D}(\mathcal{R}(\theta))(f)g.$
Theorem 2.3 ([dunkl_integralkernel]).
For all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $f,g]_{\mathcal{R}(\theta)}=[g,f]_{\mathcal{R}(\theta)}$.
Due to the symmetry of $\cdot,\cdot]_{\mathcal{R}(\theta)}$, it is straightforward to define and compute the values of $f,g]_{\mathcal{R}(\theta)}$ and $g,f]_{\mathcal{R}(\theta)}$ when $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $g\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$. When $f,g\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$, the value of $f,g]_{\mathcal{R}(\theta)}$ does not necessarily converge.
Definition 2.4.
The function $\mathcal{D}(\mathcal{R}(\theta))$ is invertible if for all $k\geq 1$, there does not exist $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ such that $f$ is homogeneous of degree $k$ and $f,g]_{\mathcal{R}(\theta)}=0$ for all $g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that is homogeneous of degree $k$. If $\mathcal{D}(\mathcal{R}(\theta))$ is not invertible, then it is singular. Let $\Theta(\mathcal{R})$ be the set of $\theta\in\theta(\mathcal{R})$ such that $\mathcal{D}(\mathcal{R}(\theta))$ is invertible.
Given $\mathcal{R}$, the paper [dunkl_singular_poly] computes all $\theta\in\theta(\mathcal{R})$ such that $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. Note that the statements of the definition of invertibility in this paper and [dunkl_singular_poly] are not the same. However, the definitions are equivalent, see Theorem˜4.5 where we consider a more general setting.
In this paper, we focus on the asymptotics of $\cdot,\cdot]_{\mathcal{R}(\theta)}$ for the root systems $A_{N-1}$, $B_{N}$, $C_{N}$, and $D_{N}$ as the number of variables $N$ increases to infinity. We use these asymptotics to determine the asymptotics of the eigenfunctions of $\mathcal{D}(\mathcal{R}(\theta))$, which we define next.
Theorem 2.5 ([dunklbessel]).
Suppose $\theta\in\Theta(\mathcal{R})$. Then, there exists a unique function $E_{a}^{\mathcal{R}(\theta)}(x)$ that is holomorphic over the domain $\mathbb{C}^{N}\times\mathbb{C}^{N}$ for $(a,x)$ and satisfies
$\begin{cases}\mathcal{D}(\mathcal{R}(\theta))(f)E_{a}^{\mathcal{R}(\theta)}(x)=f(a)E_{a}^{\mathcal{R}(\theta)}(x)&\forall f\in\mathbb{C}[x_{1},\ldots,x_{N}],\\ E_{a}^{\mathcal{R}(\theta)}(0)=1.\end{cases}$
Furthermore, $E_{a}^{\mathcal{R}(\theta)}(x)$ is holomorphic over the domain $\mathbb{C}^{N}\times\mathbb{C}^{N}\times\Theta(\mathcal{R})$ for $(a,x,\theta)$.
The following result contains some properties about the eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
Lemma 2.6 ([dunklbound]).
Suppose $\theta\in\Theta(\mathcal{R})$.
  1. (a)
    For all $h\in H(\mathcal{R})$, $E_{ha}^{\mathcal{R}(\theta)}(hx)=E_{a}^{\mathcal{R}(\theta)}(x)$.
  2. (b)
    $E_{a}^{\mathcal{R}(\theta)}(x)=E_{x}^{\mathcal{R}(\theta)}(a)$.
  3. (c)
    Suppose $c\in\mathbb{C}$. Then, $E_{a}^{\mathcal{R}(\theta)}(cx)=E_{ca}^{\mathcal{R}(\theta)}(x)$.
  4. (d)
    $\overline{E_{a}^{\mathcal{R}(\theta)}(x)}=E_{\overline{a}}^{\mathcal{R}(\overline{\theta})}(\overline{x})$.
  5. (e)
    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\\ \text{Re}(\left\langle ha,x\right\rangle))$.
We can also consider the symmetric analogue of Theorem˜2.5 after averaging over $H(\mathcal{R})$. First, we define $\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$ to be the set of $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are fixed under the action of $H(\mathcal{R})$. The following lemma is well-known.
Lemma 2.7.
Suppose $f\in\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$. Then, for $i\in[N]$, $\mathcal{D}_{i}(\mathcal{R}(\theta))f=\partial_{i}f$.
The next result is the symmetric analogue of Theorem˜2.5. See [dunkl_singular_poly] for elaboration on the proof of the result.
Theorem 2.8 ([dunklbessel]).
Suppose $\theta\in\Theta(\mathcal{R})$. Then, there exists a unique function $J_{a}^{\mathcal{R}(\theta)}(x)$ that is holomorphic over the domain $\mathbb{C}^{N}\times\mathbb{C}^{N}$ for $(a,x)$ and satisfies
$\begin{cases}\mathcal{D}(\mathcal{R}(\theta))(f)J_{a}^{\mathcal{R}(\theta)}(x)=f(a)J_{a}^{\mathcal{R}(\theta)}(x)&\forall f\in\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}],\\ J_{a}^{\mathcal{R}(\theta)}(0)=1.\end{cases}$
Furthermore, $J_{a}^{\mathcal{R}(\theta)}(x)$ is holomorphic over the domain $\mathbb{C}^{N}\times\mathbb{C}^{N}\times\Theta(\mathcal{R})$ for $(a,x,\theta)$ and
$J_{a}^{\mathcal{R}(\theta)}(x)=\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}hE_{a}^{\mathcal{R}(\theta)}(x).$
A generalization of these two results is included in Theorem˜4.5, which concerns applying a graded ring of operators to a graded vector space.
Furthermore, we let $\mathcal{D}_{H}(\mathcal{R}(\theta))$ denote the function such that for $f\in\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$, $\mathcal{D}_{H}(\mathcal{R}(\theta))(f)$ is the restriction of $\mathcal{D}(\mathcal{R}(\theta))(f)$ to $\mathbb{C}^{H(\mathcal{R})}[[x_{1},\ldots,x_{N}]]$, which is the set of $f\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$ that are fixed under the action of $H(\mathcal{R})$. Then, $\mathcal{D}_{H}(\mathcal{R}(\theta))$ is a symmetric version of $\mathcal{D}(\mathcal{R}(\theta))$.
Furthermore, for $k\geq 1$, we let $E_{a}^{\mathcal{R}(\theta)}[k](x)$ and $J_{a}^{\mathcal{R}(\theta)}[k](x)$ denote the sums of the terms of the Taylor expansions of $E_{a}^{\mathcal{R}(\theta)}(x)$ and $J_{a}^{\mathcal{R}(\theta)}(x)$, respectively, that are homogeneous of degree $k$ in $x$. This notation is used in Section˜9.
We mention that if it is clear what root system and multiplicity function we are considering, then we often do not include $\mathcal{R}(\theta)$ in the notation. For example, if this is the case then we would denote $\mathcal{D}_{i}(\mathcal{R}(\theta))$ by $\mathcal{D}_{i}$ and $\mathcal{D}(\mathcal{R}(\theta))$ by $\mathcal{D}$.
For $i\in[N]$, we let $d_{i}$ denote the operator that lowers the degree in $x_{i}$ by one. That is, $d_{i}$ maps $x_{i}^{k}$ to $x_{i}^{k-1}$ for $k\geq 1$ and $1$ to zero.
As discussed earlier, we focus on the irreducible root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$, which are subsets of $\mathbb{R}^{N}$ for $N\geq 2$. We define these root systems.
The definition of $A^{N-1}$. For $i\in[N]$, we define $e_{i}\triangleq[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}\in\mathbb{R}^{N}$. Let $A^{N-1}\triangleq\{e_{i}-e_{j}:i,j\in[N],\,i\not=j\}$. Each $\theta\in\theta(A^{N-1})$ is constant over the root system, so we let $A^{N-1}(\theta)$ for $\theta\in\mathbb{C}$ denote the choice of $A^{N-1}$ as the root system and $\theta$ as the multiplicity function.
The reflection group $H(A^{N-1})$ permutes the entries of $\mathbb{C}^{N}$. For $i\in[N]$, the associated Dunkl operator is
[…]
where $s_{ij}$ switches the $i$th and $j$th entries of an element of $\mathbb{C}^{N}$ for distinct $i,j\in[N]$.
Furthermore, $\mathbb{C}^{H(A^{N-1})}[x_{1},\ldots,x_{N}]$ is the set of symmetric functions in $\mathbb{C}[x_{1},\ldots,x_{N}]$. Equivalently, it is the span of $\{1\}\cup\{p_{\lambda}:\lambda\in\Gamma_{N}\}$ and $\{1\}\cup\{M_{\epsilon}:\epsilon\in\mathcal{P}_{N}\}$.
The definitions of $B^{N}$ and $C^{N}$. Let $B^{N}\triangleq\bigcup_{i,j\in[N],\,i<j}\{e_{i}-e_{j},e_{j}-e_{i},e_{i}+e_{j},-e_{i}-e_{j}\}\bigcup_{i\in[N]}\{e_{i},-e_{i}\}$ and $C^{N}\triangleq\bigcup_{i,j\in[N],\,i<j}\{e_{i}-e_{j},e_{j}-e_{i},e_{i}+e_{j},-e_{i}-e_{j}\}\bigcup_{i\in[N]}\{2e_{i},-2e_{i}\}$. A multiplicity function $\theta\in\theta(B^{N})$ is constant over the roots of length $\sqrt{2}$ and over the roots of length $1$. Similarly, a multiplicity function $\theta\in\theta(C^{N})$ is constant over the roots of length $\sqrt{2}$ and over the roots of length $2$.
We have that $B^{N}$ and $C^{N}$ are dual root systems such that $H(B^{N})=H(C^{N})$. Furthermore, we always have that $\mathcal{D}(B^{N}(\theta))=\mathcal{D}(C^{N}(\theta))$. Since we do not need to differentiate between these two root systems in this paper, we let $BC^{N}$ denote the root system $B^{N}$ or $C^{N}$. Furthermore, for $\theta_{0},\theta_{1}\in\mathbb{C}$, we let $BC^{N}(\theta_{0},\theta_{1})$ denote the choice of $B^{N}$ or $C^{N}$ as the root system and the function that assigns $\theta_{1}$ to the scalar multiplies of $e_{i}$ and $\theta_{0}$ to the remaining roots as the multiplicity function.
We have that $H(BC^{N})$ permutes the entries of $\mathbb{C}^{N}$ and applies sign flips to any number of entries. Furthermore, for $i\in[N]$, the associated Dunkl operator is
[…]
where $\tau_{i}$ flips the sign of the $i$th entry of an element of $\mathbb{C}^{N}$ for $i\in[N]$. When we are working in the context of type BC root systems, we refer to $\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right)$ for distinct $i,j\in[N]$ as a type 0 switch and $\theta_{1}\frac{1-\tau_{i}}{x_{i}}$ for $i\in[N]$ as a type 1 switch.
Furthermore, $\mathbb{C}^{H(BC^{N})}[x_{1},\ldots,x_{N}]$ is the set of symmetric functions in $\mathbb{C}[x_{1},\ldots,x_{N}]$ that have all even degrees. Equivalently, it is the span of $\{1\}\cup\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$.
The definition of $D^{N}$. Let $D^{N}\triangleq\bigcup_{i,j\in[N],i<j}\{e_{i}-e_{j},e_{j}-e_{i},e_{i}+e_{j},-e_{i}-e_{j}\}$. A multiplicity function $\theta\in\theta(D^{N})$ is constant over the root system, so we let $D^{N}(\theta)$ for $\theta\in\mathbb{C}$ denote the choice of $D^{N}$ as the root system and $\theta$ as the multiplicity function.
We have that $H(D^{N})$ permutes the entries of $\mathbb{C}^{N}$ and applies sign flips to an even number of entries. Furthermore, for $i\in[N]$, the associated Dunkl operator is
[…]
Additionally, $\mathbb{C}^{H(D^{N})}[x_{1},\ldots,x_{N}]$ is the set of symmetric functions in $\mathbb{C}[x_{1},\ldots,x_{N}]$ that are sums of monomials that have all degrees of the same parity. Equivalently, it is the span of $\{1\}\cup\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}\cup\{ep_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$.
2.2. Partitions
Suppose $N\geq 1$. Let $\Gamma_{N}$ denote the set of nonempty partitions $(\lambda_{1}\geq\cdots\geq\lambda_{m})\in[N]^{m}$ for $m\geq 1$ and let $\mathcal{P}_{N}$ be the set of nonempty partitions with at most $N$ parts. Furthermore, define $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}=\bigcup_{N\geq 1}\mathcal{P}_{N}$. Note that we do not assume that $\Gamma$ contains the empty partition.
For $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{m})\in\Gamma$, let $|\lambda|\triangleq\sum_{i=1}^{m}\lambda_{i}$ and $\ell(\lambda)\triangleq m$. Also, we define $\Gamma_{\text{even};\,N}$ (resp. $\Gamma_{\text{even}}$) to be the set of $\lambda\in\Gamma_{N}$ (resp. $\Gamma$) such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$.
Suppose $k\geq 1$. Define $\Gamma_{N}[k]\triangleq\{\lambda\in\Gamma_{N}:|\lambda|=k\}$ and define $\mathcal{P}_{N}[k]$, $\Gamma_{\text{even};\,N}[k]$, $\Gamma[k]$, and $\Gamma_{\text{even}}[k]$ analogously. Furthermore, for a set $S$ and $M\in S^{\Gamma\times\Gamma}$ (resp. $S^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}$), we define $M[k]\in S^{\Gamma[k]\times\Gamma[k]}$ (resp. $S^{\Gamma_{\text{even}}[k]\times\Gamma_{\text{even}}[k]}$) to be $M$ with rows and columns restricted to $\Gamma[k]$ (resp. $\Gamma_{\text{even}}[k]$).
For a positive integer $m$ and a sequence $s=(a_{1},\ldots,a_{m})$ of nonnegative integers, define $\gamma(s)$ to be the element of $\Gamma$ that is a permutation of the sequence formed from $s$ after deleting the entries that equal zero. Furthermore, for $\nu\in\Gamma$, define $\pi(\nu)$ to be $|\nu|!$ divided by the number of permutations of $\nu$ and $\nu!\triangleq\prod_{i=1}^{\ell(\nu)}\nu_{i}!$; if $\nu$ contains $n_{i}$ copies of $i$ for all $i\geq 1$, then $\pi(\nu)=\prod_{i\geq 1}n_{i}!$. We similarly define $\pi(s)\triangleq\pi(\gamma(s))$ and $s!\triangleq\gamma(s)!$. For $x\in\mathbb{C}^{m}$, we define $x^{s}\triangleq\prod_{i=1}^{m}x_{i}^{a_{i}}$.
We also consider the sums of partitions. For $\lambda_{1},\ldots,\lambda_{k}\in\Gamma$, we define $\lambda_{1}+\cdots+\lambda_{k}\triangleq\gamma((\lambda_{1},\ldots,\lambda_{k}))$ for all $k\geq 2$, where $(\lambda_{1},\ldots,\lambda_{k})$ denotes the tuple formed by combining the entries of $\lambda_{1},\ldots,\lambda_{k}$.
For $k\geq 1$, define $p_{(k)}(x_{1},\ldots,x_{N})\triangleq x_{1}^{k}+\cdots+x_{N}^{k}$ and for $\lambda\in\Gamma$, define $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}p_{(\lambda_{i})}(x_{1},\ldots,x_{N})$. Furthermore, for $\epsilon\in\mathcal{P}_{N}$, define
$M_{\epsilon}(x_{1},\ldots,x_{N})\triangleq\sum_{\begin{subarray}{c}(a_{1},\ldots,a_{N})\in\mathbb{Z}_{\geq 0}^{N},\\ \gamma((a_{1},\ldots,a_{N}))=\epsilon\end{subarray}}\prod_{i=1}^{N}x_{i}^{a_{i}}=\sum_{\begin{subarray}{c}(a_{1},\ldots,a_{N})\in\mathbb{Z}_{\geq 0}^{N},\\ \gamma((a_{1},\ldots,a_{N}))=\epsilon\end{subarray}}(x_{1},\ldots,x_{N})^{(a_{1},\ldots,a_{N})}.$
Also, define $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$.
The following two lemmas are straightforward to deduce, but are essential components of this paper.
Lemma 2.9.
Suppose $N\geq 1$ and $c_{(k)}\in\mathbb{C}$ for all $k\geq 1$. Then,
$\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}(x_{1},\ldots,x_{N})\right)=1+\sum_{\lambda\in\Gamma}\pi(\lambda)^{-1}\prod_{i=1}^{\ell(\lambda)}c_{(\lambda_{i})}p_{\lambda}(x_{1},\ldots,x_{N})$
over $\mathbb{C}[[x_{1},\ldots,x_{N}]]$.
Lemma 2.10.
Suppose $N\geq 1$ and $a,b\in\mathbb{C}^{N}$. Then,
$\exp(\left\langle a,b\right\rangle)=1+\sum_{\epsilon\in\mathbb{Z}^{N}_{\geq 0},\,\epsilon\not=0}\frac{a^{\epsilon}b^{\epsilon}}{\epsilon!}.$
For $k\geq 1$, let $P_{k}$ denote the set of $p\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are symmetric and homogeneous of degree $k$. It is clear that a basis for $P_{k}$ is $\{M_{\epsilon}(x_{1},\ldots,x_{N}):\epsilon\in\mathcal{P}_{N}[k])$. The following result is also well known.
Lemma 2.11.
Suppose $k\geq 1$. A $\mathbb{C}$-linear basis for $P_{k}$ is given by $\{p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{N}[k]\}$.
Proof.
We denote $\{p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{N}[k]\}$ by $p(\Gamma_{N}[k])$. By the fundamental theorem of symmetric polynomials and Newton’s identities, $P_{k}$ is spanned by $p(\Gamma_{N}[k])$. To finish the proof, it suffices to show that $|\Gamma_{N}[k]|=\text{dim}(P_{k})$. Thus, it suffices to show that $|\Gamma_{N}[k]|=|\mathcal{P}_{N}[k]|$.
Suppose $\epsilon\in\mathcal{P}_{N}[k]$. For $i\geq 1$, let $\lambda_{i}=|\{j\in[\ell(\epsilon)]:\epsilon_{j}\geq i\}|$ and suppose $m$ is the largest positive integer such that $\lambda_{m}>0$. Then, define $\alpha(\epsilon)\triangleq(\lambda_{1}\geq\cdots\geq\lambda_{m})\in\Gamma_{N}[k]$. Furthermore, for $\lambda\in\Gamma_{N}[k]$, let $\epsilon_{i}=|\{j\in[\ell(\lambda)]:\lambda_{j}\geq N+i-1\}|$ for $1\leq i\leq N$ and suppose $m$ is the smallest positive integer such that $\epsilon_{i}>0$. Then, define $\alpha(\lambda)\triangleq(\epsilon_{N}\geq\cdots\geq\epsilon_{m})\in\mathcal{P}_{N}[k]$. Since $\alpha$ and $\beta$ are inverses, we have that $|\Gamma_{N}[k]|=|\mathcal{P}_{N}[k]|$. ∎
For $k\geq 1$, let $P_{k}^{\text{even}}$ and $P_{k}^{\text{odd}}$ be the set of $p\in P_{k}$ that have all even and odd degrees, respectively. It is also straightforward to deduce the following result using the previous result.
Lemma 2.12.
Suppose $k\geq 1$. A $\mathbb{C}$-linear basis for $P_{2k}^{\text{even}}$ is given by $\{p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$ and a $\mathbb{C}$-linear basis for $P_{2k+N}^{\text{odd}}$ is given by $\{e(x_{1},\ldots,x_{N})p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$.
2.3. Non-crossing partitions
Suppose $k\geq 1$. Let $NC(k)$ denote the set of noncrossing partitions of $[k]$. Recall that a partition $B_{1}\sqcup\cdots\sqcup B_{m}$ of $[k]$ is noncrossing if there does not exist distinct $i,j\in[m]$, $a,c\in B_{i}$, and $b,d\in B_{j}$ such that $a<b<c<d$. Furthermore, for $k\geq 1$, define $NC^{\text{even}}(k)$ to be the set of $\pi\in NC(k)$ such that each block of $\pi$ has even size.
Suppose $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(k)$ such that the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $1\leq q\leq m-1$. For $q\in[m]$ and $i\in B_{q}$, define $b(i;\,\pi)\triangleq\mathbf{1}\{i=\min(B_{q})\}$ and
$d(i;\,\pi)\triangleq\left|\left(\bigcup_{r=1}^{q}B_{r}\right)\bigcap\{i,i+1,\ldots,k\}\right|.$
Furthermore, define $o(\pi)$ to be the number of $i\in[k]$ such that $b(i;\,\pi)=0$ and $d(i;\,\pi)$ is odd. Also, define $f(\pi)\triangleq|B_{1}|$ and $\gamma(\pi)\triangleq\gamma((|B_{1}|,\ldots,|B_{m}|))$.
[…]
Furthermore, define $NC^{\prime}(k-1)$ to be the set of $\pi$ such that $m\geq 2$ and the minimal element of $B_{2}$ is $2$.
3. An introduction to applying a graded ring of operators to a graded vector field
The paper [dunkl_singular_poly] considers the applications of operators to a graded vector field $V\triangleq\oplus_{i\geq 0}V_{i}$, where the application of an operator to an element of $V_{i}$ outputs an element of $V_{i-1}$ for $i\geq 1$. When $V=\mathbb{C}[x_{1},\ldots,x_{N}]$ and $V_{i}$ is the set of elements of $V$ that are homogeneous of degree $i$ for all $i\geq 0$, we have that $\partial_{i}$, $\mathcal{D}(\mathcal{R}(\theta))_{i}$, and $d_{i}$ for $i\in[N]$ are examples of such an operator. In this section, we introduce a similar setting that is motivated by the Dunkl bilinear form $\cdot,\cdot]_{\mathcal{R}(\theta)}$ and we continue to discuss this setting in Sections˜4 and 5. See Example˜5.19 for the application of the framework that we introduce to the context of Dunkl operators.
Let $K$ be a field. Assume that $V\triangleq\oplus_{i\geq 0}V_{i}$ is a graded vector space such that $V_{0}=K$ and $V_{i}$ is a finite dimensional $K$-vector space for $i\geq 1$. Furthermore, assume that $R\triangleq\oplus_{i\geq 0}R_{i}$ is a graded ring such that $R_{i}$ is a finite dimensional $K$-vector space for $i\geq 0$. From the definition of a graded ring, recall that $R_{i}$ is an additive abelian group for $i\geq 0$ and $R_{i_{1}}R_{i_{2}}\subset R_{i_{1}+i_{2}}$ for all $i_{1},i_{2}\geq 0$.
For $k\in K$, let $k^{*}$ denote the element $\{v\mapsto kv\}$ of $\text{End}_{K}(V)$. Let $\mathcal{L}$ be the set of $K$-linear ring homomorphisms $L:R\rightarrow\text{End}_{K}(V)$ such that for $i_{1},i_{2}\geq 0$, $f\in V_{i_{2}}$, and $g\in R_{i_{1}}$,
$\begin{cases}L(g)f\in V_{i_{2}-i_{1}}&\text{if }i_{1}\leq i_{2},\\ L(g)f=0&\text{if }i_{2}<i_{1}.\end{cases}$
Remark 3.1.
By the definition of a ring homomorphism, $L(gh)=L(g)L(h)$ and $L(g+h)=L(g)+L(h)$ for all $g,h\in R$. Furthermore, $L(0)=0^{*}$, $L(1)=1^{*}$, and the $K$-linearity condition implies that $L(kr)=kL(r)$ for all $r\in R$ and $k\in K$.
Lemma 3.2.
Suppose $L\in\mathcal{L}$.
  1. (A)
    For all $k\in K$, $L(k)=k^{*}$.
  2. (B)
    For all $r\in R$ and $k\in K$, $L(rk-kr)=0^{*}$.
Proof.
For (A), by the $K$-linearity condition, $L(k)=kL(1)=k^{*}$ for all $k\in K$. For (B), observe that
$L(rk)=L(r)k^{*}=k^{*}L(r)=L(kr)$
for all $r\in R$ and $k\in K$, because $L(r)$ is a $K$-linear endomorphism of $V$. ∎
Remark 3.3.
We do not assume that $K\subset\text{center}(R)$ despite part (B) of Lemma˜3.2.
For $i\geq 0$, let $A_{i}$ and $B_{i}$ be bases of $V_{i}$ and $R_{i}$, respectively, as $K$-vector spaces. Define the isomorphisms $a_{i}:V_{i}\rightarrow K^{A_{i}}$ by $a_{i}(r)=[\mathbf{1}\{s=r\}]_{s\in A_{i}}^{T}$ for $r\in A_{i}$ and $b_{i}:R_{i}\rightarrow K^{B_{i}}$ by $b_{i}(r)=[\mathbf{1}\{s=r\}]_{s\in B_{i}}^{T}$ for $r\in B_{i}$. Furthermore, for $A\in\text{End}_{K}(V_{i})$, define $a_{i}(A)\in K^{A_{i}\times A_{i}}$ to be the matrix with column $r$ equal to $a_{i}(Ar)$ for $r\in A_{i}$. Then, we have that
$a_{i}(Af)=a_{i}(A)a_{i}(f)$
for all $f\in A_{i}$. Note that we denote $a_{i}$ and $b_{i}$ by $a$ and $b$ if it is clear that the input is in $A_{i}$ and $B_{i}$, respectively.
Furthermore, for $L\in\mathcal{L}$ and $i\geq 0$, let $M^{i;L}\in K^{B_{i}\times A_{i}}$ denote the matrix such that
$M^{i;L}_{rs}=L(r)s$
for $r\in B_{i}$ and $s\in A_{i}$.
Lemma 3.4.
Suppose $R$ is commutative. Then, the operators $L(g)$ for $g\in R$ are commutative.
Proof.
Suppose $g,h\in R$. Then,
$L(g)L(h)=L(gh)=L(hg)=L(h)L(g).$
Lemma 3.5.
Suppose $i\geq 0$, $f\in V_{i}$, and $g\in R_{i}$. Then,
$L(g)f=b(g)^{T}M^{i;L}a(f).$
Proof.
First, observe that
$f=\sum_{r\in A_{i}}a(f)_{r}r\text{ and }g=\sum_{s\in B_{i}}b(g)_{s}s.$
We therefore have that
$L(g)f=\sum_{r\in A_{i},\,s\in B_{i}}b(g)_{s}a(f)_{r}L(s)r=\sum_{r\in A_{i},\,s\in B_{i}}b(g)_{s}M^{i;L}_{r,s}a(f)_{r}=b(g)^{T}M^{i;L}a(f).$
Definition 3.6.
Suppose $\mathcal{V}\in\text{End}_{K}(V)$. Then, $\mathcal{V}$ is degree-preserving if $\mathcal{V}V_{i}\subset V_{i}$ for all $i\geq 0$.
For $\mathcal{V}\in\text{End}_{K}(V)$ that is degree-preserving, let $\mathcal{V}_{i}$ denote its restriction to $V_{i}$ for all $i\geq 0$. Then, it is evident that $(a(\mathcal{V}_{i}))_{i\geq 0}$ is a representation for the action of $\mathcal{V}$ on $V$. A degree-preserving operator we consider is an intertwining operator, which is well studied in the context of Dunkl operators.
Definition 3.7.
Suppose $L_{1},L_{2}\in\mathcal{L}$. Then, $(L_{1},L_{2})$ intertwines with $\mathcal{V}\in\text{End}_{K}(V)$ if:
  1. (1)
    The operator $\mathcal{V}$ is degree-preserving.
  2. (2)
    For $k\in V_{0}\triangleq K$, $\mathcal{V}k=k$.
  3. (3)
    For all $f\in V$ and $g\in R$, $L_{1}(g)\mathcal{V}f=\mathcal{V}L_{2}(g)f$
Furthermore, if these conditions are satisfied, then $(L_{1},L_{2})$ is intertwining.
Remark 3.8.
If $(L_{1},L_{2})$ is intertwining, we do not necessarily have that $(L_{2},L_{1})$ is intertwining.
Let $\mathcal{F}(V)$ denote the abelian group of formal power series $(E^{i})_{i\geq 0}$ such that $E^{i}\in V_{i}$ for all $i\geq 0$. For the addition operation, we have that $(E_{1}^{i})_{i\geq 0}+(E_{2}^{i})_{i\geq 0}\triangleq(E_{1}^{i}+E_{2}^{i})_{i\geq 0}$.
Suppose $\Psi\in\text{Hom}_{K}(R,K)$, which is the set of $K$-linear ring homomorphisms from $R$ to $K$. The formal power series $E\triangleq(E^{i})_{i\geq 0}\in\mathcal{F}(V)$ is a $\Psi$-eigenvector of $L\in\mathcal{L}$ if for all $f\in R$, we have that
$L(f)E=\Psi(f)E.$
More specifically, $E$ is a $\Psi$-eigenvector of $L$ if for all $j\geq i\geq 0$ and $f\in R_{i}$,
$L(f)E^{j}=\Psi(f)E^{j-i}.$
Definition 3.9.
Suppose $\mathcal{S}\subset\text{Hom}_{K}(R,K)$. Define the kernel of $\mathcal{S}$, which we denote as $\text{ker}(\mathcal{S})$, to be the set of $r\in R$ such that $\Psi(r)=0$ for all $\Psi\in\mathcal{S}$.
In the next section, we discuss the invertibility of an element $L$ of $\mathcal{L}$; the general definition of invertibility that we give is analogous to Definition˜2.4.
4. Invertible graded rings of operators
In this section, we introduce and characterize the invertible elements of $\mathcal{L}$. The notion of invertibility generalizes the notion of singular Dunkl operators arising from certain multiplicity functions that has been studied in [dunklbessel, dunkl_singular_poly], with a singular Dunkl operator corresponding to a non-invertible element of $\mathcal{L}$. First, we define when an element of $\mathcal{L}$ is invertible; the definition is analogous to Definition˜2.4.
Definition 4.1.
The homomorphism $L\in\mathcal{L}$ is invertible if $M^{i;L}$ is an invertible square matrix with nonzero dimensions for all $i\geq 0$.
Corollary 4.2.
If $L\in\mathcal{L}$ is invertible, then $\text{dim}(R_{i})=\text{dim}(V_{i})\geq 1\Leftrightarrow|A_{i}|=|B_{i}|\geq 1$ for all $i\geq 0$.
Corollary 4.3.
Suppose $L\in\mathcal{L}$, $A_{0}=\{a_{0}\}$, and $B_{0}=\{b_{0}\}$ for $a_{0},b_{0}\in K^{\times}$. Then, $M^{0;L}=[a_{0}b_{0}]$.
Proof.
This follows from part (A) of Lemma˜3.2. ∎
[…]
Definition 4.4.
[…]
The following result generalizes ideas that have appeared previously in the study of Dunkl operators in [dunklbessel, dunkl_singular_poly, dunkl_complex]. In particular, [dunkl_singular_poly]*Section 2 discusses a similar setting regarding operators over a graded vector space and the existence of intertwiners for these operators. We extend this idea by considering a graded ring of operators that acts on a graded vector space and connecting the invertibility of an operator with the existence of an intertwiner. Furthermore, we consider the existence of eigenvectors, which leads to additional applications such as the existence of unique eigenfunctions for the complex Dunkl operators introduced in [dunkl_complex], although we do not study this direction further.
The key contribution that we discuss in this section while proving Theorem˜4.5 is the analysis of the matrices $\{M^{i;L}\}_{i\geq 0}$. To the best of our knowledge, the specific method we use has not appeared previously. It allows for straightforward proofs that are applicable in a general setting.
Theorem 4.5.
Assume that $\text{dim}(R_{i})=\text{dim}(V_{i})\geq 1$, $A_{i}=\{a_{ij}\}_{1\leq j\leq|A_{i}|}$, and $B_{i}=\{b_{ij}\}_{1\leq j\leq|A_{i}|}$ for all $i\geq 0$. The following are equivalent.
  1. (a)
    The homomorphism $L\in\mathcal{L}$ is invertible, which by definition is equivalent to the matrix $M^{i;L}$ being invertible for all $i\geq 0$.
  2. (b)
    For all $i\geq 0$, there does not exist nonzero $f\in R_{i}$ such that $L(g)f=0$ for all $g\in R_{i}$.
  3. (c)
    For all $i\geq 0$, there does not exist nonzero $g\in R_{i}$ such that $L(g)f=0$ for all $f\in R_{i}$.
  4. (d)
    The equation $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=0$ has no nonzero solutions $f\in R_{i}$ for all $i\geq 0$.
  5. (e)
    For all $i\geq 0$, there exist unique $\text{inv-row}_{ij}\in R_{i}$ for $1\leq j\leq|A_{i}|$ such that $L(\text{inv-row}_{ij_{1}})a_{ij_{2}}=\mathbf{1}\{j_{1}=j_{2}\}$ for $j_{1},j_{2}\in[|A_{i}|]$.
  6. (f)
    For all $i\geq 0$, there exist unique $\text{inv-col}_{ij}\in V_{i}$ for $1\leq j\leq|A_{i}|$ such that $L(b_{ij_{1}})\text{inv-col}_{ij_{2}}=\mathbf{1}\{j_{1}=j_{2}\}$ for $j_{1},j_{2}\in[|A_{i}|]$.
  7. (g)
    For some invertible homomorphism $L^{\prime}\in\mathcal{L}$, $(L,L^{\prime})$ is intertwining.
Assume that $K\subset\text{center}(R)$. Then, (h) is equivalent to (a).
  1. (h)
    For all homomorphisms $L^{\prime}\in\mathcal{L}$, $(L,L^{\prime})$ intertwines with a unique intertwiner.
Assume that $\text{ker}(\text{Hom}_{K}(R,K))=\{0\}$. Then, (i) and (j) are equivalent to (a).
  1. (i)
    For some $\mathcal{S}\subset\text{Hom}_{K}(R,K)$ such that $\text{ker}(\mathcal{S})=\{0\}$, there exists a $\Psi$-eigenvector $E$ with $E^{0}=1$ for all $\Psi\in\mathcal{S}$.
  2. (j)
    For all $\Psi\in\text{Hom}_{K}(R,K)$, there exists a unique $\Psi$-eigenvector $E$ such that $E^{0}=1$.
Assume that $R$ is right-spanned by $R_{1}$. Then, (k) is equivalent to (a).
  1. (k)
    There does not exist $f\in V\backslash{V_{0}}$ such that $L(g)f=0$ for all $g\in R_{1}$.
In the remaining portions of this section, we assume that $\text{dim}(R_{i})=\text{dim}(V_{i})\geq 1$ for all $i\geq 0$. Observe that $\text{dim}(R_{0})=\text{dim}(V_{0})=1$, so $R_{0}=K$. The goal of this section is to prove Theorem˜4.5. We also prove some additional results relating the singular value decompositions of the matrices $\{M^{i;L}\}_{i\geq 0}$ to eigenvectors, see Proposition˜4.22, and the structure of non-invertible operators, see Theorem˜4.27.
Lemma 4.6.
The following is true: $(a)\Leftrightarrow(b)\Leftrightarrow(c)$.
Corollary 4.7.
If $L\in\mathcal{L}$ is invertible, then $K\subset\text{center}(R)$.
Proof.
See part (B) of Lemma˜3.2 and condition (b). ∎
In fact, $K\subset\text{center}(R)$ is a necessary and sufficient condition for the existence of invertible $L\in\mathcal{L}$, see Theorem˜5.4 in addition to the previous corollary.
4.1. Statements (d), (e), and (f)
We consider (d), (e), and (f). Recall that $A_{i}=\{a_{ij}\}_{1\leq j\leq|A_{i}|}$ and $B_{i}=\{b_{ij}\}_{1\leq j\leq|A_{i}|}$ for $i\geq 0$.
Lemma 4.8.
Suppose $i\geq 0$ and $v\in K^{A_{i}}$. Define $\varphi:K^{A_{i}}\rightarrow K^{B_{i}}$ by $\varphi(a(a_{ij}))=b(b_{ij})$ for $1\leq j\leq|A_{i}|$. Let $f=a^{-1}(v)\triangleq\sum_{s\in A_{i}}v_{s}s$. Then, for $L\in\mathcal{L}$ and $\lambda\in K$, $M^{i;L}v=\lambda\varphi(v)$ if and only if $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=\lambda f$.
Proof.
We have that $M^{i;L}v=\lambda\varphi(v)$ if and only if
$L(b_{ij})f=(M^{i;L}v)_{b_{ij}}=\lambda v_{a_{ij}}\,\,\,\forall 1\leq j\leq|A_{i}|.$
This is equivalent to $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=\lambda\sum_{j=1}^{|A_{i}|}v_{a_{ij}}a_{ij}=\lambda f$. ∎
Corollary 4.9.
The conditions (a) and (d) are equivalent.
Proof.
This follows from (b) or checking that $M^{i;L}$ has no zero eigenvalues for all $i\geq 0$ using Lemma˜4.8. ∎
Lemma 4.10.
The conditions (a) and (e) are equivalent. If either is satisfied, $b(\text{inv-row}_{ij})$ is the transpose of row $b_{ij}$ of $(M^{i;L})^{-1}$ for $i\geq 0$ and $1\leq j\leq|A_{i}|$.
Proof.
Assume that $L$ is invertible. Suppose $i\geq 0$ and $1\leq j_{1}\leq|A_{i}|$. We have that $L(\text{inv-row}_{ij_{1}})a_{ij_{2}}=\mathbf{1}\{j_{1}=j_{2}\}$ for all $j_{2}\in[|A_{i}|]$ if and only if $b(\text{inv-row}_{ij_{1}})$ is the transpose of row $b_{ij_{1}}$ of $(M^{i;L})^{-1}$. If the $\text{inv-row}_{ij}$ for $1\leq j\leq|A_{i}|$ exist, then we can form the inverse of $M^{i;L}$ by setting row $b_{ij}$ of the inverse to be $b(\text{inv-row}_{ij})^{T}$ for $1\leq j\leq|A_{i}|$ to show that $L$ is invertible. ∎
Lemma 4.11.
The conditions (a) and (f) are equivalent. If either is satisfied, $a(\text{inv-col}_{ij})$ is column $a_{ij}$ of $(M^{i;L})^{-1}$ for $i\geq 0$ and $1\leq j\leq|A_{i}|$.
Proof.
We can proceed analogously as in the proof of the previous result. ∎
The results are more straightforward in the case that $V=R$, in which case we obtain the following direct implication.
Corollary 4.12.
Assume that $V=R$ and $A_{i}=B_{i}$ for all $i\geq 0$. Then, (a) is equivalent to (d’), (e’), and (f’).
  1. (d’)
    For all $i\geq 0$, the equation $\sum_{s\in B_{i}}sL(s)f=0$ has no nonzero solutions $f\in R_{i}$.
  2. (e’)
    For all $i\geq 0$, there exist unique $\text{inv-row}(r;A_{i})\in R_{i}$ for $r\in A_{i}$ such that $L(\text{inv-row}(r;A_{i}))s=\mathbf{1}\{r=s\}$ for $r,s\in A_{i}$.
  3. (f’)
    For all $i\geq 0$, there exist unique $\text{inv-col}(r;A_{i})\in R_{i}$ for $r\in A_{i}$ such that L(s)inv-col
    (r;Ai)
[…]
    for $r,s\in A_{i}$.
4.2. Statements (g) and (h)
Lemma 4.13 ($(a)\Leftrightarrow(g)$).
Assume that $L_{2}\in\mathcal{L}$ is invertible. Then, $(L_{1},L_{2})$ is intertwining if and only if $L_{1}\in\mathcal{L}$ is invertible.
Proof.
Assume that $(L_{1},L_{2})$ intertwines with $\mathcal{V}$. Then, by Lemma˜3.5, we have that for $i\geq 0$ and for all $f\in V_{i}$ and $g\in R_{i}$,
(2) $L_{1}(g)\mathcal{V}f=b(g)^{T}M^{i;L_{1}}a(\mathcal{V}_{i})a(f);$
recall that $\mathcal{V}_{i}\triangleq\mathcal{V}|_{V_{i}}$. Since $\mathcal{V}$ is the identity on $V_{0}$,
(3) $\mathcal{V}L_{2}(g)f=b(g)^{T}M^{i;L_{2}}a(f).$
We therefore have that $M^{i;L_{1}}a(\mathcal{V}_{i})=M^{i;L_{2}}$. Then, since $L_{2}$ is invertible, $L_{1}$ is invertible.
Next, assume that $L_{1}$ is invertible. For $i\geq 0$, define $C^{i}\triangleq(M^{i;L_{1}})^{-1}M^{i;L_{2}}$ and $\mathcal{V}_{i}:V_{i}\rightarrow V_{i}$ by
$a(\mathcal{V}_{i})\triangleq C^{i};$
this uniquely defines the degree-preserving operator $\mathcal{V}$. By Corollary˜4.3, $\mathcal{V}$ acts as the identity over $V_{0}$. Then, from (2) and (3), $L_{1}(g)\mathcal{V}f=\mathcal{V}L_{2}(g)f$ whenever $f\in V_{i}$ and $g\in R_{i}$.
Suppose $0\leq i_{1}<i_{2}$. We prove that for all $f\in V_{i_{2}}$ and $g\in R_{i_{1}}$,
$L_{1}(g)\mathcal{V}f=\mathcal{V}L_{2}(g)f.$
to show that $(L_{1},L_{2})$ intertwines with $\mathcal{V}$. Since $\mathcal{V}$ is a linear operator, this is evident when $i_{1}=0$ after applying Lemma˜3.2, so assume that $i_{1}\geq 1$. Note that for all $h\in R_{i_{2}-i_{1}}$,
$L_{1}(h)L_{1}(g)\mathcal{V}f=\mathcal{V}L_{2}(hg)f=L_{1}(h)\mathcal{V}L_{2}(g)f.$
Since this expression is true for all $h$, by (b), we have that $L_{1}(g)\mathcal{V}f=\mathcal{V}L_{2}(g)f$. ∎
Lemma 4.14.
If $L_{2}\in\mathcal{L}$ is invertible and $(L_{1},L_{2})$ intertwines with $\mathcal{V}$ for some $L_{1}\in\mathcal{L}$, then $\mathcal{V}$ is unique and is a bijection from $V$ to $V$.
Proof.
Suppose $i\geq 0$. We have that $M^{i;L_{1}}a(\mathcal{V}_{i})=M^{i;L_{2}}$. Since the matrices $M^{i;L_{1}}$ are invertible for $i\geq 0$ by Lemma˜4.13, $\mathcal{V}$ is unique.
To prove that $\mathcal{V}$ is a bijection, it suffices to prove that it is a bijective endomorphism of $V_{i}$, since it is degree-preserving. However, this is clear, since the action of $\mathcal{V}$ on $V_{i}$ is isomorphic to multiplying by the invertible matrix $a(\mathcal{V}_{i})$. ∎
Lemma 4.15 ($(a)\Leftrightarrow(h)$).
Assume that $K\subset\text{center}(R)$. The homomorphism $L_{1}\in\mathcal{L}$ is invertible if and only if $(L_{1},L_{2})$ intertwines with a unique operator for all $L_{2}\in\mathcal{L}$.
Proof.
If $L_{1}$ is invertible, then from the second part of the proof of Lemma˜4.13, $(L_{1},L_{2})$ is intertwining for all $L_{2}$. The uniqueness of the intertwiner follows from $M^{i;L_{1}}a(\mathcal{V}_{i})=M^{i;L_{2}}$ for $i\geq 0$. For the reverse direction, we can select $L_{2}$ that is invertible and apply Lemma˜4.13; for the existence of an invertible $L_{2}$, see Corollary˜5.7. ∎
Lemma 4.16.
Suppose $L_{2}\in\mathcal{L}$ is invertible. If $L_{1}\in\mathcal{L}$ and $(L_{1},L_{2})$ intertwines with $\mathcal{V}$, then $\mathcal{V}$ is invertible and $(L_{2},L_{1})$ intertwines with $\mathcal{V}^{-1}$.
Proof.
Since $\mathcal{V}$ is invertible by Lemma˜4.14,
$L_{2}(g)\mathcal{V}^{-1}f=\mathcal{V}^{-1}L_{1}(g)\mathcal{V}\circ\mathcal{V}^{-1}f=\mathcal{V}^{-1}L_{1}(g)f$
for all $f,g\in R$. ∎
Remark 4.17.
We do not have $(L_{2},L_{1})$ intertwining with $\mathcal{V}$ implying that $(L_{1},L_{2})$ intertwines with $\mathcal{V}^{-1}$, since the invertibility of $L_{2}$ does not imply the invertibility of $L_{1}$ and $\mathcal{V}$ in this case.
4.3. Statements (i) and (j)
Lemma 4.18 ($(j)\Rightarrow(i)$).
Assume that $\text{ker}(\text{Hom}_{K}(R,K))=\{0\}$. Then, $(j)\Rightarrow(i)$.
Lemma 4.19 ($(a)\Rightarrow(j)$).
If the homomorphism $L\in\mathcal{L}$ is invertible and $\Psi\in\text{Hom}_{K}(R,\\ K)$, then there exists a unique $\Psi$-eigenvector $E$ such that $E^{0}=1$; in particular, after using the notation from (f), for all $i\geq 0$,
$E^{i}=\sum_{j=1}^{|A_{i}|}\Psi(b_{ij})\text{inv-col}_{ij}.$
Proof.
Suppose $i\geq 0$ and $f\in R_{i}$. Observe that we have that
$L(f)E^{i}=\Psi(f)E^{0}=\Psi(f).$
Thus,
(4) $M^{i;L}a(E^{i})=[\Psi(r)]_{r\in B_{i}}^{T}.$
If $M^{i;L}$ is invertible, $a(E^{i})$ is unique, so $E$ is unique. Afterwards, we can use (4) and Lemma˜4.10 to derive the given expression for $E^{i}$.
Next, we must prove that $E$ satisfies the eigenvalue condition. Suppose $a(E^{i})=(M^{i;L})^{-1}[\Psi(r)]_{r\in B_{i}}^{T}$ for all $i\geq 0$. Then, for all $f\in R_{i}$, we have that
$L(f)E^{i}=\Psi(f)$
because $\Psi$ is additive and $K$-linear. To show that $E^{0}=1$, suppose $A_{0}=\{a_{0}\}$ and $B_{0}=\{b_{0}\}$ so that
$a(E_{0})=(M^{0;L})^{-1}\Psi(b_{0})=\frac{b_{0}}{a_{0}b_{0}}=\frac{1}{a_{0}}$
after applying Corollary˜4.3. We prove that for $1\leq i_{1}<i_{2}$ and $f\in R_{i_{1}}$, $L(f)E^{i_{2}}=\Psi(f)E^{i_{2}-i_{1}}$.
Suppose $g\in R_{i_{2}-i_{1}}$. Then, since $\Psi$ is multiplicative and $L(g)$ is $K$-linear,
$L(g)L(f)E^{i_{2}}=\Psi(g)\Psi(f)=L(g)\Psi(f)E^{i_{2}-i_{1}}.$
We therefore have that $L(g)L(f)E^{i_{2}}=L(g)\Psi(f)E^{i_{2}-i_{1}}$ for all $g\in R_{i_{2}-i_{1}}$. By (b), we have that $L(f)E^{i_{2}}=\Psi(f)E^{i_{2}-i_{1}}$. ∎
Lemma 4.20.
Suppose $\mathcal{S}\subset\text{Hom}_{K}(R,K)$ and that there exists a $\Psi$-eigenvector $E_{\Psi}$ such that $E_{\Psi}^{0}=1$ for all $\Psi\in\mathcal{S}$. For $i\geq 0$, any $r\in R_{i}$ such that $L(r)=0$ as an operator over $V_{i}$ must be an element of $\text{ker}(\mathcal{S})$.
Proof.
We have that
$\Psi(r)E_{\Psi}^{0}=L(r)E_{\Psi}^{i}=0$
for all $\Psi\in\mathcal{S}$, so $r\in\ker(\mathcal{S})$. ∎
Corollary 4.21 ($(i)\Rightarrow(a)$).
Suppose $\mathcal{S}\subset\text{Hom}_{K}(R,K)$ and that $\text{ker}(\mathcal{S})=\{0\}$. If there exists a $\Psi$-eigenvector for all $\Psi\in\mathcal{S}$, then $L$ is invertible.
Proof.
For the sake of contradiction, assume that $L$ is not invertible. Then, for some $i\geq 0$, there exists nonzero $r\in R_{i}$ such that $L(r)f=0$ for all $f\in R_{i}$ by (b). This is a contradiction to Lemma˜4.20. ∎
Proposition 4.22 (Singular value decomposition).
Suppose $L\in\mathcal{L}$ is invertible. Suppose $i\geq 0$ and $M^{i;L}$ satisfies
$M^{i;L}=U\Sigma V^{T}$
for $U\in K^{B_{i}\times B_{i}}$, $\Sigma\in K^{B_{i}\times A_{i}}$, and $V\in K^{A_{i}\times A_{i}}$ such that:
  1. (1)
    The matrices $U$ and $V$ are orthogonal.
  2. (2)
    The matrix $\Sigma$ satisfies $\Sigma_{b_{ij_{1}}a_{ij_{2}}}=0$ if $j_{1},j_{2}\in[|A_{i}|]$ are not equal.
Then, $\Sigma_{b_{ij}a_{ij}}\in K^{\times}$ for $j\in[|A|]$ and $(M^{i;L})^{-1}=V\Sigma^{-1}U^{T}$.
For $r\in B_{i}$, let $U_{r}$ denote column $r$ of $U$ and for $s\in A_{i}$, let $V_{s}$ denote column $s$ of $V$. Then, for $\Psi\in\text{Hom}_{K}(R,K)$, the unique $\Psi$-eigenvector $E\in\mathcal{F}(V)$ with $E^{0}=1$ satisfies
$E^{i}=\sum_{j=1}^{|A_{i}|}\Sigma_{b_{ij}a_{ij}}^{-1}\Psi(b^{-1}(U_{b_{ij}}))a^{-1}(V_{a_{ij}}).$
Proof.
Since $M^{i;L}$ is invertible, $\Sigma_{b_{ij}a_{ij}}\in K^{\times}$ for $j\in[|A|]$, and the formula $(M^{i;L})^{-1}=V\Sigma^{-1}U^{T}$ is evident. It is also evident that
$(M^{i;L})^{-1}=\sum_{j=1}^{|A_{i}|}\Sigma_{b_{ij}a_{ij}}^{-1}V_{a_{ij}}U_{b_{ij}}^{T}$
Using (4) gives that
$a(E^{i})=\sum_{j=1}^{|A_{i}|}\Sigma_{b_{ij}a_{ij}}^{-1}V_{a_{ij}}U_{b_{ij}}^{T}[\Psi(r)]^{T}_{r\in B_{i}}.$
Observe that for $1\leq j\leq|A_{i}|$,
$U_{b_{ij}}^{T}[\Psi(r)]^{T}_{r\in B_{i}}=\sum_{r\in B_{i}}U_{rb_{ij}}\Psi(r)=\Psi\left(\sum_{r\in B_{i}}U_{rb_{ij}}r\right)=\Psi(b^{-1}(U_{b_{ij}})),$
which concludes the proof. ∎
Furthermore, we may consider the question of whether a nonzero $\Psi$-eigenvector $E$ exists such that $E^{0}=0$. However, the existence of such an eigenvector implies that $L\in\mathcal{L}$ is not invertible.
Proposition 4.23.
Assume that $\Psi\in\text{Hom}_{K}(R,K)$ and $E$ is a nonzero $\Psi$-eigenvector such that $E^{0}=0$. Then, (a) and (k) are false.
Proof.
It is evident that setting $f=E^{i}$, where $i$ is the smallest positive integer such that $E^{i}\not=0$, is a contradiction to both (b) and (k). ∎
Remark 4.24.
In fact, the negation of (a) implies the negation of (k), see part (A) of Lemma˜4.25.
4.4. Statement (k)
Lemma 4.25.
  1. (A)
    (k) implies (a).
  2. (B)
    Suppose $R$ is right-spanned by $R_{1}$. Then, (a) implies (k).
Proof.
We prove the converse of both statements. Assume that $R_{i}$ is right-spanned by $R_{1}$. We first prove that the negation of (k) implies the negation of (a). Assume that $f\in V$ is nonconstant and $L(g)f=0$ for all $g\in R_{1}$. Suppose $i\geq 1$ and the degree $i$ part of $f$, $f_{i}$, is nonzero. Then, we have that $L(g)f_{i}=0$ for all $g\in R_{1}$. Since $R$ is right-spanned by $R_{1}$, it is then clear that $L(g)f_{i}=0$ for all $g\in R_{i}$, so $M^{i;L}$ is not invertible.
Next, we prove that the negation of (a) implies the negation of (k), without assuming that $R$ is right-spanned by $R_{1}$. Assume that $L$ is not invertible. Then, suppose $i_{\text{min}}$ is the minimal value of $i$ such that $M^{i;L}$ is not invertible. By Corollary˜4.3, $i_{\text{min}}\geq 1$.
Suppose $f\in R_{i_{\text{min}}}$ is nonzero and $L(g)f=0$ for all $g\in V_{i_{\text{min}}}$; $f$ exists since $M^{i_{\text{min}};L}$ is not invertible. Suppose $g\in R_{1}$. Then,
$L(h)L(g)f=0\,\,\,\forall h\in R_{i_{\text{min}}-1}.$
By the minimality of $i_{\text{min}}$, $M^{i_{\text{min}}-1;L}$ is invertible, so we have that $L(g)f=0$. ∎
In the following result, we analyze the structure of non-invertible $L\in\mathcal{L}$ by inspecting the invertibility of each of the matrices $M^{i;L}$ for $i\geq 0$. First, recall the following definition of the left annihilator.
Definition 4.26.
Suppose $S\subset R$. Then, the left annihilator of $S$ is $\text{Ann}_{R}(S)\triangleq\{r\in R:rs=0\,\,\,\forall s\in S\}$.
Theorem 4.27.
Assume that $L\in\mathcal{L}$ is not invertible.
  1. (A)
    Suppose $i_{\text{min}}$ is the minimal value of $i\geq 0$ such that $M^{i;L}$ is not invertible. Then, $i_{\text{min}}\geq 1$ and there exists $f\in R_{i_{\text{min}}}$ such that $L(g)f=0$ for all $g\in R_{1}$.
  2. (B)
    Assume that $R$ is left-spanned by $R_{1}$ and that $\text{Ann}_{R}(R\backslash R_{0})=\{0\}$. Then, $M^{i;L}$ is not invertible for all $i\geq i_{\text{min}}$.
  3. (C)
    Assume that $\text{Ann}_{R}(R\backslash R_{0})=\{0\}$. Then, the number of $i\geq 1$ such that $M^{i;L}$ is not invertible is infinite.
Proof.
(A): By Corollary˜4.3, $i_{\text{min}}\geq 1$. For the existence of $f\in R_{i_{\text{min}}}$ such that $L(g)f=0$ for all $g\in R_{1}$, see the proof of part (A) of Lemma˜4.25.
[…]
5. Representations of invertible graded rings of operators
In this section, the goal is to obtain a bijection between equivalence classes of invertible $L\in\mathcal{L}$ and sequences of invertible matrices. The following lemma computes the value of $L(g)f$ when the degree of $g$ is at most the degree of $f$, assuming that $L\in\mathcal{L}$ is invertible.
Lemma 5.1.
Assume that $L\in\mathcal{L}$ is invertible. Suppose $0\leq i_{1}\leq i_{2}$, $g\in R_{i_{1}}$, and $f\in R_{i_{2}}$. Then,
$a_{i_{2}-i_{1}}(L(g)f)=(M^{i_{2}-i_{1};L})^{-1}[b_{i_{2}}(rg)^{T}]_{r\in B_{i_{2}-i_{1}}}M^{i_{2};L}a_{i_{2}}(f).$
Proof.
We have that
$a_{i_{2}-i_{1}}(L(g)f)$ $=(M^{i_{2}-i_{1};L})^{-1}[L(r)L(g)f]_{r\in B_{i_{2}-i_{1}}}^{T}$
$=(M^{i_{2}-i_{1};L})^{-1}[b_{i_{2}}(rg)^{T}M^{i_{2};L}a_{i_{2}}(f)]_{r\in B_{d_{2}-d_{1}}}^{T}$
$=(M^{i_{2}-i_{1};L})^{-1}[b_{i_{2}}(rg)^{T}]_{r\in B_{i_{2}-i_{1}}}M^{i_{2};L}a_{i_{2}}(f).$
Corollary 5.2.
Suppose $L\in\mathcal{L}$ is invertible. Suppose $0\leq i_{1}\leq i_{2}$ and $g\in R_{i_{1}}$. The homomorphism $L(g):V_{i_{2}}\rightarrow V_{i_{2}-i_{1}}$ is surjective if and only if $\{r\in R_{i_{2}-i_{1}}:rg=0\}=\{0\}$.
Proof.
Since $L$ is invertible, $L(g)$ is surjective if and only if $b_{i_{2}}(rg)^{T}]_{r\in B_{i_{2}-i_{1}}}$ has full row rank by Lemma˜5.1. Afterwards, the result is straightforward to deduce. ∎
Remark 5.3.
If $\{r\in R_{i_{2}-i_{1}}:rg=0\}=\{0\}$, then the previous corollary implies that $\text{dim}(R_{i_{2}-i_{1}})\leq\text{dim}(R_{i_{2}})$, provided that an invertible element of $\mathcal{L}$ exists. However, it is evident that an invertible element of $\mathcal{L}$ exists by Theorem˜5.4, after assuming the conditions stated in the theorem.
Suppose $V\subset W$ and $R\subset S$, where $W\triangleq\oplus_{i\geq 0}W_{i}$ and $S\triangleq\oplus_{i\geq 0}S_{i}$ are defined analogously to $V$ and $R$, respectively. Furthermore, define $\mathcal{T}$ analogously to $\mathcal{L}$ after replacing $(V,R)$ with $(W,S)$. In the remaining portions of this section, we still assume that $\text{dim}(V_{i})=\text{dim}(R_{i})\geq 1$ for all $i\geq 0$, and in addition we assume that $\text{dim}(W_{i})=\text{dim}(S_{i})\geq 1$ for all $i\geq 0$.
Let $C_{i}$ and $D_{i}$ be bases of $W_{i}$ and $S_{i}$, respectively, for $i\geq 0$. Furthermore, define the isomorphisms $c_{i}:W_{i}\rightarrow K^{C_{i}}$ and $d_{i}:S_{i}\rightarrow K^{D_{i}}$ analogously to $a_{i}$ and $b_{i}$, respectively, for $i\geq 0$.
We say that $L\in\mathcal{T}$ is $(R,V)$-closed if $L(g)f\in V$ for all $g\in R$ and $f\in V$. Furthermore, we write $L_{(R,V)}:R\rightarrow\text{End}_{K}(V)$ to denote the $(R,V)$-restriction of an $(R,V)$-closed homomorphism $L$; it is clear that $L_{(R,V)}\in\mathcal{L}$.
Define $\mathcal{I}$ to be the set of $L\in\mathcal{T}$ such that $L$ is $(R,V)$-closed and $L$ and $L_{(R,V)}$ are invertible. Define the equivalence relation $\sim$ over $\mathcal{I}$ such that if $\in\mathcal{I}, ABA\sim B$ if and only if $A_{(R,V)}=B_{(R,V)}$. Define $\mathcal{I}^{\prime}$ to be the set of $L\in\mathcal{T}$ such that $L$ is $(R,V)$-closed and $L_{(R,V)}$ is invertible and define the equivalence relation $\sim$ over $\mathcal{I}^{\prime}$ similarly.
5.1. Main result
We state and prove the main result of this section. One of its direct consequences is the existence of an invertible element of $\mathcal{L}$ given that $\text{dim}(V_{i})=\text{dim}(R_{i})\geq 1$ for all $i\geq 1$, see Corollary˜5.7. However, for this to be true, we must assume that $K\subset\text{center}(R)$ based on Corollary˜4.7, otherwise no elements of $\mathcal{L}$ will be invertible.
Theorem 5.4.
Let $\mathcal{M}$ be the set of sequences $\{M^{i}\}_{i\geq 0}$ of invertible matrices $M^{i}\in K^{B_{i}\times A_{i}}$ for $i\geq 0$ such that $M^{0}=[a_{0}b_{0}]$, where $A_{0}=\{a_{0}\}$ and $B_{0}=\{b_{0}\}$. Furthermore, assume that there exists a subring $T_{i}$ of $S_{i}$ for $i\geq 0$ such that:
  1. (1)
    For all $i\geq 0$, $S_{i}=R_{i}\oplus T_{i}$.
  2. (2)
    For all $i_{1},i_{2}\geq 0$, $r\in R_{i_{1}}$, and $t\in T_{i_{2}}$, $tr\in T_{i_{1}+i_{2}}$.
Also, assume that $K\subset\text{center}(R)$. The function $\Phi:L\mapsto\{M^{i;L_{(R,V)}}\}_{i\geq 0}$ is a bijection from $\mathcal{I}/\sim$ to $\mathcal{M}$.
Proof.
First, observe that if $L\in\mathcal{I}$, then $\Phi(L)\in\mathcal{M}$ after applying Corollary˜4.3. Suppose $m\in\mathcal{M}$ and that $\Phi(L)=m$ for $L\in\mathcal{I}$. From Lemma˜5.1, we can obtain the value of $L(g)f$ for $f\in V$ and $g\in R$, so $L_{(R,V)}$ is fixed. Therefore, $\Phi$ is injective.
To finish the proof, we show that $\Phi$ is surjective. Suppose $m=\{M^{i}\}_{i\geq 0}\in\mathcal{M}$. We must construct $L\in\mathcal{I}$ such that $M^{i;L_{(R,V)}}=M^{i}$ for all $i\geq 0$. Without loss of generality, assume that $A_{i}\subset C_{i}$ and $B_{i}\subset D_{i}$ for $i\geq 0$.
Suppose $i\geq 0$. Assume that $C_{i}\backslash A_{i}=\{c_{ij}\}_{1\leq j\leq|C_{i}|-|A_{i}|}$ and $D_{i}\backslash B_{i}=\{d_{ij}\}_{1\leq j\leq|C_{i}|-|A_{i}|}$. Suppose $N^{i}\in\mathbb{C}^{D_{i}\times C_{i}}$ satisfies $N^{i}_{rv}=M^{i}_{rv}$ for $r\in B_{i}$ and $v\in A_{i}$, $N^{i}_{d_{ij}c_{ij}}=1$ for $j\in[|C_{i}|-|A_{i}|]$, and $N^{i}$ equals zero elsewhere.
Next, we define $L$ given the matrices $N^{i}$. For $0\leq i_{1}\leq i_{2}$, $f\in W_{i_{2}}$, and $g\in S_{i_{1}}$, we define
(5) $c_{i_{2}-i_{1}}(L(g)f)\triangleq(N^{i_{2}-i_{1}})^{-1}[d_{i_{2}}(rg)^{T}]_{r\in D_{i_{2}-i_{1}}}N^{i_{2}}c_{i_{2}}(f),$
which extends to a definition of $L(g)f$ for all $f\in W$ and $g\in S$. We must prove that $L\in\mathcal{I}$ and that $\Phi(L)=m$.
Step 1. For the first step, we show that $L\in\mathcal{T}$. It suffices to prove that $L$ is a $K$-linear ring homomorphism from $S$ to $\text{End}_{K}(W)$, so it suffices to prove that $L(0)=0^{*}_{W}$, $L(1)=1^{*}_{W}$, $L(ks)=kL(s)$ for all $s\in S$ and $k\in K$, and $L(g+h)=L(g)+L(h)$ and $L(gh)=L(g)L(h)$ for $g,h\in S$; note that for $k\in K$, $k^{*}_{W}$ is the element $\{w\mapsto kw\}$ of $\text{End}_{K}(W)$ and is analogous to $k^{*}$.
It is straightforward to deduce that $L(0)=0^{*}_{W}$ and $L(g+h)=L(g)+L(h)$. To show that $L(1)=1^{*}_{W}$, set $i_{1}=0$ and $g=1$ to get that
$c_{i_{2}}(L(1)f)=(N^{i_{2}})^{-1}[d_{i_{2}}(r)^{T}]_{r\in D_{i_{2}}}N^{i_{2}}c_{i_{2}}(f)=c_{i_{2}}(f).$
Furthermore, to show that $L(ks)=kL(s)$ for all $k\in K$ and $s\in S$, we have that
$c_{i_{2}-i_{1}}(L(ks)f)$ $=(N^{i_{2}-i_{1}})^{-1}[d_{i_{2}}(rks)^{T}]_{r\in D_{i_{2}-i_{1}}}N^{i_{2}}c_{i_{2}}(f)$
$=(N^{i_{2}-i_{1}})^{-1}[d_{i_{2}}(krs)^{T}]_{r\in D_{i_{2}-i_{1}}}N^{i_{2}}c_{i_{2}}(f)$
$=c_{i_{2}-i_{1}}(kL(s)f),$
where we have used $K\subset\text{center}(R)$. Next, we prove that $L(gh)=L(g)L(h)$.
Suppose $0\leq i_{1},i_{2}\leq i_{3}$ and $i_{1}+i_{2}\leq i_{3}$. We prove that for $f\in W_{i_{3}}$, $g_{1}\in S_{i_{1}}$, and $g_{2}\in S_{i_{2}}$,
$L(g_{2}g_{1})f=L(g_{2})L(g_{1})f.$
Observe that, based on (5),
$c_{i_{3}-i_{1}-i_{2}}(L(g_{2})L(g_{1})f)=$ $(N^{i_{3}-i_{1}-i_{2}})^{-1}[d_{i_{3}-i_{1}}(r_{2}g_{2})^{T}]_{r_{2}\in D_{i_{3}-i_{1}-i_{2}}}N^{i_{3}-i_{1}}$
$(N^{i_{3}-i_{1}})^{-1}[d_{i_{3}}(r_{1}g_{1})^{T}]_{r_{1}\in D_{i_{3}-i_{1}}}N^{i_{3}}c_{i_{3}}(f)$
$=$ $(N^{i_{3}-i_{1}-i_{2}})^{-1}[d_{i_{3}-i_{1}}(r_{2}g_{2})^{T}]_{r_{2}\in D_{i_{3}-i_{1}-i_{2}}}[d_{i_{3}}(r_{1}g_{1})^{T}]_{r_{1}\in D_{i_{3}-i_{1}}}N^{i_{3}}$
$c_{i_{3}}(f)$
and
$c_{i_{3}-i_{1}-i_{2}}(L(g_{2}g_{1})f)=(N^{i_{3}-i_{1}-i_{2}})^{-1}[d_{i_{3}}(rg_{2}g_{1})^{T}]_{r\in D_{i_{3}-i_{1}-i_{2}}}N^{i_{3}}c_{i_{3}}(f).$
Hence, it suffices to prove that
$d_{i_{3}-i_{1}}(r_{2}g_{2})^{T}]_{r_{2}\in D_{i_{3}-i_{1}-i_{2}}}[d_{i_{3}}(r_{1}g_{1})^{T}]_{r_{1}\in D_{i_{3}-i_{1}}}=[d_{i_{3}}(rg_{2}g_{1})^{T}]_{r\in D_{i_{3}-i_{1}-i_{2}}}.$
Note that for $x\in D_{i_{3}-i_{1}-i_{2}}$ and $y\in D_{i_{3}}$, entry $(x,y)$ of the left hand side is
$\sum_{z\in D_{i_{3}-i_{1}}}d_{i_{3}-i_{1}}(xg_{2})_{z}d_{i_{3}}(zg_{1})_{y}.$
We have that
$\sum_{y\in D_{i_{3}}}\sum_{z\in D_{i_{3}-i_{1}}}d_{i_{3}-i_{1}}(xg_{2})_{z}d_{i_{3}}(zg_{1})_{y}y=\sum_{z\in D_{i_{3}-i_{1}}}d_{i_{3}-i_{1}}(xg_{2})_{z}zg_{1}=xg_{2}g_{1},$
so row $x$ of the left hand side matches row $x$ of the right hand side.
Step 2. To finish the proof, we must verify that $L\in\mathcal{I}$. First, it is clear that $L$ is invertible and from inspecting the matrices $N^{i}$ for $i\geq 0$, if $L$ is $(R,V)$-closed, then $L_{(R,V)}$ is invertible. Hence, it suffices to prove that $L$ is $(R,V)$-closed to complete the proof.
Assume that $0\leq i_{1}\leq i_{2}$, $f\in V_{i_{2}}$ and $g\in R_{i_{1}}$. We must prove that $c_{i_{2}-i_{1}}(L(g)f)_{c_{i_{2}-i_{1},j}}=0$ for $1\leq j\leq|C_{i_{2}-i_{1}}|-|A_{i_{2}-i_{1}}|$. Using (5), we have that
$c_{i_{2}-i_{1}}(L(g)f)_{c_{i_{2}-i_{1},j}}=\sum_{\begin{subarray}{c}x\in D_{i_{2}-i_{1}},\,y\in D_{i_{2}},\\ z\in C_{i_{2}}\end{subarray}}(N^{i_{2}-i_{1}})^{-1}_{c_{i_{2}-i_{1},j}x}d_{i_{2}}(xg)_{y}N^{i_{2}}_{yz}c_{i_{2}}(f)_{z}.$
Since $f\in V_{i_{2}}$, $c_{i_{2}}(f)_{z}=0$ for $z\in C_{i_{2}}\backslash A_{i_{2}}$, so we may assume that $z\in A_{i_{2}}$. We may also assume that $y\in B_{i_{2}}$ so that $N_{yz}^{i_{2}}$ is nonzero. Since $xg\in T_{i_{2}}$ for all $x\in D_{i_{2}-i_{1}}\backslash B_{i_{2}-i_{1}}$, in order for $d_{i_{3}}(xg)_{y}$ to be nonzero, we may assume that $x\in B_{i_{2}-i_{1}}$. However, it is then clear that $(N^{i_{2}-i_{1}})^{-1}_{c_{i_{2}-i_{1},j}x}=0$. Hence, $c_{i_{2}-i_{1}}(L(g)f)_{c_{i_{2}-i_{1},j}}=0$, which finishes the proof. ∎
Example 5.5.
We give examples of tuples $(S_{i},R_{i},T_{i})$ for $i\geq 0$ that satisfy conditions (1) and (2) of theorem˜5.4. The first example is to set $S_{i}=R_{i}$ and $T_{i}=\{0\}$ for all $i\geq 0$. Another example is to set $S_{i}$ to be the ring of homogeneous polynomials in $K[x_{1},\ldots,x_{N}]$ of degree $i$, $R_{i}$ to be the ring of homogeneous polynomials in $K[x_{1},\ldots,x_{N-1}]$ of degree $i$, $T_{0}$ to be $\{0\}$, and $T_{i}$ to be $x_{N}S_{i-1}$ for all $i\geq 1$.
Corollary 5.6.
Assume that conditions (1) and (2) from Theorem˜5.4 are true and that $K\subset\text{center}(R)$. The function $\Phi$ is a bijection from $\mathcal{I}^{\prime}/\sim$ to $\mathcal{M}$.
Proof.
The proof of injectivity is the same as the proof of injectivity given in the proof of theorem˜5.4. On the other hand, surjectivity follows from Theorem˜5.4. ∎
Corollary 5.7.
Assume that $K\subset\text{center}(R)$. The function $\alpha:L\mapsto\{M^{i;L}\}_{i\geq 0}$ is a bijection from the set of invertible elements of $\mathcal{L}$ to $\mathcal{M}$.
Proof.
This follows from Theorem˜5.4 after setting $(S,W)=(R,V)$. ∎
Corollary 5.8.
Assume that conditions (1) and (2) from Theorem˜5.4 are true and that $K\subset\text{center}(R)$. The function $\beta:L\mapsto L_{(R,V)}$ is a bijection from $\mathcal{I}/\sim$ to the set of invertible elements of $\mathcal{L}$ and from $\mathcal{I}^{\prime}/\sim$ to the set of invertible elements of $\mathcal{L}$.
Proof.
The result follows from Theorem˜5.4 and Corollaries˜5.6 and 5.7 after observing that $\beta=\alpha^{-1}\circ\Phi$. ∎
5.2. Equivariance with respect to a group action and examples
Let $H$ be a finite group which acts on $W$ and $S$ such that it acts as the identity over $W_{0}=K$ and $S_{0}=K$. Assume that for $h\in H$ and $g_{1},g_{2}\in S$, $hg_{1}g_{2}=(hg_{1})(hg_{2})$. Then, let $\mathcal{H}$ be the set of $L\in\mathcal{T}$ such that for all $f\in W$, $g\in S$, and $h\in H$,
(6) $L(hg)hf=hL(g)f.$
Let $V$ and $R$ be the set of $w\in W$ such that $hw=w$ and $s\in S$ such that $hs=s$, respectively, for all $h\in H$. In particular, for all $i\geq 0$, $V_{i}$ and $R_{i}$ are the sets of $w\in W_{i}$ and $s\in S_{i}$, respectively, such that $hw=w$ and $hs=s$ for all $h\in H$. We have that $R\triangleq\oplus_{i\geq 0}R_{i}$ forms a graded ring since $hg_{1}g_{2}=(hg_{1})(hg_{2})$ for all $g_{1},g_{2}\in S$.
Lemma 5.9.
If $L\in\mathcal{H}$, then $L$ is $(R,V)$-closed.
Proof.
Suppose $f\in V$, $g\in R$, and $h\in H$. It suffices to prove that
$hL(g)f=L(g)f.$
By (6),
$hL(g)f=L(hg)hf=L(g)f,$
which finishes the proof. ∎
Lemma 5.10.
Assume that $\text{char}(K)$ does not divide $|H|$ and that $L\in\mathcal{H}$ is invertible. Then, $L_{(R,V)}$ is invertible.
Proof.
For the sake of contradiction, assume that $L_{(R,V)}$ is not invertible. Suppose $i\geq 1$ and $g\in R_{i}$ satisfies $L(g)f=0$ for all $f\in V_{i}$. Suppose $f\in W_{i}$. Then, for all $h\in H$,
$L(g)hf=L(hg)hf=hL(g)f=L(g)f$
by (6), since $L(g)f\in K$ and $H$ acts as the identity over $K$. Thus, we have that
$L(g)f=\frac{1}{|H|}\sum_{h\in H}L(g)hf=L(g)\left(\frac{1}{|H|}\sum_{h\in H}hf\right)=0,$
since $|H|$ is not divisible by $\text{char}(K)$. It follows that $L$ is not invertible, which is a contradiction. ∎
It does not seem to be generally true that the invertibility of $L_{(R,V)}$ implies the invertibility of $L$. However, in the paper [dunkl_singular_poly], this is shown to be true for Dunkl operators. We repeat the argument from Remark 4.2 of that paper in a more general setting to prove the following result.
Lemma 5.11.
Assume that $S$ is an integral domain. If $L\in\mathcal{L}$ is $(R,V)$-closed and $L_{(R,V)}$ is invertible, then $L$ is invertible.
Proof.
For the sake of contradiction, assume that $L$ is not invertible. Suppose $i\geq 1$ and $g\in S_{i}$ is nonzero and satisfies $L(g)f=0$ for all $f\in W_{i}$. Let
$g^{\prime}=\prod_{h^{\prime}\in H}h^{\prime}g.$
For all $h\in H$, $hg^{\prime}=\prod_{h^{\prime}\in H}hh^{\prime}g=g^{\prime}$ by the commutativity of $S$, so $g^{\prime}\in R_{M}$, where $M$ is the positive integer such that $g^{\prime}\in S_{M}$.
Assume that $h^{\prime}\in H$ and $h^{\prime}g=0$. Then, $(h^{\prime})^{-1}h^{\prime}g=0$, which is a contradiction to $g\not=0$. Hence, $h^{\prime}g\not=0$. It follows that $g^{\prime}\not=0$ because $S$ is an integral domain.
Suppose $f\in V_{M}$. Then, it is evident that $L(g^{\prime})f=0$ after applying $L(h^{\prime}g)$ for some sequence of |H|1|H|-1 distinct $h^{\prime}\in H\backslash\{1\}$. Thus, $L_{(R,V)}$ is not invertible, which is a contradiction. ∎
Remark 5.12.
In the previous result, we do not assume that $hS_{i}\subset S_{i}$ for all $h\in H$ and $i\geq 1$. In particular, we do not necessarily require that $M=|H|i$, although $M\geq 1$ because $g\in S_{i}$ and $i\geq 1$.
Corollary 5.13.
Assume that $S$ is an integral domain. Then, $\mathcal{I}=\mathcal{I}^{\prime}$.
Example 5.14.
We consider a more specific example. As an introduction, consider the following basic result. We omit the proof since it is straightforward.
Lemma 5.15.
Suppose $U$ is a vector space over $K$. Assume that $L:K[x_{1},\ldots,x_{N}]\rightarrow\text{End}_{K}(U)$ is a $K$-linear ring homomorphism. Then, the operators $L(x_{i})$ for $1\leq i\leq N$ are commutative. Conversely, for commutative operators $L_{i}\in\text{End}_{K}(U)$ for $1\leq i\leq N$, there exists a unique $K$-linear ring homomorphism $L:K[x_{1},\ldots,x_{N}]\rightarrow\text{End}_{K}(U)$ such that $L(x_{i})=L_{i}$ for $1\leq i\leq N$.
Suppose $N\geq 1$, $S=W=K[x_{1},\ldots,x_{N}]$, and that $S_{i}=W_{i}$ is the group of polynomials over $K$ that are homogeneous of degree $i$ for $i\geq 0$. From Lemma˜5.15, any $L\in\mathcal{T}$ is uniquely determined by $L(x_{i})$ for $1\leq i\leq N$. Let $H$ be a finite group which acts on $K^{N}$ such that for all $h\in H$ and $f\in S_{1}$, $\{x\mapsto f(hx)\}\in S_{1}$.
Lemma 5.16.
Assume that $h\in H$. Then, for $i\geq 0$ and $f\in S_{i}$, $\{x\mapsto f(hx)\}\in S_{i}$.
Proof.
This is clear since $(hx)_{i}\in S_{1}$ for all $i\in[N]$. ∎
Hence, we can define the action of $H$ over $S$ to be $hf(x)\triangleq f(hx)$; indeed, the action of $H$ over $K$ is the identity and $hf_{1}(x)f_{2}(x)=(hf_{1}(x))(hf_{2}(x))$ for $f_{1},f_{2}\in S$. Then, $R=V$ is the set of $f\in S$ such that $f(hx)=f(x)$ for all $h\in H$ and $\mathcal{H}$ is the set of $L\in\mathcal{T}$ such that
(7) $L(g(hx))f(hx)=(L(g)f)(hx).$
for all $f,g\in S$ and $h\in H$. In contrast to the proof of Theorem˜5.4, we do not assume that $B_{i}\subset D_{i}$ for $i\geq 0$.
For $x\in K^{N}$, define $\varphi_{x}:S\rightarrow K,f\mapsto f(x)$. We revisit the conditions that we introduced earlier:
  • Both $\text{ker}(\text{Hom}_{K}(R,K))$ and $\text{ker}(\text{Hom}_{K}(S,K))$ only contain 0, since we can set the homomorphism to be $\varphi_{x}$ for a fixed value of $x\in K^{N}$. See Theorem˜4.5 for the relevant implications.
  • The ring $S$ is left- and right-spanned by $S_{1}$, but the analogous condition for $R$ is not true. See Theorems˜4.5 and 4.27 for the relevant implications.
  • The left annihilators $\text{Ann}_{R}(R\backslash R_{0})$ and $\text{Ann}_{S}(S\backslash S_{0})$ only contain 0. See Theorem˜4.27 for the relevant implications.
In the following result, we discuss the $\varphi_{x}$-eigenvectors. The result generalizes Theorems˜2.5 and 2.8 since $L$ can be any invertible operator.
Corollary 5.17.
Assume that $L\in\mathcal{T}$ is invertible (resp. $L\in\mathcal{L}$ is invertible). Furthermore, suppose $x\in K^{N}$.
  1. (A)
    There does not exist a nonzero solution $E\in\mathcal{F}(S)$ (resp. $E\in\mathcal{F}(R)$) to $L(g)E=g(x)E$ for all $g\in S$ (resp. $g\in R$) such that $E^{0}=0$.
  2. (B)
    There exists a unique solution $E\in\mathcal{F}(S)$ (resp. $E\in\mathcal{F}(R)$) to $L(g)E=g(x)E$ for all $g\in S$ (resp. $g\in R$) such that $E^{0}=1$. Using the notation of Corollary˜4.12, the solution is
    $E=\sum_{r\in D_{i}}\text{inv-col}(r;D_{i})r(x)\,\,\,\left(resp.\sum_{r\in B_{i}}\text{inv-col}(r;B_{i})r(x)\right).$
Proof.
Observe that $L_{(R,V)}$ is invertible by Lemma˜5.10. For (A), see Proposition˜4.23. For (B), see condition (i) of Theorem˜4.5 with $\Psi=\varphi_{x}$ and Lemma˜4.19. ∎
Corollary 5.18.
Assume that $L\in\mathcal{H}$ is invertible and that $\text{char}(K)$ does not divide $|H|$. Then, all conclusions of Corollary˜5.17 are true (this includes the conclusions over both $(R,V)$ and $(S,W)$).
Proof.
This follows from Lemmas˜5.9 and 5.10. ∎
Example 5.19.
We describe the framework that we have developed in the context of Dunkl operators and prove a few well-known results. Consider the setting of Example˜5.14, but set $K=\mathbb{C}$ and $H$ to be $H(\mathcal{R})$, where $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system.
Suppose $u_{i}\in\mathbb{R}^{N}$ for $1\leq i\leq N$. Then, we can define $L\in\mathcal{T}$ by $L(x_{i})\triangleq\mathcal{D}_{u_{i}}(\mathcal{R}(\theta))$ after using Lemma˜5.15 and the commutativity of the Dunkl operators. As a special case, we have that $\mathcal{D}(\mathcal{R}(\theta))\in\mathcal{T}$, where we recall that $\mathcal{D}(\mathcal{R}(\theta))(x_{i})=\mathcal{D}_{i}(\mathcal{R}(\theta))$ for $i\in[N]$. For $i\geq 0$, the matrix $M^{i;\,\mathcal{D}(\mathcal{R}(\theta))}$ has entry $(r,s)$ equal to $r,s]_{\mathcal{R}(\theta)}$ for $r\in D_{i}$ and $s\in C_{i}$.
By Lemma˜2.2, it is evident that $\mathcal{D}(\mathcal{R}(\theta))\in\mathcal{H}$ since (7) is satisfied. Furthermore, $V=R=\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$ and the operator $\mathcal{D}(\mathcal{R}(\theta))_{(R,V)}\in\mathcal{L}$ is equivalent to $\mathcal{D}_{H}(\mathcal{R}(\theta))$. By Lemmas˜5.10 and 5.11, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible if and only if $\mathcal{D}_{H}(\mathcal{R}(\theta))$ is invertible.
Observe that $E_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}(\mathcal{R}(\theta))$ and $J_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}_{H}(\mathcal{R}(\theta))$ for $\Psi:r\mapsto r(a)$, where $a\in\mathbb{C}^{N}$ is fixed. Then, part (B) of Corollary˜5.17 implies that these functions are unique eigenvectors, which is included in the statements of Theorems˜2.5 and 2.8.
We do not consider intertwining operators in the Dunkl setting in this paper. However, we mention that the intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$. As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. As explained in [rosler_positivity], in terms of Theorem˜1.2, the intertwiner is given by
$Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\mu_{a}(\epsilon)$
for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $a\in\mathbb{R}^{N}$.
For the remainder of this example, we consider the setting of Proposition˜4.22. Suppose $i\geq 0$. After assuming that $A_{i}=B_{i}$, we are particularly interested about the case where a singular value decomposition with $U=V$ exits. It would follow that
$E^{i}=\sum_{r\in A_{i}}\sigma(r)^{-1}\Psi(u_{r})u_{r},$
where $\sigma(r)$ is entry $(r,r)$ of $\Sigma$ and $u_{r}$ corresponds to column $r$ of $U$ for $r\in A_{i}$.
The paper [okounkov_olshanki_2] discusses the singular value decomposition of the matrices $M^{i;\,\mathcal{D}_{H}(A^{N-1}(\theta))}$ for $i\geq 0$ and $\theta\geq 0$. The paper [baker] generalizes this result and shows that such a decomposition exists for the matrices $M^{i;\,\mathcal{D}(A^{N-1}(\theta))}$ for $i\geq 0$ and $\theta\geq 0$. Furthermore, the paper [baker2] discusses the singular value decomposition of the matrices $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$ and $\theta_{0},\theta_{1}\geq 0$. For each $i\geq 0$, in the first and third cases, $U=V$ correspond to symmetric Jack functions while in the second case, $U=V$ correspond to nonsymmetric Jack functions.
For $\theta\geq 0$ and even $i\geq 0$, it is straightforward to compute the decomposition of $M^{i;\,\mathcal{D}_{H}(D^{N}(\theta))}$ when we restrict the rows and columns to $p_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i]$ by using the decomposition of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,0))}$. An interesting question is, what is the decomposition when we restrict the rows and columns to $ep_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i-N]$, where $i\geq N$ has the same parity as $N$? Furthermore, another interesting question which generalizes the previous question is, what are the decompositions of $M^{i;\,\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$? Recall that $M^{i;\,\mathcal{D}(D^{N}(\theta))}=M^{i;\,\mathcal{D}(BC^{N}(\theta,0))}$.
6. Leading order terms for the type A Dunkl bilinear form
Based on (4), the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma}$ is essential for calculating the coefficients of $J_{a}^{A^{N-1}(\theta)}(x)$. For the continuation of this direction, see Section˜9. The following result computes the leading order terms of the matrix in the regime $|\theta N|\rightarrow\infty$.
Theorem 6.1.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}=\theta^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}N^{k+\ell(\lambda)-\ell(\nu)}+R(N,\theta).$
where $R\in\mathbb{Q}[x,y]$ satisfies:
  1. (1)
    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
  2. (2)
    The degree $x$ is at most $k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $k-\ell(\nu)$.
Remark 6.2.
In the previous result, we do not require that $\lambda,\nu\in\Gamma_{N}$.
First, we present a short proof of the theorem using the results of [limits_prob_measures]. The proof lacks some details, but it includes the main ideas.
First proof.
We deduce the result using [limits_prob_measures]*Theorem 6.13. First, replace $G_{\theta_{N}}(x_{1},\ldots,\\ x_{N};\,\mu_{N})$ with $\exp\left(\theta N\sum_{k\geq 1}c_{k}p_{k}\right)$, where $c_{k}\in\mathbb{C}$ is fixed for all $k\geq 1$. Then, the theorem implies that for all $\lambda\in\Gamma$,
$\lim_{N\rightarrow\infty}\frac{[1]\mathcal{D}(p_{\lambda})\exp\left(\theta N\sum_{k\geq 1}c_{k}p_{(k)}\right)}{(\theta N)^{|\lambda|}N^{\ell(\lambda)}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}|B|c_{|B|}.$
For $\nu\in\Gamma$, the coefficient of $p_{\nu}$ in $\exp\left(\theta N\sum_{k\geq 1}c_{k}p_{(k)}\right)$ is $\frac{(\theta N)^{\ell(\nu)}}{\pi(\nu)}\prod_{i=1}^{\ell(\nu)}c_{\nu_{i}}$. Hence, the coefficient of $\prod_{i=1}^{\ell(\nu)}c_{\nu_{i}}$ in $1]\mathcal{D}(p_{\lambda})\exp\left(\theta N\sum_{k\geq 1}c_{k}p_{(k)}\right)$ is $\frac{(\theta N)^{\ell(\nu)}}{\pi(\nu)}[p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$. By finding the coefficient of $\prod_{i=1}^{\ell(\nu)}c_{\nu_{i}}$ on the right hand side, we obtain the formula in the statement of the theorem. ∎
However, while this proof is short, we cannot use it while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems. Therefore, we outline a new proof of Theorem˜6.1 that can be easily applied to a more general setting.
6.1. Second proof of Theorem˜6.1
We have that
(8) $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}=\mathcal{D}(p_{\lambda})p_{\nu}=\prod_{i=1}^{\ell(\lambda)}\left(\sum_{j=1}^{N}\mathcal{D}_{j}^{\lambda_{i}-1}\partial_{j}\right)p_{\nu}=\sum_{j_{1},\ldots,j_{\ell(\lambda)}=1}^{N}\prod_{i=1}^{\ell(\lambda)}(\mathcal{D}_{j_{i}}-\partial_{j_{i}}+\partial_{j_{i}})^{\lambda_{i}-1}\partial_{j_{i}}p_{\nu}.$
Suppose $j_{1},\ldots,j_{\ell(\lambda)}\in[N]$ and that for each $\mathcal{D}_{j_{i}}$ operator, we select either $\mathcal{D}_{j_{i}}-\partial_{j_{i}}$ (which only consists of switches) or $\partial_{j_{i}}$; the total number of such choices that we make is $k-\ell(\lambda)$. Let $d$ be the number of times we choose $\partial_{j_{i}}$.
Assume that $d<\ell(\nu)-\ell(\lambda)$. Then, the total number of derivatives is less than $\ell(\nu)$. Recall that when applying a derivative to a product of terms, we select one of the terms and apply the derivative to it, and we iterate over all such selections. Because the number of derivatives is less than $\ell(\nu)$, for every path, we will never apply a derivative to $p_{\nu_{l}}$ for some $l\in[\ell(\nu)]$. Then, we have that we will not be able to eliminate the symmetric term $p_{\nu_{l}}$ without eliminating the entire expression. This is straightforward to deduce, since we must eliminate the symmetric term using a switch. Thus, the total contribution in this case is zero.
Next, assume that $d\geq\ell(\nu)-\ell(\lambda)$. The total number of switches is $k-\ell(\lambda)-d$. The number of ways to select these switches is therefore $(N-1)^{k-\ell(\lambda)-d}$. Furthermore, the number of ways to select $j_{1},\ldots,j_{\ell(\lambda)}$ is $N^{\ell(\lambda)}$. Note that for each selection of switches and $j_{1},\ldots,j_{\ell(\lambda)}$, any term of $p_{\nu}$ which contains $x_{i}$ for some $i$ that is not included in a switch or $\{j_{1},\ldots,j_{\ell(\lambda)}\}$ will have a contribution of zero. Hence, the sums of the coefficients of the terms of $p_{\nu}$ which may have nonzero contribution is at most $(k-d)^{\ell(\nu)}\leq k^{\ell(\nu)}$. Furthermore, for each sequence of switches and derivatives, the derivatives will increase the absolute value of the coefficient by at most $k^{\ell(\lambda)+d}$ while the switches will increase it by at most $(k|\theta|)^{k-\ell(\lambda)-d}$. Hence, for each selection of the switches and $j_{1},\ldots,j_{\ell(\lambda)}$, the magnitude of the total contribution is at most $k^{k+\ell(\nu)}|\theta|^{k-\ell(\lambda)-d}$. Then, the magnitude of the total contribution is less than
$k^{k+\ell(\nu)}|\theta|^{k-\ell(\lambda)-d}N^{k-d}.$
To achieve the leading order term, we must have that $d=\ell(\nu)-\ell(\lambda)$. Furthermore, each of the indices $j_{1},\ldots,j_{\ell(\lambda)}$ must be distinct, and each switch must be from some $j_{i}$ to a distinct index which does not appear in $\{j_{1},\ldots,j_{\ell(\lambda)}\}$. Then, the leading order term is $N^{k-\ell(\nu)+\ell(\lambda)}$ multiplied by a constant that depends on $\theta$.
It is not challenging to determine that the remainder term $R$ is a polynomial with rational coefficients that satisfies conditions (1) and (2). To see that $R$ is a polynomial with rational coefficients, we perform case work on the choices of $j_{1},\ldots,j_{\ell(\lambda)}$ and the switches. For condition (1), we note that after selecting $j_{1},\ldots,j_{\ell(\lambda)}$, the only way to obtain a factor of $N$ is with a switch, which will in turn add a factor of $\theta$. For condition (2), we note that we have separated the leading order term in $x$, and that $d\geq\ell(\nu)-\ell(\lambda)$ so the number of switches is at most $k-\ell(\nu)$ to obtain the maximal degree in $y$. For more details of a similar argument, see [limits_prob_measures]*Section 6.
The final step is to compute the coefficient of $N^{k+\ell(\lambda)-\ell(\nu)}$. Recall that the indices $j_{1},\ldots,j_{\ell(\lambda)}$ are distinct and that the switches are from some $j_{i}$ to a distinct index not in $\{j_{1},\ldots,j_{\ell(\lambda)}\}$. The total number of choices for the indices is therefore
$N(N-1)\cdots(N-k-\ell(\lambda)+\ell(\nu)+1)=(1+o_{N}(1))N^{k+\ell(\lambda)-\ell(\nu)},$
so the coefficient of $N^{k+\ell(\lambda)-\ell(\nu)}$ is the sum of the contributions of the sequences after we fix the indices.
In particular, let $\mathcal{S}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-\ell(\nu)$. Note that the operators $s_{i}$ are ordered from $i=1$ to $|\lambda|$, meaning that $s_{1}$ appears first and $s_{|\lambda|}$ appears last.
Then, the coefficient of $N^{k+\ell(\lambda)-\ell(\nu)}$ is
(9) $\sum_{s\in\mathcal{S}}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu}.$
Recall that in order to have a nonzero contribution, we must apply at least one derivative to each of the $p_{\nu_{l}}$. In this case, the number of derivatives is exactly $\ell(\nu)$, so we apply exactly one derivative to each of the $p_{\nu_{l}}$.
Furthermore, note that when we apply the $j$th switch from $i$ to $\ell(\lambda)+j$, $x_{\ell(\lambda)+j}$ does not appear before we apply the switch, other than among the $p_{\nu_{l}}$ that have yet to be assigned to. Also, any terms that contain $x_{\ell(\lambda)+j}$ after applying the switch will have a contribution of zero, because none of the remaining derivatives or switches will be able to remove the variable $x_{\ell(\lambda)+j}$. Note that applying the switch from $i$ to $\ell(\lambda)+j$ to $x_{i}^{e}$ outputs $\theta(x_{i}^{e-1}+x_{i}^{e-2}x_{\ell(\lambda)+j}+\cdots+x_{\ell(\lambda)+j}^{e-1})$, and the only resulting term that does not contain $x_{\ell(\lambda)+j}$ is $\theta x_{i}^{e-1}$. It follows that the switch from $i$ to $\ell(\lambda)+j$ is equivalent to $\theta d_{i}$.
Remark 6.3.
The derivatives multiply the leading order coefficient by $\prod_{l=1}^{\ell(\nu)}\nu_{l}$ and the factors of $\theta$ from the switches multiply the coefficient by $\theta^{k-\ell(\nu)}$. These expressions appear in the statement of Theorem˜6.1.
To compute (9), we must choose the locations of the $\ell(\nu)-\ell(\lambda)$ unallocated derivatives and then assign each of the $\ell(\nu)$ derivatives to a distinct symmetric term $p_{\nu_{l}}$. In particular, for $s\in\mathcal{S}$, define $\mathcal{P}(s)\triangleq\{i\in[|\lambda|]:s_{i}\text{ is a derivative}\}$ and let $\mathcal{H}(s)$ denote the set of bijections $\zeta:\mathcal{P}(s)\rightarrow[\ell(\nu)]$. Then for $s\in\mathcal{S}$ and $\zeta\in\mathcal{H}(s)$, the pair $(s,\zeta)$ corresponds to applying the sequence $s$ of operators and for $i\in\mathcal{P}(s)$, allocating the derivative $s_{i}$ to $p_{\nu_{\zeta(i)}}$.
Consider the following process to compute the contribution after choosing $s$ and $\zeta$. We start with the coefficient $c=1$ and iterate over the operators of $s$ from $s_{1}$ to $s_{|\lambda|}$. If we apply $\partial_{j_{i}}=\partial_{i}$ to $p_{\nu_{l}}$ at a step, where $l$ is determined by $\zeta$, then we multiply $c$ by $\partial_{i}p_{\nu_{l}}$. Otherwise, if we apply the $j$th switch from $i$ to $\ell(\lambda)+j$, then we apply $\theta d_{i}$ to $c$. Based on the previous discussion, the final value of $c$ will be the contribution from $(s,\zeta)$, and we denote this value of $c$ by $c(s,\zeta)$. We then have that
(10) $\sum_{s\in\mathcal{S}}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu}=\sum_{s\in\mathcal{S},\,\zeta\in\mathcal{H}(s)}c(s,\zeta).$
Suppose $s\in\mathcal{S}$, $\zeta\in\mathcal{H}(s)$, and $i\in[\ell(\lambda)]$. Define
$S_{i}(s,\zeta)\triangleq\{\zeta(j):j\in[1+\lambda_{1}+\cdots+\lambda_{i-1},\lambda_{1}+\cdots+\lambda_{i}]\cap\mathcal{P}(s)\};$
that is, $S_{i}$ is the set of $l$ such that $p_{\nu_{l}}$ is assigned to by some $\partial_{i}$.
Lemma 6.4.
Suppose $s\in\mathcal{S}$, $\zeta\in\mathcal{H}(s)$, and $c(s,\zeta)\not=0$. Then
$\sum_{l\in S_{i}(s,\zeta)}|\nu_{l}|=\lambda_{i}$
for all $i\in[\ell(\lambda)]$.
Proof.
Consider $i=1$. If $\sum_{l\in S_{1}(s,\zeta)}|\nu_{l}|<\lambda_{1}$, then $c(s,\zeta)=0$ will be zero. This is because among the first $\lambda_{1}$ operators, we apply $\lambda_{1}-|S_{1}(s,\zeta)|$ switches to $c$ and we multiply $c$ by the polynomials $\partial_{1}p_{\nu_{l}}$ for each $l\in S_{1}(s,\zeta)$. Since we start at $c=1$ and the total degree of the multiplied polynomials is less than the number of switches, $c(s,\zeta)=0$. Furthermore, if $\sum_{l\in S_{1}(s,\zeta)}|\nu_{l}|>\lambda_{1}$, then $c(s,\zeta)=0$. This is because after the first $\lambda$ operators, $c(s,\zeta)$ will be a multiple of $x_{1}$. We will not be able to remove the factor of $x_{1}$ using later operators, so the final contribution will be zero. Therefore, for $c(s,\zeta)$ to be nonzero, $\sum_{l\in S_{1}(s,\zeta)}|\nu_{l}|=\lambda_{1}$. In particular, after applying the first $\lambda_{1}$ operators, $c$ will be a constant.
We can apply the same argument using induction to deduce that $\sum_{l\in S_{i}(s,\zeta)}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. ∎
Let $\mathcal{G}$ denote the set of $(s,\zeta)$ such that $\sum_{l\in S_{i}(s,\zeta)}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. Based on the proof of Lemma˜6.4, note that for each block of $\lambda_{i}$ operators, we multiply $c$ by some constant factor. For $(s,\zeta)\in\mathcal{G}$, we let $C_{i}(s,\zeta)$ denote the factor we multiply $c$ by when applying the block of $\lambda_{i}$ operators $\{s_{j}:j\in[1+\lambda_{1}+\cdots+\lambda_{i-1},\lambda_{1}+\cdots+\lambda_{i}]\}$ associated with $j_{i}=i$. Then, we have that for $(s,\zeta)\in\mathcal{G}$,
(11) $c(s,\zeta)=\prod_{i=1}^{\ell(\lambda)}C_{i}(s,\zeta),$
Afterwards, it is evident that if we condition on the values $S_{i}(s,\zeta)$, then the blocks of $\lambda_{i}$ operators will be independent, since we can allocate the locations of the derivatives and assign the derivatives to the $p_{\nu_{l}}$ independently within each block. Suppose $\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}$ such that $\sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. Then, conditioned on $S_{i}(s,\gamma)=S_{i}$ for all $i\in[\ell(\lambda)]$, the total contribution will be the product of the contributions from each of the blocks. In particular, the main idea for the next step is that the contribution to the leading order coefficient of $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ when we condition on the sets $S_{i}$, $1\leq i\leq\ell(\lambda)$ will be the leading order coefficient of
$\prod_{i=1}^{\ell(\lambda)}\left[p_{\lambda_{i}},\prod_{j\in S_{i}}p_{\nu_{j}}\right]_{A^{N-1}(\theta)}.$
So, we only need to consider the case where $\ell(\lambda)=1$ and then sum over the partitions $\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}$. We justify this idea rigorously in the following lemma.
Lemma 6.5.
For $\nu^{\prime}\in\Gamma$, let $c_{\nu^{\prime}}$ denote the quantity $\sum_{s\in\mathcal{S},\,\zeta\in\mathcal{H}(s)}c(s,\zeta)$ from (10) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$. Suppose $\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}$ such that $\sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. Then,
$\sum_{\begin{subarray}{c}(s,\zeta)\in\mathcal{G},\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}c(s,\zeta)=\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:\,l\in S_{i}))}.$
Remark 6.6.
Note that $c_{\nu^{\prime}}$ is simply the leading order coefficient of $p_{(|\nu^{\prime}|)},p_{\nu^{\prime}}]_{A^{N-1}(\theta)}$.
Proof of Lemma˜6.5.
For $s\in\mathcal{S}$ and $i\in[\ell(\lambda)]$, define $s[i]\triangleq\{s_{j}:j\in[1+\lambda_{1}+\cdots+\lambda_{i-1},\lambda_{1}+\cdots+\lambda_{i}]\}$ and $\mathcal{P}_{i}(s)\triangleq\mathcal{P}(s)\cap[1+\lambda_{1}+\cdots+\lambda_{i-1},\lambda_{1}+\cdots+\lambda_{i}]$.
Let $\mathcal{T}$ denote the set of $s\in\mathcal{S}$ such that $|\mathcal{P}_{i}(s)|=|S_{i}|$ for all $i\in[\ell(\lambda)]$ and for $i\in[\ell(\lambda)]$, define $\mathcal{T}_{i}\triangleq\{s[i]:s\in\mathcal{T}\}$. Observe that
$\mathcal{T}\cong\mathcal{T}_{1}\otimes\cdots\otimes\mathcal{T}_{\ell(\lambda)},$
since the number of switches in each element of $\mathcal{T}_{i}$ is fixed at $\lambda_{i}-|S_{i}|$; the bijection is given by $s\mapsto s[1]\times\cdots\times s[\ell(\lambda)]$.
Suppose $i\in[\ell(\lambda)]$. Define the function $\mathcal{P}_{i}^{\prime}$ over $\mathcal{T}_{i}$ by $\mathcal{P}_{i}^{\prime}(s[i])\triangleq\mathcal{P}_{i}(s)$ for $s\in\mathcal{T}$; this function is well defined. Furthermore, for $s\in\mathcal{T}$, define $\mathcal{H}_{i}(s[i])$ to be the set of bijections $\zeta:\mathcal{P}_{i}^{\prime}(s[i])\rightarrow S_{i}$. Then,
$\{(s,\zeta):s\in\mathcal{T},\,\zeta\in\mathcal{H}(s),\,S_{i}(s,\zeta)=S_{i}\,\forall i\in[\ell(\lambda)]\}\cong$
$\bigotimes_{i=1}^{\ell(\lambda)}\{(s,\zeta):s\in\mathcal{T}_{i},\,\zeta\in\mathcal{H}_{i}(s)\},$
where the bijection is given by $(s,\zeta)\mapsto(s[1],\zeta|_{\mathcal{P}_{1}(s)})\times\cdots\times(s[\ell(\lambda)],\zeta|_{\mathcal{P}_{\ell(\lambda)}(s)})$. Furthermore, define the function $C_{i}^{\prime}$ over $\{(s,\zeta):s\in\mathcal{T}_{i},\,\zeta\in\mathcal{H}_{i}(s)\}$ by $C_{i}^{\prime}(s[i],\zeta|_{\mathcal{P}_{i}(s)})\triangleq C_{i}(s,\zeta)$ for $s\in\mathcal{T}$ and $\zeta\in\mathcal{H}(s)$ such that $(s,\zeta)\in\mathcal{G}$ and $S_{i}(s,\zeta)=S_{i}$.
After using (11), it is evident that
$\sum_{\begin{subarray}{c}(s,\zeta)\in\mathcal{G},\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}c(s,\zeta)$ $=\sum_{\begin{subarray}{c}(s,\zeta)\in\mathcal{G},\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}\prod_{i=1}^{\ell(\lambda)}C_{i}(s,\zeta)$
$=\sum_{\begin{subarray}{c}s\in\mathcal{T},\zeta\in\mathcal{H}(s),\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}\prod_{i=1}^{\ell(\lambda)}C_{i}(s,\zeta)$
$=\sum_{\begin{subarray}{c}\mathbf{s}_{i}\in\mathcal{T}_{i},\,\zeta_{i}\in\mathcal{H}_{i}(\mathbf{s}_{i}),\\ 1\leq i\leq\ell(\lambda)\end{subarray}}\prod_{i=1}^{\ell(\lambda)}C_{i}^{\prime}(\mathbf{s}_{i},\zeta_{i})$
$=\prod_{i=1}^{\ell(\lambda)}\sum_{\mathbf{s}_{i}\in\mathcal{T}_{i},\,\zeta_{i}\in\mathcal{H}_{i}(\mathbf{s}_{i})}C_{i}^{\prime}(\mathbf{s}_{i},\zeta_{i}).$
However, observe that for $i\in[\ell(\lambda)]$, $\sum_{\mathbf{s}_{i}\in\mathcal{T}_{i},\,\zeta_{i}\in\mathcal{H}_{i}(\mathbf{s}_{i})}C_{i}^{\prime}(\mathbf{s}_{i},\zeta_{i})$ is the same as the quantity $\sum_{s\in\mathcal{S},\zeta\in\mathcal{H}(s)}c(s,\zeta)$ from (10) for when $\lambda$ is set as $(\lambda_{i})$ and $\nu$ is set as $\gamma((\nu_{l}:l\in S_{i}))$. Then, we have that
$\sum_{\begin{subarray}{c}(s,\zeta)\in\mathcal{G},\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}\prod_{i=1}^{\ell(\lambda)}C_{i}(s,\zeta)=\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:l\in S_{i}))}.$
Since using (10) and Lemma˜6.5 gives that
(12) $\begin{split}\sum_{s\in\mathcal{S}}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu}&=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\sum_{\begin{subarray}{c}(s,\zeta)\in\mathcal{G},\\ S_{i}(s,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}c(s,\zeta)\\ &=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:l\in S_{i}))},\end{split}$
it suffices to compute the values of $c_{\nu^{\prime}}$.
Lemma 6.7.
For $\nu^{\prime}\in\Gamma$, let $n_{\nu^{\prime}}$ equal the number of noncrossing partitions of $|\nu^{\prime}|]$ with block size given by $\nu^{\prime}$. Then, $c_{\nu^{\prime}}=\theta^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\pi(\nu^{\prime})n_{\nu^{\prime}}$
Remark 6.8.
The $\pi(\nu^{\prime})$ factor arises from the fact that we must order the blocks of the same size.
Proof of Lemma˜6.7.
In this proof, we use the same notation for $\mathcal{S}$, $\mathcal{H}(s)$ for $s\in\mathcal{S}$, and $c(s,\zeta)$ for $s\in\mathcal{S}$ and $\zeta\in\mathcal{H}(s)$ that is mentioned earlier in the subsection after we set $\lambda$ to be $(|\nu^{\prime}|)$ and $\nu$ to be $\nu^{\prime}$. First, recall that by definition,
$c_{\nu^{\prime}}=\sum_{s\in\mathcal{S},\,\zeta\in\mathcal{H}(s)}c(s,\zeta).$
The number of derivatives $\partial_{1}$ in $s\in\mathcal{S}$ is $\ell(\nu^{\prime})$ and the number of switches, which are each equivalent to $\theta d_{1}$, is $|\nu^{\prime}|-\ell(\nu^{\prime})$. Suppose $n^{\prime}_{\nu^{\prime}}$ is the number of ways to choose the locations of the $\ell(\nu^{\prime})$ derivatives of $s$ and assign them to distinct $p_{\nu^{\prime}_{l}}$ for $l\in[\ell(\nu^{\prime})]$ such that the resulting contribution $c(s,\zeta)$ is nonzero. Then $c_{\nu^{\prime}}=\theta^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}n^{\prime}_{\nu^{\prime}}$ by the argument in Remark˜6.3. Hence, it suffices to show that $n^{\prime}_{\nu^{\prime}}=\pi(\nu^{\prime})n_{\nu^{\prime}}$.
Suppose $s\in\mathcal{S}$ and $s=\{s_{i}\}_{1\leq i\leq|\nu^{\prime}|}$; recall that $s_{1}=\partial_{1}$. Furthermore, suppose $\zeta\in\mathcal{H}(s)$. Assume that $c(s,\zeta)\not=0$, or equivalently that $c(s,\zeta)=\theta^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}$. Then, we construct the noncrossing partition $\alpha(s,\zeta)$ of $|\nu^{\prime}|]$ using the following procedure.
  1. (1)
    Initiate $n=1$. Let $\mathcal{A}$ be a stack which is initially empty.
  2. (2)
    Suppose $s_{n}=\partial_{1}$ and $\zeta(n)=l$ so that $s_{n}$ is assigned to $\nu^{\prime}_{l}$. Then, we add $(\{n\},\nu^{\prime}_{l})$ to the top of $\mathcal{A}$.
  3. (3)
    Suppose $s_{n}$ is a switch. Then, iterate through $\mathcal{A}$ from top to bottom. For the first element $(S,L)$ such that $|S|<L$, add $n$ to $S$.
  4. (4)
    Increment $n$ by one.
  5. (5)
    If $n=|\nu^{\prime}|+1$, then for the noncrossing partition, output the union of the sets $S$ for $(S,L)\in\mathcal{A}$.
  6. (6)
    If $n\leq|\nu^{\prime}|$, then return to step (2).
Since $c(s,\zeta)\not=0$, it is straightforward to deduce that we can always preform step (3) by counting the degree of $x_{1}$. Then, the output partition of $|\nu^{\prime}|]$ will have block sizes given by $\nu^{\prime}$ and will be noncrossing due to the structure of $\mathcal{A}$.
To reverse the algorithm, suppose $\sigma$ is a noncrossing partition of $\lambda_{i}$ with blocks sizes given by $\nu^{\prime}$. Suppose $\sigma=B_{1}\sqcup\cdots\sqcup B_{\ell(\nu^{\prime})}$, where $B_{q}$ starts before $B_{q+1}$ for $1\leq q\leq\ell(\nu^{\prime})-1$. Suppose $q\in[\ell(\nu^{\prime})]$ and that the first element of $B_{q}$ is $b$. Then, set $s_{b}$ to be $\partial_{1}$ and assign $s_{b}$ to $p_{\nu^{\prime}_{l}}$ for some $l\in[\ell(\nu^{\prime})]$ such that $p_{\nu^{\prime}_{l}}$ has not been previously assigned to and $|\nu^{\prime}_{l}|=|B_{q}|$; note that we then have that $\zeta(b)=l$. Afterwards, we set the remaining operators to be switches based on condition (3) in the definition of $\mathcal{S}$.
It is evident that the resulting pair $(s,\zeta)$ satisfies $\alpha(s,\zeta)=\sigma$ and that the number of choices for $(s,\zeta)$ is $\pi(\nu^{\prime})$. This establishes that $n^{\prime}_{\nu^{\prime}}=\pi(\nu^{\prime})n_{\nu^{\prime}}$. ∎
Using (12) and Lemma˜6.7, it then follows that the leading order coefficient is
$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\theta^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))(\#\,\text{noncrossing partitions of }[\lambda_{i}]\text{ with shape }\gamma((\nu_{j}:j\in S_{i})))$
$=\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\theta^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}(\#\,\text{noncrossing partitions of }[\lambda_{i}]\text{ with shape }\gamma_{i})$
$=\theta^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}.$
6.2. Proof of part (A) of Theorem˜1.1
We are now prepared to prove part (A) of Theorem˜1.1.
Define the infinite dimensional matrix $\mathcal{M}^{A}\in\mathbb{Z}_{\geq 0}^{\Gamma\times\Gamma}$ such that for $\lambda,\nu\in\Gamma$,
$\mathcal{M}^{A}_{\lambda\nu}\triangleq\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}.$
The following lemma is straightforward to deduce.
Lemma 6.9.
Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$.
  1. (A)
    If $\lambda=\nu$, then $\mathcal{M}^{A}_{\lambda\nu}=\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)$.
  2. (B)
    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{A}_{\lambda\nu}=0$.
The next theorem generalizes part (A) of Theorem˜1.1. As we mentioned previously, we focus on the less general version stated in Corollary˜6.11 because it is more relevant for analyzing the convergence of Bessel functions.
Theorem 6.10.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\\ |\theta N|=\infty$. The following are equivalent.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\sum_{\lambda\in\Gamma}\mathcal{M}^{A}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
Proof.
Suppose $\nu\in\Gamma$ and $N\geq 1$ satisfies $\nu\in\Gamma_{N}$. We have that
$p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\sum_{\lambda\in\Gamma_{N}}[p_{\lambda}]F_{N}\cdot[p_{\nu},p_{\lambda}]_{A^{N-1}(\theta)}.$
Then, using Theorem˜6.1 after noting that each summand of $R$ is $o_{N}(|\theta N|^{|\nu|}N^{\ell(\nu)})$, we get that
(13) $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}.$
We justify this equation. For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)<\ell(\nu)$, we have that
$p_{\nu},p_{\lambda}]_{A^{N-1}(\theta)}=O_{N}(|\theta N|^{|\nu|-\ell(\nu)}N^{\ell(\lambda)})=o_{N}(|\theta N|^{|\nu|-\ell(\lambda)}N^{\ell(\nu)}),$
which then corresponds to the term $\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1)$ since $\mathcal{M}^{A}_{\nu\lambda}=0$. For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)\geq\ell(\nu)$, it is straightforward to obtain
$p_{\nu},p_{\lambda}]_{A^{N-1}(\theta)}=(\theta N)^{|\nu|-\ell(\lambda)}N^{\ell(\nu)}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))$
using Theorem˜6.1.
(a) implies (b): It is straightforward to recover the equation from (b) using (13).
(b) implies (a): We proceed with induction on $|\lambda|$.
Assume that the result is true when $|\lambda|\leq k-1$ for $k\geq 1$. We prove that the result is true when $|\lambda|=k$. First, using (b) and (13) gives that for all $\nu\in\Gamma$ with $|\nu|=k$,
$\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[k]}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}=\sum_{\lambda\in\Gamma}\mathcal{M}^{A}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
After deleting the terms of lower order by the inductive hypothesis, we have that
$\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[k]}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=\sum_{\lambda\in\Gamma}\mathcal{M}^{A}_{\nu\lambda}c_{\lambda}.$
Hence,
$\lim_{N\rightarrow\infty}(\mathcal{M}^{A}[k]+o_{N}(1))\left[\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}-c_{\lambda}\right]_{\lambda\in\Gamma[k]}^{T}=0.$
However, observe that $\mathcal{M}^{A}[k]$ is upper-triangular with a positive diagonal from Lemma˜6.9. Thus, $\mathcal{M}^{A}[k]$ is invertible, which proves (a). ∎
Corollary 6.11.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\\ |\theta N|=\infty$. The following are equivalent.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
Proof.
It suffices to prove that for all $\nu\in\Gamma$,
$\sum_{\lambda\in\Gamma}\mathcal{M}^{A}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right)=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}$
when $c_{\gamma}=0$ for all $\gamma\in\Gamma$ with $\ell(\gamma)\geq 2$. Observe that by the definition of $\mathcal{M}^{A}$, the left hand side equals
$\sum_{\lambda\in\Gamma}\prod_{l=1}^{\ell(\lambda)}\lambda_{l}\pi(\lambda)\frac{\prod_{l=1}^{\ell(\lambda)}c_{(\lambda_{l})}}{\pi(\lambda)}\left[\prod_{l=1}^{\ell(\lambda)}x_{\lambda_{l}}\right]\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}x_{|B|}$
$=\sum_{\lambda\in\Gamma}\prod_{l=1}^{\ell(\lambda)}\lambda_{l}c_{(\lambda_{l})}\left[\prod_{l=1}^{\ell(\lambda)}x_{\lambda_{l}}\right]\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}x_{|B|}$
$=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
From Newton’s identities, it is straightforward to verify that when we write the degree $k$ elementary symmetric polynomial as a linear combination of power sums, the coefficient of $p_{(k)}$ is $\frac{(-1)^{k-1}}{k}$. Using the previous result as well as the results of [limits_prob_measures], we obtain the following generalization.
Corollary 6.12.
Suppose $\epsilon\in\mathcal{P}$. Assume that
$M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda},$
where $c_{\epsilon\lambda}\in\mathbb{R}$ for $\lambda\in\Gamma$ are unique. Then, $c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
Remark 6.13.
When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}\\ p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables. We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
Proof of Corollary˜6.12.
Consider [limits_prob_measures]*Theorem 1.6 with $c=0$, $\theta_{N}=\theta>0$ for all $N\geq 1$, and $G_{\theta_{N}}(x_{1},\ldots,x_{N};\,\mu_{N})=\exp(Nm_{\epsilon})$. Note that it is not necessarily true that there exists $\mu_{N}$ with this particular Bessel generating function; however, the proof of the theorem implies that
$\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)},$
where $c_{(k)}=0$ for $k\geq 1$ such that $k\not=|\epsilon|$ and $c_{(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{\theta^{-1}|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
By Corollary˜6.11, we have that $\lim_{N\rightarrow\infty}\frac{c_{(|\epsilon|)}(N)}{\theta N}=(-1)^{\ell(\epsilon)-1}\frac{\theta^{-1}|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$, where $G_{\theta_{N}}(x_{1},\ldots,\\ x_{N};\,\mu_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$. However, observe that
$c_{(|\epsilon|)}(N)=Nc_{\epsilon,(|\epsilon|)}\Rightarrow c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}.$
6.3. Leading order terms of the type A Dunkl bilinear form in the $\theta N\rightarrow c\in\mathbb{C}$ regime
Next, we compute the leading order terms of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma}$ in the regime $\theta N\rightarrow c\in\mathbb{C}$.
Suppose $k\geq 1$ and $\pi\in NC(k)$. Define the polynomial $W^{A}(\pi)\in\mathbb{Z}[x]$ by
$W^{A}(\pi)(x)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(x+d(i;\,\pi)).$
Theorem 6.14.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{A}(\pi)(N\theta)\prod_{B\in\pi}x_{|B|}+R(N,\theta),$
where $R\in\mathbb{Q}[x,y]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
Similarly to how Theorem˜6.1 can be proved using the results of [limits_prob_measures], this theorem can be proved using Theorems 3.8 and 3.10 of the paper [matrix], although the weight function that the paper uses is written differently than how we write $W^{A}(\pi)$. However, we include the proof that follows the method of the second proof of Theorem˜6.1. This proof can be used while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems.
Note that Theorem˜6.17 generalizes the results of the paper [matrix], since it allows $c_{\lambda}$ to be nonzero if $\ell(\lambda)\geq 2$. However, Corollary˜6.18 has already been proved in the paper.
Proof of Theorem˜6.14.
We can use the same argument as the proof of Theorem˜6.1. The main difference is that we are including all terms such that the degree of $N$ is $\ell(\lambda)$ greater than the degree of $\theta$ in the leading order term. This means that we still have that $j_{i}$ for $1\leq i\leq\ell(\lambda)$ are distinct to obtain the factor of $N^{\ell(\lambda)}$ and that the switches are to distinct indices so that the degrees of $N$ and $\theta$ are the same, after removing the factor of $N^{\ell(\lambda)}$. However, we no longer require that $d=\ell(\nu)-\ell(\lambda)$, since we are not assuming that the order of $N\theta$ is greater than the order of $1$. In particular, $d\geq\ell(\nu)-\ell(\lambda)$ and the number of switches is $k-\ell(\lambda)-d$. This essentially means that the derivatives do not all have to be assigned to a $p_{\nu_{l}}$ to contribute to the leading order term.
Furthermore, for the remainder term $R$, we have that each term which is $N^{\ell(\lambda)}$ multiplied by a power of $N\theta$ is already included in the leading order term, so it is straightforward to deduce that it follows the given condition. We alter the definition of $\mathcal{S}$ as follows.
Let $\mathcal{S}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq n(s)$, where $n(s)$ denotes the number of switches of $s$.
Then, the leading order term is
(14) $N^{\ell(\lambda)}\sum_{s\in\mathcal{S}}N^{n(s)}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu},$
where the factor of $N^{\ell(\lambda)}$ arises from the choices of the $j_{i}$ and the factor of $N^{n(s)}$ arises from selecting the indices of the switches. For simplicity, we compute
(15) $\sum_{s\in\mathcal{S}}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu},$
and then multiply by $N^{\ell(\lambda)}$ and replace $\theta$ with $\theta N$ to obtain (14).
We have that the analogous results in the proof of Theorem˜6.1 are true, other than Lemma˜6.7, which we must alter since the derivatives have the same weight as $\theta N$. In other words, as we mentioned earlier, the derivatives in $s$ do not all have to be assigned to a $p_{\nu_{l}}$ to contribute to the leading order term.
Lemma 6.15.
Suppose $\nu^{\prime}\in\Gamma$. Let $c_{\nu^{\prime}}$ be the value of (15) when $\lambda$ is set as $(|\nu^{\prime}|)$ and $\nu$ is set as $\nu^{\prime}$ and define
$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\,\gamma(\pi)=\nu^{\prime}\end{subarray}}W_{\pi}^{A}(x).$
Then, $c_{\nu^{\prime}}=\theta^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\pi(\nu^{\prime})n_{\nu^{\prime}}(\theta)$.
Proof.
The proof is essentially the same as the proof of Lemma˜6.7. The only difference is that at each position where we previously apply a switch, we can also choose to apply a derivative. The derivative would multiply the coefficient by the current degree of $x_{1}$, which corresponds to the $d(i;\,\pi)$ factor in the formula for $W_{\pi}^{A}$. ∎
Using (12), (15), and Lemma˜6.15, and then replacing $\theta$ with $\theta N$ and multiplying by a factor of $N^{\ell(\lambda)}$, it follows that the leading order coefficient is
$N^{\ell(\lambda)}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W_{\pi}^{A}(N\theta)$
$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}W_{\pi}^{A}(N\theta)$
$=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}W_{\pi}^{A}(N\theta).$
We now prove the analogue Theorem˜6.17 of Theorem˜6.10. The method we use is the same, however we include it for completeness.
Define the infinite dimensional matrix $\mathcal{W}^{A}\in\mathbb{Z}_{\geq 0}^{\Gamma\times\Gamma}[y]$ such that for $\lambda,\nu\in\Gamma$,
$\mathcal{W}^{A}_{\lambda\nu}(y)\triangleq\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{A}(\pi)(y)\prod_{B\in\pi}x_{|B|}.$
The following lemma is straightforward to deduce.
Lemma 6.16.
Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$.
  1. (A)
    If $\lambda=\nu$, then $\mathcal{W}^{A}(y)_{\lambda\nu}=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}W^{A}([\nu_{l}])(y)$.
  2. (B)
    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{A}_{\lambda\nu}=0$.
Theorem 6.17.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\\ \theta N=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\sum_{\lambda\in\Gamma}\mathcal{W}^{A}(c)_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
Then, (a) implies (b), and if $c$ is not a negative integer, then (b) implies (a).
Proof.
We follow the method of the proof of Theorem˜6.10. Suppose $\nu\in\Gamma$. Then, using Theorem˜6.14 after noting that each summand of $R$ is $o_{N}(N^{\ell(\nu)})$, we get that
(16) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}.$
For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)<\ell(\nu)$, we have that
$p_{\nu},p_{\lambda}]_{A^{N-1}(\theta)}=O_{N}(N^{\ell(\lambda)})=o_{N}(N^{\ell(\nu)}),$
which then corresponds to the term $(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))$ since $\mathcal{W}^{A}(c)_{\nu\lambda}=0$. For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)\geq\ell(\nu)$, it is straightforward to obtain from Theorem˜6.14 that
$p_{\nu},p_{\lambda}]_{A^{N-1}(\theta)}=N^{\ell(\nu)}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1)).$
(a) implies (b): It is straightforward to recover the equation from (b) using (16).
(b) implies (a): Assume that $c$ is not a negative integer. We proceed with induction on $|\lambda|$.
Assume that the result is true when $|\lambda|\leq k-1$ for $k\geq 1$. We prove that the result is true when $|\lambda|=k$. First, using (b) and (16) gives that for all $\nu\in\Gamma[k]$,
$\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[k]}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}=\sum_{\lambda\in\Gamma}\mathcal{W}^{A}(c)_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
After deleting the terms of lower order by the inductive hypothesis, we have that
$\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[k]}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))c_{\lambda}(N)=\sum_{\lambda\in\Gamma}\mathcal{W}^{A}(c)_{\nu\lambda}c_{\lambda}.$
Hence,
$\lim_{N\rightarrow\infty}(\mathcal{W}^{A}(c)[k]+o_{N}(1))\left[c_{\lambda}(N)-c_{\lambda}\right]_{\lambda\in\Gamma[k]}^{T}=0.$
However, observe that $\mathcal{W}^{A}(c)[k]$ is upper-triangular with a nonzero diagonal from
Lemma˜6.16, since $c$ is not a negative integer. Thus, $\mathcal{W}^{A}(c)[k]$ is invertible, which proves (a). ∎
Corollary 6.18 ([matrix]*Theorems 3.8 and 3.10).
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\\ \left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}c_{\lambda}(N)=0$ if $\ell(\lambda)>1$
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}W^{A}(\pi)(c)\prod_{B\in\pi}|B|c_{(|B|)}.$
Then, (a) implies (b), and if $c$ is not a negative integer, then (b) implies (a).
Proof.
The same method as the proof of Corollary˜6.11 can be used. ∎
7. Leading order terms for the type BC Dunkl bilinear form
We transition to the type BC root system and first prove the analogues of the results of Section˜6 for the $|\theta_{0}N|\rightarrow\infty$ regime and later prove the analogues of the results for the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$ regime. Note that we always have that $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ and that the asymptotic invertibility of $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ will depend on $c_{0}$ and $c_{1}$.
7.1. Proof of part (B) of Theorem˜1.1
We first consider the $|\theta_{0}N|\rightarrow\infty$ regime.
Theorem 7.1.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $(2\theta_{0})^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}\left(1+\frac{\theta_{1}}{N\theta_{0}}\right)^{o(\pi)}\times$
$N^{k+\ell(\lambda)-\ell(\nu)}+R(N,\theta_{0},\theta_{1}).$
where $R\in\mathbb{Q}[x,y,z]$ satisfies:
  1. (1)
    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
  2. (2)
    The degree of $x$ is at most $k+\ell(\lambda)-\ell(\nu)-1$ and the sum of the degrees of $y$ and $z$ in each summand is at most $k-\ell(\nu)$.
  3. (3)
    In no summand is the degree of $x$ $\ell(\lambda)$ greater than the degree of $y$ while the degrees of $y$ and $z$ add to $k-\ell(\nu)$.
Proof.
We follow the proof of Theorem˜6.1 given in Subsection˜6.1. The main difference is that we regard $\theta_{1}$ as having weight $N\theta_{0}$, while $N$ is considered to be the same as before and $2\theta_{0}$ is considered to be the same as $\theta$. However, we still regard $N\theta_{0}$ and $\theta_{1}$ as being higher order than a constant. So, we still have that the $j_{i}$ for $i\in[\ell(\lambda)]$ are distinct to obtain the factor of $N^{\ell(\lambda)}$. Afterwards, we have that the type 0 switches are to distinct indices, so that the degrees of $\theta_{0}$ and $N$ are the same after removing the factor of $N^{\ell(\lambda)}$. In contrast with the proof of Theorem˜6.14, we still have that $d=\ell(\nu)-\ell(\lambda)$.
It is not challenging to see that the remainder term $R$ is a polynomial with rational coefficients that satisfies conditions (1), (2), and (3). For condition (1), we have that after selecting $j_{1},\ldots,j_{\ell(\lambda)}$, to obtain an additional factor of $N$, we must include a type 0 switch, which will also add a factor of $\theta_{0}$. For condition (2), we note that we have already identified the leading order term, so the degree of $x$ is at most $k+\ell(\lambda)-\ell(\nu)-1$. Furthermore, the number of switches is at most $k-\ell(\nu)$ since $d\geq\ell(\nu)-\ell(\lambda)$, so the degrees of $y$ and $z$ in each summand add to at most $k-\ell(\nu)$. For condition (3), we note that any term that satisfies this condition is included in the leading order term.
Let $\mathcal{S}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    The $j$th type 0 switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-\ell(\nu)-n(s)$, where $n(s)$ is the number of type 1 switches of $s$.
Then, the leading order term is
(17) $N^{\ell(\lambda)}\sum_{s\in\mathcal{S}}N^{k-\ell(\nu)-n(s)}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu},$
where $N^{\ell(\lambda)}$ corresponds to the number of choices for $j_{1},\ldots,j_{\ell(\lambda)}$ and $N^{k-\ell(\nu)-n(s)}$ corresponds to the number of choices for the indices of the type 0 switches. For simplicity, we compute
(18) $\sum_{s\in\mathcal{S}}s_{|\lambda|}\circ\cdots\circ s_{1}p_{\nu}.$
Afterwards, we replace $\theta_{0}$ with $N\theta_{0}$ and multiply by a factor of $N^{\ell(\lambda)}$ to compute (17).
We have that the analogous results in the proof of Theorem˜6.14 are true, other than Lemma˜6.7, which we must alter to account for the type 1 switches.
Lemma 7.2.
Suppose $\nu^{\prime}\in\Gamma_{\text{even}}$. Let $c_{\nu^{\prime}}$ be the value of (18) when $\lambda$ is set as $(|\nu^{\prime}|)$ and $\nu$ is set as $\nu^{\prime}$ and define
$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\\ \gamma(\pi)=\nu^{\prime}\end{subarray}}(1+x)^{o(\pi)}.$
Then, $c_{\nu^{\prime}}=(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\pi(\nu^{\prime})n_{\nu^{\prime}}\left(\frac{\theta_{1}}{\theta_{0}}\right)$.
Proof.
We can follow the proof of Lemma˜6.7 with a few differences. First, we replace $\theta$ by $2\theta_{0}$. The only other difference is that when the degree of $x_{1}$ is odd, we can apply either the type 0 or type $1$ switch, which would be equivalent to $2\theta_{0}d_{1}$ and $2\theta_{1}d_{1}$, respectively. To account for these choices, for each noncrossing partition $\pi$, we must multiply by $(1+\frac{\theta_{1}}{\theta_{0}})^{o(\pi)}$, since $o(\pi)$ counts the number of locations at which there is no $\partial_{1}$ and the degree of $x_{1}$ is odd. ∎
Recall that to compute the leading order term, we must replace $\theta_{0}$ with $N\theta_{0}$ and multiply by a factor of $N^{\ell(\lambda)}$. Then, using (12), (18), and Lemma˜7.2, the leading order term is
$N^{\ell(\lambda)}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}(2N\theta_{0})^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}\left(1+\frac{\theta_{1}}{N\theta_{0}}\right)^{o(\pi)}$
$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}(2N\theta_{0})^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}\left(1+\frac{\theta_{1}}{N\theta_{0}}\right)^{o(\pi)}$
$=N^{\ell(\lambda)}(2N\theta_{0})^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}\left(1+\frac{\theta_{1}}{N\theta_{0}}\right)^{o(\pi)}.$
Define the infinite dimensional matrix $\mathcal{M}^{BC}\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}[y]$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$,
$\mathcal{M}^{BC}_{\lambda\nu}(y)\triangleq 2^{|\nu|-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}(1+y)^{o(\pi)}.$
Lemma 7.3.
Suppose $\lambda,\nu\in\Gamma_{\text{even}}$ and $|\lambda|=|\nu|$.
  1. (A)
    If $\lambda=\nu$, then $\mathcal{M}^{BC}_{\lambda\nu}(y)=\pi(\nu)\prod_{l=1}^{\ell(\nu)}2^{\nu_{l}-1}\nu_{l}(1+y)^{o([\nu_{l}])}$.
  2. (B)
    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{BC}_{\lambda\nu}(y)=0$.
Theorem 7.4.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N;even}}c_{\lambda}(N)p_{\lambda}\right)$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Suppose $c_{\lambda}\in\mathbb{C}$ for all $\lambda\in\Gamma_{\text{even}}$. Consider the following statements:
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta_{0}N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma}\mathcal{M}^{BC}_{\nu\lambda}(c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
Then, (a) implies (b), and if $c\not=-1$, then (b) implies (a).
Proof.
The same method as the proof of Theorem˜6.10 can be used. For the implication of (a) from (b), we note that $\mathcal{M}^{BC}(c)$ is invertible when $c\not=-1$ by Lemma˜7.3. ∎
Corollary 7.5.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
$\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda(N)}(N)}{\theta_{0}N}=c_{\lambda}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta_{0}N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
Then, (a) implies (b), and if $c\not=-1$, then (b) implies (a).
Proof.
See the proof of Corollary˜6.11. ∎
7.2. Leading order terms of the type BC Dunkl bilinear form in the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime
Next, we consider the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime. Define the polynomial $W^{BC}(\pi)\in\mathbb{Z}[x,y]$ by
$W^{BC}(\pi)(x,y)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi)).$
Theorem 7.6.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in\pi}x_{|B|}$
$+R(N,\theta_{0},\theta_{1}),$
where $R\in\mathbb{Q}[x,y,z]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
Similarly to Theorems˜6.1 and 7.1, we can prove Theorem˜7.6 using Theorems 4.8 and 5.5 of the paper [rectangularmatrix], although the weight function that the paper uses is written differently than how we write $W^{BC}(\pi)$. We include a different proof of Theorem˜7.6 as well. Furthermore, we similarly have that Theorem˜7.8 generalizes the results of the paper while Corollary˜7.9 has already been proved in the paper.
Proof of Theorem˜7.6.
The idea is the same as the proof of Theorem˜6.1, except we add the modifications from the proofs of Theorems˜6.14 and 7.1. The leading order term consists of $N^{\ell(\lambda)}$ multiplied by a power of $N\theta_{0}$ and a power of $\theta_{1}$. Then, we still have that the $j_{i}$ are distinct and the type 0 switches are to distinct indices. However, now we have that $d\geq\ell(\nu)-\ell(\lambda)$ rather than $d=\ell(\nu)-\ell(\lambda)$.
For the modification of Lemma˜6.7, at each position which is not at the start of a block, we can apply a type 0 switch, the type 1 switch, or the derivative. These correspond to the terms $x$, $\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y$, and $d(i;\,\pi)$, respectively, in the formula for $W^{BC}(x,y)$. ∎
Define the infinite dimensional matrix $\mathcal{W}^{BC}\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}[y,z]$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$,
$\mathcal{W}^{BC}_{\lambda\nu}(y,z)\triangleq\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(y,z)\prod_{B\in\pi}x_{|B|}.$
Lemma 7.7.
Suppose $\lambda,\nu\in\Gamma_{\text{even}}$ and $|\lambda|=|\nu|$.
  1. (A)
    If $\lambda=\nu$, then $\mathcal{W}^{BC}_{\lambda\nu}(y,z)=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}W^{BC}([\nu_{l}])(y,z)$.
  2. (B)
    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}_{\lambda\nu}(y,z)=0$.
Theorem 7.8.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
$\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(2c_{0},2c_{1})[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
Then, (a) implies (b), and if $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer, then (b) implies (a).
Proof.
See the proof of Theorem˜6.17. For the implication of (a) from (b), consider the formula for $W^{BC}([l])(2c_{0},2c_{1})$ for some $l\in 2\mathbb{N}$. If $d(i;\,\pi)$ is odd, then the term $2c_{0}+2c_{1}+d(i;\,\pi)$ is nonzero. Moreover, if $d(i;\,\pi)$ is even and at least two, then the term $2c_{0}+d(i;\,\pi)$ is nonzero; note that if $d(i;\,\pi)=0$, then we must have that $b(i;\,\pi)=1$. Therefore, $W^{BC}([l])(2c_{0},2c_{1})\not=0$, so $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]$ is invertible for all $k\geq 1$ by Lemma˜7.7. ∎
Corollary 7.9 ([rectangularmatrix]*Theorems 4.8 and 5.5).
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}c_{\lambda}(N)=0$ if $\ell(\lambda)\geq 2$.
  2. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}.$
Then, (a) implies (b), and if $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer, then (b) implies (a).
Proof.
See the proof of Corollary˜6.18. ∎
8. Leading order terms of the type D Dunkl bilinear form
8.1. Orthogonality results
First, we prove some orthogonality results which are based on the presence of sign flips in $H(\mathcal{R})$. These results are relevant for separating terms with all even degrees and all odd degrees with respect to the Dunkl bilinear form for the $D^{N}$ root system. The following theorem expresses the Dunkl bilinear form as an integral when the multiplicity function is nonnegative.
Theorem 8.1 ([dunkl_integralkernel]*Theorem 3.10).
Assume that $\theta\in\theta(\mathcal{R})$ and $\theta\geq 0$. Then, for $p,q\in\mathbb{C}[x_{1},\ldots,x_{N}]$,
$p,q]_{\mathcal{R}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}(\mathcal{D}(e^{-\frac{p_{(2)}}{2}})p)(\mathcal{D}(e^{-\frac{p_{(2)}}{2}})q)h^{2}wdx,$
where:
  • The function $h:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $h(x)\triangleq\prod_{r\in\mathcal{R}^{+}}|\left\langle x,r\right\rangle|^{\theta(r)}$.
  • The function $w:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $w(x)\triangleq\frac{e^{-\left\lVert x\right\rVert_{2}^{2}}}{(2\pi)^{\frac{N}{2}}}$.
  • The constant $c_{N}$ is defined as $c_{N}\triangleq\int_{\mathbb{R}^{N}}h^{2}wdx$.
Remark 8.2.
The operator $\mathcal{D}(e^{-\frac{p_{(2)}}{2}})$ is equivalent to $\sum_{k=0}^{\infty}\left(-\frac{1}{2}\right)^{k}\frac{\mathcal{D}(p_{(2)}^{k})}{k!}$. Furthermore, [dunkl_integralkernel] gives a formula for $c_{N}$.
In order to analyze whether two monomials are orthogonal with respect to the Dunkl bilinear form $\cdot,\cdot]_{\mathcal{R}(\theta)}$, we can analyze the parities of the degrees of $x_{1},\ldots,x_{N}$ as well as the presence of sign flips in $H(\mathcal{R})$. We deduce the following result using the equivariance property of the Dunkl operators given in Lemma˜2.2.
Theorem 8.3.
Suppose $N\geq 1$ and $\lambda,\nu\in\mathbb{Z}_{\geq 0}^{N}$. Assume that for some $i\in[N]$, $\lambda_{i}$ and $\nu_{i}$ do not have the same parity and that $H(\mathcal{R})$ contains the reflection that flips the sign of $x_{i}$. Then, $\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}]_{\mathcal{R}(\theta)}=0$.
Proof.
First, assume that $\theta\geq 0$. By Theorem˜8.1,
(19) $\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{\mathcal{R}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}\left(\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\lambda_{i}}\right)\left(\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right)h^{2}wdx$
Assume $i\in[N]$ such that $\lambda_{i}$ and $\nu_{i}$ do not have the same sign and that $\sigma\in H$ flips the sign of $i$. Without loss of generality, assume that $\lambda_{i}$ is odd and $\nu_{i}$ is even. Then, by Lemma˜2.2, we have that
$\sigma\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\lambda_{i}}=-\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\lambda_{i}}$
and
$\sigma\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\nu_{i}}=\mathcal{D}\left(e^{-\frac{p_{(2)}}{2}}\right)\prod_{i=1}^{N}x_{i}^{\nu_{i}}.$
Then, because $h$ is $H$-invariant and $\sigma w=w$, if we apply $\sigma$ to the integrand of (19), we will flip its sign. It is then clear that (19) evaluates to zero.
Next, observe that $\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{\mathcal{R}(\theta)}$ is a polynomial in $\theta$ with all real coefficients. Thus, since it evaluates to zero for all choices of nonnegative real-valued $\theta$, it must evaluate to zero for all choices of complex-valued $\theta$. ∎
Corollary 8.4.
Suppose $N\geq 1$ and $\lambda,\nu\in\mathbb{Z}_{\geq 0}^{N}$. Assume that for some $i\in[N]$, $\lambda_{i}$ and $\nu_{i}$ do not have the same parity. Then, $\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{D^{N}(\theta)}=0$ for all $\theta\in\mathbb{C}$.
Proof.
We cannot directly apply Theorem˜8.3 since $H(D^{N}(\theta))$ only contains reflections that flip an even number of signs. However, we have that
$\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{D^{N}(\theta)}=\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{BC^{N}(\theta,0)},$
and we can apply Theorem˜8.3 to deduce that $\left[\prod_{i=1}^{N}x_{i}^{\lambda_{i}},\prod_{i=1}^{N}x_{i}^{\nu_{i}}\right]_{BC^{N}(\theta,0)}=0$. ∎
Corollary 8.5.
Suppose $N\geq 1$ and $H(\mathcal{R})$ contains the reflection that flips the sign of $x_{i}$ for some $i\in[N]$. Then, $E_{a}^{\mathcal{R}(\theta)}(x)$ is a linear combination of $r(a)s(x)$ for monomials $r$ and $s$ such that the degree of $a_{i}$ in $r(a)$ and the degree of $x_{i}$ in $s(x)$ have the same parity.
Proof.
Recall the formula (4), where $\Psi:r\mapsto r(a)$ after $a\in\mathbb{C}^{N}$ is fixed. By Theorem˜8.3, the matrix $M^{k;\,\mathcal{D}}$ has two orthogonal components corresponding to when the degrees of $x_{i}$ are even and odd for all $k\geq 1$. Then, it is straightforward to deduce the result by computing the inverse of $M^{k;\,\mathcal{D}}$, which has the same two orthogonal components. ∎
[…]
Corollary 8.6.
Suppose $\theta\in\mathbb{C}$ such that $D^{N}(\theta)$ is invertible.
  1. (A)
    Both $E_{a}^{D^{N}(\theta)}(x)$ and $J_{a}^{D^{N}(\theta)}(x)$ are linear combinations of $r(a)s(x)$ for monomials $r$ and $s$ such that for all $i\in[N]$, the degree of $a_{i}$ in $r(a)$ and the degree of $x_{i}$ in $s(x)$ have the same parity. In particular, $J_{a}^{D^{N}(\theta)}(x)$ is a linear combination of terms $r(a)s(x)$ such that $r$ and $s$ both have all even degrees or all odd degrees.
  2. (B)
    The function $E_{a}^{D^{N}(\theta)}(x)$ equals $E_{a}^{BC^{N}(\theta,0)}(x)$ and the function
    $\frac{1}{2}\left(J_{a}^{D^{N}(\theta)}(x_{1},x_{2},\ldots,x_{N})+J_{a}^{D^{N}(\theta)}(-x_{1},x_{2},\ldots,x_{N})\right)$
    equals $J_{a}^{BC^{N}(\theta,0)}(x)$.
Proof.
(A): The result for $E^{D^{N}(\theta)}_{a}(x)$ follows from the proof of Corollary˜8.5 after using Corollary˜8.4 rather than Theorem˜8.3. For $J_{a}^{D^{N}(\theta)}(x)$, we note that Corollary˜8.4 implies that for $k\geq 1$, the matrix $M^{2k;\,\mathcal{D}_{H}(D^{N}(\theta))}$ has two orthogonal components corresponding to $\{p_{\lambda}:\lambda\in\Gamma_{\text{even}}[2k]\}$ and $\{ep_{\lambda}:\lambda\in\Gamma_{\text{even}}[2k-N]\}$ assuming that $N$ is even. Then, the result is straightforward to obtain after using (4), computing the inverse of the matrix, and dividing it into the same two orthogonal components. If $N$ is odd, then the proof is more straightforward, since the row and column sets of $M^{2k;\,\mathcal{D}_{H}(D^{N}(\theta))}$ equal $\{p_{\lambda}:\lambda\in\Gamma_{\text{even}}[2k]\}$ while the row and column sets of $M^{2k-1;\,\mathcal{D}_{H}(D^{N}(\theta))}$ equal $\{ep_{\lambda}:\lambda\in\Gamma_{\text{even}}[2k-1-N]\}$ for $k\geq 1$.
(B): It is straightforward to deduce that $BC^{N}(\theta,0)$ is invertible if and only if $D^{N}(\theta)$ is invertible and to show that $E_{a}^{D^{N}(\theta)}(x)$ equals $E_{a}^{BC^{N}(\theta,0)}(x)$. Afterwards, the goal is to prove that the sum of the summands of $J_{a}^{D^{N}(\theta)}(x)$ with all even degrees in $a$ and $x$ is $J_{a}^{BC^{N}(\theta,0)}(x)$. Suppose $k\geq 1$. The result follows from the fact that when we restrict the matrix $M^{2k;\,\mathcal{D}_{H}(D^{N}(\theta))}$ to have row and column sets equal to $\{p_{\lambda}:\lambda\in\Gamma_{\text{even}}[2k]\}$, we obtain the matrix $M^{2k;\,\mathcal{D}_{H}(BC^{N}(\theta,0))}$. ∎
Corollary 8.7.
Assume that $\theta\in\mathbb{C}$ and $\mathcal{D}(D^{N}(\theta))$ is invertible. Suppose $a\in\mathbb{C}^{N}$. The coefficients of terms of $J_{a}^{D^{N}(\theta)}(x_{1},\ldots,x_{N})$ with all odd degrees in $x_{1},\ldots,x_{N}$ are all zero if and only if $a_{1}\cdots a_{N}=0$.
Proof.
Each coefficient is a multiple of $a_{1}\cdots a_{N}$ by Corollary˜8.6, so it is straightforward to deduce that the coefficients are all zero if $a_{1}\cdots a_{N}=0$. Next, assume that the coefficients are all zero. By (4) and Corollary˜8.4, we have that since the coefficients of terms with all odd degrees are all zero and $\mathcal{D}_{H}(D^{N}(\theta))$ is invertible, $ep_{\lambda}(a)=0$ for all $\lambda\in\Gamma_{\text{even}}$. It is then evident $a_{1}\cdots a_{N}=0$. ∎
The following result shows that $e$ is a harmonic function with respect to $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ for all $\theta_{0},\theta_{1}\in\mathbb{C}$.
Lemma 8.8.
For all $\theta_{0},\theta_{1}\in\mathbb{C}$, $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))(p_{(2)})e=0$.
Proof.
First, observe that
$\mathcal{D}_{1}e=(1+2\theta_{1})x_{2}\cdots x_{N}.$
Afterwards, it is straightforward to deduce that $\mathcal{D}_{1}^{2}e=0$. ∎
Corollary 8.9.
For all $\theta\in\mathbb{C}$, $\mathcal{D}(D^{N}(\theta))(p_{(2)})e=0$.
Proof.
Set $\theta_{1}=0$ in Lemma˜8.8. ∎
8.2. The leading order terms of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and an introduction to the leading order terms of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma_{\text{even}}$
Next, we prove the results analogous to those of Section˜6 for the type D root system. However, we must now compute both $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma$. First, we consider the computation of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$. As we will see, these values are asymptotically equivalent to the values of $p_{\lambda},p_{\nu}]_{A^{N-1}(2\theta)}$ in both the $|\theta N|\rightarrow\infty$ and $\theta N\rightarrow c\in\mathbb{C}$ regimes.
First, we consider the $|\theta N|\rightarrow\infty$ regime.
Theorem 8.10.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then, Theorem˜6.1 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $\theta$ replaced by $2\theta$ is true.
Proof.
We can use the same method as the proof of Theorem˜6.1. The reason for this is that because the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$. This is because after applying $\frac{1-\sigma_{j_{i}j}}{x_{j_{i}}-x_{j}}$ or $\frac{1-\tau_{j_{i}}\tau_{j}s_{j_{i}j}}{x_{j_{i}}+x_{j}}$ to $x_{j_{i}}^{d}$ for some $d\geq 1$, the only resulting term without $x_{j}$ is $\theta x_{j_{i}}^{d-1}$ in both cases. Hence, we only need to replace $\theta$ with $2\theta$ in Theorem˜6.1 to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$. ∎
Define the infinite dimensional matrix $\mathcal{M}^{D}\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$.
Next, we consider the $\theta N\rightarrow c\in\mathbb{C}$ regime.
Theorem 8.11.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Then, statements (a) and (c) of Theorem˜8.21 are equivalent.
Proof.
This follows from Theorem˜8.10 and the proof of Theorem˜6.10 after $\Gamma$ and $\mathcal{M}^{A}$ are replaced with $\Gamma_{\text{even}}$ and $\mathcal{M}^{D}$, respectively. ∎
Theorem 8.12.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then, Theorem˜6.14 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $N\theta$ replaced by $2N\theta$ is true.
Proof.
Similarly to the proof of Theorem˜8.10, since the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$. Then, to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$, we only need to replace $N\theta$ with $2N\theta$ in Theorem˜6.14. ∎
Theorem 8.13.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Then, in Theorem˜8.26, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
Proof.
This follows from Theorem˜8.12 and the proof of Theorem˜6.17 after $\Gamma$ and $c$ replaced with $\Gamma_{\text{even}}$ and $2c$, respectively. ∎
Remark 8.14.
Theorems 8.10, 8.11, 8.12, and 8.13 also follow from the analogous results for the type BC root system after setting $\theta_{1}=0$.
Next, we consider the computation of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$, which is considerably more complicated. First, we exhibit a more direct computation in the case where $p_{\lambda},p_{\nu}=p_{(2)}^{k}$ and the multiplicity function is nonnegative by using Theorem˜8.1.
Lemma 8.15.
Suppose $k\geq 0$. Then, for all $\theta\in\mathbb{R}_{\geq 0}$,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$
Proof.
Suppose $\theta\geq 0$. By Theorem˜8.1,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})h^{2}wdx,$
where $c_{N}$, $h$, and $w$ are defined with respect to the type $D$ root system with constant multiplicity $\theta$.
By Lemma˜8.8, $e$ is harmonic. Then, using [dunkl_integralkernel]*Proposition 3.9 gives that
$\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k}=(-1)^{k}k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)e,$
where $\gamma=N(N-1)\theta$ and $L_{n}^{(\alpha)}$ denotes the generalized Laguerre polynomial. Hence,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)\right)^{2}(eh)^{2}wdx.$
For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$. Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent. We similarly have that
$\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}=(c^{\prime}_{N})^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{N}{2}+\gamma^{\prime}-1\right)}\left(\frac{p_{2}}{2}\right)\right)^{2}(h^{\prime})^{2}wdx$
However, it is evident that $\gamma^{\prime}=\gamma+N$ and $(h^{\prime})^{2}=(eh)^{2}$, if we assume that the positive roots for the type B root system are $\{[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}:i\in[N]\}$ in addition to the positive roots for the type D root system. Then, we have that
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}$ $=\frac{c^{\prime}_{N}}{c_{N}}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}$
$=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-1}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$
We compute the leading order terms of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ in the following theorem. The leading order terms are obviously not polynomials, but can be expressed as a product involving gamma functions and polynomials. In particular, the product $\prod_{i=1}^{N-k}(1+2(i-1))\theta$ which appears in the following theorem can also be expressed as $(2\theta)^{N-k}\frac{\Gamma(\frac{1}{2\theta}+N-k)}{\Gamma(\frac{1}{2\theta})}$.
Theorem 8.16.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=\prod_{i=1}^{N-k}(1+2(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(2N\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}+R(N,\theta)\bigg),$
where $R\in\mathbb{Q}[x,y]$ satisfies:
  1. (1)
    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
  2. (2)
    The degree $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$.
8.3. Proof of Theorem˜8.16
We follow the framework of the proof of Theorem˜6.1. Observe that
$ep_{\lambda},ep_{\nu}]=\mathcal{D}(p_{\lambda})\mathcal{D}(e)ep_{\nu}=\sum_{j_{1},\ldots,j_{\ell(\lambda)}=1}^{N}\prod_{i=1}^{\ell(\lambda)}(\mathcal{D}_{j_{i}}-\partial_{j_{i}}+\partial_{j_{i}})^{\lambda_{i}-1}\partial_{j_{i}}\mathcal{D}_{1}\cdots\mathcal{D}_{N}ep_{\nu}.$
First, we consider case work on the operators in $\mathcal{D}(p_{\lambda})$. Assume that $\{j_{i}\}_{1\leq i\leq\ell(\lambda)}=[m_{1}]$. Furthermore, assume that the terms $\mathcal{D}_{j_{i}}-\partial_{j_{i}}$ have switches from $j_{i}$ to some $j\in[m_{2}]$ for some $m_{2}\geq m_{1}$, and that each element in $m_{2}]\backslash[m_{1}]$ appears in a switch. Moreover, we can replace the sets $m_{1}]$ and $m_{2}]\backslash[m_{1}]$ with any disjoint subsets of $[N]$ with size $m_{1}$ and $m_{2}-m_{1}$, respectively, and obtain the same result.
Next, we consider the operators $\mathcal{D}_{j}$ in $\mathcal{D}_{1}\cdots\mathcal{D}_{N}$. First, observe that by symmetry, we assume that the switches in $\mathcal{D}_{j}$ are from $j$ to some element of $\{j+1,\ldots,N\}$ for $j\in[N]$. This is because prior to applying $\mathcal{D}_{j}$, we have applied $\mathcal{D}_{j+1},\ldots,\mathcal{D}_{N}$, so we are applying $\mathcal{D}_{j}$ to a polynomial which is invariant under switches between two distinct variables in $\{x_{1},\ldots,x_{j}\}$. To justify this, using Lemma˜2.2 gives that for $h\in H$,
$h\mathcal{D}(x_{j+1}\cdots x_{N})ep_{\nu}=\mathcal{D}(h(x_{j+1}\cdots x_{N}))h(ep_{\nu})=\mathcal{D}(h(x_{j+1}\cdots x_{N}))ep_{\nu}.$
If $h$ corresponds to a switch between two distinct variables in $\{x_{1},\ldots,x_{j}\}$, then $h(x_{j+1}\cdots\\ x_{N})=x_{j+1}\cdots x_{N}$, so we would have that $\mathcal{D}(x_{j+1}\cdots x_{N})ep_{\nu}$ is invariant under the action of $h$.
Furthermore, each operator $\mathcal{D}_{j}$ either deletes the factor of $x_{j}$ in $e$ or applies $\partial_{j}$ to some $p_{\nu_{l}}$. Assume that there exists $j\in\{m_{2}+1,\ldots,N\}$ such that $\mathcal{D}_{j}$ applies $\partial_{j}$ to some $p_{\nu_{l}}$. Let $j_{\text{max}}$ be the largest such $j$. After applying $\mathcal{D}_{j_{\text{max}}}$, the exponent of $x_{j_{\text{max}}}$ will be positive. Then, we have that we will never be able to remove the factor of $x_{j_{\text{max}}}$. When applying $\mathcal{D}_{j}$ for some $j\in\{1,\ldots,j_{\text{max}}-1\}$, we will not be able to remove the factor with a switch since the exponent of $x_{j}$ is also positive. Furthermore, there are no switches to $j_{\text{max}}$ in the operators in $\mathcal{D}(p_{\lambda})$ by assumption. Then, for the contribution to be nonzero, we delete $x_{j}$ for each $j\in\{m_{2}+1,\ldots,N\}$, which results in multiplying by a factor of
$\prod_{j=1}^{N-m_{2}}(1+2(j-1)\theta)=\prod_{j=N-k}^{N-m_{2}}(1+2(j-1)\theta)\prod_{j=1}^{N-k}(1+2(j-1)\theta);$
note that $m_{2}\leq k$. The leading order term of this expression is
(20) $(2\theta N)^{k-m_{2}}\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
Afterwards, the expression we must evaluate is
$\mathcal{D}(p_{\lambda})\mathcal{D}_{1}\cdots\mathcal{D}_{m_{2}}x_{1}\cdots x_{m_{2}}p_{\nu}.$
Let $d_{1}$ be the number of derivatives in $\mathcal{D}(p_{\lambda})$, which does not include the $\ell(\lambda)$ derivatives at the starts of the blocks. Furthermore, let $d_{2}$ be the number of derivatives in $\mathcal{D}_{j}$ for $j\in[m_{1}]$ and let $d_{3}$ be the number of derivatives in $\mathcal{D}_{j}$ for $j\in[m_{2}]\backslash[m_{1}]$. The total number of derivatives is $d_{1}+d_{2}+d_{3}+\ell(\lambda)$; recall that this quantity must be at least $\ell(\nu)$ for the contribution to be nonzero.
We have that the number of choices for the set $m_{2}]$ is $O(N^{m_{2}})$; afterwards, the number of ways to choose the operators in $\mathcal{D}(p_{\lambda})$ is $O(1)$. Note that if we choose a new set to replace $m_{2}]$, then we can reorder the operators in $\mathcal{D}_{1}\cdots\mathcal{D}_{N}$ accordingly using the commutativity property of the Dunkl operators.
The total number of switches is $k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)$, which will multiply the contribution a factor of $(2\theta)^{k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)}$. Furthermore, the operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ will contribute an additional factor of $O(N^{m_{2}-d_{2}-d_{3}})$ for deciding the indices of their switches. Thus, using (20) gives that the total contribution is, without including $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$,
$O(N^{2m_{2}-d_{2}-d_{3}}\theta^{k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k-m_{2}})=O(N^{m_{2}-d_{2}-d_{3}}\theta^{k-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k}).$
Additionally, observe that the number of switches in $\mathcal{D}(p_{\lambda})$ is $k-d_{1}-\ell(\lambda)$, so
$m_{2}-m_{1}\leq k-d_{1}-\ell(\lambda).$
Therefore, the total contribution is
$O(N^{m_{1}}(\theta N)^{k-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k}).$
Since $m_{1}\leq\ell(\lambda)$ and $d_{1}+d_{2}+d_{3}+\ell(\lambda)\geq\ell(\nu)$, the total contribution is $O(N^{\ell(\lambda)}(\theta N)^{2k-\ell(\nu)})$.
To achieve the leading order term, we must have that $m_{1}=\ell(\lambda)$, meaning that the $j_{i}$ are distinct, and that $d_{1}+d_{2}+d_{3}+\ell(\lambda)=\ell(\nu)$, meaning that the total number of derivatives is $\ell(\nu)$ and therefore that each derivative is assigned to a distinct $p_{\nu_{l}}$.
In addition, we must have have that for $j\in[m_{2}]$, if $\mathcal{D}_{j}$ is not $\partial_{j}$, then it is equivalent to $2\theta Nd_{j}$; this is based on the factor of $O(N^{m_{2}-d_{2}-d_{3}})$ mentioned previously. That is, each of these $\mathcal{D}_{j}$ will delete the factor of $x_{j}$ in $e$, which will multiply the leading order term of the contribution by $2\theta N$. It is also possible that $\mathcal{D}_{j}$ is the switch from $j$ to $j^{\prime}$ such that $\mathcal{D}_{j^{\prime}}$ is a derivative which has already been assigned to some $p_{\nu_{l}}$ so that the exponent of $x_{j^{\prime}}$ is positive before applying $\mathcal{D}_{j}$. Then, $\mathcal{D}_{j}$ will not delete the factor of $x_{j}$, however the number of choices for $j^{\prime}$ is $O(1)$ so this case will not contribute to the leading order term.
Furthermore, we must have that $m_{2}-m_{1}=k-d_{1}-\ell(\lambda)$, which means that the switches in $\mathcal{D}(p_{\lambda})$ are to distinct indices which are also not elements of $\{j_{i}\}_{1\leq i\leq\ell(\lambda)}$. Since $m_{1}=\ell(\lambda)$, this is equivalent to $m_{2}=k-d_{1}$.
Furthermore, for the remainder term $R$, recall that to delete the variables $x_{j}$ for $j\in\{m_{2}+1,\ldots,N\}$, we multiply by a factor of
$\prod_{j=N-k+1}^{N-m_{2}}(1+2(j-1)\theta)\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
In particular, after we remove the term $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$, we obtain a polynomial with degree that is independent of $N$. Afterwards, we can deduce that the remainder term $R$ is a polynomial in $\mathbb{Q}[x,y]$ by using casework. Furthermore, for (1), we still have that the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$, by the argument that after selecting $j_{1},\ldots,j_{\ell(\lambda)}$ and deleting the variables $x_{j}$ for $j\in\{m_{2}+1,\ldots,N\}$, each factor of $N$ must be accompanied by a factor of $\theta$. For (2), we also have that in $R$, the degree of $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$, since we have removed the leading order term which is of order $N^{\ell(\lambda)}(\theta N)^{2k-\ell(\nu)}$.
Let $\mathcal{S}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
  1. (1)
[…]
  2. (2)
    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-d_{1}(s)-\ell(\lambda)$, where $d_{1}(s)\leq\ell(\nu)-\ell(\lambda)$ is the number of derivatives in $s$ other than the ones that appear at $1+\lambda_{1}+\cdots+\lambda_{i-1}$ for some $i\in[\ell(\lambda)]$.
The sequence $s\in\mathcal{S}$ records the locations of the derivatives, since the indices of the switches are fixed after the locations are selected. Then, we have that the leading order term is
$\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{k}\sum_{s\in\mathcal{S}}N^{-d_{1}(s)}(2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{1}\circ\cdots\circ\mathcal{D}_{k-d_{1}(s)}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
$=\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{k}\sum_{s\in\mathcal{S}}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
Note that we have reversed the order of $\mathcal{D}_{1},\ldots,\mathcal{D}_{k-d_{1}(s)}$ using the commutativity of the Dunkl operators. The purpose of this is to simplify the notation; the actual content of the proof does not change.
We explain how we arrive at this formula. First, observe that $m_{2}=k-d_{1}(s)$. Then, recall that the term $(2\theta N)^{d_{1}(s)}$ is included in (20) as $(2\theta N)^{k-m_{2}}$ and that the term $N^{k-d_{1}(s)}$ is from the number of choices for $m_{2}]=[k-d_{1}(s)]$, which can be replaced by any ordered list of $k-d_{1}(s)$ distinct elements of $[N]$. It suffices to compute
(21) $\sum_{s\in\mathcal{S}}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
For $s\in\mathcal{S}$, let $\mathcal{R}^{\prime}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
  1. (1)
    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or the switch from $i$ to an element of $\{k-d_{1}(s)+1,\ldots,N\}\cup[i-1]$.
  2. (2)
    The total number of derivatives in $r$ is $\ell(\nu)-\ell(\lambda)-d_{1}(s)$.
Then, the leading order term of (21) is
$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}^{\prime}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
Recall that for $r\in\mathcal{R}(s)$, if $r_{i}$ is a switch, then we may regard it as $2\theta Nd_{i}$ so that $s$ and $r$ contribute to the leading order term. Therefore, we modify the definition of $\mathcal{R}^{\prime}(s)$ as follows.
For $s\in\mathcal{S}$, let $\mathcal{R}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
  1. (1)
    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or $2\theta d_{i}$.
  2. (2)
    The total number of derivatives in $r$ is $\ell(\nu)-\ell(\lambda)-d_{1}(s)$.
Definition 8.17.
For $r\in\mathcal{R}(s)$, let $d_{2}(r)$ denote the number of derivatives among the operators $r_{i}$ for $i\in[\ell(\lambda)]$ and $d_{3}(r)$ denote the number of derivatives among the operators $r_{i}$ for $\ell(\lambda)+1\leq i\leq k-d_{1}(s)$. Furthermore, let $d(r)\triangleq d_{2}(r)+d_{3}(r)$.
We have that $d(r)+d_{1}(s)=\ell(\nu)-\ell(\lambda)$ for $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$. The leading order term of (21) is
$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}N^{k-d_{1}(s)-d(r)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
$=\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}N^{k-\ell(\nu)+\ell(\lambda)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
Note that the term $N^{k-d_{1}(s)-d(r)}$ accounts for the factors of $N$ arising from the switches among the operators of $r$, the number of which is $k-d_{1}(s)-d(r)$. Furthermore, we have that $k-d_{1}(s)-d(r)=k-\ell(\nu)+\ell(\lambda)$. Therefore, we compute
(22) $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
For $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$, we let $\mathcal{P}(s,r)$ denote the pair $(S,R)$, where $S$ consists of the set of $i\in[\ell(\lambda)]$ such that $s_{i}$ is a derivative and $R$ consists of the set of $i\in[k-d_{1}(s)]$ such that $r_{i}$ is a derivative; note that $|\mathcal{P}(s,r)|=\ell(\nu)$. Let $\mathcal{H}(s,r)$ denote the set of bijections $\zeta:\mathcal{P}(s,r)\rightarrow[\ell(\nu)]$; for $\zeta\in\mathcal{H}(s,r)$, we let $\zeta_{S}$ and $\zeta_{R}$ denote $\zeta$ restricted to the first and second elements of $(S,R)$, respectively.
For $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(s,r)$, we similarly define $c(s,r,\zeta)$. We start with $c=1$. We apply $r_{i}$ from $i=1$ to $k-d_{1}(s)$ and then $s_{i}$ from $i=1$ to $|\lambda|$. Suppose the current operator is in $r$. If the operator is $\partial_{i}$ and is assigned to $p_{\nu_{l}}$ by $\zeta_{R}$, then we multiply $c$ by $x_{i}\partial_{i}p_{\nu_{l}}$. Otherwise, if the operator is $2\theta d_{i}$, then we multiply $c$ by $2\theta$. Suppose the operator is in $s$. If the operator is $\partial_{i}$ and is assigned to $p_{\nu_{l}}$ by $\zeta_{S}$, then we multiply $c$ by $\partial_{i}p_{\nu_{l}}$. Otherwise, if the operator is the switch from $i$ to $j$, then we apply the operator to $c$. The final value of $c$ is $c(s,r,\zeta)$. Afterwards, (22) equals
(23) $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s),\,\zeta\in\mathcal{H}(s,r)}(2\theta)^{d_{1}(s)}c(s,r,\zeta).$
Suppose $i\in[\ell(\lambda)]$. Define $A_{i}(s)$ to be the set of $j$ such that there is a switch from $i$ to $j$ in $s$. Then,
$\ell(\lambda)+1,\ldots,k-d_{1}(s)]=\bigsqcup_{i=1}^{\ell(\lambda)}A_{i}(s).$
Furthermore, define $S_{i}(s,r,\zeta)$ to be the set of $l\in[\ell(\nu)]$ such that some $\partial_{i}$ is assigned to $p_{\nu_{l}}$ or $\partial_{j}$ is assigned to $p_{\nu_{l}}$ for some $j\in A_{i}(s)$; recall that we may have that $r_{i}=\partial_{i}$.
Lemma 8.18.
Suppose $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(r,s)$. If $c(s,r,\zeta)\not=0$, then
$\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}=\lambda_{i}.$
for all $i\in[\ell(\lambda)]$.
Proof.
Suppose $i\in[\ell(\lambda)]$. Note that the only way to remove a factor of $x_{i}$ or a factor of $x_{j}$ for $j\in A_{i}(s)$ is with the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$. Based on the algorithm to compute $c$, the total degree in the variables $x_{i}$ and $x_{j}$ for $j\in A_{i}(s)$ that we increase $c$ by is $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$. The total degree that the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ lower $c$ by is $\lambda_{i}$, so $\lambda_{i}\geq\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$.
Conversely, the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ will only lower the degrees of the variables $x_{i}$ and $x_{j}$ for $j\in A_{i}(s)$. Since the operators lower $c$ by a total degree of $\lambda_{i}$ and the total degree in the variables that we add to $c$ is $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$, we must also have that $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}\geq\lambda_{i}$ for the contribution to be nonzero. ∎
Let $\mathcal{G}$ be the set of $(s,r,\zeta)$ for $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(r,s)$ such that $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. Suppose $(s,r,\zeta)\in\mathcal{G}$. Then, we define $C_{i}(s,r,\zeta)$ for $i\in[\ell(\lambda)]$ as the factor that $c$ is multiplied by when we apply
$s_{\lambda_{1}+\cdots+\lambda_{i}}\circ\cdots\circ s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}\circ\prod_{j\in A_{i}(s)}r_{j}\circ r_{i},$
where we apply the $r_{j}$s in $\prod_{j\in A_{i}(s)}r_{j}$ starting from lower values of $j$. Note that the blocks $s_{\lambda_{1}+\cdots+\lambda_{i}}\circ\cdots\circ s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}$ and $\prod_{j\in A_{i}(s)}r_{j}$ are consecutive in $s$ and $r$, respectively, although these two blocks and $r_{i}$ are not consecutive in (22). However, we can regard them as consecutive because the factors that operators associated with different $i\in[\ell(\lambda)]$ multiply $c$ by are independent. In particular, we have that
$c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}C_{i}(s,r,\zeta).$
For $i\in[\ell(\lambda)]$, define $d_{1}(s;\,i)$ to be the number of derivatives among $s_{\lambda_{1}+\cdots+\lambda_{i-1}+2},\ldots,\\ s_{\lambda_{1}+\cdots+\lambda_{i}}$; note that we do not include the derivative at $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}$. Then, we have that
(24) $(2\theta)^{d_{1}(s)}c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}(2\theta)^{d_{1}(s;\,i)}C_{i}(s,r,\zeta).$
Similarly to Lemma˜6.5, we have that conditioned on the values of $S_{i}(s,r,\zeta)$, we can factorize the contributions.
Lemma 8.19.
For $\nu^{\prime}\in\Gamma_{\text{even}}$, let $c_{\nu^{\prime}}$ denote the quantity $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s),\,\zeta\in\mathcal{H}(r,s)}(2\theta)^{d_{1}(s)}c(s,r,\\ \zeta)$ from (23) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$. Suppose $\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}$ such that $\sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$. Then,
$\sum_{\begin{subarray}{c}(s,r,\zeta)\in\mathcal{G},\\ S_{i}(s,r,\zeta)=S_{i},\,1\leq i\leq\ell(\lambda)\end{subarray}}(2\theta)^{d_{1}(s)}c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:\,l\in S_{i}))}.$
Proof.
We describe the general idea of the proof, which is the same as that of Lemma˜6.5. Suppose that for $i\in[\ell(\lambda)]$, we are given a sequence $(s_{i},r_{i},\zeta_{i})$ which has the structure for $(s,r,\zeta)$ when $\lambda$ is set as $(\lambda_{i})$ and $\nu$ is set as $\gamma((\nu_{l}:l\in S_{i}))$. Then, we can combine the $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ for $i\in[\ell(\lambda)]$ to uniquely form an ordered tuple $(s,r,\zeta)\in\mathcal{G}$ such that $S_{i}(s,r,\zeta)=S_{i}$ for $i\in[\ell(\lambda)]$; we combine the $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ based on the conditions in the definition of $\mathcal{S}$. We can also uniquely decompose the ordered tuple $(s,r,\zeta)\in\mathcal{G}$ such that $S_{i}(s,r,\zeta)=S_{i}$ for $i\in[\ell(\lambda)]$ into ordered tuples $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ for $i\in[\ell(\lambda)]$. In particular, $\mathbf{s}_{i}$ would correspond to $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ and $\mathbf{r}_{i}$ would correspond to $r_{j}$ for $j\in A_{i}(s)$ and $r_{i}$. Afterwards, using (24) and the fact that $C_{i}(s,r,\zeta)=c(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ finishes the proof. ∎
Thus, (22) equals
(25) $\begin{split}&\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}\\ &=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\sum_{\begin{subarray}{c}(s,r,\zeta)\in\mathcal{G},\\ S_{i}(s,r,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}(2\theta)^{d_{1}(s)}c(s,r,\zeta)\\ &=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:\,l\in S_{i}))}.\end{split}$
Lemma 8.20.
For $\nu^{\prime}\in\Gamma_{\text{even}}$,
$c_{\nu^{\prime}}=(2\theta)^{2|\nu^{\prime}|-\ell(\nu^{\prime})}\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC^{\prime}(|\nu^{\prime}|),\,b(\pi)=\nu^{\prime}}\mathbf{1}\{z(\pi)=0\}.$
Proof.
First, we compute the exponent of $2\theta$ in $(2\theta)^{d_{1}(s)}c(s,r,\zeta)$. The number of switches in $s$ is $|\nu^{\prime}|-d_{1}(s)-1$ and the number of switches in $r$ is $|\nu^{\prime}|-d_{1}(s)-d(r)=|\nu^{\prime}|-\ell(\nu^{\prime})+1$. The total number of switches is then $2|\nu^{\prime}|-\ell(\nu^{\prime})-d_{1}(s)$. Hence, the exponent of $2\theta$ is $2|\nu^{\prime}|-\ell(\nu^{\prime})$.
Following this, the general idea is the same as that of Lemma˜6.7, however there are two differences. The first difference is due to the possibility that $r_{1}=\partial_{1}$. This case would correspond to the first block $B_{1}$ having length $1+\nu_{l}$; note that the length is $1+\nu_{l}$ rather than $\nu_{l}$ since there is an additional factor of $x_{1}$, so after applying $\partial_{1}$ to $p_{\nu_{l}}$, we would obtain the term $\nu_{l}x_{1}^{\nu_{l}}$ rather than $\nu_{l}x_{1}^{\nu_{l}-1}$. However, since $s_{1}=\partial_{1}$, the second block $B_{2}$ would start at $2$. On the other hand, if $r_{1}=2\theta d_{1}$, then that would correspond to the first block $B_{1}$ having length $1$; in this case, the remaining blocks would be a noncrossing partition of $\{2,\ldots,|\nu^{\prime}|+1\}$.
The second difference is due to the operators $r_{j}$ for $j\in A_{1}(s)$ which are derivatives. If $j\in A_{1}(s)$ and $r_{j}=\partial_{j}$ is assigned to $p_{\nu_{l}}$, after applying $r_{j}$ to $p_{\nu_{l}}$ we would obtain a factor of $\nu_{l}x_{j}^{\nu_{l}}$, where the exponent is $\nu_{l}$ rather than $\nu_{l}-1$ since there is an additional factor of $x_{j}$. Afterwards, when we apply the switch from $1$ to $j$, we will need to delete the power of $x_{j}$. However, note that the exponent of $x_{1}$ must be zero when we apply this switch, otherwise we will not be able to delete the power of $x_{j}$ and the contribution will be 0. Furthermore, in order to delete the power of $x_{j}$ when we apply the switch, we will replace $x_{j}^{\nu_{l}}$ with $2\theta x_{1}^{\nu_{l}-1}$; this is straightforward to deduce because $\nu_{l}$ is even. The outcome is then equivalent to applying $2\theta\partial_{1}$ to $p_{\nu_{l}}$.
We explain how this impacts the bijective correspondence. Suppose $\pi\in NC^{\prime}([|\nu^{\prime}|])$ and $i\in[|\nu^{\prime}|+1]$ such that $i\geq 3$ and $i$ is the minimal value of a block of $\pi$ of length $L$; recall that $i=1$ and $2$ correspond to $r_{1}$ and $s_{1}$, respectively. Then, we would have that $s_{i-1}$ is either a $\partial_{1}$ which is assigned to $p_{\nu_{l}}$ or a switch from $1$ to $j$ such that $r_{j}=\partial_{j}$ is assigned to $p_{\nu_{l}}$ for some $\nu_{l}$ such that $|\nu_{l}|=L$.
However, note that for $c(s,r,\zeta)$ to be nonzero, we can only choose to set $r_{j}=\partial_{j}$ for $j\in A_{1}(s)$ such that the exponent of $x_{1}$ is zero when applying the switch from $1$ to $j$. That is, we must select a location $i\in\{3,\ldots,|\nu^{\prime}|+1\}$ for the switch in $\pi$ such that $i$ is the minimal element of the block that contains it and all earlier blocks do not contain elements greater than $i$; in other words, we must have that $b(i;\,\pi)=1$ and $d(i;\,\pi)$ equal to the size of the block that contains it. If such a location $i$ in $\pi$ exists, then for $s_{i-1}$, we must choose $\partial_{1}$ or the switch from $1$ to $j$ and set $r_{j}=\partial_{j}$, where $j$ is determined by the previous switches. However, the latter option is equivalent to applying $2\theta\partial_{1}$ rather than $\partial_{1}$. We have observed that after multiplying by $(2\theta)^{d_{1}(s)}$, the exponent of $2\theta$ is constant. Therefore, when summing the contributions from selecting $2\theta\partial_{1}$ and $\partial_{1}$, we will obtain zero. Hence, any $\pi\in NC^{\prime}(|\nu^{\prime}|)$ with $z(\pi)\geq 1$ will have a contribution of zero.
On the other hand, if $\pi\in NC^{\prime}(|\nu^{\prime}|)$ satisfies $z(\pi)=0$, then the contribution corresponding to it is $(2\theta)^{2|\nu^{\prime}|-\ell(\nu^{\prime})}\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}$. ∎
Using (25) and Lemma˜8.20, it suffices to compute
$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}(2\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))(\#\,\pi\in NC^{\prime}(\lambda_{i})\text{ such that }z(\pi)=0\text{ and }\gamma(b(\pi))=\gamma((\nu_{j}:j\in S_{i})))$
$=\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}(2\theta)^{2k-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\prod_{i=1}^{\ell(\lambda)}(\#\,\pi\in NC^{\prime}(\lambda_{i})\text{ such that }z(\pi)=0\text{ and }\gamma(b(\pi))=\gamma_{i})$
$=(2\theta)^{2k-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i}),\,z(\pi)=0}\prod_{B\in b(\pi)}x_{|B|}.$
Suppose $k\geq 2$ and $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC^{\prime}(k)$, where the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $q\in[m-1]$. Then, we show that $z(\pi)=0$ if and only if $k+1\in B_{1}\cup B_{2}$. If $k+1\in B_{1}$, then clearly $z(\pi)=0$; if $k+1\in B_{2}$, then $B_{1}=\{1\}$ and it is also the case that $z(\pi)=0$. Next, assume that $k+1\notin B_{1}\cup B_{2}$. If $B_{1}\not=\{1\}$, let $M$ be the maximal element of $B_{1}$. Then, $M<k+1$ and it is evident that $b(M+1;\,\pi)=0$ and $d(M+1;\,\pi)$ equals the size of the block that contains $M+1$. Therefore, $z(\pi)\geq 1$. Next, assume that $B_{1}=\{1\}$. If we let $M$ be the maximal element of $B_{2}$, then we can similarly deduce that $b(M+1;\,\pi)=0$ and $d(M+1;\,\pi)$ equals the size of the block that contains $M+1$ to conclude that $z(\pi)\geq 1$.
Afterwards, we observe that $b:\{\pi\in NC^{\prime}(k):z(\pi)=0\}\rightarrow NC(\{2,\ldots,k+1\})$ is a bijection. To show that $b$ is surjective, suppose $\pi^{\prime}=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(\{2,\ldots,k+1\})$, the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $q\in[m-1]$, and $k+1\in B_{i}$ for $i\in[m]$. If $i=1$, then let $\pi=\{1\}\sqcup\pi^{\prime}$. If $i>1$, then let $\pi$ be $\pi^{\prime}$ with $B_{i}$ replaced by $\{1\}\cup B_{i}$. In both cases, we have that $\pi\in NC^{\prime}(k)$, $z(\pi)=0$, and $b(\pi)=\pi^{\prime}$.
Next, we show that $b$ is injective. We use the same notation for $\pi^{\prime}$. Suppose $\pi\in NC^{\prime}(k)$, $z(\pi)=0$, and $b(\pi)=\pi^{\prime}$. If $i=1$, then we must have that the block containing $1$ in $\pi$ is $\{1\}$ because the block containing $2$ also contains $k+1$. Hence, there is a unique choice for $\pi$. Next, assume that $i>1$. Then, we must have that the block that contains $1$ in $\pi$ also contains $k+1$. Therefore, the block that contains $1$ in $\pi$ is $B_{i}\cup\{1\}$ and we similarly have that there is a unique choice for $\pi$.
Hence, the value of (25) is
$(2\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}.$
After multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{2k-\ell(\nu)+\ell(\lambda)}$, we obtain the formula in the statement of the theorem.
8.4. Proof of part (C) of Theorem˜1.1
We prove part (C) of Theorem˜1.1, see Corollary˜8.22, as well as a generalization, see Theorem˜8.21. Recall that $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$ for $\lambda,\nu\in\Gamma_{\text{even}}$.
Theorem 8.21.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=d_{\lambda}\in\mathbb{C}$.
  3. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{D^{N}(\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{M}^{D}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
  4. (d)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{(\theta N)^{|\nu|}N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{M}^{D}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
Then, (a) and (c) are equivalent.
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) and (d) are equivalent.
Proof.
See Theorem˜8.11 for the equivalence of (a) and (c). For the equivalence of (b) and (d), we can use Theorem˜8.16 and follow the method of the proof of Theorem˜6.10, after dividing by the nonzero quantity $\prod_{j=1}^{N}(1+2(j-1)\theta)$. ∎
Corollary 8.22.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
  2. (b)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
  3. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{D^{N}(\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
  4. (d)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}(\theta N)^{|\nu|}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
Then, (a) and (c) are equivalent.
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) and (d) are equivalent.
8.5.
Next, we consider the asymptotics of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ in the $\theta N\rightarrow c\in\mathbb{C}$ regime. Afterwards, we arrive at results similar to those of the previous subsection for this regime, see Theorem˜8.26 and Corollary˜8.27.
For $\pi\in NC(k+1)$, define
$W^{D;\,\text{odd}}(\pi)(x)\triangleq$ $(1+\mathbf{1}\{b(2;\,\pi)=0\}(d(2;\,\pi)-1))\times$
$\prod_{i\in[k+1],\,i\geq 3,\,b(i;\,\pi)=0}(x+d(i;\,\pi))\left(\frac{1}{1+x}\right)^{z(\pi)+\mathbf{1}\{f(\pi)>1\}}.$
Theorem 8.23.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=\prod_{i=1}^{N-k}(1+2(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(1+2N\theta)^{k}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(2N\theta)\prod_{B\in b(\pi)}x_{|B|}$
$+R(N,\theta)\bigg),$
where $R\in\mathbb{Q}[x,y]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
Proof.
We follow the same framework as the proof of Theorem˜8.16. The only difference is that in this case, we view $\theta N$ as having the same weight as a constant to compute the leading order term.
We use the same notation for $m_{1}$, $m_{2}$, $d_{1}$, $d_{2}$, and $d_{3}$. Define $d_{1}^{\prime}$, $d_{2}^{\prime}$, and $d_{3}^{\prime}$ to be the number of derivatives in $\mathcal{D}(p_{\lambda})$ other than the $\ell(\lambda)$ derivatives at the starts of the blocks, $\mathcal{D}_{j}$ for $j\in[m_{1}]$, and $\mathcal{D}_{j}$ for $j\in[m_{2}]\backslash[m_{1}]$, respectively, that we assign to some $p_{\nu_{l}}$. Then, we have that $d^{\prime}_{1}+d^{\prime}_{2}+d^{\prime}_{3}=\ell(\nu)-\ell(\lambda)$ and $d_{i}^{\prime}\leq d_{i}$ for $i\in[3]$.
The factor of $\prod_{j=1}^{N-m_{2}}(1+2(j-1)\theta)$ remains the same. However, its leading order term is now
(26) $(1+2\theta N)^{k-m_{2}}\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
The number of ways to select $m_{2}]$ is $O(N^{m_{2}})$ and the number of ways to choose the operators in $\mathcal{D}(p_{\lambda})$ is $O(1)$. Furthermore the switches in $\mathcal{D}(p_{\lambda})$ multiply the contribution by a factor of $(2\theta)^{k-d_{1}-\ell(\lambda)}$. The operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ which are not a derivative assigned to some $p_{\nu_{l}}$ will contribute a factor of $(1+2N\theta)^{m_{2}-d_{2}^{\prime}-d_{3}^{\prime}}$. Therefore, using (26) gives that the total contribution is, without including $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$,
$O(N^{m_{2}}\theta^{k-d_{1}-\ell(\lambda)}(1+2N\theta)^{k-d_{2}^{\prime}-d_{3}^{\prime}}).$
We still have that $m_{2}-m_{1}\leq k-d_{1}-\ell(\lambda)$, so the total contribution is
$O(N^{m_{1}}(\theta N)^{k-d_{1}-\ell(\lambda)}(1+2N\theta)^{k-d_{2}^{\prime}-d_{3}^{\prime}}).$
Furthermore, $m_{1}\leq\ell(\lambda)$, so multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$ gives the leading order term, which is $O(N^{\ell(\lambda)})\prod_{j=1}^{N-k}(1+2(j-1)\theta)$.
To achieve the leading order term, we must have that $m_{2}-m_{1}=k-d_{1}-\ell(\lambda)$, which means that the switches in $\mathcal{D}(p_{\lambda})$ are to distinct indices not in $m_{1}]$, and that $m_{1}=\ell(\lambda)$, which means that the $j_{i}$ are distinct. Furthermore, each of the operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ which is not a derivative assigned to some $p_{\nu_{l}}$ should delete $x_{j}$ by either applying $\partial_{j}$ or $2N\theta d_{j}$.
Let $\mathcal{S}^{\prime}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
  1. (1)
[…]
  2. (2)
    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-d_{1}(s)-\ell(\lambda)$, where $d_{1}(s)$ is the number of derivatives in $s$ other than the ones that appear at $1+\lambda_{1}+\cdots+\lambda_{i-1}$ for some $i\in[\ell(\lambda)]$.
The sequence $s\in\mathcal{S}^{\prime}$ records the locations of the derivatives, since the indices of the switches are fixed after the locations are selected. Then, we have that the leading order term is
$\prod_{j=1}^{N-k}(1+2(j-1)\theta)\sum_{s\in\mathcal{S}^{\prime}}N^{k-d_{1}(s)}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ$
$\mathcal{D}_{1}\circ\cdots\circ\mathcal{D}_{k-d_{1}(s)}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
$=\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}\sum_{s\in\mathcal{S}^{\prime}}N^{k-d_{1}(s)-\ell(\lambda)}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ$
$\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
Recall that since $m_{2}=k-d_{1}(s)$, the term $(1+2\theta N)^{d_{1}(s)}$ is included in (20) as $(1+2\theta N)^{k-m_{2}}$ and the term $N^{k-d_{1}(s)}$ is from the number of choices for $m_{2}]=[k-d_{1}(s)]$, which can be replaced by any ordered list of $k-d_{1}(s)$ distinct elements of $[N]$.
Define $\mathcal{S}$ to be the set of $s\in\mathcal{S}^{\prime}$ with each switch multiplied by a factor of $N$; this is equivalent to replacing $\theta$ with $\theta N$. Then, since the number of switches in $s\in\mathcal{S}^{\prime}$ is $k-d_{1}(s)-\ell(\lambda)$, the leading order term is
$\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}\sum_{s\in\mathcal{S}}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
For $s\in\mathcal{S}$, let $\mathfrak{R}^{\prime}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
  1. (1)
    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or $(1+2\theta N)d_{i}$.
  2. (2)
    If $1\leq i\leq k-d_{1}(s)$ and $r_{i}=\partial_{i}$, then $r_{i}$ must be assigned to some $p_{\nu_{l}}$.
Note that we use a slight abuse of notation, since condition (2) should be a condition on the injective function $\zeta$ which assigns the derivatives of $s\in\mathcal{S}$ and $r\in\mathfrak{R}^{\prime}(s)$ to the $p_{\nu_{l}}$ rather than a condition on $r\in\mathfrak{R}^{\prime}(s)$. However, we avoid using the notation for $\zeta$ for simplicity.
Then, the leading order term, without including $\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}$, is
(27) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}^{\prime}(s)}(1+2N\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
The reasoning behind conditions (1) and (2) is that if $r_{i}$ is not a derivative which is assigned to some $p_{\nu_{l}}$, then to contribute to the leading order term, it must delete the variable $x_{i}$ and multiply the contribution by a factor of $1+2N\theta$. In particular, we have that each derivative in $r\in\mathfrak{R}^{\prime}(s)$ must be assigned to some $p_{\nu_{l}}$, while this is not necessarily true for a derivative in $s\in\mathcal{S}$.
For $s\in\mathcal{S}$, let $\mathfrak{R}(s)$ denote the set of $r\in\mathfrak{R}^{\prime}(s)$ with each of its operators multiplied by a factor of $\frac{1}{1+2\theta N}$. After using (27), the leading order term, without including $\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}$, is
$(1+2N\theta)^{k}\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
Therefore, we compute
(28) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
To finish the proof, we can use the framework as the proof of Theorem˜8.16. In particular, for $\nu^{\prime}\in\Gamma_{\text{even}}$, let $c_{\nu^{\prime}}$ denote the quantity (28) when $\lambda$ is set as $(|\nu^{\prime}|)$ and $\nu$ is set as $\nu^{\prime}$. We similarly have that
(29) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:l\in S_{i}))}.$
Lemma 8.24.
For $\nu^{\prime}\in\Gamma_{\text{even}}$,
$c_{\nu^{\prime}}=\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC(|\nu^{\prime}|+1),\,\gamma(b(\pi))=\nu^{\prime}}W^{D;\,\text{odd}}(\pi)(2N\theta).$
Proof.
The proof follows the method of the proof of Lemma˜8.20. First, we state the bijective correspondence. Suppose $s\in\mathcal{S}$, $r\in\mathfrak{R}(s)$, and $(s,r,\zeta)$ corresponds to $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(k+1)$, where the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $1\leq q\leq m-1$; we describe $\pi$ based on $(s,r,\zeta)$. We have not defined $\zeta$ yet; in this setting, $\zeta$ assigns each derivative in $r$ and some of the derivatives in $s$ to a $p_{\nu^{\prime}_{l}}$ so that each $p_{\nu^{\prime}_{l}}$ for $l\in[\ell(\nu^{\prime})]$ is assigned to once.
The block $B_{1}$ corresponds to the operator $r_{1}$; if $r_{1}=\frac{1}{1+2N\theta}\partial_{1}$, then $r_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$ and $B_{1}$ has size $|\nu^{\prime}_{l}|+1$. Otherwise, $r_{1}=d_{1}$ and $B_{1}$ has size $1$. Afterwards, for $1\leq i\leq|\nu^{\prime}|$, when $s_{i}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$ or when $s_{i}$ is a switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to some $p_{\nu^{\prime}_{l}}$, then there is a block in $\pi$ with size $\nu^{\prime}_{l}$ and minimal element $i+1$. Note that in contrast with Lemma˜8.20, we use $NC(k+1)$ rather than $NC^{\prime}(k)$. The reason for this is that $s_{1}=\partial_{1}$ is not necessarily assigned to a $p_{\nu^{\prime}_{l}}$, so $B_{2}$ does not necessarily have minimal element $2$.
Next, we justify the weight $W^{D;\,\text{odd}}(\pi)(2N\theta)$ for each $\pi\in NC(k+1)$. First, the reasoning for the coefficient of $\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}$ is the same.
If $r_{1}=\frac{1}{1+2N\theta}\partial_{1}$, then we must multiply the contribution by a factor of $\frac{1}{1+2N\theta}$; this is the reasoning for the factor $(1+2N\theta)^{-\mathbf{1}\{f(\pi)>1\}}$. Otherwise, if $r_{1}=d_{1}$, then we do not need to multiply by an additional factor.
Suppose $i\geq 3$ and $b(i;\,\pi)=0$. Then, $s_{i-1}$ is a switch or an unassigned derivative. If $s_{i-1}$ is the switch from $1$ to $j$, then because $b(i;\,\pi)=0$, $r_{j}=d_{j}$. In this case, we would multiply the contribution by a factor of $2N\theta+d(i;\,\pi)$, where the term $2N\theta$ is due to $s_{i-1}$.
Suppose $b(2;\,\pi)=0$. Then, because $s_{1}=\partial_{1}$ is unassigned, we would multiply the contribution by a factor of $d(2;\,\pi)$.
Next, suppose $i\geq 2$ and $b(i;\,\pi)=1$. Suppose $d(i;\,\pi)$ is greater than the size of the block that contains $i$. Then, we would have that $s_{i-1}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$, so we would multiply the contribution by a factor of $1$. Note that we could also have that $s_{i-1}$ is the switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to some $p_{\nu^{\prime}_{l}}$, but because $d(i;\,\pi)$ is greater than the size of the block that contains $i$, the contribution in this case would be zero. As mentioned earlier, this is due to the fact that the degree of $x_{1}$ before applying $s_{i-1}$ is positive, so we will not be able to eliminate the factor of $x_{j}$ arising from $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$.
Next, suppose $d(i;\,\pi)$ equals the size of the block that contains it; recall that this is equivalent to all earlier blocks having no elements greater than $i$. Then, if $s_{i-1}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$, we would similarly multiply the contribution by a factor of $1$. On the other hand, if $s_{i-1}$ is the switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to $p_{\nu^{\prime}_{l}}$, then we would multiply the contribution by a factor of $\frac{2N\theta}{1+2N\theta}$. Recall that in this case, $s_{i-1}$ must modify $x_{j}^{\nu^{\prime}_{l}}$ to a multiple of $x_{i}^{\nu^{\prime}_{l}-1}$; in order to do so, we will obtain $2N\theta x_{i}^{\nu^{\prime}_{l}-1}$ since $\nu^{\prime}_{l}$ is even. Since $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ has a factor of $\frac{1}{1+2N\theta}$, we obtain a contribution of $\frac{2N\theta}{1+2N\theta}$. Then, the total contribution in this case is $1-\frac{2N\theta}{1+2N\theta}=\frac{1}{1+2N\theta}$, which is the reasoning for the factor $(1+2N\theta)^{-z(\pi)}$. ∎
Using (29) and Lemma˜8.24, it suffices to compute
$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}+1),\\ \gamma(b(\pi))=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W^{D;\,\text{odd}}(\pi)(2N\theta)$
$=\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}+1),\\ \gamma(b(\pi))=\gamma_{i}\end{subarray}}W^{D;\,\text{odd}}(\pi)(2N\theta)$
$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(2N\theta)\prod_{B\in b(\pi)}x_{|B|}.$
After multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)(1+2N\theta)^{k}N^{\ell(\lambda)}$, we obtain the formula in the statement of the theorem. ∎
Define the infinite dimensional matrix $\mathcal{W}^{D;\,\text{odd}}(x)\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}[x]$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$,
$\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)\triangleq(1+x)^{|\nu|}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(x)\prod_{B\in b(\pi)}x_{|B|}$
The following lemma is straightforward to deduce.
Lemma 8.25.
Suppose $\lambda,\nu\in\Gamma_{\text{even}}$ and $|\lambda|=|\nu|$.
  1. (A)
    If $\lambda=\nu$, then $\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)=(1+x)^{|\nu|}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}W^{D;\,\text{odd}}(\{1\}\cup\{2,\ldots,\nu_{l}+1\})(x)$.
  2. (B)
    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)=0$.
Theorem 8.26.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
  3. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{D^{N}(\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{A}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
  4. (d)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
Proof.
For the implications involving (a) and (c), see Theorem˜8.13. Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Using Theorem˜8.23 gives that $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ is divisible by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$. Then, after dividing by this nonzero quantity, we can use Theorem˜8.23 and the same analysis as in the proof of Theorem˜6.17 to deduce that (b) implies (d). Furthermore, we similarly have that if $2c$ is not a negative integer, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]$ is invertible for all $k\geq 1$, so (d) implies (b) if this is the case. ∎
Corollary 8.27.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}c_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
  2. (b)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}d_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
  3. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{D^{N}(\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{A}(\pi)(2c)\prod_{B\in\pi}|B|c_{(|B|)}.$
  4. (d)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}$
    $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(|\nu|+1)}W^{D;\,\text{odd}}(\pi)(2c)\prod_{B\in b(\pi)}(1+2c)^{|B|}|B|d_{(|B|)},$
    where $d_{(k)}=0$ for odd positive integers $k$.
Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
Remark 8.28.
As mentioned in Remark˜8.14, the equivalences between (a) and (c) of Theorems˜8.21 and 8.26 and Corollaries˜8.22 and 8.27 follow from the analogous results for the type BC root system after setting $\theta_{1}=0$.
9. Asymptotics of the coefficients of Bessel functions
In this section, we determine the asymptotics of the coefficients of Bessel functions in each of the regimes mentioned in Subsection˜1.2. Note that we determine the asymptotics for the coefficients of the terms with a fixed degree and the analyses for different degrees are separate. Therefore, additional assumptions are required to deduce the convergence of the Bessel functions themselves rather than only their coefficients. For more discussion about this direction, see Subsection˜10.3.
The basic idea of the proofs is to determine the asymptotics of the inverse of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma[k]}$ and the analogous matrices for the $BC^{N}$ and $D^{N}$ root systems so that we can apply (4).
9.1. Coefficients for the $A^{N-1}$ root system
First, we consider the regime $|\theta N|\rightarrow\infty$.
Lemma 9.1.
Assume that $|\theta N|\rightarrow\infty$. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}\\ )_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}(\mathcal{M}^{A}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{M}^{A}[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{M}^{A}[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{A}[k]_{\lambda\nu}^{-1}=0$.
Proof.
(a): Define the diagonal matrices $D_{1}$ and $D_{2}$ with diagonal entry $(\gamma,\gamma)$ given by $N^{\ell(\gamma)}$ and $(\theta N)^{\ell(\gamma)}$, respectively, for all $\gamma\in\Gamma[k]$. Based on Theorem˜6.1, it is evident that
$M=(\theta N)^{k}D_{1}\left(\mathcal{M}^{A}[k]+o_{N}(1)\right)D_{2}^{-1}.$
To obtain that, consider
$\mathcal{M}^{\prime}=(\theta N)^{-k}D_{1}^{-1}MD_{2}.$
For $\lambda^{\prime},\nu^{\prime}\in\Gamma$, entry $(\lambda^{\prime},\nu^{\prime})$ of this matrix is $(\theta N)^{-k}N^{-\ell(\lambda^{\prime})}(\theta N)^{\ell(\nu^{\prime})}M_{\lambda^{\prime}\nu^{\prime}}$. If $\ell(\lambda^{\prime})\leq\ell(\nu^{\prime})$, then by Theorem˜6.1, the entry is $\mathcal{M}^{A}_{\lambda^{\prime}\nu^{\prime}}+o_{N}(1)$. Otherwise, if $\ell(\lambda^{\prime})>\ell(\nu^{\prime})$, the entry is
$O_{N}(N^{2(\ell(\nu^{\prime})-\ell(\lambda^{\prime}))}\theta^{\ell(\nu^{\prime})-\ell(\lambda^{\prime})})=O_{N}(N^{\ell(\nu^{\prime})-\ell(\lambda^{\prime})}(\theta N)^{\ell(\nu^{\prime})-\ell(\lambda^{\prime})})=o_{N}(1).$
Hence, $\mathcal{M}^{\prime}=\mathcal{M}^{A}[k]+o_{N}(1)$. It follows that
$M^{-1}=(\theta N)^{-k}D_{2}\left(\mathcal{M}^{A}[k]^{-1}+o_{N}(1)\right)D_{1}^{-1}.$
With this expression, we can approximate $M^{-1}_{\lambda\nu}$ and therefore $M_{\nu\lambda}^{-1}$ as well.
(b) and (c): Suppose the elements of $I\subset\Gamma[k]$ are consecutive based on the ordering of the rows and columns of $\mathcal{M}^{A}$; note that the rows and columns are ordered increasingly by length. Then, we have that
$\mathcal{M}^{A}[k][I,I]^{-1}=\mathcal{M}^{A}[k]^{-1}[I,I],$
where $\mathcal{M}^{A}[k][I,I]$ and $\mathcal{M}^{A}[k]^{-1}[I,I]$ denote $\mathcal{M}^{A}[k]$ and $\mathcal{M}^{A}[k]^{-1}$ with their rows and columns restricted to $I$, respectively. It is straightforward to prove (b) and (c) using this result. ∎
Corollary 9.2.
If $|\theta N|\rightarrow\infty$, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{A^{N-1}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma[k]}(\theta N)^{-k}\theta^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}\right)^{-1}(1+o_{N}(1))r(a)r(x)$
$+\sum_{r,s\in\Gamma[k],\,r\not=s}(\theta N)^{-k}(\theta N)^{\min(\ell(r),\ell(s))}N^{-\max(\ell(r),\ell(s))}\times$
$(\mathcal{M}^{A}[k]^{-1}_{rs}+\mathcal{M}^{A}[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
Proof.
See (4), Lemma˜6.9, and Lemma˜9.1. ∎
Next, we consider the regime $\theta N\rightarrow c\in\mathbb{C}$. Recall from Theorem˜6.17 that in order for the invertibility of $\mathcal{D}(A^{N-1}(\theta))$ to hold in an asymptotic sense, we must have that $c$ is not a negative integer. We require the invertibility of $\mathcal{D}(A^{N-1}(\theta))$ to compute the Bessel function, so we assume that $c$ is not a negative integer.
Lemma 9.3.
Assume that $\theta N\rightarrow c\in\mathbb{C}$ and that $c$ is not a negative integer. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}(\mathcal{W}^{A}(c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{A}(c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{A}(c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{A}(c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
(a): We can use the same argument as the proof of Lemma˜9.1 and Theorem˜6.14 to obtain that
$M=D(\mathcal{W}^{A}(c)[k]+o_{N}(1))$
where $D$ is a diagonal matrix with diagonal entry $(\gamma,\gamma)$ given by $N^{\ell(\gamma)}$ for $\gamma\in\Gamma[k]$. Hence,
$M^{-1}=D^{-1}(\mathcal{W}^{A}(c)[k]^{-1}+o_{N}(1)),$
which completes the proof since we can approximate $M^{-1}_{\lambda\nu}$ and therefore $M^{-1}_{\nu\lambda}$ as well. ∎
Corollary 9.4.
If $\theta N\rightarrow c\in\mathbb{C}$ and $c$ is not a negative integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{A^{N-1}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{A}([r_{l}])(c)\right)^{-1}(1+o_{N}(1))r(a)r(x)$
$+\sum_{r,s\in\Gamma[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}(\mathcal{W}^{A}(c)[k]^{-1}_{rs}+\mathcal{W}^{A}(c)[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
Proof.
See (4), Lemma˜6.16, and Lemma˜9.3. ∎
9.2. Coefficients for the $BC^{N}$ root system
The framework of this subsection is analogous to that of the $A^{N-1}$ root system.
Lemma 9.5.
Suppose $|\theta_{0}N|\rightarrow\infty$ and $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ such that $c\not=-1$. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $(\theta_{0}N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}(\mathcal{M}^{BC}(c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{M}^{BC}(c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{M}^{BC}(c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{BC}(c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜7.1. ∎
Corollary 9.6.
If $|\theta_{0}N|\rightarrow\infty$, $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$, and $c\not=-1$, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}(\theta_{0}N)^{-k}\theta_{0}^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}2^{r_{l}-1}r_{l}(1+c)^{o([r_{l}])}\right)^{-1}\times$
$(1+o_{N}(1))r(a)r(x)+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}(\theta_{0}N)^{-k}(\theta_{0}N)^{\min(\ell(r),\ell(s))}N^{-\max(\ell(r),\ell(s))}\times$
$(\mathcal{M}^{BC}(c)[k]^{-1}_{rs}+\mathcal{M}^{BC}(c)[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
Proof.
See (4), Lemma˜7.3, and Lemma˜9.5. ∎
Lemma 9.7.
Suppose $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$, $c_{0}$ is not a negative integer, and $2c_{0}+2c_{1}$ is not a negative odd integer. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}\\ )_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}(\mathcal{W}^{BC}(2c_{0},2c_{1})[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜7.6. ∎
Corollary 9.8.
If $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$, $c_{0}$ is not a negative integer, and $2c_{0}+2c_{1}$ is not a negative odd integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{BC}([r_{l}])(2c_{0},2c_{1})\right)^{-1}\times$
$(1+o_{N}(1))r(a)r(x)+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}\times$
$(\mathcal{W}^{BC}(c)[k]^{-1}_{rs}+\mathcal{W}^{BC}(c)[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
Proof.
See (4), Lemma˜7.7, and Lemma˜9.7. ∎
9.3. Coefficients for the $D^{N}$ root system
The framework of this subsection is analogous to that of the $A^{N-1}$ root system.
Lemma 9.9.
Suppose $|\theta N|\rightarrow\infty$. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{M}^{D}[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{M}^{D}[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{D}[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.10. ∎
Lemma 9.10.
Suppose $|\theta N|\rightarrow\infty$. Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$. The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
$(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)^{-1}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
Proof.
See Theorem˜8.16. ∎
Corollary 9.11.
If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}(\theta N)^{-k}\theta^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}2^{r_{l}-1}r_{l}\right)^{-1}\times$
$\left(1+o_{N}(1)+(1+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\right)r(a)r(x)$
$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}(\theta N)^{-k}(\theta N)^{\min(\ell(r),\ell(s))}N^{-\max(\ell(r),\ell(s))}\times$
$\bigg[\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1)$
$+(\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
Proof.
See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10. Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$. ∎
Lemma 9.12.
Suppose $\theta N\rightarrow c\in\mathbb{C}$ and $2c$ is not a negative integer. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}(\mathcal{W}^{A}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{A}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{A}(2c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{A}(2c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.12. ∎
Lemma 9.13.
Suppose $\theta N\rightarrow c\in\mathbb{C}$, $2c$ is not a negative integer, and assume that $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large. Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)^{-1}(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.23. ∎
Corollary 9.14.
If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}\right)^{-1}\bigg[\prod_{l=1}^{\ell(r)}W^{A}([r_{l}])(2c)^{-1}+o_{N}(1)$
$+\left(\prod_{l=1}^{\ell(r)}W^{D;\,\text{odd}}(\{1\}\cup\{2,\ldots,r_{l}+1\})(2c)^{-1}+o_{N}(1)\right)\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)r(x)$
$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}\bigg[\mathcal{W}^{A}(2c)[k]^{-1}_{rs}+\mathcal{W}^{A}(2c)[k]^{-1}_{sr}+o_{N}(1)$
$+(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{rs}+\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
Proof.
10. Applications
In this section, we discuss applications of the main results of this paper. Some settings that we consider include the weak convergence of the measures in ˜1.4 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
10.1. Weak convergence to the free convolution
The results of this subsection are based on the assumption of ˜1.4. Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems. In particular, we prove Corollary˜1.5. For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.7.
The following result describes the support of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$. As we mentioned earlier, [trimeche] shows that there exists a solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that is signed and is supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$. If we assume the conjecture, then we can reduce the domain of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ further, see Corollary˜10.1. We conjecture that there always exists a signed solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that has the same support as what is stated in the corollary.
Corollary 10.1.
Assume that ˜1.4 is true and that $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$. Then, for any choice of $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
Proof.
Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.2. Then, we have that
$\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}(\epsilon_{i}^{1}+\epsilon_{i}^{2})}d\mu_{a_{1}}^{\text{sym}}(\epsilon^{1})d\mu_{a_{2}}^{\text{sym}}(\epsilon^{2})=\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)d\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}(a).$
for all $x\in\mathbb{C}^{N}$, where we have used the fact that $\mu_{a}^{\text{sym}}$ is compactly supported for all $a\in\mathbb{R}^{N}$ to obtain the left hand side.
By considering $x\in i\mathbb{R}^{N}$, we have the characteristic functions of the random variables $\epsilon^{1}+\epsilon^{2}$ where $\epsilon^{1}\sim\mu_{a_{1}}^{\text{sym}}$ and $\epsilon^{2}\sim\mu_{a_{2}}^{\text{sym}}$ and $\epsilon$ where $a\sim\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ and $\epsilon\sim\mu_{a}^{\text{sym}}$ are equal. Therefore, these two random variables have the same distribution.
Suppose $u\in\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ and $r>0$. We show that the probability that $\epsilon$ is in $B(u,r)$ is positive to show that $u$ is in the support of $\epsilon$. First, observe that
$\Pr[\epsilon\in B(u,r)]\geq\Pr_{a\sim\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right].$
For the sake of contradiction, assume that $\Pr_{a\sim\mu_{a_{1},a_{2}},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right]=0$. Then, there exists a sequence $\{a_{i}\}_{i\geq 1}$ of elements of $B(u,r)$ such that
$\lim_{i\rightarrow\infty}\mu_{a_{i}}^{\text{sym}}\left[B(a_{i},\frac{r}{2})\right]=0,$
since $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}\left[B(u,\frac{r}{2})\right]>0$. Let $a$ be an element of the closed ball $\overline{B}(u,r)$ that is the limit of a subsequence $\{a_{i_{j}}\}_{j\geq 1}$ of $\{a_{i}\}_{i\geq 1}$.
Observe that for all $x\in\mathbb{R}^{N}$,
$\lim_{j\rightarrow\infty}J_{a_{i_{j}}}^{\mathcal{R}(\theta)}(ix)=J_{a}^{\mathcal{R}(\theta)}(ix).$
Hence, it follows that the characteristic functions of $\mu^{\text{sym}}_{a_{i_{j}}}$ converge pointwise to the characteristic function of $\mu^{\text{sym}}_{a}$. By Lévy’s continuity theorem, $\mu^{\text{sym}}_{a_{i_{j}}}$ converges weakly to $\mu^{\text{sym}}_{a}$. It then follows that
$\mu^{\text{sym}}_{a}\left(B(a,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a_{i_{j}},\frac{r}{2})\right)=0,$
which is a contradiction since $a\in\text{supp}(\mu_{a}^{\text{sym}})$ by Theorem˜1.2. Thus, $u$ is in the support of $\epsilon$.
However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$. Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.2, the result follows. ∎
In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$. Afterwards, it is straightforward to deduce Corollaries˜1.5 and 1.7.
Corollary 10.2.
Assume that ˜1.4 is true. Suppose that for $N\geq 1$, $a(N),\\ b(N)\in\mathbb{R}^{N}$. For all $N\geq 1$, assume that $\theta,\theta_{0},\theta_{1}\geq 0$.
  • (A)
    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{d}^{a}$ and $m_{d}^{b}$ are real numbers such that
    $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}b(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{b}$
    for $d\geq 1$. Let $a_{(k)}$ and $b_{(k)}$ for $k\geq 1$ be real numbers that solve
    $m_{d}^{a}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|a_{(|B|)}\text{ and }m_{d}^{b}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|b_{(|B|)}$
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{d}\right]}{(\theta N)^{d}N}=\sum_{\pi\in NC(d)}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
  • (B)
    Suppose $c\geq 0$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c$. Suppose $m_{2d}^{a}$ and $m_{2d}^{b}$ are real numbers such that
    $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{2d}}{(\theta_{0}N)^{2d}N}=m_{2d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}b(N)_{i}^{2d}}{(\theta_{0}N)^{2d}N}=m_{2d}^{b}$
    for $d\geq 1$. Let $a_{(2k)}$ and $b_{(2k)}$ for $k\geq 1$ be real numbers that solve
    $m_{2d}^{a}=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|a_{(|B|)}\text{ and}$
    $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta_{0}N)^{2d}N}$
    $=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
  • (C)
    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{2d}^{a}$ and $m_{2d}^{b}$ are real numbers such that
    $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{2d}}{(\theta_{0}N)^{2d}N}=m_{2d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}b(N)_{i}^{2d}}{(\theta_{0}N)^{2d}N}=m_{2d}^{b}$
    for $d\geq 1$. Let $a_{(2k)}$ and $b_{(2k)}$ for $k\geq 1$ be real numbers that solve
    $m_{2d}^{a}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|a_{(|B|)}\text{ and }$
    $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta N)^{2d}N}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
Remark 10.3.
We note that the proof of Corollary˜10.2 is straightforward and that the framework of the proof has appeared previously, for example see [matrix]*Proof of Theorem 1.5.
Proof of Corollary˜10.2.
(A): Suppose $N\geq 1$. Suppose $\Omega_{N}\triangleq B(0,r_{N})\subset\mathbb{C}^{N}$ for $r_{N}$ sufficiently small such that $J_{a(N)}^{A^{N-1}(\theta)}$ and $J_{b(N)}^{A^{N-1}(\theta)}$ are nonzero over $\Omega_{N}$. Note that since
$J_{a(N)}^{A^{N-1}(\theta)}(0)=J_{b(N)}^{A^{N-1}(\theta)}(0)=1,$
$\Omega_{N}$ exists. Then, we can express $J_{a(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}a_{\lambda}(N)p_{\lambda}\right)$ and $J_{b(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}b_{\lambda}(N)p_{\lambda}\right)$ over $\Omega_{N}$.
For $N\geq 1$, $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Corollary˜10.1 and is therefore exponentially decaying, see Definition˜10.15. Thus, part (C) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
(30) $\begin{split}\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]&=\left[p_{\lambda},J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}\right]_{A^{N-1}(\theta)}\\ &=\left[p_{\lambda},\exp\left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)\right]_{A^{N-1}(\theta)};\end{split}$
the second equality follows from the fact that $J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}=\exp\\ \left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)$ over $\Omega_{N}$ and $0\in\Omega_{N}$.
By Corollary˜6.11, we have that
$\lim_{N\rightarrow\infty}\frac{a_{(k)}(N)}{\theta N}=a_{(k)}\text{ and }\lim_{N\rightarrow\infty}\frac{b_{(k)}(N)}{\theta N}=b_{(k)}$
for $k\geq 1$ and
$\lim_{N\rightarrow\infty}\frac{a_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=\lim_{N\rightarrow\infty}\frac{b_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$
for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$. Using Corollary˜6.11 again and (30) then gives that for all $\lambda\in\Gamma$,
$\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]}{N^{\ell(\lambda)}(\theta N)^{|\lambda|}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}(a_{(|B|)}+b_{(|B|)}).$
(B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.22, respectively, and follow the same idea as the proof of (A). ∎
With this result, we are prepared to prove Corollary˜1.5.
Proof of Corollary˜1.5.
(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta(\frac{a_{i}}{\theta N})\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition. Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$. It follows from [billingsley]*Theorem 30.2 that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\right\}_{N\geq 1}$ converges in distribution to $\mu$.
(B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$. It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero. Using part (C) of Corollary˜10.2, we can then deduce that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(\frac{a_{i}}{2\theta N}\right)\right]\right\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$. Afterwards, we can follow the steps of (A). ∎
We can use the same framework to prove Corollary˜1.7.
Proof of Corollary˜1.7.
[…]
Suppose $d\geq 1$ and $\pi\in NC^{\text{even}}(2d)$. Recall that $o(\pi)$ is the number of $i\in[2d]$ such that $i$ is not the minimal element of a block of $\pi$ and $d(i;\,\pi)$ is odd. Suppose $\pi=B_{1}\sqcup\cdots\sqcup B_{m}$ such that $\min(B_{j})<\min(B_{j+1})$ for $1\leq j\leq m-1$. Suppose $i\in B_{j}$ for $j\in[m]$. We have that
$d(i;\,\pi)=\sum_{j^{\prime}=1}^{j}|B_{j^{\prime}}|-(i-1)\equiv i+1\pmod{2}$.
Hence, $d(i;\,\pi)$ is odd if and only if $i$ is even. It follows that $o(\pi)$ equals $d$ minus the number of even $i\in[2d]$ such that $b(i;\,\pi)=1$. If $E(\pi)$ denotes the number of even $i\in[2d]$ such that $b(i;\,\pi)=1$, then,
$(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|a_{(|B|)}=(1+c)^{-E(\pi)}\prod_{B\in\pi}(1+c)^{|B|}2^{|B|-1}|B|a_{(|B|)};$
[…]
10.2. Upper bounds of the magnitudes of Bessel functions
In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.2.
Lemma 10.4.
Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a,x\in\mathbb{C}^{N}$. Then,
$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}\leq\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
Proof.
First, observe that from Theorem˜1.2,
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
for a measure $\mu_{a}$ over $\mathbb{R}^{N}$ which is supported on the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$. In particular,
$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}d\mu_{a}(\epsilon).$
Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ and $\sum_{h\in H}c_{h}=1$. Then, for $h^{\prime}\in H$,
$e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}=e^{\sum_{i=1}^{N}\text{Re}(x_{i}\sum_{h\in H}c_{h}h^{\prime}ha_{i})}\leq\sum_{h\in H}c_{h}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}$
by Jensen’s inequality. It follows that
$\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}\leq\sum_{h\in H}\frac{c_{h}}{|H|}\sum_{h^{\prime}\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}=\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
This completes the proof. ∎
Lemma 10.5.
Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$. Suppose $R$ and $I$ are nonnegative real numbers. Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
  1. (A)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq I^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{R\sum_{i=1}^{N}|\text{Re}(x_{i})|+I\sum_{i=1}^{N}|\text{Im}(x_{i})|}.$
  2. (B)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq d!R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq d!I^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\left(\sum_{d=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d}\sum_{d=0}^{\infty}\left(2I\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{d}\right)^{\frac{1}{2}}.$
Proof.
(A): We use Theorem˜1.2. Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$. First, observe that
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle-\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)$
$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{-2\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}$
$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}.$
Next, note that since $\mu_{a}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{C}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(2|\text{Re}(x)|)M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ satisfy $\sum_{h\in H}c_{h}=1$. Furthermore, suppose $\nu\in\Gamma$. By Muirhead’s inequality and the triangle inequality,
$\frac{M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)}\leq\frac{\sum_{i=1}^{N}|\text{Re}(\epsilon_{i})|^{|\nu|}}{N}\leq\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}.$
By Jensen’s inequality, we have that
$\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}\leq\frac{1}{N}\sum_{h\in H}c_{h}\sum_{i=1}^{N}|\text{Re}(ha_{i})|^{|\nu|}=\frac{\sum_{i=1}^{N}|\text{Re}(a_{i})|^{|\nu|}}{N}\leq R^{|\nu|}.$
Hence,
(31) $\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\nu!}=e^{2R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
Similarly,
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq e^{2I\sum_{i=1}^{N}|\text{Im}(x)_{i}|}.$
Afterwards, it is straightforward to conclude the proof.
(B): The equation (31) becomes
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\pi(\nu)}=\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{k}.$
Similarly,
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{k},$
which completes the proof. ∎
Lemma 10.6.
Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$. Suppose $M$ is a nonnegative real number. Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
  1. (A)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq M^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}}.$
  2. (B)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq d!M^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\sum_{d=0}^{\infty}\left(M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}\right)^{d}.$
Proof.
For $z\in\mathbb{C}^{N}$, we let $\tilde{z}$ denote $(\left\lVert z_{i}\right\rVert_{2})_{i\in[N]}$. First, observe that for $a,b\in\mathbb{C}^{N}$,
$\left\lVert e^{\left\langle a,b\right\rangle}\right\rVert_{2}=e^{\text{Re}(\left\langle a,b\right\rangle)}\leq e^{\left\lVert\left\langle a,b\right\rangle\right\rVert_{2}}\leq e^{\left\langle\tilde{a},\tilde{b}\right\rangle}$.
Therefore, after referring to Theorem˜1.2, we can deduce that
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\tilde{x},\tilde{\epsilon}\right\rangle}d\mu_{a}(\epsilon).$
Afterwards, we can apply the same method used in the proof of Lemma˜10.5. ∎
Parts (A) and (B) of the following lemma are from [rank_infinity], where they are proved using results from [baker2, baker, okounkov_olshanki, bessel_convolution, forrester]. We state a proof of part (A) which does not involve Jack polynomials. Furthermore, we use a continuity argument to remove the requirement that the multiplicity function is nonnegative, which extends upon previous results.
Lemma 10.7.
Suppose $N\geq 2$, $a\in\mathbb{C}^{N}$, and $z\in\mathbb{C}$. For $k\geq 1$, suppose $c_{k}$ is the homogeneous degree $k$ polynomial such that $c_{k}(a)$ is the coefficient of $z^{k}$ in $\exp\\ \left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$.
  1. (A)
    Assume that $\theta\in\mathbb{C}$ such that $\mathcal{D}(A^{N-1}(\theta))$ is invertible. The value of $J_{a}^{A^{N-1}(\theta)}(z,0,\\ \ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a)}{(\theta N)_{k}}z^{k}$.
  2. (B)
    Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ such that $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ is invertible. The value of $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0\ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}z^{2k}$.
  3. (C)
    Assume that $\theta\in\mathbb{C}$ such that $\mathcal{D}(D^{N}(\theta))$ is invertible. The value of $J_{a}^{D^{N}(\theta)}(z,0,\\ \ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}((N-1)\theta+\frac{1}{2})_{k}(\theta N)_{k}}z^{2k}$.
Proof.
(A): Suppose $k\geq 1$. Let $r_{k}$ be the homogeneous degree $k$ polynomial such that $r_{k}(a)$ is the coefficient of $z^{k}$ in $J_{a}^{A^{N-1}(\theta)}(z,0,\ldots,0)$. Then,
$J_{a}^{A^{N-1}(\theta)}(z,0\ldots,0)=1+\sum_{k=1}^{\infty}r_{k}(a)z^{k}.$
The goal is to show that $r_{k}=\frac{c_{k}}{(\theta N)_{k}}$.
Observe that $r_{k}$ is the unique solution to $r_{k},m_{\epsilon}]_{A^{N-1}(\theta)}=0$ for $\epsilon\in\mathcal{P}_{N}[k]$ such that $\ell(\epsilon)\geq 2$ and $r_{k},p_{(k)}]_{A^{N-1}(\theta)}=1$. Hence, it suffices to show that $r_{k}=\frac{c_{k}}{(\theta N)_{k}}$ satisfies these properties.
First, observe that $\mathcal{D}_{2}\mathcal{D}_{1}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)=0$; note that the Dunkl operators are applied to the variable $a$. This is because
$\mathcal{D}_{1}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)=\theta\sum_{m=1}^{\infty}a_{1}^{m-1}z^{m}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right).$
Applying $\mathcal{D}_{2}$ to this expression gives
$\theta^{2}\left[\left(\sum_{m=1}^{\infty}a_{1}^{m-1}z^{m}\right)\left(\sum_{m=1}^{\infty}a_{2}^{m-1}z^{m}\right)-\sum_{m=1}^{\infty}z^{m}\sum_{i=0}^{m-2}a_{1}^{m-2-i}a_{2}^{i}\right]\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)=0.$
Hence, $c_{k},m_{\epsilon}]_{A^{N-1}(\theta)}=0$ for $\epsilon\in\mathcal{P}_{N}[k]$ such that $\ell(\epsilon)\geq 2$.
To finish the proof, we must show that $c_{k},p_{(k)}]_{A^{N-1}(\theta)}=(\theta N)_{k}$. Equivalently, we must show that
$\left[p_{(k)}(a),\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)\right]_{A^{N-1}(\theta)}=(\theta N)_{k}z^{k},$
since the left hand side equals $c_{k},p_{(k)}]_{A^{N-1}(\theta)}z^{k}$. However, note that this expression equals
$\mathcal{D}(p_{(k)})\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$ $=\left(\sum_{i=1}^{N}\mathcal{D}_{i}\right)^{k}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$
$=\left(\sum_{i=1}^{N}\partial_{i}\right)^{k}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$
Hence, $c_{k},p_{(k)}]_{A^{N-1}(\theta)}z^{k}=[1]\left(\sum_{i=1}^{N}\partial_{i}\right)^{k}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$.
First, note that
$1]e^{\sum_{i=1}^{N}\partial_{i}}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)=\exp\left(N\theta\sum_{m=1}^{\infty}\frac{z^{m}}{m}\right).$
Hence, the value of $1]\left(\sum_{i=1}^{N}\partial_{i}\right)^{k}\exp\left(\theta\sum_{m=1}^{\infty}p_{(m)}(a)\frac{z^{m}}{m}\right)$ is $k!$ times the term of degree $k$ in $z$ of $\exp\left(N\theta\sum_{m=1}^{\infty}\frac{z^{m}}{m}\right)$. That is, $c_{k},p_{(k)}]_{A^{N-1}(\theta)}=\partial_{z}^{k}\exp\left(N\theta\sum_{m=1}^{\infty}\frac{z^{m}}{m}\right)|_{z=0}$.
It is straightforward to deduce that this quantity is $(\theta N)_{k}$. Initiate $c=1$. At each of $k$ steps, we can decide to either multiply $c$ by $N\theta\sum_{m=1}^{\infty}z^{m}$ or differentiate $c$. Label the terms that we multiply $c$ by starting from $1$. We must multiply $c$ by $N\theta\sum_{m=1}^{\infty}z^{m}$ at the first step, and we label this term as $1$. If we do not multiply $c$ by $N\theta\sum_{m=1}^{\infty}z^{m}$ at a step, then we must replace one of the previously labeled terms with its derivative. Therefore, we always have that $c$ is the product of the labeled terms, which may have been differentiated.
Suppose we are currently at step $i\in[k]$ such that $i\geq 2$. Suppose $L\geq 1$ terms have been labeled. For $j\in[L]$, let $d_{j}$ be the number of times that term $j$ has been differentiated; note that $d_{j}$ does not include the first derivative applied to $N\theta\sum_{m=1}^{\infty}\frac{z^{m}}{m}$. Then, we have that
$\sum_{j\in L}d_{j}+L=i-1.$
Furthermore, the constant term of term $j$ is $d_{j}!$ for all $j\in[L]$ so the constant term of $c$ is $\prod_{j=1}^{L}d_{j}!$. For step $j$, we can decide to either differentiate term $j$ for some $j\in[L]$ or multiply $c$ by $N\theta\sum_{m=1}^{\infty}z^{m}$. If we differentiate term $j$ for some $j\in[L]$, we will multiply the constant term of $c$ by $d_{j}+1$. Otherwise, if we multiply $c$ by $N\theta\sum_{m=1}^{\infty}z^{m}$, we will multiply the constant term by $N\theta$. By adding these choices, we will multiply the constant term of $c$ by
$\sum_{j=1}^{L}(d_{j}+1)+N\theta=N\theta+i-1.$
When $i=1$, we multiply the constant term by $N\theta$. Therefore,
$c_{k},p_{(k)}]_{A^{N-1}(\theta)}=\partial_{z}^{k}\exp\left(N\theta\sum_{m=1}^{\infty}\frac{z^{m}}{m}\right)\bigg|_{z=0}=\prod_{i=1}^{k}(N\theta+i-1)=(N\theta)_{k}.$
(B): We follow the same method as (A). Suppose $k\geq 1$. Let $r_{2k}$ be the homogeneous degree $k$ polynomial such that $r_{2k}(a)$ is the coefficient of $z^{2k}$ in $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0,\ldots,0)$. It suffices to show that $r_{2k}(a)=\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}$.
It suffices to show that $c_{k}(a^{2}),m_{\epsilon}(a^{2})]_{BC^{N}(\theta_{0},\theta_{1})}=0$ for all $\epsilon\in\mathcal{P}_{N}[k]$ such that $\ell(\epsilon)\geq 2$ and that $c_{k}(a^{2}),p_{(2k)}(a)]_{BC^{N}(\theta_{0},\theta_{1})}=4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}$. For the first statement, we similarly have that $\mathcal{D}_{2}\mathcal{D}_{1}\exp\left(\theta_{0}\sum_{m=1}^{\infty}p_{(2m)}(a)\frac{z^{m}}{m}\right)=0$. Observe that
$\mathcal{D}_{1}\exp\left(\theta_{0}\sum_{m=1}^{\infty}p_{(2m)}(a)\frac{z^{m}}{m}\right)=2\theta_{0}\left(\sum_{m=1}^{\infty}a_{1}^{2m-1}z^{m}\right)\exp\left(\theta_{0}\sum_{m=1}^{\infty}p_{(2m)}(a)\frac{z^{m}}{m}\right).$
Applying $\mathcal{D}_{2}$ to this expression gives
$4\theta_{0}^{2}\left[\left(\sum_{m=1}^{\infty}a_{1}^{2m-1}z^{m}\right)\left(\sum_{m=1}^{\infty}a_{2}^{2m-1}z^{m}\right)-\sum_{m=1}^{\infty}z^{m}\sum_{i=0}^{m-2}a_{1}^{2m-3-2i}a_{2}^{2i+1}\right]$
$\exp\left(\theta_{0}\sum_{m=1}^{\infty}p_{(2m)}(a)\frac{z^{m}}{m}\right)=0.$
Hence, it suffices to show that $c_{k}(a^{2}),p_{(2k)}(a)]_{BC^{N}(\theta_{0},\theta_{1})}=4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}$. It should be possible to do so using a similar argument as in (A). However, we use a different argument that proves the identity for all $\theta_{0},\theta_{1}\in\mathbb{C}$ to avoid the computational details. First, we note that $c_{k}(a^{2}),p_{(2k)}(a)]_{BC^{N}(\theta_{0},\theta_{1})}$ is a polynomial in $\theta_{0}$ and $\theta_{1}$ with real coefficients, where we assume that $N$ is fixed. Since the expression is true whenever $\theta_{0},\theta_{1}\geq 0$ from [baker2] or [rank_infinity], it must be true for all $\theta_{0},\theta_{1}\in\mathbb{C}$.
(C): We note that the terms with all odd degrees do not contribute to the value of $J_{a}^{D^{N}(\theta)}(z,0,\ldots,0)$. For $k\geq 1$, these terms are also orthogonal to $c_{k}(a^{2})$ by Corollary˜8.4 or the argument in (B). Thus, $c_{k}(a^{2})$ is orthogonal to $m_{\epsilon}(a^{2})$ for all $\epsilon\in\mathcal{P}_{N}[k]$ and $e(a)m_{\epsilon}(a^{2})$ for all $\epsilon\in\mathcal{P}_{N}[k-\frac{N}{2}]$. Hence, we obtain the formula by setting $\theta_{1}=0$ in (B). ∎
10.3. The uniform convergence of Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime
First, we introduce the following well-known result; we include the proof, which is also well-known, for completeness. A similar argument is used in [rank_infinity].
Lemma 10.8.
Suppose $r\geq 1$ and that $\Omega\subset\mathbb{C}^{r}$ is open and simply connected. Let $\{f_{N}\}_{N\geq 1}$ be a sequence of holomorphic functions over $\Omega$.
  1. (A)
    Assume that $f:\Omega\rightarrow\mathbb{C}$ and $\{f_{N}\}_{N\geq 1}$ converges to $f$ uniformly over compact subsets of $\Omega$. Then, $f$ is holomorphic over $\Omega$ and the partial derivatives of $\{f_{N}\}_{N\geq 1}$ converge to the partial derivatives of $f$ uniformly over compact subsets of $\Omega$.
  2. (B)
    Assume that $\{f_{N}\}_{N\geq 1}$ is uniformly bounded over compact subsets of $\Omega$. Suppose $z\in\Omega$, $\lim_{N\rightarrow\infty}f_{N}(z)=f(z)$, and the limits of the partial derivatives of $\{f_{N}\}_{N\geq 1}$ evaluated at $z$ equal the partial derivatives of $f$ evaluated at $z$. Then, $\{f_{N}\}_{N\geq 1}$ converges to $f$ uniformly over compact subsets of $\Omega$.
Proof.
(A): To show that $f$ is holomorphic, note that because $\Omega$ is simply connected, we may apply Cauchy’s integral theorem and Morera’s theorem. Afterwards, we can apply Cauchy’s integral formula to deduce that the partial derivatives of $\{f_{N}\}_{N\geq 1}$ converge to the partial derivatives of $f$ uniformly over compact subsets of $\Omega$.
Suppose the closed ball $\overline{B}(a,b)$ is a subset of $\Omega$. For the sake of contradiction, assume that for all $\epsilon>0$, the closed ball $\overline{B}(a,b+\epsilon)$ is not a subset of $\Omega$. Then, for $\epsilon>0$, let $x_{\epsilon}$ be an element of $\overline{B}(a,b)$ such that $\overline{B}(x_{\epsilon},\epsilon)\not\subset\Omega$. By the compactness of $\overline{B}(a,b)$, $\{x_{\epsilon}\}_{\epsilon>0}$ has a convergent subsequence as $\epsilon\rightarrow 0$. Assume that the limit is $x$, which is an element of $\overline{B}(a,b)$ because the set is closed. Then, for all $\delta>0$, $B(x,\delta)\not\subset\Omega$, because there exists sufficiently small $\epsilon$ such that $\overline{B}(x_{\epsilon},\epsilon)\subset B(x,\delta)$. This is a contradiction to $\Omega$ being open.
Assume that $\epsilon>0$ and $\overline{B}(a,b+\epsilon)\subset\Omega$. Since $\{f_{N}\}_{N\geq 1}$ is uniformly converging over $\overline{B}(a,b+\epsilon)$, by applying the Cauchy integral formula with the contour set as the boundary of $\overline{B}(a,b+\epsilon)$, we have that $\{\partial_{i}f_{N}\}_{N\geq 1}$ is uniformly converging over $\overline{B}(a,b)$ to $\partial_{i}f$ for all $i\in[r]$. Then, we can proceed with induction to obtain that all partial derivatives of $\{f_{N}\}_{N\geq 1}$ are uniformly converging over $\overline{B}(a,b)$.
Next, let $K\subset\Omega$ be a compact set. By the same argument that we mentioned earlier, there exists $\epsilon>0$ such that $\overline{B}(x,\epsilon)\subset\Omega_{r}$ for all $x\in K$. Then, $K\subset\bigcup_{x\in K}B(x,\epsilon)$; note that the union is over open balls rather than closed balls. Let $F$ be a finite subcover so that $K\subset\bigcup_{x\in F}B(x,\epsilon)$. Then, the partial derivatives of $\{f_{N}\}_{N\geq 1}$ are uniformly converging over $B(x,\epsilon)$ for each $x\in F$, so the partial derivatives are uniformly converging over $K$ to the partial derivatives of $f$.
(B): Suppose a subsequence of $\{f_{N}\}_{N\geq 1}$ converges uniformly to $g$ over compact subsets of $\Omega$. By (A), $g$ is holomorphic and the partial derivatives of the subsequence converge to the partial derivatives of $g$ uniformly over compact subsets of $\Omega$. Therefore, the partial derivatives of $f$ and $g$ evaluated at $z$ are equal. Furthermore, since $\lim_{N\rightarrow\infty}f_{N}(z)=f(z)$, we have that $f(z)=g(z)$. This implies that $f=g$ over $\Omega$ by the identity theorem.
Afterwards, applying Montel’s theorem implies that $\{f_{N}\}_{N\geq 1}$ converges to $f$ uniformly over compact subsets of $\Omega$. For the sake of contradiction, assume that $K$ is a compact subset of $\Omega$ such that $\{f_{N}\}_{N\geq 1}$ does not converge uniformly to $f$ over $K$. Then, suppose $\epsilon>0$ and $\{N_{j}\}_{j\geq 1}$ is an increasing sequence of positive integers such that for all $j\geq 1$, $\sup_{z\in K}|f_{N_{j}}(z)-f(z)|\geq\epsilon$. By Montel’s theorem, since $\{f_{N_{j}}\}_{j\geq 1}$ is uniformly bounded over compact subsets of $\Omega$, it has a uniformly converging subsequence over compact subsets of $\Omega$. However, this subsequence must uniformly converge to $f$ over $K$, which is a contradiction to $\sup_{z\in K}|f_{N_{j}}(z)-f(z)|\geq\epsilon$ for all $j\geq 1$. ∎
We use the previous result to justify the uniform convergence of the Bessel functions $J_{a(N)}^{\mathcal{R}(\theta)}(z_{1},\ldots,z_{r},0,\ldots,0)$ as $N\rightarrow\infty$, similarly to what is proved in the papers [assiotis, rank_infinity]. While the papers consider when $\theta\in\mathbb{R}_{\geq 0}$ is fixed, we consider when $\theta N\rightarrow c\in\mathbb{R}_{\geq 0}$. Furthermore, we focus on when we know some uniform bounds on the moments of $\{a(N)\}_{N\geq 1}$. In particular, if the bounds are strong enough, then we can deduce uniform convergence over compact subsets of $\mathbb{C}^{r}$.
Theorem 10.9.
Suppose $m_{d}\in\mathbb{C}$ for $d\geq 1$. Assume that for $N\geq 1$, $a(N)\in\mathbb{C}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N}=m_{d}$ for all $d\geq 1$. Suppose $M$ is a positive real number. Consider the following two conditions.
  1. (C1)
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq M^{d}$.
  2. (C2)
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq d!M^{d}$.
Assume that $\theta,\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$. Suppose $r\geq 1$ and define $\Omega_{r}\triangleq\{z\in\mathbb{C}:\sum_{i=1}^{r}\left\lVert z_{i}\right\rVert_{2}<\frac{1}{M}\}$.
  • (A)
    Suppose that as $N\rightarrow\infty$, $\theta N\rightarrow c\in\mathbb{R}_{\geq 0}$. Also, suppose $c_{(k)}$ for $k\geq 1$ solve
    $m_{d}=\sum_{\pi\in NC(d)}W^{A}(\pi)(c)\prod_{B\in\pi}|B|c_{(|B|)}$
    for $d\geq 1$.
    As $N\rightarrow\infty$, $J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
  • (B)
    Suppose that as $N\rightarrow\infty$, $\theta_{0}N\rightarrow c_{0}\in\mathbb{R}_{\geq 0}$ and $\theta_{1}\rightarrow c_{1}\in\mathbb{R}_{\geq 0}$. Also, suppose $c_{(2k)}$ for $k\geq 1$ solve
    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}$
    for $d\geq 1$.
    As $N\rightarrow\infty$, $J^{BC^{N}(\theta_{0},\theta_{1})}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}c_{(2k)}p_{(2k)}\right)$ uniformly over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
  • (C)
    Suppose that as $N\rightarrow\infty$, $\theta N\rightarrow c\in\mathbb{R}_{\geq 0}$. Also, suppose $c_{(2k)}$ for $k\geq 1$ solve
    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{A}(\pi)(2c)\prod_{B\in\pi}|B|c_{(|B|)}$
    for $d\geq 1$.
    As $N\rightarrow\infty$, $J^{D^{N}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}c_{(2k)}p_{(2k)}\right)$ uniformly over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
Proof.
(A): By Lemma˜10.6, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
Suppose $k\geq 1$ and $r\in\Gamma[k]$. We compute the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ as $N\rightarrow\infty$. First, note that for $s\in\Gamma$,
$\lim_{N\rightarrow\infty}N^{-\ell(s)}s(a(N))=\prod_{i=1}^{\ell(s)}m_{s_{i}}.$
Then, using Corollary˜9.4 gives that it is evident that the coefficient of $r(x)$ as $N\rightarrow\infty$ is
$\sum_{s\in\Gamma[k]}\mathcal{W}^{A}(c)[k]^{-1}_{rs}\prod_{i=1}^{\ell(s)}m_{s_{i}}.$
To show that this quantity is $\frac{\prod_{i=1}^{\ell(r)}c_{(r_{i})}}{\pi(r)}$, it suffices to show that
$\mathcal{W}^{A}(c)[k]\left[\frac{\prod_{i=1}^{\ell(r)}c_{(r_{i})}}{\pi(r)}\right]_{r\in\Gamma[k]}^{T}=\left[\prod_{i=1}^{\ell(r)}m_{r_{i}}\right]_{r\in\Gamma[k]}^{T},$
since $\mathcal{W}^{A}(c)[k]$ is invertible. However, this is evident based on the formula for $\mathcal{W}^{A}(c)$ and the definition of the $c_{(k)}$; for example, see the proof of Corollary˜6.11. Thus, the $N\rightarrow\infty$ limit of the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ is the coefficient of $r(x)$ in $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$.
Afterwards, by part (B) of Lemma˜10.8, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
(B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and Corollary˜9.14 for (C). For (C), observe that the terms with all odd degrees are eliminated over $\Omega$. ∎
10.4. The uniform convergence of Bessel functions for Vershik-Kerov
sequences
In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system. The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.14. First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
Theorem 10.10.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{|\nu|}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\theta^{|\nu|-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}[p_{\nu}]\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right).$
Then, (a) implies (b) and if $\theta\not=0$, then (b) implies (a).
Proof.
See Theorem˜6.1. The main idea is that for $\lambda,\nu\in\Gamma$, the leading order term of the expansion of $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ given in the theorem is nonzero if and only if $\lambda=\nu$. ∎
Theorem 10.11.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=(2\theta_{0})^{|\nu|-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(1+c)^{o([\nu_{l}])}[p_{\nu}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
Then, (a) implies (b) and if $\theta_{0}\not=0$ and $c\not=-1$, then (b) implies (a).
Proof.
See Theorem˜7.1. ∎
Theorem 10.12.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}c_{\lambda}(N)=c_{\lambda}\in\mathbb{C}$.
  2. (b)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
  3. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{|\nu|}}[p_{\nu},F_{N}]_{D^{N}(\theta)}=(2\theta)^{|\nu|-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}[p_{\nu}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
  4. (d)
    For all $\nu\in\Gamma_{\text{even}}$
    $\lim_{N\rightarrow\infty}\frac{1}{\prod_{j=1}^{N}(1+2(j-1)\theta)N^{|\nu|}}[ep_{\nu},F_{N}]_{D^{N}(\theta)}$
    $=(2\theta)^{|\nu|-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}[p_{\nu}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
Then, (a) implies (c) and if $\theta\not=0$, then (c) implies (a).
Assume that $0\notin\{1+2j\theta:j\in\mathbb{N}\}$. Then, (b) implies (d) and if $\theta\not=0$, then (d) implies (b).
Proof.
See Theorems˜8.10 and 8.16. ∎
Theorem˜10.13 is an example of an application of the results of this paper. Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity]. Part (C) has not appeared previously, although it is straightforward to deduce. See Remark˜10.14 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
Theorem 10.13.
Suppose $m_{d}\in\mathbb{R}$ for $d\geq 1$. Assume that for $N\geq 1$, $a(N)\in\mathbb{R}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$. Assume that $M$ is a positive real number such that for sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N^{d}}\left|\sum_{i=1}^{N}a(N)_{i}^{d}\right|\leq M^{d}$. Also, assume that $\theta,\theta_{0}>0$ are fixed and that $\theta_{1}\geq 0$ for all $N\geq 1$. Suppose $r\geq 1$.
  • (A)
    As $N\rightarrow\infty$, $J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$ uniformly over compact subsets of $\left\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\right\}$.
  • (B)
    Suppose $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\geq 0$. As $N\rightarrow\infty$, $J^{BC^{N}(\theta_{0},\theta_{1})}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}\frac{m_{2k}p_{(2k)}(2\theta_{0})^{-2k+1}(1+c)^{-k}}{2k}\right)$ uniformly over compact subsets of $\left\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{2\theta_{0}\sqrt{1+c}}{rM}\,\,\forall i\in[r]\right\}$.
  • (C)
    As $N\rightarrow\infty$, $J^{D^{N}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)$ converges to $\exp\left(\sum_{k\geq 1}\frac{m_{2k}p_{(2k)}(2\theta)^{-2k+1}}{2k}\right)$ uniformly over compact subsets of $\left\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{2\theta}{rM}\,\,\forall i\in[r]\right\}$.
Proof.
(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.2 and Hölder’s inequality gives that
$\left|J^{A^{N-1}(\theta)}_{a(N)}(x_{1},\ldots,x_{r},0,\ldots,0)\right|$ $=\left|\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{r}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)\right|$
$\leq\prod_{i=1}^{r}\left|\int_{\mathbb{R}^{N}}e^{r\text{Re}(x_{i}\epsilon_{i})}d\mu_{a}(\epsilon)\right|^{\frac{1}{r}}$
$=\prod_{i=1}^{r}J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)^{\frac{1}{r}}.$
Afterwards, using Lemma˜10.7 gives that
$J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)$ $\leq 1+\sum_{k=1}^{\infty}\frac{|c_{k}(a)|}{(\theta N)^{k}}|r\text{Re}(x_{i})|^{k}$
$=1+\sum_{k=1}^{\infty}\left|c_{k}\left(\frac{a}{\theta N}\right)r\text{Re}(x_{i})\right|^{k}$
$\leq\exp\left(\theta\sum_{m=1}^{\infty}\left|p_{(m)}\left(\frac{a}{N}\right)\right|\frac{|r\text{Re}(x_{i})|^{m}}{\theta^{m}m}\right)$
$\leq\exp\left(\theta\sum_{m=1}^{\infty}M^{m}\frac{|r\text{Re}(x_{i})|^{m}}{\theta^{m}m}\right).$
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$. We can then apply part (B) of Lemma˜10.8 to conclude the proof.
(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.7 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.7 and Corollary˜9.11. ∎
Remark 10.14.
From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered. Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.13 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$. This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
10.5. Bessel generating functions for exponentially decaying probability measures
Rather than consider the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ for a single value of $a$, we can consider the average of $J_{a}^{\mathcal{R}(\theta)}(x)$ over a distribution of $a$. The resulting function is referred to as the Bessel generating function and has been studied previously in [gaussianfluctuation, matrix, limits_prob_measures, rectangularmatrix]. This notion is also related to the Dunkl transform discussed in [dunklbound].
Suppose $\mu$ is a probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
$G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)].$
However, we must place restrictions on $\mu$ so that the Bessel generating function converges. In particular, we consider the class of exponentially decaying probability measures studied in [limits_prob_measures], which is a modification of the class of measures studied in [matrix].
Definition 10.15 (Definition 1.2 of [limits_prob_measures]).
A Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$ if $\int_{\mathbb{C}^{N}}e^{R\left\lVert a\right\rVert_{2}}d\mu(a)$ is finite.
If $\mu$ is exponentially decaying, then we have that $G_{\mu}^{\mathcal{R}(\theta)}$ is holomorphic in a neighborhood of the origin, which is required to apply the results of this paper. This implication as well as other essential implications are included in the following lemma.
Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (C) is [matrix]*Proposition 2.11).
Suppose the Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ satisfies $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$.
  1. (A)
    The Bessel generating function $G_{\mu}^{\mathcal{R}(\theta)}(x)$ converges and is holomorphic over the closed ball of radius $R$ centered at the origin.
  2. (B)
    There exist unique constants $c_{\lambda}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(\mu)p_{\lambda}\right)$
    in a neighborhood of the origin.
  3. (C)
    For all $\lambda\in\Gamma$,
    $1]\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)=\operatorname{\mathbb{E}}_{a\sim\mu}[p_{\lambda}(a)].$
Proof.
For (B), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$. ∎
Remark 10.17.
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (C) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
Using Lemma˜10.16 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system. It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems. The corollary generalizes the results of [limits_prob_measures].
Corollary 10.18.
Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 1$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
Define $c_{\lambda}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ as in part (B) of Lemma˜10.16. Then, the following are equivalent.
  1. (a)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (b)
    For all $\nu\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
Remark 10.19.
Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
11. Computing the Dunkl bilinear form with combinatorics
For any $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ and $\lambda,\nu\in\Gamma$, we set
$p_{\lambda},p_{\nu}]_{A^{0}(\theta)},\,[p_{\lambda},p_{\nu}]_{BC^{1}(\theta_{0},\theta_{1})},\,[p_{\lambda},p_{\nu}]_{D^{1}(\theta)}\triangleq[1]\partial_{1}^{\lambda}x_{1}^{\nu}=\mathbf{1}\{|\lambda|=\nu|\}|\lambda|!.$
Note that $A^{0}$, $B^{1}$, $C^{1}$, and $D^{1}$ are not actually root systems; however, we use this notation for simplicity.
The following result computes combinatorial expressions for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$, $p_{\lambda},\\ p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}$, and $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ in terms of the values of the quantities for small values of $N$, including $N=1$.
Theorem 11.1.
Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$. For (B) and (C), assume that $\lambda,\nu\in\Gamma_{\text{even}}$. Let $k=|\lambda|$ and suppose $N\geq 1$. Furthermore, assume that $k+\ell(\lambda)-\ell(\nu)<N$. For all $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$, the following expressions are true:
$(A)\,\,\,[p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ $=\sum_{i=1}^{k+\ell(\lambda)-\ell(\nu)}(-1)^{k+\ell(\lambda)-\ell(\nu)-i}[p_{\lambda},p_{\nu}]_{A^{i-1}(\theta)}\times$
$\binom{N}{i}\binom{N-i-1}{k+\ell(\lambda)-\ell(\nu)-i}$
$(B)\,\,\,[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}$ $=\sum_{i=1}^{k+\ell(\lambda)-\ell(\nu)}(-1)^{k+\ell(\lambda)-\ell(\nu)-i}[p_{\lambda},p_{\nu}]_{BC^{i}(\theta_{0},\theta_{1})}\times$
$\binom{N}{i}\binom{N-i-1}{k+\ell(\lambda)-\ell(\nu)-i}$
$(C)\,\,\,[p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ $=\sum_{i=1}^{k+\ell(\lambda)-\ell(\nu)}(-1)^{k+\ell(\lambda)-\ell(\nu)-i}[p_{\lambda},p_{\nu}]_{D^{i}(\theta)}\times$
$\binom{N}{i}\binom{N-i-1}{k+\ell(\lambda)-\ell(\nu)-i}$
Proof.
We confirm expression (A). Expressions (B) and (C) follow similarly. Let $\Delta=k+\ell(\lambda)-\ell(\nu)$ and $k=|\lambda|$. Suppose $\mathcal{S}$ is the set of sequences of operators appearing in $\mathcal{D}(A^{N-1}(\theta))(p_{\lambda})$, so that
$p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}=\sum_{s\in\mathcal{S}}sp_{\nu},$
where for $s\in\mathcal{S}$, $sp_{\nu}\triangleq s_{k}\circ\cdots\circ s_{1}p_{\nu}$.
For $s\in\mathcal{S}$, let $i(s)$ be the number of distinct indices in $s$. Observe that if $s\in\mathcal{S}$ and $s_{k}\circ\cdots\circ s_{1}p_{\nu}\not=0$, then we must have that $i(s)\leq k+\ell(\lambda)-\ell(\nu)$, by the argument that $s$ must have at least $\ell(\nu)$ derivatives. Thus, let $\mathcal{S}^{\prime}$ be the set of $s\in\mathcal{S}$ such that $i(s)\leq k+\ell(\lambda)-\ell(\nu)$, so that
$p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}=\sum_{s\in\mathcal{S}^{\prime}}sp_{\nu}.$
Note that for $i\geq 1$, $p_{\lambda},p_{\nu}]_{A^{i-1}(\theta)}$ corresponds to picking a sequence of operators in $\mathcal{D}(A^{N-1}(\theta))(p_{\lambda})$ with all indices in $[i]$. Then, $p_{\lambda},p_{\nu}]_{A^{i-1}(\theta)}\binom{N}{i}$ corresponds to picking this sequence as well as replacing $[i]$ by any subset of $[N]$ with size $i$.
Suppose $s\in\mathcal{S}^{\prime}$. Assume that the set of distinct indices in $s$ is $[i(s)]$. Then, it is clear that
$sp_{\nu}(x_{1},\ldots,x_{N})=sp_{\nu}(x_{1},\ldots,x_{i(s)},0,\ldots,0).$
For $i\geq i(s)$, the number of times that $sp_{\nu}$ is counted in $p_{\lambda},p_{\nu}]_{A^{i-1}(\theta)}\binom{N}{i}$ is $\binom{N-i(s)}{i-i(s)}$. Hence, the total number of times that $sp_{\nu}$ is counted is
$\sum_{i=i(s)}^{k+\ell(\lambda)-\ell(\nu)}(-1)^{k+\ell(\lambda)-\ell(\nu)-i}\binom{N-i(s)}{i-i(s)}\binom{N-i-1}{k+\ell(\lambda)-\ell(\nu)-i}$.
It suffices to show that this quantity equals one. Equivalently, it suffices to show that
(32) $\sum_{i=0}^{\Delta-z}(-1)^{i}\binom{N-z}{\Delta-z-i}\binom{N-\Delta-1+i}{i}=1$
for $\Delta\geq 1$, $z\in[\Delta]$, and $N\in\mathbb{N}$.
We prove (32) using induction on $z$ from $z=\Delta$ to $1$. The base case $z=\Delta$ is clear. Assume that the inductive hypothesis is true for $z=m$, where $m\in\{2,\ldots,\Delta\}$. We prove that it is true for z=m1z=m-1. We have that
$\sum_{i=0}^{\Delta-m+1}(-1)^{i}\binom{N-m+1}{\Delta-m+1-i}\binom{N-\Delta-1+i}{i}$
$=\sum_{i=0}^{\Delta-m+1}(-1)^{i}\left(\binom{N-m}{\Delta-m-i}+\binom{N-m}{\Delta-m+1-i}\right)\binom{N-\Delta-1+i}{i}$.
Next, observe that
$\sum_{i=0}^{\Delta-m+1}(-1)^{i}\binom{N-m}{\Delta-m-i}\binom{N-\Delta-1+i}{i}$
$=\sum_{i=0}^{\Delta-m}(-1)^{i}\binom{N-m}{\Delta-m-i}\binom{N-\Delta-1+i}{i}=1$
by the inductive hypothesis for $z=m$. Hence, it suffices to show that
$\sum_{i=0}^{\Delta-m+1}(-1)^{i}\binom{N-m}{\Delta-m+1-i}\binom{N-\Delta-1+i}{i}=0.$
This expression evaluates to
$\frac{(N-m)!}{(N-\Delta-1)!}\sum_{i=0}^{\Delta-m+1}\frac{(-1)^{i}}{(\Delta-m+1-i)!i!}=0.$
Remark 11.2.
The previous result is not applicable to when $k+\ell(\lambda)-\ell(\nu)\geq N$, but if this is the case, then we must consider when the number of distinct indices is $N$. Then, to obtain a summation formula for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ in (A), we must set $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ to be one of the summands. Similarly, we cannot obtain analogous formulas for (B) and (C).
Furthermore, for (B) and (C) we require that $\lambda,\nu\in\Gamma_{\text{even}}$ so that $i(s)\leq k+\ell(\lambda)-\ell(\nu)$. If this is not the case, then $s$ does not necessarily need to have at least $\ell(\nu)$ derivatives and the operators at the locations $1+\lambda_{1}+\cdots+\lambda_{i}$ for $0\leq i\leq\ell(\lambda)-1$ do not necessarily have to be derivatives, so we can have that $i(s)>k+\ell(\lambda)-\ell(\nu)$. Afterwards, we cannot ensure that the order of the Dunkl bilinear form is $N^{k+\ell(\lambda)-\ell(\nu)}$, if we assume that $\lambda$ and $\nu$ are fixed.
Assuming that $\theta$ is fixed, the order of the expressions in Theorem˜11.1 matches the order of the expressions in Theorems˜6.1, 7.1 and 8.10, which is $N^{k+\ell(\lambda)-\ell(\nu)}$. However, the leading order coefficients are not apparent from the formulas.
12. Leading order terms of the type A and BC Dunkl bilinear forms after multiplying by $x_{1}\cdots x_{N}$
In this section, we prove analogues of Theorems˜8.16 and 8.23. First, we prove the analogues of Theorem˜8.16 for the $A^{N-1}$ and $BC^{N}$ root systems.
Theorem 12.1.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{A^{N}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(N\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}+R(N,\theta)\bigg),$
where $R\in\mathbb{Q}[x,y]$ satisfies:
  1. (1)
    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
  2. (2)
    The degree $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$.
Proof.
The same method as the proof of Theorem˜8.16 can be used. ∎
Corollary 12.2.
Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$. For (B), assume that $\lambda,\nu\in\Gamma_{\text{even}}$. Suppose $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
$(A)\,\,\,[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+(i-1)\theta)[p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$
$(B)\,\,\,[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+2(i-1)\theta)[p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$
Proof.
(A) follows from Theorems˜6.1 and 12.1 and (B) follows from Theorems˜8.10 and 8.16. ∎
Theorem 12.3.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})\bigg(N^{2k+\ell(\lambda)-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}(2\theta_{0})^{k-\ell(\nu)}\times$
$\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}+R(N,\theta_{0},\theta_{1})\bigg),$
where $R\in\mathbb{Q}[x,y,z]$ satisfies:
  1. (1)
    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
  2. (2)
    The degree of $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the sum of the degrees of $y$ and $z$ in each summand is at most $2k-\ell(\nu)$.
  3. (3)
    In no summand is the degree of $x$ $\ell(\lambda)$ greater than the degree of $y$ while the degrees of $y$ and $z$ add to $2k-\ell(\nu)$.
Proof.
First, we have that $h\mathcal{D}(e)ep_{\nu}=\mathcal{D}(he)hep_{\nu}=\mathcal{D}(e)ep_{\nu}$ for all $h\in H(BC^{N}(\theta_{0},\theta_{1}))$. Hence, $\mathcal{D}(e)ep_{\nu}$ has all even degrees. It follows that we can replace $\mathcal{D}(ep_{\lambda})$ with
$\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}\mathcal{D}_{j}^{\lambda_{i}-1}\partial_{j}\mathcal{D}(e),$
as we have done previously.
After noting that $\theta_{1}$ has order $\theta_{0}N$, the same method as the proof of Theorem˜8.16 can be used. In this case, we can view the switch from $i$ to $j$ as
$\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right).$
We replace $\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}$ with $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}$ to compute the leading order term. Thus, we can view the switch from $i$ to $j$ as
(33) $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right);$
afterwards, we follow the same method. The goal is to compute
$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu},$
where $\mathcal{S}$ is defined in the same way except that the switch from $i$ to $j$ is replaced with (33) and $\mathcal{R}(s)$ is defined in the same way except that $2\theta d_{i}$ is replaced by $(2\theta_{0}+\frac{2\theta_{1}}{N})d_{i}$. We get that we can similarly reduce to the case that $\ell(\lambda)=1$. Hence, let $c_{\nu^{\prime}}$ be the value of (22) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$ for $\nu^{\prime}\in\Gamma_{\text{even}}$.
The only difference is in Lemma˜8.20. Assume that $\pi\in NC^{\prime}(|\nu^{\prime}|)$ and $b(\pi)=\nu^{\prime}$. For $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$ that are associated with $\pi$, we have that the number of switches in $s$ which are not associated with a derivative in $r$ is $|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1$ and the number of switches in $r$ is $|\nu^{\prime}|-\ell(\nu^{\prime})+1$; recall Definition˜8.17 for the definition of $d_{3}(r)$. The switches in $r$ contribute a factor of $\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}$. On the other hand, the switches in $s$ which are not associated with a derivative in $r$ contribute a factor of $(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}(1+\frac{\theta_{1}}{\theta_{0}N})^{o(\pi)}$. The switches in $s$ which are associated with a derivative in $r$ contribute a factor of $(-2\theta_{0})^{d_{3}(r)}$. Hence, the total contribution is
$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}(-2\theta_{0})^{d_{3}(r)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-d_{3}(r)}(-1)^{d_{3}(r)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)+d_{3}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)},$
where we have used $d_{1}(s)+d_{2}(r)+d_{3}(r)=\ell(\nu^{\prime})-1$. Observe that the number of locations for a switch in $s$ that is associated with a derivative in $r$ is $z(\pi)$. For each of these locations, we can decide to place the derivative in $s$ or in $r$. After summing over these choices, we obtain
$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-z(\pi)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}-2\theta_{0}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-d_{2}(r)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
Thus, we obtain that
$c_{\nu^{\prime}}=$ $\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC^{\prime}(|\nu^{\prime}|),\,b(\pi)=\nu^{\prime}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\times$
$\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
It follows that
$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{j}:j\in S_{i}))}$
$=\prod_{l=1}^{\ell(\nu)}\nu_{l}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{\lambda_{i}-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\times$
$(2\theta_{0})^{\lambda_{i}-\ell(b(\pi))}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma_{i}\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}.$
After multiplying by $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})N^{2k+\ell(\lambda)-\ell(\nu)}$, we obtain the desired result. ∎
Next, we prove the analogues of Theorem˜8.23 for the $A^{N-1}$ and $BC^{N}$ root systems.
Theorem 12.4.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(1+N\theta)^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(N\theta)\prod_{B\in\pi}x_{|B|}$
$+R(N,\theta)\bigg),$
where $R\in\mathbb{Q}[x,y]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
Proof.
We can prove this in the same way as we prove Theorem˜8.23. ∎
For $\pi\in NC(k+1)$, define
$W^{BC;\,\text{odd}}(\pi)(x,y)\triangleq$ $(1+\mathbf{1}\{b(2;\,\pi)=0\}(d(2;\,\pi)-1))\times$
$\prod_{i\in[k+1],i\geq 3,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi))\left(\frac{1+y}{1+x+y}\right)^{z(\pi)}\times$
$\left(\frac{1}{1+x+y}\right)^{\mathbf{1}\{f(\pi)>1\}}.$
Theorem 12.5.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2\theta_{1}+2(i-1)\theta_{0})\times$
$\bigg(N^{\ell(\lambda)}(1+2\theta_{1}+2N\theta_{0})^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{BC;\text{odd}}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in b(\pi)}x_{|B|}$
$+R(N,\theta_{0},\theta_{1})\bigg),$
where $R\in\mathbb{Q}[x,y,z]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
Proof.
We can prove this theorem in the same way as we prove Theorem˜8.23. The main difference is that in the last step of the proof of Lemma˜8.24, the total contribution is
$1-\frac{2N\theta_{0}}{1+2\theta_{1}+2N\theta_{0}}=\frac{1+2\theta_{1}}{1+2\theta_{1}+2N\theta_{0}},$
which would correspond to $\frac{1+y}{1+x+y}$ rather than $\frac{1}{1+x+y}$. ∎

### v3

@@
 We prove a generalized version of this result in Corollary˜10.18.
-Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix].
+Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]; see Lemma˜1.2 for a more direct connection with the question that is posed.
 In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$.
@@
 See Subsection˜1.6, Section˜8, and Section˜11 for more discussion regarding this direction.
+We can also restate Theorem˜1.1 in terms of the Taylor series expansions over the standard basis $M_{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ and the coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$.
+See Subsection˜2.2 for the definitions of these notions.
+Lemma 1.2.
+Repeat the statement of Theorem˜1.1. Suppose
+$F_{N}^{A}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m^{A}_{\epsilon}(N)M_{\epsilon})\right)\in\mathbb{C}[[x_{1},\ldots,x_{N}]],$
+$F_{N}^{BC}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{BC}_{\epsilon}(N)M_{\epsilon}\right)\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]],$
+$F_{N}^{D}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{D}_{\epsilon}(N)M_{\epsilon}\right)+e\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}l^{D}_{\epsilon}(N)M_{\epsilon}\right)$
+$\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]].$
+Consider the following statements.
+  1. (a’)
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  2. (a”)
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  3. (c’)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  4. (c”)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  5. (e’)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  6. (e”)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  7. (f’)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  8. (f”)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+Then, (a’) and (a”), (c’) and (c”), (e’) and (e”), and (f’) and (f”) are equivalent to (a), (c), (e), and (f), respectively.
+Remark 1.3.
+The partial derivative $\partial^{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ denotes $\prod_{i=1}^{\ell(\epsilon)}\partial_{i}^{\epsilon_{i}}$.
+We can clearly replace it with $\prod_{i=1}^{\ell(\epsilon)}\partial_{\pi(i)}^{\epsilon_{i}}$ for any injection $\pi:[\ell(\epsilon)]\rightarrow[N]$ since the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are symmetric.
+Furthermore, in (f”), $(1,\ldots,1)$ consists of $N$ copies of $1$ and $(\epsilon+(1,\ldots,1))!=\prod_{i=1}^{\ell(\epsilon)}(\epsilon_{i}+1)$.
+Proof of Lemma˜1.2.
+The equivalences involving (a’), (c’), (e’), and (f’) are straightforward by the definition of the coefficients $c_{\epsilon\lambda}$.
+The equivalences involving (a”), (c”), (e”), and (f”) follow from $m_{\epsilon}^{A}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}$, $m_{\epsilon}^{BC}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}$, $m_{\epsilon}^{D}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}$, and $l_{\epsilon}^{D}(N)=((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}$.
+∎
 1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
@@
 The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well.
-Theorems that we prove over regimes 3 and 4 are Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix].
+Theorems that we prove over regimes 3 and 4 include Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix].
+We also prove Theorem˜8.26 for regime 3 and the type D root system.
 1.3. Related works
@@
 The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8.
-In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.2.
+In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.4.
 We state a well-known conjecture related to an integral representation of the product of two Bessel functions; it can easily be extended to the product of an arbitrary number of Bessel functions.
@@
 Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$.
-There exists a nonnegative probability measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ over $\mathbb{R}^{N}$ such that
+There exists a nonnegative Borel probability measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ over $\mathbb{R}^{N}$ such that
 $J^{\mathcal{R}(\theta)}_{a_{1}}(x)J^{\mathcal{R}(\theta)}_{a_{2}}(x)=\int_{\mathbb{R}^{N}}J^{\mathcal{R}(\theta)}_{a}(x)d\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}(a)$
@@
 The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$.
-By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in ˜1.4 is satisfied.
+By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.6 is satisfied.
 It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
-The paper [rosler_product] shows that ˜1.4 holds in the context of radially symmetric Bessel functions.
+The paper [rosler_product] shows that Conjecture˜1.6 holds in the context of radially symmetric Bessel functions.
 Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
-As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in ˜1.4. For the proofs of the corollaries, see Subsection˜10.1.
+As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.6.
+For the proofs of the corollaries, see Subsection˜10.1.
 Corollary 1.7.
-Assume that ˜1.4 is true.
-Furthermore, assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$.
-Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$.
+Assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$.
+Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$.
 Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
@@
   1. (A)
-    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
+    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
+Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
   2. (B)
-    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\tilde{\mu}$ as $N\rightarrow\infty$.
-Remark 1.6.
-By “always”, we mean that the statement is true for any choices of the measures $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ and $\mu_{a(N),b(N)}^{D^{N}(\theta)}$, since there may be multiple choices.
+    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
+Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in distribution and in terms of moments.
+Conjecture˜1.6 is known for $\theta\in\{\frac{1}{2},1,2\}$.
+By [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$.
+When $\theta\in\{\frac{1}{2},1,2\}$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively.
+Then, by [voiculescu], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
 For the type BC root system, we can deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
 Corollary 1.8.
-Assume that ˜1.4 is true.
-Furthermore, assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
-Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$.
-Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
+Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\\ \frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
+Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$.
+Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\\ \delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
 Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$.
 Let $\mu$ be the rectangular free convolution with $\lambda$ set as $\frac{1}{1+c}$ of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ as defined in [rectangular_free_convolution]*Proposition-Definition 2.1.
-The distribution $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
-We also note that analogues of part (A) of Corollary˜1.5 and Corollary˜1.7 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix].
-We can similarly establish the analogue of part (B) of Corollary˜1.5 in the $\theta N\rightarrow c\in\mathbb{C}$ regime; its statement is almost the same as that of the analogue of part (A) of the corollary.
+Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6.
+Then, $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),\,b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
+We also note that analogues of part (A) of Corollary˜1.7 and Corollary˜1.8 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix].
+We can similarly establish the analogue of part (B) of Corollary˜1.7 in the $\theta N\rightarrow c\in\mathbb{C}$ regime; its statement is almost the same as that of the analogue of part (A) of the corollary.
+As expected, we can recover the free convolution from the convolution for the $\theta N\rightarrow c\in\mathbb{C}$ regime by taking $|c|\rightarrow\infty$, see [matrix]*Theorem 8.7.
+Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic.
+In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.21.
+We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
 1.5. Uniform convergence of the Bessel functions
@@
 In order to prove uniform convergence over compact subsets of an open and simply connected domain, we follow the argument of [rank_infinity] and first prove that the Bessel functions are uniformly bounded and then apply Montel’s theorem, see Subsections˜10.3 and 10.4.
-To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.2.
+To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.4.
 An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
@@
 This framework and especially the content of Section˜4 are useful for analyzing the Bessel functions.
+1.8. Equivalence conditions for non-exponential holomorphic functions
+An exponential formal power series is a formal power series $\exp(F_{N})$ for some $F_{N}\in\mathbb{C}[[x_{1},\ldots,x_{N}\\ ]]$.
+Analyzing exponential formal power series is useful since the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ are positive when $a,x\in\mathbb{R}^{N}$ and $\theta$ is nonnegative.
+Then, the Bessel generating function $\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)]$, assuming that $\mu$ is supported over $\mathbb{R}^{N}$ and the integral converges, is a positive holomorphic function and can be expressed as an exponential formal power series.
+We have established equivalence conditions for exponential formal power series.
+However, it is also straightforward to deduce analogous results for non-exponential formal power series.
+We describe these results in the following list.
+  • In Theorems˜6.10 and 6.17, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}$ and replace $\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}$.
+  • In Theorems˜7.4 and 7.8, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}$ and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$.
+  • In Theorems˜8.21 and 8.26, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
 1.9. Paper organization
@@
 In Section˜12, we analyze a special case of the leading order terms of the Dunkl bilinear form for the $A^{N-1}$ and $BC^{N}$ root systems.
+1.10. Acknowledgments
+The author would like to thank Vadim Gorin for giving comments on the paper.
 2. Basic definitions and notation
@@
 Lemma 2.2 ([dunkloperators]).
-Suppose $h\in H(\mathcal{R})$. Then, for all $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $h\mathcal{D}(\mathcal{R}(\theta))(f)h^{-1}=\mathcal{D}(\mathcal{R}(\theta))(hf)$.
+Suppose $h\in H(\mathcal{R})$. Then, for all $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $h\mathcal{D}(\mathcal{R}(\theta))\\ (f)h^{-1}=\mathcal{D}(\mathcal{R}(\theta))(hf)$.
 By the previous lemma, we have that for all $h\in H(\mathcal{R})$, $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, and $g\in\mathbb{C}[[x_{1},\ldots,x_{N}]]$, $h\mathcal{D}(\mathcal{R}(\theta))(f)g=\mathcal{D}(\mathcal{R}(\theta))(hf)hg$.
@@
   5. (e)
-    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\\ \text{Re}(\left\langle ha,x\right\rangle))$.
+    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\text{Re}(\\ \left\langle ha,x\right\rangle))$.
 We can also consider the symmetric analogue of Theorem˜2.5 after averaging over $H(\mathcal{R})$.
@@
 $\Gamma$) such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$.
+Similarly, define $\mathcal{P}_{\text{even};\,N}$ to be the set of $\lambda\in\mathcal{P}_{N}$ such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$.
 Suppose $k\geq 1$.
@@
 A $\mathbb{C}$-linear basis for $P_{2k}^{\text{even}}$ is given by $\{p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$ and a $\mathbb{C}$-linear basis for $P_{2k+N}^{\text{odd}}$ is given by $\{e(x_{1},\ldots,x_{N})p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$.
+Additionally, define the unique coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$ such that
+$M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}.$
+Remark 2.13.
+When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables.
+We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
 2.3. Non-crossing partitions
@@
 As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible.
-As explained in [rosler_positivity], in terms of Theorem˜1.2, the intertwiner is given by
+As explained in [rosler_positivity], in terms of Theorem˜1.4, the intertwiner is given by
 $Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\mu_{a}(\epsilon)$
@@
 ∎
-However, while this proof is short, we cannot use it while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems.
+Although this proof is short, we cannot use it while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems.
 Therefore, we outline a new proof of Theorem˜6.1 that can be easily applied to a more general setting.
@@
 Then, using Theorem˜6.1 after noting that each summand of $R$ is $o_{N}(|\theta N|^{|\nu|}N^{\ell(\nu)})$, we get that
-(13) $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}.$
+(13) $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[|\nu|]}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}.$
 We justify this equation. For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)<\ell(\nu)$, we have that
@@
 Corollary 6.12.
-Suppose $\epsilon\in\mathcal{P}$. Assume that
-$M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda},$
-where $c_{\epsilon\lambda}\in\mathbb{R}$ for $\lambda\in\Gamma$ are unique. Then, $c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
+Suppose $\epsilon\in\mathcal{P}$. Then, $c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
 Remark 6.13.
-When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}\\ p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables.
-We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
+See Subsection˜2.2 for the definition of $c_{\epsilon,\,(|\epsilon|)}$.
 Proof of Corollary˜6.12.
@@
 We follow the method of the proof of Theorem˜6.10. Suppose $\nu\in\Gamma$. Then, using Theorem˜6.14 after noting that each summand of $R$ is $o_{N}(N^{\ell(\nu)})$, we get that
-(16) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}.$
+(16) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[|\nu|]}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}.$
 For $\lambda\in\Gamma_{N}$ such that $\ell(\lambda)<\ell(\nu)$, we have that
@@
 $\lim_{N\rightarrow\infty}(\mathcal{W}^{A}(c)[k]+o_{N}(1))\left[c_{\lambda}(N)-c_{\lambda}\right]_{\lambda\in\Gamma[k]}^{T}=0.$
-However, observe that $\mathcal{W}^{A}(c)[k]$ is upper-triangular with a nonzero diagonal from
-Lemma˜6.16, since $c$ is not a negative integer. Thus, $\mathcal{W}^{A}(c)[k]$ is invertible, which proves (a). ∎
+However, observe that $\mathcal{W}^{A}(c)[k]$ is upper-triangular with a nonzero diagonal from Lemma 6.16, since $c$ is not a negative integer.
+Thus, $\mathcal{W}^{A}(c)[k]$ is invertible, which proves (a).
+∎
 Corollary 6.18 ([matrix]*Theorems 3.8 and 3.10).
@@
 Corollary 7.5.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
-$\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$.
+Assume that $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
+Consider the following statements.
   1. (a)
@@
 Theorem 7.8.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
-$\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$. Consider the following statements.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$.
+Assume that $\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$.
+Consider the following statements.
   1. (a)
@@
 For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$.
-Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent.
+Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B^{N}(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent.
 We similarly have that
@@
 In this section, we discuss applications of the main results of this paper.
-Some settings that we consider include the weak convergence of the measures in ˜1.4 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
+Some settings that we consider include the weak convergence of the measures in Conjecture˜1.6 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
 10.1. Weak convergence to the free convolution
-The results of this subsection are based on the assumption of ˜1.4.
+The results of this subsection are based on the assumption of Conjecture˜1.6.
 Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems.
-In particular, we prove Corollary˜1.5.
-For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.7.
+In particular, we prove Corollary˜1.7.
+For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.8.
 The following result describes the support of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$.
 As we mentioned earlier, [trimeche] shows that there exists a solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that is signed and is supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$.
-If we assume the conjecture, then we can reduce the domain of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ further, see Corollary˜10.1.
+If we assume the conjecture, then we can reduce the domain of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ further, see Lemma˜10.1.
 We conjecture that there always exists a signed solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that has the same support as what is stated in the corollary.
-Corollary 10.1.
-Assume that ˜1.4 is true and that $\theta\in\theta(\mathcal{R})$ is nonnegative.
-Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$.
-Then, for any choice of $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
+Lemma 10.1.
+Assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
+Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
+Then, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
 Proof.
-Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.2. Then, we have that
+Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.4. Then, we have that
 $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}(\epsilon_{i}^{1}+\epsilon_{i}^{2})}d\mu_{a_{1}}^{\text{sym}}(\epsilon^{1})d\mu_{a_{2}}^{\text{sym}}(\epsilon^{2})=\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)d\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}(a).$
@@
 $\Pr[\epsilon\in B(u,r)]\geq\Pr_{a\sim\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right].$
-For the sake of contradiction, assume that $\Pr_{a\sim\mu_{a_{1},a_{2}},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right]=0$.
-Then, there exists a sequence $\{a_{i}\}_{i\geq 1}$ of elements of $B(u,r)$ such that
-$\lim_{i\rightarrow\infty}\mu_{a_{i}}^{\text{sym}}\left[B(a_{i},\frac{r}{2})\right]=0,$
+For the sake of contradiction, assume that $\Pr_{a\sim\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right]=0$.
+Then, there exists a sequence $\{z_{i}\}_{i\geq 1}$ of elements of $B(u,r)$ such that
+$\lim_{i\rightarrow\infty}\mu_{z_{i}}^{\text{sym}}\left[B(z_{i},\frac{r}{2})\right]=0,$
 since $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}\left[B(u,\frac{r}{2})\right]>0$.
-Let $a$ be an element of the closed ball $\overline{B}(u,r)$ that is the limit of a subsequence $\{a_{i_{j}}\}_{j\geq 1}$ of $\{a_{i}\}_{i\geq 1}$.
+Let $z$ be an element of the closed ball $\overline{B}(u,r)$ that is the limit of a subsequence $\{z_{i_{j}}\}_{j\geq 1}$ of $\{z_{i}\}_{i\geq 1}$.
 Observe that for all $x\in\mathbb{R}^{N}$,
-$\lim_{j\rightarrow\infty}J_{a_{i_{j}}}^{\mathcal{R}(\theta)}(ix)=J_{a}^{\mathcal{R}(\theta)}(ix).$
-Hence, it follows that the characteristic functions of $\mu^{\text{sym}}_{a_{i_{j}}}$ converge pointwise to the characteristic function of $\mu^{\text{sym}}_{a}$.
-By Lévy’s continuity theorem, $\mu^{\text{sym}}_{a_{i_{j}}}$ converges weakly to $\mu^{\text{sym}}_{a}$.
+$\lim_{j\rightarrow\infty}J_{z_{i_{j}}}^{\mathcal{R}(\theta)}(ix)=J_{z}^{\mathcal{R}(\theta)}(ix).$
+Hence, it follows that the characteristic functions of $\mu^{\text{sym}}_{z_{i_{j}}}$ converge pointwise to the characteristic function of $\mu^{\text{sym}}_{z}$.
+By Lévy’s continuity theorem, $\mu^{\text{sym}}_{z_{i_{j}}}$ converges weakly to $\mu^{\text{sym}}_{z}$.
 It then follows that
-$\mu^{\text{sym}}_{a}\left(B(a,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a_{i_{j}},\frac{r}{2})\right)=0,$
-which is a contradiction since $a\in\text{supp}(\mu_{a}^{\text{sym}})$ by Theorem˜1.2. Thus, $u$ is in the support of $\epsilon$.
+$\mu^{\text{sym}}_{z}\left(B(z,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z_{i_{j}},\frac{r}{2})\right)=0,$
+which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.4. Thus, $u$ is in the support of $\epsilon$.
 However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$.
-Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.2, the result follows.
+Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.4, the result follows.
 ∎
 In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$.
-Afterwards, it is straightforward to deduce Corollaries˜1.5 and 1.7.
+Afterwards, it is straightforward to deduce Corollaries˜1.7 and 1.8.
 Corollary 10.2.
-Assume that ˜1.4 is true. Suppose that for $N\geq 1$, $a(N),\\ b(N)\in\mathbb{R}^{N}$. For all $N\geq 1$, assume that $\theta,\theta_{0},\theta_{1}\geq 0$.
+Suppose that for $N\geq 1$, $a(N),b(N)\in\mathbb{R}^{N}$. For all $N\geq 1$, assume that $\theta,\theta_{0},\theta_{1}\geq 0$.
   • (A)
@@
     $m_{d}^{a}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|a_{(|B|)}\text{ and }m_{d}^{b}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|b_{(|B|)}$
-    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
+    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
     $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{d}\right]}{(\theta N)^{d}N}=\sum_{\pi\in NC(d)}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
     $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
-    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
+    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
     $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta_{0}N)^{2d}N}$
@@
     $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
-    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
+    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
     $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta N)^{2d}N}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
 Remark 10.3.
-We note that the proof of Corollary˜10.2 is straightforward and that the framework of the proof has appeared previously, for example see [matrix]*Proof of Theorem 1.5.
+We note that the proof of Corollary˜10.2 is straightforward and that its framework has appeared previously, for example see [matrix]*Proof of Theorem 1.5.
+The main contribution of the proof is the application of Corollary˜6.11.
 Proof of Corollary˜10.2.
@@
 Then, we can express $J_{a(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}a_{\lambda}(N)p_{\lambda}\right)$ and $J_{b(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}b_{\lambda}(N)p_{\lambda}\right)$ over $\Omega_{N}$.
-For $N\geq 1$, $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Corollary˜10.1 and is therefore exponentially decaying, see Definition˜10.15.
-Thus, part (C) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
+For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15.
+Thus, part (E) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
 (30) $\begin{split}\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]&=\left[p_{\lambda},J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}\right]_{A^{N-1}(\theta)}\\ &=\left[p_{\lambda},\exp\left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)\right]_{A^{N-1}(\theta)};\end{split}$
@@
 (B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.22, respectively, and follow the same idea as the proof of (A). ∎
-With this result, we are prepared to prove Corollary˜1.5.
-Proof of Corollary˜1.5.
-(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta(\frac{a_{i}}{\theta N})\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$.
+With this result, we are prepared to prove Corollary˜1.7.
+Proof of Corollary˜1.7.
+(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$.
 Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition.
 Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$.
-It follows from [billingsley]*Theorem 30.2 that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\right\}_{N\geq 1}$ converges in distribution to $\mu$.
+It follows from [billingsley]*Theorem 30.2 that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in distribution to $\mu$.
 (B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$.
 It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero.
-Using part (C) of Corollary˜10.2, we can then deduce that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(\frac{a_{i}}{2\theta N}\right)\right]\right\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$.
+Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$.
 Afterwards, we can follow the steps of (A).
 ∎
-We can use the same framework to prove Corollary˜1.7.
-Proof of Corollary˜1.7.
+We can use the same framework to prove Corollary˜1.8.
+Proof of Corollary˜1.8.
 […]
@@
 10.2. Upper bounds of the magnitudes of Bessel functions
-In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.2.
+In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.4.
 Lemma 10.4.
@@
 Proof.
-First, observe that from Theorem˜1.2,
+First, observe that from Theorem˜1.4,
 $J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
@@
 Proof.
-(A): We use Theorem˜1.2.
+(A): We use Theorem˜1.4.
 Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$.
@@
 $\left\lVert e^{\left\langle a,b\right\rangle}\right\rVert_{2}=e^{\text{Re}(\left\langle a,b\right\rangle)}\leq e^{\left\lVert\left\langle a,b\right\rangle\right\rVert_{2}}\leq e^{\left\langle\tilde{a},\tilde{b}\right\rangle}$.
-Therefore, after referring to Theorem˜1.2, we can deduce that
+Therefore, after referring to Theorem˜1.4, we can deduce that
 $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\tilde{x},\tilde{\epsilon}\right\rangle}d\mu_{a}(\epsilon).$
@@
     Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ such that $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ is invertible.
-The value of $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0\ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}z^{2k}$.
+The value of $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0,\ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}z^{2k}$.
   3. (C)
@@
 (A): Suppose $k\geq 1$. Let $r_{k}$ be the homogeneous degree $k$ polynomial such that $r_{k}(a)$ is the coefficient of $z^{k}$ in $J_{a}^{A^{N-1}(\theta)}(z,0,\ldots,0)$. Then,
-$J_{a}^{A^{N-1}(\theta)}(z,0\ldots,0)=1+\sum_{k=1}^{\infty}r_{k}(a)z^{k}.$
+$J_{a}^{A^{N-1}(\theta)}(z,0,\ldots,0)=1+\sum_{k=1}^{\infty}r_{k}(a)z^{k}.$
 The goal is to show that $r_{k}=\frac{c_{k}}{(\theta N)_{k}}$.
@@
 Theorem 10.10.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$. Consider the following statements.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$.
+Assume that $\theta\in\mathbb{C}$ is fixed.
+Consider the following statements.
   1. (a)
@@
 Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$.
-Assume that $\theta_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
+Assume that $\theta_{0}\in\mathbb{C}$ is fixed and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
 Consider the following statements.
@@
 Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$.
-Assume that $\theta\in\mathbb{C}$.
+Assume that $\theta\in\mathbb{C}$ is fixed.
 Consider the following statements.
@@
 Proof.
-(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.2 and Hölder’s inequality gives that
+(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.4 and Hölder’s inequality gives that
 $\left|J^{A^{N-1}(\theta)}_{a(N)}(x_{1},\ldots,x_{r},0,\ldots,0)\right|$ $=\left|\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{r}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)\right|$
@@
 It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$.
-Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$.
+Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity].
 We can then apply part (B) of Lemma˜10.8 to conclude the proof.
@@
 This notion is also related to the Dunkl transform discussed in [dunklbound].
-Suppose $\mu$ is a probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
+Suppose $\mu$ is a Borel probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
 $G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)].$
@@
 This implication as well as other essential implications are included in the following lemma.
-Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (C) is [matrix]*Proposition 2.11).
+Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (E) is [matrix]*Proposition 2.11).
 Suppose the Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$.
@@
   2. (B)
-    There exist unique constants $c_{\lambda}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
-    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(\mu)p_{\lambda}\right)$
+    There exist unique constants $c_{\lambda}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
+    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}^{\mathcal{R}(\theta)}(\mu)p_{\lambda}\right)$
     in a neighborhood of the origin.
   3. (C)
+    There exist unique constants $m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ such that
+    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)M_{\epsilon}\right)$
+    in a neighborhood of the origin.
+  4. (D)
+    For all $\lambda\in\Gamma$, the function $\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and equals
+    $\int_{\mathbb{C}^{N}}p_{\lambda}(a)J_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
+    over the open ball of radius $R$ centered at the origin.
+  5. (E)
     For all $\lambda\in\Gamma$,
-    $1]\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)=\operatorname{\mathbb{E}}_{a\sim\mu}[p_{\lambda}(a)].$
+    $\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(0)=\operatorname{\mathbb{E}}_{a\sim\mu}[p_{\lambda}(a)].$
 Proof.
-For (B), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$. ∎
+For (B) and (C), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
+(D): We include a detailed proof of this statement for completeness.
+We prove this result for the nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
+It is straightforward to deduce the result for the Bessel function afterwards using the expression relating the two functions in Theorem˜2.8.
+Note that some of the ideas of this proof have previously appeared in [limits_prob_measures]*Remark 4.6.
+Suppose $x\in\overline{B}_{R}(0)$, which is the closed ball of radius $R$ centered at the origin. By part (e) of Lemma˜2.6, the integral
+$\int_{\mathbb{C}^{N}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
+converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$.
+Furthermore, we can show that the integral is holomorphic over $\overline{B}_{R}(0)$ by following [matrix]*Proof of Lemma 2.9.
+Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
+The goal is to prove using induction that for any sequence $\{i_{j}\}_{1\leq j\leq m}$ of elements of $[N]$, we have that for $x\in B_{R}(0)$,
+$\prod_{j=1}^{m}\mathcal{D}_{i_{j}}\int_{\mathbb{C}^{N}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
+and both sides of the equation converge to a holomorphic function over $B_{R}(0)$, which is the open ball of radius $R$ centered at the origin.
+We have already shown that this statement is true when the sequence is empty.
+Assume that the statement is true for the sequence $\{i_{j}\}_{1\leq j\leq m}$.
+We show that the statement is true for the sequence $\{i_{j}\}_{1\leq j\leq m+1}$.
+It suffices to prove that for $x\in B_{R}(0)$,
+(32) $\mathcal{D}_{i_{j+1}}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m+1}a_{i_{j+1}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
+and that the right hand side is holomorphic over $B_{R}(0)$. Note that $\left\lVert x\right\rVert_{2}<R$, so applying part (e) of Lemma˜2.6 gives that
+$\left|\prod_{j=1}^{m+1}a_{i_{j+1}}E_{a}^{\mathcal{R}(\theta)}(x)\right|\leq\sqrt{|H(\mathcal{R})}|\prod_{j=1}^{m+1}a_{i_{j+1}}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2}),$
+which implies that the right hand side converges.
+Afterwards, we can use the method used in [matrix]*Proof of Lemma 2.9 to show that the right hand side is holomorphic.
+To complete the proof, it suffices to show that (32) is true.
+Without loss of generality, assume that $i_{m+1}=1$.
+Suppose $x\in B_{R}(0)$. Observe that
+(33) $\partial_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\lim_{\epsilon\rightarrow 0}\int_{\mathbb{C}^{N}}\frac{\prod_{j=1}^{m}a_{i_{j}}(E_{a}(x_{1}+\epsilon,x_{2},\ldots,x_{N})-E_{a}(x))}{\epsilon}d\mu(a).$
+By [dunklbound]*Lemma 3.5,
+$\frac{|E_{a}(x_{1}+\epsilon,x_{2},\ldots,x_{N})-E_{a}(x)|}{\epsilon}\leq e\sqrt{|H(\mathcal{R})|}\left\lVert a\right\rVert_{2}\exp(\left\lVert a\right\rVert_{2}(\left\lVert x\right\rVert_{2}+\epsilon)).$
+If we assume that $\epsilon<R-\left\lVert x\right\rVert_{2}$, then the absolute value of the integrand of the right hand side of (33) is upper bounded by an integrable function, since $\mu$ is exponentially decaying at rate $R$.
+Then, we can apply the dominated convergence theorem to deduce that
+$\partial_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\partial_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}(x)d\mu(a).$
+Suppose $\alpha\in\mathcal{R}^{+}$. Then, it is clear that
+$\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)-\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(r_{\alpha}x)d\mu(a)$
+$=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}(E_{a}^{\mathcal{R}(\theta)}(x)-E_{a}^{\mathcal{R}(\theta)}(r_{\alpha}x))d\mu(a)$
+since $r_{\alpha}x\in B_{R}(0)$. After using the expression given in $\mathcal{D}_{1}$ in Subsection˜2.1, we deduce that
+$\mathcal{D}_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\mathcal{D}_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}a_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a),$
+which completes the induction.
+(E): This is straightforward to deduce from (D). ∎
 Remark 10.17.
 The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system.
-However, (A) and (C) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
+However, (A) and (E) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
+Furthermore, observe that (A) implies (B) and (C).
+However, we require the exponentially decaying condition to prove (D) and (E).
 Using Lemma˜10.16 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system.
@@
 Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 1$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-Define $c_{\lambda}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ as in part (B) of Lemma˜10.16. Then, the following are equivalent.
+Define $G_{\mu}^{A^{N-1}(\theta)}$, $c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$, and $m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in Lemma˜10.16.
+Then, the following are equivalent.
   1. (a)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  2. (a’)
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  3. (a”)
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   4. (b)
@@
     $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
-Remark 10.19.
 Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
+When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers.
+Furthermore, we can generalize Corollary˜1.7 and Corollary˜1.8 to when $a(N)$ and $b(N)$ are randomly distributed. We first state the following generalization of Lemma˜10.1.
+Lemma 10.19.
+Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative.
+Suppose $\mu_{1}$, $\mu_{2}$, and $\mu$ are Borel probability measures over $\mathbb{R}^{N}$ such that $G_{\mu}^{\mathcal{R}(\theta)}=G_{\mu_{1}}^{\mathcal{R}(\theta)}G_{\mu_{2}}^{\mathcal{R}(\theta)}$ over an open subset of $i\mathbb{R}^{N}$.
+Then, every element of $\text{supp}(\mu)$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$ for $i\in\{1,2\}$.
+Proof.
+We follow the method of the proof of Lemma˜10.1.
+We have that for $x$ in an open subset of $i\mathbb{R}^{N}$, the characteristic function of the random variables $\epsilon_{1}+\epsilon_{2}$ where $a_{i}\sim\mu_{i}$ and $\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}$ for $i\in\{1,2\}$ and $\epsilon$ where $a\sim\mu$ and $\epsilon\sim\mu_{a}^{\text{sym}}$.
+Then, since the characteristic function is defined based on its values in an open set, the characteristic functions of the two random variables are equal everywhere, so the two random variables have the same distribution.
+We can similarly deduce that for all $\in\text{supp}(\mu), uu$ is in the support of $\epsilon$. Additionally, the support of $\epsilon$ is the same as the support of $\epsilon^{1}+\epsilon^{2}$.
+Suppose $i\in\{1,2\}$, $u\in\text{supp}(\epsilon_{i})$, and $r>0$. Observe that
+(34) $\int_{\mathbb{R}^{N}}\Pr_{\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}}[\epsilon_{i}\in B(u,r)]d\mu_{i}(a_{i})=\Pr[\epsilon_{i}\in B(u,r)]>0.$
+For the sake of contradiction, assume that the distance between $u$ and $\text{supp}(\mu_{a_{i}}^{\text{sym}})$ for any element $a_{i}\in\text{supp}(\mu_{i})$ is greater than $2r$.
+Then, for all $a_{i}\in\text{supp}(\mu_{i})$, $\Pr_{\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}}[\epsilon_{i}\in B(u,r)]=0$, which is a contradiction to (34).
+Recall from Theorem˜1.4 that for $a\in\mathbb{R}^{N}$, the support of $\mu_{a}^{\text{sym}}$ is a subset of the convex hull of $H(\mathcal{R})a$.
+By considering when $r\rightarrow 0$, we have that $u$ must be in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$.
+∎
+Corollary 10.20.
+Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative.
+  1. (A)
+    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are exponentially decaying Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ such that:
+    1. 1.
+      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{R}$ in terms of moments.
+    2. 2.
+      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{R}$ in terms of moments.
+    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
+Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
+  2. (B)
+    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
+Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
+Proof.
+(A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution.
+Instead of using Theorem˜1.1, we use Corollary˜10.18.
+(B): We use Lemma˜10.19 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying.
+Afterwards, we follow the proof of (A) for the remaining steps of the proof.
+∎
+We can also state the analogues of Corollary˜10.20 for the type BC and D root systems, which are similar to Corollary˜1.8 and part (B) of Corollary˜1.7, respectively.
+To determine what the differences are, it suffices to compare Corollary˜10.20 with part (A) of Corollary˜1.7.
+Example 10.21.
+As an example, we consider the Dyson Brownian motion with a random initial condition.
+For $\theta>0$, the Dyson Brownian motion $DBM(\theta)$ over $\mathbb{R}^{N}$ with initial values $(a_{1}(0),\ldots,a_{N}(0))$ is the unique strong solution to
+[…]
+where $B_{i}$, $1\leq i\leq N$ are independent standard Brownian motions.
+For $t\geq 0$, let $\mu_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
+(36) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=B_{a}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
+for all $x\in\mathbb{C}^{N}$.
+Suppose $\eta$ is a probability measure over $\mathbb{R}^{N}$.
+Let $\mu_{t}(\eta)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ if we assume that the initial values $(a_{1}(0),\ldots,a_{N}(0))$ are distributed according to $\eta$ independently of the Brownian motions.
+In particular, we define
+$\mu_{t}(\eta)\triangleq\operatorname{\mathbb{E}}_{a\sim\eta}[\mu_{t}(a)].$
+Lemma 10.22.
+Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then,
+$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
+over $\overline{B}_{R}(0)$ for all $t\geq 0$.
+Proof.
+See Lemma˜10.16 and (36). ∎
+Furthermore, for $\theta>0$, define Hermite $\theta$-ensemble to be the random variable $H_{\theta}$ over $\mathbb{R}^{N}$ with density at $a$ proportional to $\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{2\theta}\prod_{i=1}^{N}e^{-\frac{\theta a_{i}^{2}}{2}}$.
+By [betaprocess]*Proposition 4.2,
+$G_{H_{\theta}}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{1}{2\theta}\sum_{i=1}^{N}x_{i}^{2}\right)$
+over $\mathbb{C}^{N}$.
+For $t>0$, let $H_{\theta}^{t}$ be the pushforward of $H_{\theta}$ with respect to multiplication by $t$, so that $H_{\theta}^{t}$ is a random variable over $\mathbb{R}^{N}$ with density at $a$ proportional to $\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{2\theta}\prod_{i=1}^{N}e^{-\frac{\theta a_{i}^{2}}{2t^{2}}}$.
+This random variable is studied in the paper [airy_beta] and Lemma 3.7 of the paper computes that
+$G_{H_{\theta}^{t}}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t^{2}}{2\theta}\sum_{i=1}^{N}x_{i}^{2}\right)$
+over $\mathbb{C}^{N}$.
+Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.22 implies that
+(37) $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
+over a neighborhood of the origin for exponentially decaying Borel probability measures $\eta$. This example satisfies Conjecture˜1.6 for randomly distributed $a_{1}$ and $a_{2}$.
+Corollary 10.23.
+Suppose $\theta>0$ and that for $N\geq 1$, $\eta_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{R}^{N}$ such that as $N\rightarrow\infty$, $\operatorname{\mathbb{E}}_{a\sim\eta_{N}}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the distribution $\eta$ over $\mathbb{R}$.
+Then, for $\alpha\geq 0$, $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the free convolution of $\eta$ and the semicircular law multiplied by $\sqrt{\frac{\alpha}{\theta}}$ as $N\rightarrow\infty$.
+If $\eta$ is compactly supported, then $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges weakly to the free convolution as $N\rightarrow\infty$.
+Proof.
+Assume that $\alpha>0$.
+We can follow the method of the proof of Corollary˜10.20.
+First, since condition (b) of Corollary˜10.18 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well.
+The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$.
+This does not follow from (37) and Lemma˜10.19 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
+Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $\alpha N$ conditioned on the initial value $a$ is at most
+$\exp\left(-\frac{(\left\lVert a\right\rVert_{2}-\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}\right)y^{O(1)},$
+where we have applied the upper bound $B_{a}\left(\frac{y}{\alpha N}\right)\leq\sqrt{N!}\exp\left(\frac{\left\lVert a\right\rVert_{2}\left\lVert y\right\rVert_{2}}{\alpha N}\right)$ that we deduce from part (e) of Lemma˜2.6.
+Assume that $\eta_{N}$ exponentially decays at rate $R$. Then, we have that for all $\epsilon>0$,
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
+and
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
+This implies that $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$.
+Then, because
+$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=G_{\eta_{N}}^{A^{N-1}(\theta)}(x)\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
+over $\overline{B}_{R}(0)$ by Lemma˜10.22, applying Corollary˜10.18 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$.
+Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$.
+If $\eta$ is compactly supported, then the free convolution is compactly supported and therefore uniquely determined by its moments by Carleman’s condition, which implies that the sequence of probability measures also converges weakly to the free convolution after applying [billingsley]*Theorem 30.2.
+If $\alpha=0$, then the convergence in terms of moments is clear and the weak convergence follows from [billingsley]*Theorem 30.2. ∎
 11. Computing the Dunkl bilinear form with combinatorics
@@
 Note that $A^{0}$, $B^{1}$, $C^{1}$, and $D^{1}$ are not actually root systems; however, we use this notation for simplicity.
-The following result computes combinatorial expressions for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$, $p_{\lambda},\\ p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}$, and $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ in terms of the values of the quantities for small values of $N$, including $N=1$.
+The following result computes combinatorial expressions for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$, $p_{\lambda},p_{\nu}\\ ]_{BC^{N}(\theta_{0},\theta_{1})}$, and $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ in terms of the values of the quantities for small values of $N$, including $N=1$.
 Theorem 11.1.

### v4

@@
 (C): Suppose
-$F_{N}^{D}(x_{1},\ldots,x_{N})$ $=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$
+$F_{N}^{D}(x_{1},\ldots,x_{N})$ $=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)++ed_{0}(N)+e\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$
 $\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]$
@@
   2. (f)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
   3. (g)
@@
   4. (h)
-    For all $\nu\in\Gamma_{\text{even}}$,
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))F_{N}^{D}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
+    and for all $\nu\in\Gamma_{\text{even}}$,
     $\lim_{N\rightarrow\infty}\frac{[1]\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}F_{N}^{D}}{N^{\ell(\nu)}(\theta N)^{|\nu|}\prod_{j=1}^{N}(1+2(j-1)\theta)}$
@@
 Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin.
-We prove a generalized version of this result in Corollary˜10.18.
+We prove a generalized version of this result in Corollary˜10.17.
 Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]; see Lemma˜1.2 for a more direct connection with the question that is posed.
@@
   1. (a’)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  2. (a”)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   2. (c’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  4. (c”)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   3. (e’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  6. (e”)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   4. (f’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  8. (f”)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-Then, (a’) and (a”), (c’) and (c”), (e’) and (e”), and (f’) and (f”) are equivalent to (a), (c), (e), and (f), respectively.
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+Then, (a’), (c’), (e’), and (f’) are equivalent to (a), (c), (e), and (f), respectively.
 Remark 1.3.
@@
 Proof of Lemma˜1.2.
-The equivalences involving (a’), (c’), (e’), and (f’) are straightforward by the definition of the coefficients $c_{\epsilon\lambda}$.
-The equivalences involving (a”), (c”), (e”), and (f”) follow from $m_{\epsilon}^{A}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}$, $m_{\epsilon}^{BC}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}$, $m_{\epsilon}^{D}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}$, and $l_{\epsilon}^{D}(N)=((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}$.
-∎
+The equivalences are straightforward after using the definition of the coefficients $c_{\epsilon\lambda}$. ∎
 1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
@@
 Corollary 1.8.
-Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\\ \frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
+Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}\\ =c\in\mathbb{C}$.
 Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$.
-Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\\ \delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
+Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
 Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$.
@@
 Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic.
-In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.21.
+In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.26.
 We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
@@
   5. (e)
-    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\text{Re}(\\ \left\langle ha,x\right\rangle))$.
+    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\\ \text{Re}(\left\langle ha,x\right\rangle))$.
 We can also consider the symmetric analogue of Theorem˜2.5 after averaging over $H(\mathcal{R})$.
@@
   • The function $h:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $h(x)\triangleq\prod_{r\in\mathcal{R}^{+}}|\left\langle x,r\right\rangle|^{\theta(r)}$.
-  • The function $w:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $w(x)\triangleq\frac{e^{-\left\lVert x\right\rVert_{2}^{2}}}{(2\pi)^{\frac{N}{2}}}$.
+  • The function $w:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $w(x)\triangleq\frac{e^{-\frac{\left\lVert x\right\rVert_{2}^{2}}{2}}}{(2\pi)^{\frac{N}{2}}}$.
   • The constant $c_{N}$ is defined as $c_{N}\triangleq\int_{\mathbb{R}^{N}}h^{2}wdx$.
@@
 Theorem 8.21.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
 Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
@@
   2. (b)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=d_{\lambda}\in\mathbb{C}$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=d_{\lambda}\in\mathbb{C}$.
   3. (c)
@@
   4. (d)
-    For all $\nu\in\Gamma_{\text{even}}$,
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1.$
+    Furthermore, for all $\nu\in\Gamma_{\text{even}}$,
     $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{(\theta N)^{|\nu|}N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{M}^{D}_{\nu\lambda}[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
@@
 For the equivalence of (b) and (d), we can use Theorem˜8.16 and follow the method of the proof of Theorem˜6.10, after dividing by the nonzero quantity $\prod_{j=1}^{N}(1+2(j-1)\theta)$.
-∎
+In particular, we have that $e,F_{N}]_{D^{N}(\theta)}=(d_{0}(N)+1)[e,e]_{D^{N}(\theta)}=(d_{0}(N)+1)\prod_{j=1}^{N}(1+2(j-1)\theta))$, so we can deduce that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}$ and
+$\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
+are equivalent. ∎
 Corollary 8.22.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
 Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
@@
   2. (b)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
   3. (c)
@@
   4. (d)
-    For all $\nu\in\Gamma_{\text{even}}$,
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1.$
+    Furthermore, for all $\nu\in\Gamma_{\text{even}}$,
     $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}(\theta N)^{|\nu|}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
@@
 Theorem 8.26.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
 Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$.
@@
   2. (b)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
   3. (c)
@@
   4. (d)
-    For all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
+    and for all $\nu\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
+    $=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
 Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
+Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
 Proof.
 For the implications involving (a) and (c), see Theorem˜8.13.
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$.
+Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$.
 Using Theorem˜8.23 gives that $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ is divisible by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$.
@@
 Corollary 8.27.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
 Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$.
@@
   2. (b)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}d_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}d_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
   3. (c)
@@
   4. (d)
-    For all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}$
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
+    and for all $\nu\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
     $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(|\nu|+1)}W^{D;\,\text{odd}}(\pi)(2c)\prod_{B\in b(\pi)}(1+2c)^{|B|}|B|d_{(|B|)},$
@@
 Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
+Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
 Remark 8.28.
@@
 Corollary 9.11.
-If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
+If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then
+$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
+and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
 $J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}(\theta N)^{-k}\theta^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}2^{r_{l}-1}r_{l}\right)^{-1}\times$
@@
 Proof.
-See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10. Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$. ∎
+See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10.
+Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$.
+Furthermore, we obtain the term $\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$ in $J_{a}^{D^{N}(\theta)}$ from $e,J_{a}^{D^{N}(\theta)}]_{D^{N}(\theta)}=e(a)$.
+∎
 Lemma 9.12.
@@
 Corollary 9.14.
-If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
+If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then
+$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
+and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
 $J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}\right)^{-1}\bigg[\prod_{l=1}^{\ell(r)}W^{A}([r_{l}])(2c)^{-1}+o_{N}(1)$
@@
 Then, we can express $J_{a(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}a_{\lambda}(N)p_{\lambda}\right)$ and $J_{b(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}b_{\lambda}(N)p_{\lambda}\right)$ over $\Omega_{N}$.
-For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15.
-Thus, part (E) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
+For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.14.
+Thus, part (D) of Lemma˜10.15 implies that for all $\lambda\in\Gamma$,
 (30) $\begin{split}\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]&=\left[p_{\lambda},J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}\right]_{A^{N-1}(\theta)}\\ &=\left[p_{\lambda},\exp\left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)\right]_{A^{N-1}(\theta)};\end{split}$
-the second equality follows from the fact that $J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}=\exp\\ \left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)$ over $\Omega_{N}$ and $0\in\Omega_{N}$.
+the second equality follows from the fact that $J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}=\exp\big(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+\\ b_{\lambda}(N))p_{\lambda}\big)$ over $\Omega_{N}$ and $0\in\Omega_{N}$.
 By Corollary˜6.11, we have that
@@
 $(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|a_{(|B|)}=(1+c)^{-E(\pi)}\prod_{B\in\pi}(1+c)^{|B|}2^{|B|-1}|B|a_{(|B|)};$
-[…]
+the equation also holds after replacing $a_{(|B|)}$ with $b_{(|B|)}$.
+By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\\ \left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\mu$.
+∎
 10.2. Upper bounds of the magnitudes of Bessel functions
@@
 Lemma 10.4.
-Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a,x\in\mathbb{C}^{N}$. Then,
-$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}\leq\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
+Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a\in\mathbb{R}^{N}$ and $x\in\mathbb{C}^{N}$. Then,
+$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}\leq\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})ha_{i}}.$
 Proof.
@@
 for a measure $\mu_{a}$ over $\mathbb{R}^{N}$ which is supported on the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$. In particular,
-$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}d\mu_{a}(\epsilon).$
+$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon).$
 Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ and $\sum_{h\in H}c_{h}=1$. Then, for $h^{\prime}\in H$,
-$e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}=e^{\sum_{i=1}^{N}\text{Re}(x_{i}\sum_{h\in H}c_{h}h^{\prime}ha_{i})}\leq\sum_{h\in H}c_{h}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}$
+$e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}=e^{\sum_{i=1}^{N}\text{Re}(x_{i})\sum_{h\in H}c_{h}h^{\prime}ha_{i}}\leq\sum_{h\in H}c_{h}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}ha_{i}}$
 by Jensen’s inequality. It follows that
-$\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}\leq\sum_{h\in H}\frac{c_{h}}{|H|}\sum_{h^{\prime}\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}=\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
+$\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}\leq\sum_{h\in H}\frac{c_{h}}{|H|}\sum_{h^{\prime}\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}ha_{i}}=\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})ha_{i}}.$
 This completes the proof. ∎
 Lemma 10.5.
-Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$.
-Suppose $R$ and $I$ are nonnegative real numbers.
+Suppose $N\geq 1$, $a\in\mathbb{R}^{N}$, and $x\in\mathbb{C}^{N}$.
+Suppose $R$ is a nonnegative real number.
 Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$.
@@
   1. (A)
-    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq I^{d}$. Then,
-    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{R\sum_{i=1}^{N}|\text{Re}(x_{i})|+I\sum_{i=1}^{N}|\text{Im}(x_{i})|}.$
+    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a_{i}|^{d}\leq R^{d}$. Then,
+    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{R\sum_{i=1}^{N}|\text{Re}(x_{i})|}.$
   2. (B)
-    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq d!R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq d!I^{d}$. Then,
-    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\left(\sum_{d=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d}\sum_{d=0}^{\infty}\left(2I\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{d}\right)^{\frac{1}{2}}.$
+    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a_{i}|^{d}\leq d!R^{d}$. Then,
+    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d}.$
 Proof.
@@
 First, observe that
-$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle-\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)$
-$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{-2\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}$
-$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}.$
+$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\mu_{a}(\epsilon)\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon).$
 Next, note that since $\mu_{a}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
-$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{C}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(2|\text{Re}(x)|)M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
+$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
 Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ satisfy $\sum_{h\in H}c_{h}=1$. Furthermore, suppose $\nu\in\Gamma$. By Muirhead’s inequality and the triangle inequality,
-$\frac{M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)}\leq\frac{\sum_{i=1}^{N}|\text{Re}(\epsilon_{i})|^{|\nu|}}{N}\leq\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}.$
+$\frac{M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)}\leq\frac{\sum_{i=1}^{N}|\epsilon_{i}|^{|\nu|}}{N}\leq\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|ha_{i}|\right)^{|\nu|}}{N}.$
 By Jensen’s inequality, we have that
-$\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}\leq\frac{1}{N}\sum_{h\in H}c_{h}\sum_{i=1}^{N}|\text{Re}(ha_{i})|^{|\nu|}=\frac{\sum_{i=1}^{N}|\text{Re}(a_{i})|^{|\nu|}}{N}\leq R^{|\nu|}.$
+$\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|ha_{i}|\right)^{|\nu|}}{N}\leq\frac{1}{N}\sum_{h\in H}c_{h}\sum_{i=1}^{N}|ha_{i}|^{|\nu|}=\frac{\sum_{i=1}^{N}|a_{i}|^{|\nu|}}{N}\leq R^{|\nu|}.$
 Hence,
-(31) $\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\nu!}=e^{2R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
-Similarly,
-$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq e^{2I\sum_{i=1}^{N}|\text{Im}(x)_{i}|}.$
-Afterwards, it is straightforward to conclude the proof.
+(31) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
 (B): The equation (31) becomes
-$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\pi(\nu)}=\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{k}.$
-Similarly,
-$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{k},$
+$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
 which completes the proof. ∎
-Lemma 10.6.
-Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$.
-Suppose $M$ is a nonnegative real number.
-Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$.
-Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
-  1. (A)
-    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq M^{d}$. Then,
-    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}}.$
-  2. (B)
-    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq d!M^{d}$. Then,
-    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\sum_{d=0}^{\infty}\left(M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}\right)^{d}.$
-Proof.
-For $z\in\mathbb{C}^{N}$, we let $\tilde{z}$ denote $(\left\lVert z_{i}\right\rVert_{2})_{i\in[N]}$. First, observe that for $a,b\in\mathbb{C}^{N}$,
-$\left\lVert e^{\left\langle a,b\right\rangle}\right\rVert_{2}=e^{\text{Re}(\left\langle a,b\right\rangle)}\leq e^{\left\lVert\left\langle a,b\right\rangle\right\rVert_{2}}\leq e^{\left\langle\tilde{a},\tilde{b}\right\rangle}$.
-Therefore, after referring to Theorem˜1.4, we can deduce that
-$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\tilde{x},\tilde{\epsilon}\right\rangle}d\mu_{a}(\epsilon).$
-Afterwards, we can apply the same method used in the proof of Lemma˜10.5. ∎
 Parts (A) and (B) of the following lemma are from [rank_infinity], where they are proved using results from [baker2, baker, okounkov_olshanki, bessel_convolution, forrester].
@@
     Assume that $\theta\in\mathbb{C}$ such that $\mathcal{D}(D^{N}(\theta))$ is invertible.
-The value of $J_{a}^{D^{N}(\theta)}(z,0,\\ \ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}((N-1)\theta+\frac{1}{2})_{k}(\theta N)_{k}}z^{2k}$.
+The value of $J_{a}^{D^{N}(\theta)}(z,0,\ldots,\\ 0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}((N-1)\theta+\frac{1}{2})_{k}(\theta N)_{k}}z^{2k}$.
 Proof.
@@
 Theorem 10.8.
-Suppose $m_{d}\in\mathbb{C}$ for $d\geq 1$.
-Assume that for $N\geq 1$, $a(N)\in\mathbb{C}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N}=m_{d}$ for all $d\geq 1$.
+Suppose $m_{d}\in\mathbb{R}$ for $d\geq 1$.
+Assume that for $N\geq 1$, $a(N)\in\mathbb{R}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}a(N)_{i}^{d}=m_{d}$ for all $d\geq 1$.
 Suppose $M$ is a positive real number.
@@
   1. (C1)
-    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq M^{d}$.
+    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a(N)_{i}|^{d}\leq M^{d}$.
   2. (C2)
-    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq d!M^{d}$.
+    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a(N)_{i}|^{d}\leq d!M^{d}$.
 Assume that $\theta,\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$. Suppose $r\geq 1$ and define $\Omega_{r}\triangleq\{z\in\mathbb{C}:\sum_{i=1}^{r}\left\lVert z_{i}\right\rVert_{2}<\frac{1}{M}\}$.
@@
 Proof.
-(A): By Lemma˜10.6, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
+(A): By Lemma˜10.5, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
 Suppose $k\geq 1$ and $r\in\Gamma[k]$. We compute the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ as $N\rightarrow\infty$. First, note that for $s\in\Gamma$,
@@
 Thus, the $N\rightarrow\infty$ limit of the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ is the coefficient of $r(x)$ in $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$.
-Afterwards, by part (B) of Lemma˜10.8, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
+Afterwards, by part (B) of Lemma˜10.7, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
 (B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and Corollary˜9.14 for (C).
@@
 In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system.
-The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.14.
+The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.13.
 First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
@@
 Theorem 10.11.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$.
+Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$.
 Assume that $\theta\in\mathbb{C}$ is fixed.
@@
   2. (b)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
   3. (c)
@@
   4. (d)
-    For all $\nu\in\Gamma_{\text{even}}$
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{1}{\prod_{j=1}^{N}(1+2(j-1)\theta)}[e,F_{N}]_{D^{N}(\theta)}=d_{0}+1$
+    and for all $\nu\in\Gamma_{\text{even}}$
     $\lim_{N\rightarrow\infty}\frac{1}{\prod_{j=1}^{N}(1+2(j-1)\theta)N^{|\nu|}}[ep_{\nu},F_{N}]_{D^{N}(\theta)}$
@@
 See Theorems˜8.10 and 8.16. ∎
-Theorem˜10.13 is an example of an application of the results of this paper.
+Theorem˜10.12 is an example of an application of the results of this paper.
 Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity].
 Part (C) has not appeared previously, although it is straightforward to deduce.
-See Remark˜10.14 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
+See Remark˜10.13 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
 Theorem 10.12.
@@
 $=\prod_{i=1}^{r}J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)^{\frac{1}{r}}.$
-Afterwards, using Lemma˜10.7 gives that
+Afterwards, using Lemma˜10.6 gives that
 $J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)$ $\leq 1+\sum_{k=1}^{\infty}\frac{|c_{k}(a)|}{(\theta N)^{k}}|r\text{Re}(x_{i})|^{k}$
@@
 Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity].
-We can then apply part (B) of Lemma˜10.8 to conclude the proof.
-(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.7 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.7 and Corollary˜9.11. ∎
+We can then apply part (B) of Lemma˜10.7 to conclude the proof.
+(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.6 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.6 and Corollary˜9.11. ∎
 Remark 10.13.
 From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered.
-Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.13 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$.
+Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.12 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$.
 This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
@@
 This implication as well as other essential implications are included in the following lemma.
-Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (E) is [matrix]*Proposition 2.11).
+Lemma 10.15 ((A) is [limits_prob_measures]*Lemma 1.4 and (D) is [matrix]*Proposition 2.11).
 Suppose the Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$.
@@
   1. (A)
-    The Bessel generating function $G_{\mu}^{\mathcal{R}(\theta)}(x)$ converges and is holomorphic over the closed ball of radius $R$ centered at the origin.
+    The Bessel generating function $G_{\mu}^{\mathcal{R}(\theta)}(x)$ converges over the closed ball of radius $R$ centered at the origin and is holomorphic over the interior of this domain.
   2. (B)
-    There exist unique constants $c_{\lambda}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
-    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}^{\mathcal{R}(\theta)}(\mu)p_{\lambda}\right)$
+    Let $\{1\}\cup\mathcal{B}$ be a basis of $\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$.
+There exist unique constants $c_{b}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $b\in\mathcal{B}$ such that
+    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
     in a neighborhood of the origin.
   3. (C)
-    There exist unique constants $m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ such that
-    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)M_{\epsilon}\right)$
-    in a neighborhood of the origin.
-  4. (D)
     For all $\lambda\in\Gamma$, the function $\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and equals
@@
     over the open ball of radius $R$ centered at the origin.
-  5. (E)
+  4. (D)
     For all $\lambda\in\Gamma$,
@@
 Proof.
-For (B) and (C), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
-(D): We include a detailed proof of this statement for completeness.
+For (B), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
+(C): We include a detailed proof of this statement for completeness.
 We prove this result for the nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
@@
 converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$.
-Furthermore, we can show that the integral is holomorphic over $\overline{B}_{R}(0)$ by following [matrix]*Proof of Lemma 2.9.
+Furthermore, we can show that the integral is holomorphic over $B_{R}(0)$ by following [matrix]*Proof of Lemma 2.9.
 Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
@@
 which completes the induction.
-(E): This is straightforward to deduce from (D). ∎
+(D): This is straightforward to deduce from (C). ∎
 Remark 10.16.
 The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system.
-However, (A) and (E) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
-Furthermore, observe that (A) implies (B) and (C).
-However, we require the exponentially decaying condition to prove (D) and (E).
-Using Lemma˜10.16 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system.
+However, (A) and (D) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
+Furthermore, observe that (A) implies (B).
+However, we require the exponentially decaying condition to prove (C) and (D).
+Using Lemma˜10.15 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system.
 It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems.
@@
 Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 1$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-Define $G_{\mu}^{A^{N-1}(\theta)}$, $c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$, and $m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in Lemma˜10.16.
+For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{N}\}$, define the coefficients $c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in part (B) of Lemma˜10.15.
 Then, the following are equivalent.
   1. (a)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   2. (a’)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  3. (a”)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   3. (b)
@@
     $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
-Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
+Condition (b) of Corollary˜10.17 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
 When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers.
+Next, we state the analogous result for the $BC^{N}$ root system.
+Corollary 10.18.
+Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$.
+Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ have nonnegative real part for all $N\geq 1$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
+For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$, define the coefficients $c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.15.
+Then, the following are equivalent.
+  1. (a)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  2. (a’)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  3. (b)
+    For all $\nu\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta_{0}N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
+For the type D analogue of Corollary˜10.17, we adjust the notation used to match the notation of part (C) of Theorem˜1.1.
+Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.15.
+Lemma 10.19.
+Suppose the Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$. Furthermore, assume that $\theta\in\mathbb{C}$ has nonnegative real part.
+Let $\{1\}\cup\{\mathcal{B}\}$ be a basis of $\mathbb{C}^{H(BC^{N})}[x_{1},\ldots,x_{N}]$.
+There exist unique constants $d_{0}^{D^{N}(\theta)}(\mu)\\$ $\in\mathbb{C}$ and $c_{b}^{D^{N}(\theta)}(\mu),d_{b}^{D^{N}(\theta)}(\mu)\in\mathbb{C}$ for $b\in\mathcal{B}$ such that
+$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
+over a neighborhood of the origin.
+Proof.
+We have that $\frac{1}{2}(G_{\mu}^{D^{N}(\theta)}(x)+G_{\mu}^{D^{N}(\theta)}(-x))$ is even and evaluates to one at the origin, from which the uniqueness of $c_{b}^{D^{N}(\theta)}(\mu)$ for $b\in\mathcal{B}$ follows.
+Furthermore, we have that $\frac{1}{2}(G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x))$ is odd and $\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}$ is even.
+It is clear that $d_{0}^{D^{N}(\theta)}(\mu)=\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}-1$.
+Next, we have that $\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}-d_{0}^{D^{N}(\theta)}(\mu)$ evaluates to one at the origin, from which the uniqueness of $d_{b}^{D^{N}(\theta)}(\mu)$ for $b\in\mathcal{B}$ follows.
+∎
+Corollary 10.20.
+Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$.
+Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 1$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
+Define the coefficient $d_{0}^{D^{N}(\theta)}(\mu)$, the coefficients $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, and the coefficients $c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N}),d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$ as they are defined in Lemma˜10.19.
+Consider the following statements.
+  1. (a)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}\\ =0$ if $\ell(\lambda)\geq 2$.
+  2. (a’)
+    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  3. (b)
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  4. (b’)
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+  5. (c)
+    For all $\nu\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
+  6. (d)
+    It is the case that
+    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[e(a)\right]}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
+    and for all $\nu\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[e(a)p_{\nu}\left(\frac{a}{\theta N}\right)\right]}{N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
 Furthermore, we can generalize Corollary˜1.7 and Corollary˜1.8 to when $a(N)$ and $b(N)$ are randomly distributed. We first state the following generalization of Lemma˜10.1.
@@
 Corollary 10.22.
-Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative.
   1. (A)
-    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are exponentially decaying Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ such that:
+[…]
     1. 1.
-      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{R}$ in terms of moments.
+      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
     2. 2.
-      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{R}$ in terms of moments.
+      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
     For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
@@
   2. (B)
+    Suppose $\theta\in\mathbb{C}$ is positive.
 Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
-For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
+For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
 Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
@@
 (A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution.
-Instead of using Theorem˜1.1, we use Corollary˜10.18.
-(B): We use Lemma˜10.19 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying.
+Instead of using Theorem˜1.1, we use Corollary˜10.17.
+(B): We use Lemma˜10.21 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying.
 Afterwards, we follow the proof of (A) for the remaining steps of the proof.
 ∎
-We can also state the analogues of Corollary˜10.20 for the type BC and D root systems, which are similar to Corollary˜1.8 and part (B) of Corollary˜1.7, respectively.
-To determine what the differences are, it suffices to compare Corollary˜10.20 with part (A) of Corollary˜1.7.
+In the following two corollaries, we use notation defined in Corollary˜1.8.
+Corollary 10.23.
+[…]
+  1. (A)
+    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are exponentially decaying Borel probability measures over $\mathbb{C}^{N}$ for $N\geq 1$ such that:
+    1. 1.
+      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
+    2. 2.
+      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
+    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
+    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
+  2. (B)
+    Suppose $\theta_{0}$ and $\theta_{1}$ are nonnegative.
+Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
+    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
+Remark 10.24.
+Note that in part (A) of the previous corollary, the rectangular free convolution is not defined in [rectangular_free_convolution] when $c$ is complex.
+In this case, we define it using part (B) of Corollary˜10.2.
+Corollary 10.25.
+  1. (A)
+[…]
+    1. 1.
+      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
+    2. 2.
+      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
+    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin.
+Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
+  2. (B)
+    Suppose $\theta\in\mathbb{C}$ is positive.
+Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin.
+Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
 Example 10.26.
@@
 For $t\geq 0$, let $\mu_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
-(36) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=B_{a}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
+(36) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
 for all $x\in\mathbb{C}^{N}$.
@@
 Lemma 10.27.
-Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then,
-$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
-over $\overline{B}_{R}(0)$ for all $t\geq 0$.
+Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$.
+Then, for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$ and
+$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
+over $B_{R}(0)$.
 Proof.
-See Lemma˜10.16 and (36). ∎
+First, we show that for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
+Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $t$ conditioned on the initial value $a$ is at most
+$\exp\left(-\frac{(\left\lVert a\right\rVert_{2}-\left\lVert y\right\rVert_{2})^{2}}{2t}\right)y^{O(1)},$
+where we have applied the upper bound $B_{a}\left(\frac{y}{\alpha N}\right)\leq\sqrt{N!}\exp\left(\frac{\left\lVert a\right\rVert_{2}\left\lVert y\right\rVert_{2}}{\alpha N}\right)$ that we deduce from part (e) of Lemma˜2.6.
+Then, we have that for all $\epsilon>0$,
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
+and
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
+This implies that $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$.
+For the computation of the Bessel generating function $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}$, we have that
+$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}(\eta)(a).$
+for $x\in\mathbb{C}^{N}$. Because $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
+$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}(a)(a(t))\right)d\eta(a).$
+Afterwards, applying (36) and part (A) of Lemma˜10.15 implies that this expression equals
+$\int_{\mathbb{R}^{N}}J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)d\eta(a)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x).$
 Furthermore, for $\theta>0$, define Hermite $\theta$-ensemble to be the random variable $H_{\theta}$ over $\mathbb{R}^{N}$ with density at $a$ proportional to $\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{2\theta}\prod_{i=1}^{N}e^{-\frac{\theta a_{i}^{2}}{2}}$.
@@
 over $\mathbb{C}^{N}$.
-Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.22 implies that
+Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.27 implies that
 (37) $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
@@
 Assume that $\alpha>0$.
-We can follow the method of the proof of Corollary˜10.20.
-First, since condition (b) of Corollary˜10.18 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well.
+We can follow the method of the proof of Corollary˜10.22.
+First, since condition (b) of Corollary˜10.17 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well.
 The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$.
-This does not follow from (37) and Lemma˜10.19 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
-Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $\alpha N$ conditioned on the initial value $a$ is at most
-$\exp\left(-\frac{(\left\lVert a\right\rVert_{2}-\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}\right)y^{O(1)},$
-where we have applied the upper bound $B_{a}\left(\frac{y}{\alpha N}\right)\leq\sqrt{N!}\exp\left(\frac{\left\lVert a\right\rVert_{2}\left\lVert y\right\rVert_{2}}{\alpha N}\right)$ that we deduce from part (e) of Lemma˜2.6.
-Assume that $\eta_{N}$ exponentially decays at rate $R$. Then, we have that for all $\epsilon>0$,
-$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
-$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
-and
-$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
-$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
-This implies that $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$.
-Then, because
-$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=G_{\eta_{N}}^{A^{N-1}(\theta)}(x)\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
-over $\overline{B}_{R}(0)$ by Lemma˜10.22, applying Corollary˜10.18 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$.
+This does not follow from (37) and Lemma˜10.21 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
+Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.27, $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$. Then, because
+$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$
+over $B_{R}(0)$ by Lemma˜10.27, applying Corollary˜10.17 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$.
 Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$.

### v5

@@
 Abstract.
-For the type A, BC, and D root systems, we determine equivalent conditions between the coefficients of an exponential holomorphic function and the asymptotic values taken by the Dunkl bilinear form when one of its entries is the function.
+We determine equivalent conditions between the asymptotic coefficients of the Bessel generating functions of a sequence of probability measures and the asymptotic expected values of power sums when their inputs are sampled from these measures.
 We establish these conditions over the $|\theta N|\rightarrow\infty$ regime for the type A and D root systems and over the $|\theta_{0}N|\rightarrow\infty,\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ regime for the type BC root system.
-We also generalize existing equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type A root system and over the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ regime for the type BC root system and prove new equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type D root system.
+We also establish equivalent conditions over the $\theta N\rightarrow c\in\mathbb{C}$ regime for the type A and D root systems and over the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ regime for the type BC root system that generalize existing results.
 Furthermore, we determine the asymptotics of the coefficients of the Bessel functions over the regimes that we have mentioned.
@@
 For positive integers $N\geq 2$, we study the asymptotics of the Bessel functions associated to the irreducible root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$.
-We denote the Bessel function associated to a finite root system $\mathcal{R}$ and multiplicity function $\theta$ as $J^{\mathcal{R}(\theta)}_{a}(x)$, where $a,x\in\mathbb{C}^{N}$, provided that the function exists.
+We denote the Bessel function associated to a finite root system $\mathcal{R}$ and multiplicity function $\theta$ by $J^{\mathcal{R}(\theta)}_{a}(x)$, where $a,x\in\mathbb{C}^{N}$, provided that the function exists.
 It is an eigenfunction of the Dunkl operators associated to $\mathcal{R}$ and $\theta$, which are introduced in [dunkloperators].
 Furthermore, the function is symmetric with respect to the reflection group generated by $\mathcal{R}$ in the variables $a$ and $x$.
-For the definitions of root systems, multiplicity functions, and the Bessel functions, see Subsection˜2.1.
 For certain choices of $\theta$, the paper [dunklbessel] shows that $J^{\mathcal{R}(\theta)}_{a}(x)$ exists, is unique, and is holomorphic over $a,x\in\mathbb{C}^{N}$, see Theorem˜2.8.
-The paper also discusses a nonsymmetric eigenfunction, see Theorem˜2.5; the formula relating the two eigenfunctions is given in Theorem˜2.8.
+The paper also discusses a nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$, see Theorem˜2.5; we have that $J^{\mathcal{R}(\theta)}_{a}(x)$ is the average of $E_{ga}^{\mathcal{R}(\theta)}(x)$ when $g$ is selected uniformly at random from the finite reflection group generated by $\mathcal{R}$.
 In this paper, we analyze the coefficients of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is fixed and $\mathcal{R}$ is one of $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$.
 The multiplicity function $\theta$ varies with $N$ and should be viewed as a function of $N$, although we denote it as $\theta$ rather than $\theta(N)$ for brevity.
-We outline the notation we use to record the root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$ and the corresponding multiplicity functions.
+To compute the coefficients of $J_{a}^{\mathcal{R}(\theta)}$, it is equivalent to compute the values of
+$f(\mathcal{D}_{1}(\mathcal{R}(\theta)),\ldots,\mathcal{D}_{N}(\mathcal{R}(\theta)))g(x_{1},\ldots,x_{N})$
+for all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are homogeneous, have the same degree, and symmetric with respect to the reflection group generated by $\mathcal{R}$; for an explanation of why this is the case, see Lemma˜4.19.
+We compute the asymptotics of these values in Theorem˜1.2.
+1.1. Notation
+We introduce some notation required for stating the main result. For additional definitions and notation, see Section˜2.
+For $N\geq 1$, $\Gamma_{N}$ consists of partitions with all parts of size at most $N$, $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}$, and $\Gamma_{\text{even}}$ and $\Gamma_{\text{even};\,N}$ are the elements of $\Gamma$ and $\Gamma_{N}$, respectively, with all parts of even size; for $\lambda\in\Gamma$, $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$.
+Additionally, $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$.
+Furthermore, for $k\geq 1$, $NC(k)$ denotes the set of noncrossing partitions of $[k]$ and $NC^{\text{even}}(2k)$ denotes the set of noncrossing partitions of $[2k]$ with all even block sizes.
+We outline the notation we use to denote the root systems $A^{N-1}$, $B^{N}$, $C^{N}$, and $D^{N}$ and the corresponding multiplicity functions.
 Note that $e_{i}\triangleq[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}$ for $i\in[N]$.
@@
 for $i\in[N]$, where $s_{ij}$ switches the $i$th and $j$th entries of an element of $\mathbb{C}^{N}$ for distinct $i,j\in[N]$ and $\tau_{i}$ flips the sign of the $i$th entry of an element of $\mathbb{C}^{N}$ for $i\in[N]$.
-To compute the coefficients of $J_{a}^{\mathcal{R}(\theta)}$, it is equivalent to compute the values of
-$f(\mathcal{D}_{1}(\mathcal{R}(\theta)),\ldots,\mathcal{D}_{N}(\mathcal{R}(\theta)))g(x_{1},\ldots,x_{N})$
-for all $f,g\in\mathbb{C}[x_{1},\ldots,x_{N}]$ that are homogeneous, have the same degree, and symmetric with respect to the reflection group generated by $\mathcal{R}$; for the explanation of why this is the case, see Lemma˜4.19.
-We compute the asymptotics of these values in Theorem˜1.1.
+1.2. Bessel generating functions and related works
+In addition to considering the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ for a single value of $a$, we can consider the average of $J_{a}^{\mathcal{R}(\theta)}(x)$ when $a$ is randomly sampled.
+The resulting function is referred to as the Bessel generating function and has been studied previously in [gaussianfluctuation, matrix, airy_beta, limits_prob_measures, rectangularmatrix].
+This notion is also related to the Dunkl transform introduced in [dunklbound].
+We discuss results related to Bessel generating functions in Subsection˜10.5.
+Definition 1.1.
+Suppose $\mu$ is a Borel probability measure over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
+$G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\mathbb{E}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)],$
+provided that the expected value converges.
+For $k\geq 1$, the type A Bessel function satisfies
+$\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}J_{a}^{A^{N-1}(\theta)}(x)=\sum_{i=1}^{N}a_{i}^{k}J_{a}^{A^{N-1}(\theta)}(x).$
+Therefore, after assuming regularity conditions on $\mu$, we have that
+$1]\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}G_{\mu}^{A^{N-1}(\theta)}(x)=\mathbb{E}_{a\sim\mu}\left[\sum_{i=1}^{N}a_{i}^{k}\right],$
+see Lemma˜10.16 for more details.
+As $N\rightarrow\infty$, in order to analyze the moments of $\mu$, it suffices to analyze $1]\sum_{i=1}^{N}\mathcal{D}_{i}(A^{N-1}(\theta))^{k}G_{\mu}^{\mathcal{R}(\theta)}(x)$.
+For the type A root system, the paper [gaussianfluctuation] addresses this question for $\theta=1$ and [matrix] does so for the $\theta N\rightarrow c\geq 0$ regime.
+For the type BC root system, [rectangularmatrix] does so for the $\theta_{0}N\rightarrow c_{0}\geq 0$ and $\theta_{1}\rightarrow c_{1}\geq 0$ regime.
+We address this question for both regimes and the type A, BC, and D root systems.
+The paper [gaussianfluctuation] defines the Bessel generating function to be the average of $\frac{J_{a}^{\mathcal{R}(\theta)}(x)}{J_{a}^{\mathcal{R}(\theta)}(\rho)}$ when $a$ is sampled from a probability ensemble; we consider when $\rho=0$.
+In order to determine the moments of the probability ensemble, the paper computes the value of the Bessel generating function after applying Dunkl operators when $x=\rho$ and as $N\rightarrow\infty$, $\rho$ converges to a fixed distribution after rescaling.
+For the $|\theta N|\rightarrow\infty$ regime, the paper [limits_prob_measures] introduces a framework for analyzing the asymptotic moments when the coefficients of the logarithm of the type A Bessel generating functions with two or more variables can have nonzero $N\rightarrow\infty$ limits after scaling by $(\theta N)^{-1}$.
+The paper [domino_tiling] studies Schur generating functions and considers the same setting.
+The converse statement is that we can determine the asymptotics of the coefficients of $G_{\mu}^{\mathcal{R}(\theta)}$ as $N\rightarrow\infty$ from the values of $p_{\nu}(\mu)$ for $\nu\in\Gamma$.
+The paper [matrix] establishes this statement for the type A root system and the $\theta N\rightarrow c\geq 0$ regime and asks whether it is true for the $|\theta N|\rightarrow\infty$ regime.
+We answer this question and establish the converse statement for both regimes and the type A, BC, and D root systems.
+Furthermore, the papers [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov, domino_tiling] similarly analyze the coefficients of Jack and Schur generating functions to establish a law of large numbers.
+The papers [unitary, llnclt, jackgenfunc, rectangularmatrix] also prove converse statements and [fluctuations, unitary, llnclt, gaussianfluctuation, domino_tiling] consider the central limit theorem in addition to the law of large numbers.
 1.3. Main result
-For the notation regarding partitions and noncrossing partitions in the following result, see Subsections˜2.2 and 2.3, respectively.
-We briefly mention that for $N\geq 1$, $\Gamma_{N}$ consists of partitions with all parts of size at most $N$, $\Gamma\triangleq\bigcup_{N\geq 1}\Gamma_{N}$, and $\Gamma_{\text{even}}$ and $\Gamma_{\text{even};\,N}$ are the elements of $\Gamma$ and $\Gamma_{N}$, respectively, with all parts of even size; for $\lambda\in\Gamma$, $p_{\lambda}(x_{1},\ldots,x_{N})\triangleq\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$.
-Additionally, $e(x_{1},\ldots,x_{N})\triangleq x_{1}\cdots x_{N}$.
-Furthermore, for $k\geq 1$, $NC(k)$ denotes the set of of noncrossing partitions of $[k]$ and $NC^{\text{even}}(k)$ denotes the set of noncrossing partitions of $[k]$ with all even block sizes.
 We state the main result of this paper, which is over the regime $|\theta N|\rightarrow\infty$ for the $A^{N-1}$ and $D^{N}$ root systems and the regime $|\theta_{0}N|\rightarrow\infty$, $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c\in\mathbb{C}$ for the $BC^{N}$ root system.
@@
     $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
-Then, (c) implies (d), and if $c\not=-1$, then (d) implies (c).
+Then, (c) implies (d) and if $c\not=-1$, then (d) implies (c).
 (C): Suppose
@@
 Proof.
-See Corollaries˜6.11, 7.5 and 8.22 for the proofs of (A), (B), and (C), respectively. ∎
-First, we note that we actually prove the generalizations where the $N\rightarrow\infty$ limits of $\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$, $\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}$, and $\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$ can be nonzero when $\ell(\lambda)>1$, see Theorems˜6.10, 7.4 and 8.21.
+See Corollaries˜6.11, 7.5 and 8.10 for the proofs of (A), (B), and (C), respectively. ∎
+First, we note that we actually prove the generalizations where the $N\rightarrow\infty$ limits of $\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$, $\frac{c_{\lambda}(N)}{(\theta_{0}N)^{\ell(\lambda)}}$, and $\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}$ can be nonzero when $\ell(\lambda)>1$, see Theorems˜6.10, 7.4 and 8.9.
 We present this corollary because it is more applicable to the setting where the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are set as $J_{a(N)}^{A^{N-1}(\theta)}$, $J_{a(N)}^{BC^{N}(\theta_{0},\theta_{1})}$, and $J_{a(N)}^{D^{N}(\theta)}$ in some sufficiently small neighborhood of the origin where the Bessel functions are nonzero so that their logarithms are holomorphic over the neighborhood; $a(N)\in\mathbb{C}^{N}$ varies with $N$.
@@
 Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 2}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin.
-We prove a generalized version of this result in Corollary˜10.17.
-Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]; see Lemma˜1.2 for a more direct connection with the question that is posed.
-In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$.
-The main contribution of this paper is the set of equivalence relations in the theorem where some conditions are based on scaling the coefficients of the logarithms of the input functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ by varying powers of $\theta N$.
-The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed is mentioned in the appendix of [matrix].
-However, in [limits_prob_measures], the coefficients are only scaled by $(\theta N)^{-1}$ and it is similarly the case that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-As we mention in Section˜6, a method to prove part (A) of Theorem˜1.1 is to prove Theorem˜6.1 using the results of [limits_prob_measures].
+We prove a generalized version of this result in Corollary˜10.18.
+As we mentioned previously, Theorem˜1.2 resolves the question posed in the appendix of the paper [matrix].
+The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed that we use is mentioned in the paper.
+In particular, we generalize the results of [limits_prob_measures], where the coefficients are only scaled by $(\theta N)^{-1}$.
+As we mention in Section˜6, a method to prove part (A) of Theorem˜1.2 is to prove Theorem˜6.1 using the results of [limits_prob_measures].
 However, this method is not applicable to parts (B) and (C), so we develop a new approach which can be used in these settings in Subsection˜6.1.
 This approach is applicable to the $\theta N\rightarrow c\in\mathbb{C}$ regime as well.
-Interestingly, in part (C) of Theorem˜1.1, we also consider the asymptotics of coefficients of terms with degree increasing to infinity.
-See Subsection˜1.6, Section˜8, and Section˜11 for more discussion regarding this direction.
-We can also restate Theorem˜1.1 in terms of the Taylor series expansions over the standard basis $M_{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ and the coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$.
-See Subsection˜2.2 for the definitions of these notions.
-Lemma 1.2.
-Repeat the statement of Theorem˜1.1. Suppose
-$F_{N}^{A}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m^{A}_{\epsilon}(N)M_{\epsilon})\right)\in\mathbb{C}[[x_{1},\ldots,x_{N}]],$
-$F_{N}^{BC}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{BC}_{\epsilon}(N)M_{\epsilon}\right)\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]],$
-$F_{N}^{D}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{D}_{\epsilon}(N)M_{\epsilon}\right)+e\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}l^{D}_{\epsilon}(N)M_{\epsilon}\right)$
-$\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]].$
-Consider the following statements.
-  1. (a’)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  2. (c’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  3. (e’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  4. (f’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-Then, (a’), (c’), (e’), and (f’) are equivalent to (a), (c), (e), and (f), respectively.
-Remark 1.3.
-The partial derivative $\partial^{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ denotes $\prod_{i=1}^{\ell(\epsilon)}\partial_{i}^{\epsilon_{i}}$.
-We can clearly replace it with $\prod_{i=1}^{\ell(\epsilon)}\partial_{\pi(i)}^{\epsilon_{i}}$ for any injection $\pi:[\ell(\epsilon)]\rightarrow[N]$ since the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are symmetric.
-Furthermore, in (f”), $(1,\ldots,1)$ consists of $N$ copies of $1$ and $(\epsilon+(1,\ldots,1))!=\prod_{i=1}^{\ell(\epsilon)}(\epsilon_{i}+1)$.
-Proof of Lemma˜1.2.
-The equivalences are straightforward after using the definition of the coefficients $c_{\epsilon\lambda}$. ∎
-1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
-Observe that in Theorem˜1.1, we consider the following regimes:
+1.4. The high temperature regime
+Observe that in Theorem˜1.2, we consider the following regimes:
   1. 1.
@@
     The $N\rightarrow\infty$ limit of $\theta_{0}N$ is $c_{0}\in\mathbb{C}$ and the $N\rightarrow\infty$ limit of $\theta_{1}$ is $c_{1}\in\mathbb{C}$ for $BC^{N}(\theta_{0},\theta_{1})$.
-Note that [matrix] considers regime 3 for the root system $A^{N-1}$ while [rectangularmatrix] considers regime 4.
-The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well.
+The approach that we use to prove Theorem˜1.2 over regimes 1 and 2 is adaptable to proving similar results over regimes 3 and 4 as well.
 Theorems that we prove over regimes 3 and 4 include Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix].
-We also prove Theorem˜8.26 for regime 3 and the type D root system.
-1.3. Related works
-Some related works that we have already mentioned are [matrix, limits_prob_measures, rectangularmatrix].
-These works are based on analyzing the Bessel generating functions of exponentially decaying measures to compute the asymptotic moments of the measures.
-While we focus on a broader setting, we discuss the applications of this paper’s results to exponentially decaying measures in Subsection˜10.5.
-Another work that is based on analyzing Bessel generating functions to compute asymptotic moments is [gaussianfluctuation], which considers when $\theta=1$.
-Furthermore, the papers [representations, fluctuations, unitary, llnclt, qmk, jackgenfunc, perelomov-popov] consider similar results by analyzing the coefficients of Jack and Schur generating functions.
-The paper [limits_prob_measures] introduces a framework for analyzing the asymptotic moments when the coefficients of terms of the Taylor series of the type A Bessel generating functions with two or more variables can have nonzero $N\rightarrow\infty$ limits after scaling by $(\theta N)^{-1}$.
-In fact, we are able to apply the results of the paper to obtain a straightforward proof of part (A) of Theorem˜1.1, as we mentioned earlier.
-In addition, we present a new proof method that is applicable to other settings, such as the $BC^{N}$ and $D^{N}$ root systems and regimes 3 and 4 that we described in Subsection˜1.2.
-1.4. Weak convergence to the free convolution
-Note that the definitions of $\theta(\mathcal{R})$ and $H(\mathcal{R})$ are included in Subsection˜2.1.
-First, we discuss an integral representation of the nonsymmetric eigenfunction for the Dunkl operators. This integral representation assumes that the multiplicity function is nonnegative.
+We also prove Theorem˜8.11 for regime 3 and the type D root system.
+1.5. Proof method
+For the proof of Theorem˜1.2 for the type A root system, the main idea is that when $|\theta N|\rightarrow\infty$ and $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is applied to $p_{\lambda}$ for some $\lambda\in\Gamma$, only certain sequences of operators contribute to the leading order term.
+Initialize $S$ to be $\ell(\lambda)]$.
+In order to contribute to the leading order term, we must have that the first operator of $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is $\partial_{i}$ applied to $p_{(\lambda_{j})}$ for some $j\in S=[\ell(\lambda)]$, since $p_{\lambda}$ is symmetric.
+Afterwards, we remove $j$ from $S$ and replace $p_{\lambda}$ with $\lambda_{j}x_{i}^{\lambda_{j}-1}\frac{p_{\lambda}}{p_{(\lambda_{j})}}$.
+For subsequent operators, we either diminish the degree of $x_{i}$ by one and multiply by $\theta N$ or apply $\partial_{i}$ to $p_{(\lambda_{j})}$ for some $j\in S$ and remove $j$ from $S$.
+Afterwards, we can deduce the result by constructing a bijection with noncrossing partitions.
+Let $d_{i}$ denote the operator which diminishes the degree of $x_{i}$ by one.
+Then, to contribute to the leading order term, $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is equivalent to $(\partial_{i}+\theta Nd_{i})^{k-1}\partial_{i}$, where $\partial_{i}$ is only applied to $p_{(\lambda_{j})}$ for $j\in S$.
+[…]
+In the $\theta N\rightarrow c\in\mathbb{C}$ regime, we have that $\mathcal{D}_{i}(A^{N-1}(\theta))^{k}$ is equivalent to $(\partial_{i}+cd_{i})^{k-1}\partial_{i}$.
+In contrast, we do not impose the constraint that $\partial_{i}$ is only applied to $p_{(\lambda_{j})}$ for $j\in S$.
+Similarly, in the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$ and $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime, $\mathcal{D}_{i}(BC^{N}(\theta_{0},\theta_{1}))^{k}$ is equivalent to $\left(\partial_{i}+2c_{0}d_{i}+c_{1}\frac{1-\tau_{i}}{x_{i}}\right)^{k-1}\partial_{i}$.
+Afterwards, we similarly construct a bijection with noncrossing partitions.
+1.6. Probabilistic applications
+As examples of applications, we prove that an observation of the DBM converges to a fixed measure as $\theta N\rightarrow\infty$ and determine the measure; it is straightforward to prove an analogous result in the $\theta N\rightarrow c\geq 0$ regime.
+We can similarly establish the law of large numbers for the $\beta$-Hermite ensemble as $\theta N\rightarrow\infty$, see [limits_prob_measures]; this ensemble is equivalent to the DBM started at zero.
+For the $\beta$-Laguerre ensemble, we also discuss the law of large numbers as $\theta N\rightarrow\infty$ and reprove a result of [betalaguerre].
+Although we only consider the type A DBM in this paper, we can prove analogous results for the type BC and type D DBMs using the same method.
+For the definitions of these Brownian motions, see [demni_sde].
+The paper [airy_beta] also uses Dunkl operators to analyze the type A DBM. The paper shows that multiple observations of the largest particles of the DBM converge to the Airyβ line ensemble.
+1.7. Integral representations of the Bessel functions
+First, we introduce the intertwining measure, which is a representation of the Bessel function as an integral of exponential functions.
+For a finite root system $\mathcal{R}$, $H(\mathcal{R})$ denotes the finite reflection group generated by $r_{\alpha}$ for $\alpha\in\mathcal{R}$ and for $a\in\mathbb{R}^{N}$, $H(\mathcal{R})a\triangleq\{ha:h\in H(\mathcal{R})\}$.
+Moreover, $\theta(\mathcal{R})$ is the set of multiplicity functions over $\mathcal{R}$.
 Theorem 1.3 ([rosler_positivity]).
@@
 Suppose $a\in\mathbb{R}^{N}$.
-There exists a unique Borel probability measure $\mu_{a}$ whose support is contained in the convex hull of $H(\mathcal{R})a$ such that
-$E_{a}^{\mathcal{R}}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
-for all $x\in\mathbb{C}^{N}$. Furthermore, $\text{supp}(\mu_{a})\cap H(\mathcal{R})a$ is nonempty and the Borel probability measure
-$\mu_{a}^{\text{sym}}\triangleq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}\mu_{ha}$
+There exists a unique Borel probability measure $\nu^{\mathcal{R}(\theta)}_{a}$ whose support is contained in the convex hull of $H(\mathcal{R})a$ such that
+$E_{a}^{\mathcal{R}}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu^{\mathcal{R}(\theta)}_{a}(\epsilon)$
+for all $x\in\mathbb{C}^{N}$. Furthermore, $\text{supp}(\nu_{a}^{\mathcal{R}(\theta)})\cap H(\mathcal{R})a$ is nonempty and the Borel probability measure
+$\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}\triangleq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}\nu_{ha}^{\mathcal{R}(\theta)}$
 is invariant with respect to the action of $H(\mathcal{R})$ and satisfies
-$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)$
+$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}(\epsilon)$
 for all $x\in\mathbb{C}^{N}$.
@@
 The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8.
-In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.4.
-We state a well-known conjecture related to an integral representation of the product of two Bessel functions; it can easily be extended to the product of an arbitrary number of Bessel functions.
-The conjecture also assumes that the multiplicity function is nonnegative.
+In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.3.
+The intertwining measure $\nu_{a}^{\mathcal{R}(\theta)}$ is equivalent to the intertwining operator between the Dunkl operator associated to $\mathcal{R}$ and $\theta$ and $\partial$, the derivative operator which is equivalent to setting $\theta=0$.
+For the definition of an intertwining operator, see Definition˜3.7 and for the explanation of this equivalence, see Example˜5.18.
+We study intertwiners in a more general setting in Theorem˜4.5.
+In addition, we can analyze the product of two Bessel functions.
+We state a well-known conjecture about the existence of an integral representation of such products; it can easily be extended to the product of an arbitrary number of Bessel functions.
+The conjecture similarly assumes that the multiplicity function is nonnegative.
 Conjecture 1.5.
@@
 The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$.
-By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.6 is satisfied.
-It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
-The paper [rosler_product] shows that Conjecture˜1.6 holds in the context of radially symmetric Bessel functions.
-Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
-As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.6.
+For simplicity, we may assume that it is symmetric.
+By [trimeche], there exists a distribution $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.5 is satisfied.
+It remains to determine whether the symmetric version of this distribution is nonnegative.
+Conjecture˜1.5 is established for $A^{N-1}(\theta)$ when $\theta\in\{\frac{1}{2},1,2\}$ as well as for additional choices of the root system and multiplicity function, see the paper [rosler_product].
+The paper also shows that the conjecture holds in the context of radially symmetric Bessel functions.
+In addition, analogues of Conjecture˜1.5 have been established for the Schur functions, Hall-Littlewood polynomials, zonal polynomials, quaternionic zonal polynomials, and Schur’s $Q$-functions, see [macdonald].
+However, the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators is false, see [convolution_maximal].
+As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.5.
 For the proofs of the corollaries, see Subsection˜10.1.
+First, we define the notion of convergence in terms of power sums, which implies convergence in terms of moments.
+Definition 1.6.
+Suppose $\eta$ is a distribution over $\mathbb{R}$ with finite moments.
+For $N\geq 1$, suppose $\eta_{N}$ is a Borel probability measure over $\mathbb{R}^{N}$.
+Then, $\eta_{N}$ converges to $\eta$ in terms of power sums if for all $\nu\in\Gamma$,
+(1) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\mathbb{E}_{a\sim\eta_{N}}\left[p_{\nu}(a)\right]=\prod_{i=1}^{\ell(\eta)}\mathbb{E}_{a\sim\eta}[a^{\nu_{i}}].$
+Converging in terms of power sums is similar to requiring that $\eta_{N}\approx\otimes^{N}\eta$.
+Conversely, we have that if $\eta_{N}=\otimes^{N}\eta$ and $\eta$ has finite moments, then $\eta_{N}$ converges to $\eta$ in terms of power sums as $N\rightarrow\infty$.
+This is because for all $\nu\in\Gamma$,
+$\mathbb{E}_{a\sim\eta_{N}}\left[p_{\nu}(a)\right]=\frac{N!\prod_{i=1}^{\ell(\nu)}\mathbb{E}_{a\sim\mu}[a^{\nu_{i}}]}{(N-\ell(\nu))!}+O(N^{\ell(\nu)-1}),$
+since $\eta$ has finite moments.
+Moreover, observe that when $\eta_{N}$ is distributed as a point mass for all $N\geq 1$, convergence in terms of power sums is equivalent to convergence in terms of moments.
+Converging in terms of moments corresponds to requiring that (1) is satisfied when $\nu\in\Gamma$ has length one.
+Another weaker condition that is often studied is to require that (1) is satisfied when $\nu\in\Gamma$ has length one or consists of two equal parts.
+This would imply that $\frac{1}{N^{\ell(\nu)}}p_{\nu}(\eta_{N})$ converges to $\prod_{i=1}^{\ell(\nu)}\mathbb{E}_{a\sim\eta}[a^{\nu_{i}}]$ in probability for all $\nu\in\Gamma$.
+Furthermore, if $\eta$ is determined by its moments, then after applying [billingsley]*Theorem 30.2, convergence in terms of power sums implies that $\mathbb{E}_{a\sim\eta_{N}}\left[\frac{1}{N}\sum_{i=1}^{N}\delta_{(a_{i})}\right]$ converges to $\eta$ in distribution as $N\rightarrow\infty$.
+In particular, Carleman’s condition implies that $\eta$ is determined by its moments when it is compactly supported.
 Corollary 1.7.
 Assume that $\theta\geq 0$ for all $N\geq 2$ and $\lim_{N\rightarrow\infty}\theta N=\infty$.
-Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$.
-Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
-Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$ as defined in [nica_speicher_2006]*Definition 12.1.
+Let $\mu_{a}$ and $\mu_{b}$ be Borel probability measures over $\mathbb{R}$ with finite moments.
+Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 2$ such that $\delta_{a(N)}(x\mapsto\frac{x}{\theta N})\rightarrow\mu_{a}$ and $\delta_{b(N)}(x\mapsto\frac{x}{\theta N})\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
+Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$.
 Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$.
@@
   1. (A)
-    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
-Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
+    Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.5.
+Then, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}(x\mapsto\frac{x}{\theta N})$ converges to $\mu$ as $N\rightarrow\infty$ in terms of power sums.
   2. (B)
-    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
-Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in distribution and in terms of moments.
-Conjecture˜1.6 is known for $\theta\in\{\frac{1}{2},1,2\}$.
-By [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$.
-When $\theta\in\{\frac{1}{2},1,2\}$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively.
-Then, by [voiculescu], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
-For the type BC root system, we can deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
+    Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.5.
+Then, $\frac{1}{2}\mu_{a(N),\,b(N)}^{D^{N}(\theta)}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{D^{N}(\theta)}(x\mapsto-\frac{x}{\theta N})$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in terms of power sums.
+The free convolution is introduced in [voiculescu1] and is defined for pairs of measures with finite variance in [maassen].
+However, the results of this paper require $\mu_{a}$ and $\mu_{b}$ to have finite moments.
+For a discussion of the concepts of free probability, see [nica_speicher_2006].
+If we assume that there is at most one measure with free cumulants given by adding the free cumulants of $\mu_{a}$ and $\mu_{b}$, we have that this corollary proves the existence of the free convolution as a probability measure when $\mu_{a}$ and $\mu_{b}$ have finite moments.
+This follows from considering when $\theta\in\{\frac{1}{2},1,2\}$ so that Conjecture˜1.5 is known to be true.
+When $\theta\in\{\frac{1}{2},1,2\}$, by [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$.
+Moreover, when $\theta=\frac{1}{2}$, $1$, and $2$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively.
+Then, by [voiculescu2], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
+For the type BC root system, we deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
 Corollary 1.8.
 Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 2$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}\\ =c\in\mathbb{C}$.
-Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$.
-Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
+Let $\mu_{a}$ and $\mu_{b}$ be Borel probability measures over $\mathbb{R}$ with finite moments.
+Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 2$ such that $\delta_{a(N)}(x\mapsto\frac{x}{\theta_{0}N})\rightarrow\mu_{a}$ and $\delta_{b(N)}(x\mapsto\frac{x}{\theta_{0}N})\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
 Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$.
 Let $\mu$ be the rectangular free convolution with $\lambda$ set as $\frac{1}{1+c}$ of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ as defined in [rectangular_free_convolution]*Proposition-Definition 2.1.
-Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6.
-Then, $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),\,b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
+Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.5.
+Then, $\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to $\mu$ as $N\rightarrow\infty$ in terms of power sums.
 We also note that analogues of part (A) of Corollary˜1.7 and Corollary˜1.8 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix].
@@
 Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic.
-In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.26.
-We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
+In particular, we use this framework to analyze the Dyson Brownian motion (DBM) with random initial values in Example˜10.28.
+We establish that if the average of the scaled empirical distribution of the initial values converge in terms of power sums to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of power sums to a measure determined by $\eta$ and $\alpha$.
 1.8. Uniform convergence of the Bessel functions
@@
 Additionally, we require that $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer.
-When $|\theta N|\rightarrow\infty$ and $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N(\theta N)^{k}}$ exists for all $k\in\mathbb{N}$, we can no longer compute the asymptotic coefficients; for example, in part (A) of Theorem˜1.1, $c_{\lambda}(N)$ has order $(\theta N)^{\ell(\lambda)}$ for $\lambda\in\Gamma$.
+When $|\theta N|\rightarrow\infty$ and $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N(\theta N)^{k}}$ exists for all $k\in\mathbb{N}$, we can no longer compute the asymptotic coefficients; for example, in part (A) of Theorem˜1.2, $c_{\lambda}(N)$ has order $(\theta N)^{\ell(\lambda)}$ for $\lambda\in\Gamma$.
 In this case, we can still approximate the coefficients, although they will not converge.
@@
 In order to prove uniform convergence over compact subsets of an open and simply connected domain, we follow the argument of [rank_infinity] and first prove that the Bessel functions are uniformly bounded and then apply Montel’s theorem, see Subsections˜10.3 and 10.4.
-To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.4.
+To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.3.
 An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
-1.6. Coefficients of terms with all odd degrees in the type D Bessel function
-The type D Bessel function is a symmetric linear combination of terms with all even and all odd degrees.
-To analyze the coefficients of the terms with all odd degrees of these functions, we compute the asymptotics of
-(1) $\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}\prod_{j=1}^{N}x_{j}\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$
-for $\nu,\lambda\in\Gamma_{\text{even}}$.
-These quantities are mentioned in part (h) of Theorem˜1.1 and are interesting because they are no longer polynomials in $N$ and $\theta$; however, they can be expressed as the product of a polynomial and an expression involving gamma functions, see Theorem˜8.16.
-In fact, we can consider analogous results for $A^{N-1}$ and $BC^{N}$, although terms with all odd degrees are not particularly important for analyzing the Bessel functions for these root systems so the results are not used while computing their coefficients, see Theorems˜12.1 and 12.3.
-Perhaps more interesting are the asymptotics of (1) in the regime $\theta N\rightarrow c\in\mathbb{C}$ for $A^{N-1}$ and $D^{N}$ and the regime $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ for $BC^{N}$.
-For the corresponding leading order terms for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, see Theorems 12.4, 12.5, and 8.23, respectively.
-Of course, we are most interested in the $D^{N}$ case since that is when summands which are multiples of $\prod_{j=1}^{N}x_{j}$ are present in the Bessel function.
-However, it is interesting that we can apply the methods that we develop for the $D^{N}$ case to the $A^{N-1}$ and $BC^{N}$ cases.
 1.9. Applying a graded ring of operators to a graded vector field
@@
 We extend this notion by considering a graded ring of operators.
-The results that we obtain are relevant to Dunkl operators and in particular the Dunkl bilinear form, see Example˜5.19; for the definition of the Dunkl bilinear form, see Subsection˜2.1.
-We also study the notion of invertible operators, which is the focus of [dunkl_singular_poly].
-See Section˜4 for the definition of invertibility.
-This framework and especially the content of Section˜4 are useful for analyzing the Bessel functions.
+The results that we obtain are relevant to Dunkl operators and in particular the Dunkl bilinear form, see Example˜5.18; for the definition of the Dunkl bilinear form, see Subsection˜2.1.
+We also study the notion of invertible operators, which is the focus of [dunkl_singular_poly]; see Definition˜4.1 for the definition of invertibility.
+In particular, we prove Theorem˜4.5, which establishes many equivalent conditions for the invertibility of an operator.
+By the equivalence of conditions (a) and (i) of the theorem, for a fixed multiplicity function, the Bessel function exists if and only if the Dunkl operator is invertible.
+Due to this, the content of Section˜4 is useful for analyzing the Bessel functions.
+In particular, in Section˜9, we obtain the asymptotic coefficients of the Bessel functions by computing the inverses of the matrices that store the values of the Dunkl bilinear form.
 1.10. Equivalence conditions for non-exponential holomorphic functions
@@
   • In Theorems˜7.4 and 7.8, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}$ and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$.
-  • In Theorems˜8.21 and 8.26, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
+  • In Theorems˜8.9 and 8.11, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
 1.11. Paper organization
@@
 Afterwards, in Section˜5, we discuss a representation of invertible graded rings of operators as sequences of invertible matrices and connect the framework introduced in Section˜3 to the Dunkl operators.
-In Sections 6, 7, and 8, we discuss the leading order terms of the Dunkl bilinear form and prove Theorem˜1.1 for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, respectively.
+In Sections 6, 7, and 8, we discuss the leading order terms of the Dunkl bilinear form and prove Theorem˜1.2 for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, respectively.
 Following this, in Section˜9, we determine the asymptotics of the coefficients of the terms of the Bessel functions that are homogeneous with a fixed degree.
-In Section˜10, we discuss applications of the results of this paper and in Section˜11, we present combinatorial expressions for the Dunkl bilinear form.
-In Section˜12, we analyze a special case of the leading order terms of the Dunkl bilinear form for the $A^{N-1}$ and $BC^{N}$ root systems.
+In Section˜10, we discuss applications of the results of this paper.
+Following this, in Section˜11, we present combinatorial expressions for the Dunkl bilinear form.
 1.12. Acknowledgments
@@
 Furthermore, $E_{a}^{\mathcal{R}(\theta)}(x)$ is holomorphic over the domain $\mathbb{C}^{N}\times\mathbb{C}^{N}\times\Theta(\mathcal{R})$ for $(a,x,\theta)$.
-The following result contains some properties about the eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
+The following lemma contains some properties about the eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
+See Lemma˜10.4 for an upper bound on the symmetric eigenfunction that is similar to part (e) of the lemma.
 Lemma 2.6 ([dunklbound]).
@@
 A $\mathbb{C}$-linear basis for $P_{2k}^{\text{even}}$ is given by $\{p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$ and a $\mathbb{C}$-linear basis for $P_{2k+N}^{\text{odd}}$ is given by $\{e(x_{1},\ldots,x_{N})p_{\lambda}(x_{1},\ldots,x_{N}):\lambda\in\Gamma_{\text{even};\,N}[2k]\}$.
-Additionally, define the unique coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$ such that
+Additionally, define the unique coefficients $c_{\epsilon\lambda}\in\mathbb{Q}$ for $\epsilon,\lambda\in\Gamma$ such that
 $M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}.$
+We define $c_{\lambda\epsilon}^{\text{inv}}$ for $\epsilon,\lambda\in\Gamma$ such that
+$p_{\lambda}=\sum_{\epsilon\in\Gamma}c_{\lambda\epsilon}^{\text{inv}}M_{\epsilon}.$
 Remark 2.13.
@@
 We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
+This is similarly the case when we write $p_{\lambda}=\sum_{\epsilon\in\Gamma}c_{\lambda\epsilon}^{\text{inv}}M_{\epsilon}$.
+It is clear that $c_{\lambda\lambda}=\frac{1}{\pi(\lambda)}$ and $c_{\lambda\lambda}^{\text{inv}}=\pi(\lambda)$.
 2.3. Non-crossing partitions
@@
 Recall that a partition $B_{1}\sqcup\cdots\sqcup B_{m}$ of $[k]$ is noncrossing if there does not exist distinct $i,j\in[m]$, $a,c\in B_{i}$, and $b,d\in B_{j}$ such that $a<b<c<d$.
-Furthermore, for $k\geq 1$, define $NC^{\text{even}}(k)$ to be the set of $\pi\in NC(k)$ such that each block of $\pi$ has even size.
+Furthermore, for $k\geq 1$, define $NC^{\text{even}}(2k)$ to be the set of $\pi\in NC(2k)$ such that each block of $\pi$ has even size.
 Suppose $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(k)$ such that the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $1\leq q\leq m-1$.
@@
 Also, define $f(\pi)\triangleq|B_{1}|$ and $\gamma(\pi)\triangleq\gamma((|B_{1}|,\ldots,|B_{m}|))$.
-[…]
-Furthermore, define $NC^{\prime}(k-1)$ to be the set of $\pi$ such that $m\geq 2$ and the minimal element of $B_{2}$ is $2$.
+2.4. Additional notation
+For a measure $\mu$ over $\mathbb{R}^{N}$ and measurable $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$, we let $\mu(f)$ denote the pushforward of $\mu$ with respect to the map $f$.
+For example, for a point mass $\delta_{a}$ located at $a\in\mathbb{R}^{N}$, $\delta_{a}(f)$ denotes the point mass $\delta_{f(a)}$.
 3. An introduction to applying a graded ring of operators to a graded vector field
@@
 In this section, we introduce a similar setting that is motivated by the Dunkl bilinear form $\cdot,\cdot]_{\mathcal{R}(\theta)}$ and we continue to discuss this setting in Sections˜4 and 5.
-See Example˜5.19 for the application of the framework that we introduce to the context of Dunkl operators.
+See Example˜5.18 for the application of the framework that we introduce to the context of Dunkl operators.
 Let $K$ be a field.
@@
 The key contribution that we discuss in this section while proving Theorem˜4.5 is the analysis of the matrices $\{M^{i;L}\}_{i\geq 0}$.
-To the best of our knowledge, the specific method we use has not appeared previously.
-It allows for straightforward proofs that are applicable in a general setting.
+This method allows for straightforward proofs that are applicable in a general setting.
 Theorem 4.5.
@@
   2. (b)
-    For all $i\geq 0$, there does not exist nonzero $f\in R_{i}$ such that $L(g)f=0$ for all $g\in R_{i}$.
+    For all $i\geq 0$, there does not exist nonzero $f\in V_{i}$ such that $L(g)f=0$ for all $g\in R_{i}$.
   3. (c)
-    For all $i\geq 0$, there does not exist nonzero $g\in R_{i}$ such that $L(g)f=0$ for all $f\in R_{i}$.
+    For all $i\geq 0$, there does not exist nonzero $g\in R_{i}$ such that $L(g)f=0$ for all $f\in V_{i}$.
   4. (d)
-    The equation $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=0$ has no nonzero solutions $f\in R_{i}$ for all $i\geq 0$.
+    The equation $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=0$ has no nonzero solutions $f\in V_{i}$ for all $i\geq 0$.
   5. (e)
@@
 since $|H|$ is not divisible by $\text{char}(K)$. It follows that $L$ is not invertible, which is a contradiction. ∎
-It does not seem to be generally true that the invertibility of $L_{(R,V)}$ implies the invertibility of $L$.
-However, in the paper [dunkl_singular_poly], this is shown to be true for Dunkl operators.
+The paper [dunkl_singular_poly] shows that in the setting of Dunkl operators, the invertibility of $L_{(R,V)}$ implies the invertibility of $L$.
 We repeat the argument from Remark 4.2 of that paper in a more general setting to prove the following result.
@@
 $g^{\prime}=\prod_{h^{\prime}\in H}h^{\prime}g.$
-For all $h\in H$, $hg^{\prime}=\prod_{h^{\prime}\in H}hh^{\prime}g=g^{\prime}$ by the commutativity of $S$, so $g^{\prime}\in R_{M}$, where $M$ is the positive integer such that $g^{\prime}\in S_{M}$.
+For all $h\in H$, $hg^{\prime}=\prod_{h^{\prime}\in H}hh^{\prime}g=g^{\prime}$ by the commutativity of $S$, so $g^{\prime}\in R$.
 Assume that $h^{\prime}\in H$ and $h^{\prime}g=0$.
@@
 It follows that $g^{\prime}\not=0$ because $S$ is an integral domain.
+We can express $g^{\prime}$ as a unique linear combination of elements of $R_{j}$ for $j\geq 0$.
+Let $m$ be the smallest nonnegative integer $j$ such that the component of $R_{j}$ in this linear combination is nonzero.
+Then, $m\geq i\geq 1$ because $g\in S_{i}$ and $g^{\prime}\not=0$.
 Suppose $f\in V_{m}$.
-Then, it is evident that $L(g^{\prime})f=0$ after applying $L(h^{\prime}g)$ for some sequence of |H|1|H|-1 distinct $h^{\prime}\in H\backslash\{1\}$.
-Thus, $L_{(R,V)}$ is not invertible, which is a contradiction.
+Then, $L(g^{\prime})f\in K$.
+Suppose $f^{\prime}=\prod_{h^{\prime}\in H\backslash\{1\}}L(h^{\prime}g)f$.
+Then, $L(g)f^{\prime}\in K$, so if $f^{\prime\prime}$ is the projection of $f^{\prime}$ onto $W_{0}\oplus\cdots\oplus W_{i}$, we have that $L(g^{\prime})f=L(g)f^{\prime}=L(g)f^{\prime\prime}$.
+However, $L(g)f^{\prime\prime}=0$, which is a contradiction to the invertibility of $L_{(R,V)}$.
 ∎
-Remark 5.12.
-In the previous result, we do not assume that $hS_{i}\subset S_{i}$ for all $h\in H$ and $i\geq 1$.
-In particular, we do not necessarily require that $M=|H|i$, although $M\geq 1$ because $g\in S_{i}$ and $i\geq 1$.
 Corollary 5.12.
@@
 Suppose $N\geq 1$, $S=W=K[x_{1},\ldots,x_{N}]$, and that $S_{i}=W_{i}$ is the group of polynomials over $K$ that are homogeneous of degree $i$ for $i\geq 0$.
-From Lemma˜5.15, any $L\in\mathcal{T}$ is uniquely determined by $L(x_{i})$ for $1\leq i\leq N$.
+From Lemma˜5.14, any $L\in\mathcal{T}$ is uniquely determined by $L(x_{i})$ for $1\leq i\leq N$.
 Let $H$ be a finite group which acts on $K^{N}$ such that for all $h\in H$ and $f\in S_{1}$, $\{x\mapsto f(hx)\}\in S_{1}$.
@@
 Assume that $L\in\mathcal{H}$ is invertible and that $\text{char}(K)$ does not divide $|H|$.
-Then, all conclusions of Corollary˜5.17 are true (this includes the conclusions over both $(R,V)$ and $(S,W)$).
+Then, all conclusions of Corollary˜5.16 are true (this includes the conclusions over both $(R,V)$ and $(S,W)$).
 Proof.
@@
 We describe the framework that we have developed in the context of Dunkl operators and prove a few well-known results.
-Consider the setting of Example˜5.14, but set $K=\mathbb{C}$ and $H$ to be $H(\mathcal{R})$, where $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system.
+Consider the setting of Example˜5.13, but set $K=\mathbb{C}$ and $H$ to be $H(\mathcal{R})$, where $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system.
 Suppose $u_{i}\in\mathbb{R}^{N}$ for $1\leq i\leq N$.
-Then, we can define $L\in\mathcal{T}$ by $L(x_{i})\triangleq\mathcal{D}_{u_{i}}(\mathcal{R}(\theta))$ after using Lemma˜5.15 and the commutativity of the Dunkl operators.
+Then, we can define $L\in\mathcal{T}$ by $L(x_{i})\triangleq\mathcal{D}_{u_{i}}(\mathcal{R}(\theta))$ after using Lemma˜5.14 and the commutativity of the Dunkl operators.
 As a special case, we have that $\mathcal{D}(\mathcal{R}(\theta))\in\mathcal{T}$, where we recall that $\mathcal{D}(\mathcal{R}(\theta))(x_{i})=\mathcal{D}_{i}(\mathcal{R}(\theta))$ for $i\in[N]$.
@@
 Observe that $E_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}(\mathcal{R}(\theta))$ and $J_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}_{H}(\mathcal{R}(\theta))$ for $\Psi:r\mapsto r(a)$, where $a\in\mathbb{C}^{N}$ is fixed.
-Then, part (B) of Corollary˜5.17 implies that these functions are unique eigenvectors, which is included in the statements of Theorems˜2.5 and 2.8.
-We do not consider intertwining operators in the Dunkl setting in this paper.
-However, we mention that the intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$.
+Then, part (B) of Corollary˜5.16 implies that these functions are unique eigenvectors, which is included in the statements of Theorems˜2.5 and 2.8.
+The intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$.
 As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible.
-As explained in [rosler_positivity], in terms of Theorem˜1.4, the intertwiner is given by
-$Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\mu_{a}(\epsilon)$
-for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $a\in\mathbb{R}^{N}$.
+As explained in [rosler_positivity], in terms of Theorem˜1.3, the intertwiner is given by
+$Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\nu^{\mathcal{R}(\theta)}_{a}(\epsilon)$
+for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $a\in\mathbb{R}^{N}$. Since the Bessel function is
+$J_{a}^{\mathcal{R}(\theta)}(x)=\frac{1}{|H(\mathcal{R})|}V\sum_{g\in H(\mathcal{R})}e^{\left\langle gx,\cdot\right\rangle}(a),$
+the characteristic function of $\frac{1}{|H(\mathcal{R})|}\sum_{g\in H(\mathcal{R})}\nu_{ga}^{\mathcal{R}(\theta)}=\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}$ is given by $J_{a}^{\mathcal{R}(\theta)}(ix)$.
 For the remainder of this example, we consider the setting of Proposition˜4.22.
@@
 For $\theta\geq 0$ and even $i\geq 0$, it is straightforward to compute the decomposition of $M^{i;\,\mathcal{D}_{H}(D^{N}(\theta))}$ when we restrict the rows and columns to $p_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i]$ by using the decomposition of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,0))}$.
-An interesting question is, what is the decomposition when we restrict the rows and columns to $ep_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i-N]$, where $i\geq N$ has the same parity as $N$?
-Furthermore, another interesting question which generalizes the previous question is, what are the decompositions of $M^{i;\,\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$?
+By (20), the decomposition for the remaining entries is given by that of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,1))}$.
+Furthermore, another interesting question is, what are the decompositions of $M^{i;\,\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$?
 Recall that $M^{i;\,\mathcal{D}(D^{N}(\theta))}=M^{i;\,\mathcal{D}(BC^{N}(\theta,0))}$.
@@
 $=\theta^{k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}.$
-6.2. Proof of part (A) of Theorem˜1.1
-We are now prepared to prove part (A) of Theorem˜1.1.
+6.2. Proof of part (A) of Theorem˜1.2
 Define the infinite dimensional matrix $\mathcal{M}^{A}\in\mathbb{Z}_{\geq 0}^{\Gamma\times\Gamma}$ such that for $\lambda,\nu\in\Gamma$,
@@
     If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{A}_{\lambda\nu}=0$.
-The next theorem generalizes part (A) of Theorem˜1.1.
+The next theorem generalizes part (A) of Theorem˜1.2.
 As we mentioned previously, we focus on the less general version stated in Corollary˜6.11 because it is more relevant for analyzing the convergence of Bessel functions.
@@
 $c_{(|\epsilon|)}(N)=Nc_{\epsilon,(|\epsilon|)}\Rightarrow c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}.$
-6.3. Leading order terms of the type A Dunkl bilinear form in the $\theta N\rightarrow c\in\mathbb{C}$ regime
-Next, we compute the leading order terms of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma}$ in the regime $\theta N\rightarrow c\in\mathbb{C}$.
+6.3. Leading order terms of the type A Dunkl bilinear form in the high temperature regime
+Next, we compute the leading order terms of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}\\ )_{\lambda,\nu\in\Gamma}$ in the regime $\theta N\rightarrow c\in\mathbb{C}$.
 Suppose $k\geq 1$ and $\pi\in NC(k)$. Define the polynomial $W^{A}(\pi)\in\mathbb{Z}[x]$ by
@@
 This proof can be used while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems.
-Note that Theorem˜6.17 generalizes the results of the paper [matrix], since it allows $c_{\lambda}$ to be nonzero if $\ell(\lambda)\geq 2$.
-However, Corollary˜6.18 has already been proved in the paper.
 Proof of Theorem˜6.14.
@@
 Suppose $\nu^{\prime}\in\Gamma$. Let $c_{\nu^{\prime}}$ be the value of (15) when $\lambda$ is set as $(|\nu^{\prime}|)$ and $\nu$ is set as $\nu^{\prime}$ and define
-$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\,\gamma(\pi)=\nu^{\prime}\end{subarray}}W_{\pi}^{A}(x).$
+$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\,\gamma(\pi)=\nu^{\prime}\end{subarray}}W^{A}(\pi)(x).$
 Then, $c_{\nu^{\prime}}=\theta^{|\nu^{\prime}|-\ell(\nu^{\prime})}\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\pi(\nu^{\prime})n_{\nu^{\prime}}(\theta)$.
@@
 $N^{\ell(\lambda)}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
-$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W_{\pi}^{A}(N\theta)$
-$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}W_{\pi}^{A}(N\theta)$
-$=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}W_{\pi}^{A}(N\theta).$
+$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W^{A}(\pi)(N\theta)$
+$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}W^{A}(\pi)(N\theta)$
+$=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}W^{A}(\pi)(N\theta).$
 We now prove the analogue Theorem˜6.17 of Theorem˜6.10. The method we use is the same, however we include it for completeness.
@@
 ∎
-Corollary 6.18 ([matrix]*Theorems 3.8 and 3.10).
+Note that Theorem˜6.17 generalizes the results of the paper [matrix], since it allows $c_{\lambda}$ to be nonzero if $\ell(\lambda)\geq 2$.
+However, Corollary˜6.18 has already been proved in the paper.
+Corollary 6.18.
 Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\\ \left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$.
@@
 Note that we always have that $\frac{\theta_{1}}{\theta_{0}N}\rightarrow c_{1}\in\mathbb{C}$ and that the asymptotic invertibility of $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ will depend on $c_{0}$ and $c_{1}$.
-7.1. Proof of part (B) of Theorem˜1.1
+7.1. Proof of part (B) of Theorem˜1.2
 We first consider the $|\theta_{0}N|\rightarrow\infty$ regime.
@@
 See the proof of Corollary˜6.11. ∎
-7.2. Leading order terms of the type BC Dunkl bilinear form in the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime
+7.2. Leading order terms of the type BC Dunkl bilinear form in the high temperature regime
 Next, we consider the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime. Define the polynomial $W^{BC}(\pi)\in\mathbb{Z}[x,y]$ by
-$W^{BC}(\pi)(x,y)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi)).$
+$W^{BC}(\pi)(x,y)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(2x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}2y+d(i;\,\pi)).$
 Theorem 7.6.
 Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in\pi}x_{|B|}$
+$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(N\theta_{0},\theta_{1})\prod_{B\in\pi}x_{|B|}$
 $+R(N,\theta_{0},\theta_{1}),$
@@
     For all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(2c_{0},2c_{1})[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
+    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(c_{0},c_{1})[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
 Then, (a) implies (b), and if $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer, then (b) implies (a).
@@
 See the proof of Theorem˜6.17.
-For the implication of (a) from (b), consider the formula for $W^{BC}([l])(2c_{0},2c_{1})$ for some $l\in 2\mathbb{N}$.
+For the implication of (a) from (b), consider the formula for $W^{BC}([l])(c_{0},c_{1})$ for some $l\in 2\mathbb{N}$.
 If $d(i;\,\pi)$ is odd, then the term $2c_{0}+2c_{1}+d(i;\,\pi)$ is nonzero.
 Moreover, if $d(i;\,\pi)$ is even and at least two, then the term $2c_{0}+d(i;\,\pi)$ is nonzero; note that if $d(i;\,\pi)=0$, then we must have that $b(i;\,\pi)=1$.
-Therefore, $W^{BC}([l])(2c_{0},2c_{1})\not=0$, so $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]$ is invertible for all $k\geq 1$ by Lemma˜7.7.
+Therefore, $W^{BC}([l])(c_{0},c_{1})\not=0$, so $\mathcal{W}^{BC}(c_{0},c_{1})[k]$ is invertible for all $k\geq 1$ by Lemma˜7.7.
 ∎
-Corollary 7.9 ([rectangularmatrix]*Theorems 4.8 and 5.5).
+We similarly have that Theorem˜7.8 generalizes the results of the paper [rectangularmatrix] while Corollary˜7.9 has already been proved in the paper.
+Corollary 7.9.
 Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$.
@@
     For all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}.$
+    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(c_{0},c_{1})\prod_{B\in\pi}|B|c_{(|B|)}.$
 Then, (a) implies (b), and if $c_{0}$ is not a negative integer and $2c_{0}+2c_{1}$ is not a negative odd integer, then (b) implies (a).
@@
 Suppose $a\in\mathbb{C}^{N}$.
-The coefficients of terms of $J_{a}^{D^{N}(\theta)}(x_{1},\ldots,x_{N})$ with all odd degrees in $x_{1},\ldots,x_{N}$ are all zero if and only if $a_{1}\cdots a_{N}=0$.
+The coefficients of the terms of $J_{a}^{D^{N}(\theta)}(x_{1},\ldots,x_{N})$ with all odd degrees in $x_{1},\ldots,x_{N}$ are all zero if and only if $a_{1}\cdots a_{N}=0$.
 Proof.
@@
 ∎
-The following result shows that $e$ is a harmonic function with respect to $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ for all $\theta_{0},\theta_{1}\in\mathbb{C}$.
-Lemma 8.8.
-For all $\theta_{0},\theta_{1}\in\mathbb{C}$, $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))(p_{(2)})e=0$.
-Proof.
-First, observe that
-$\mathcal{D}_{1}e=(1+2\theta_{1})x_{2}\cdots x_{N}.$
-Afterwards, it is straightforward to deduce that $\mathcal{D}_{1}^{2}e=0$. ∎
-Corollary 8.9.
-For all $\theta\in\mathbb{C}$, $\mathcal{D}(D^{N}(\theta))(p_{(2)})e=0$.
-Proof.
-Set $\theta_{1}=0$ in Lemma˜8.8. ∎
-8.2.
-The leading order terms of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and an introduction to the leading order terms of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma_{\text{even}}$
-Next, we prove the results analogous to those of Section˜6 for the type D root system.
-However, we must now compute both $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma$.
-First, we consider the computation of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$.
-As we will see, these values are asymptotically equivalent to the values of $p_{\lambda},p_{\nu}]_{A^{N-1}(2\theta)}$ in both the $|\theta N|\rightarrow\infty$ and $\theta N\rightarrow c\in\mathbb{C}$ regimes.
-First, we consider the $|\theta N|\rightarrow\infty$ regime.
-Theorem 8.10.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$.
-Then, Theorem˜6.1 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $\theta$ replaced by $2\theta$ is true.
-Proof.
-We can use the same method as the proof of Theorem˜6.1.
-The reason for this is that because the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$.
-This is because after applying $\frac{1-\sigma_{j_{i}j}}{x_{j_{i}}-x_{j}}$ or $\frac{1-\tau_{j_{i}}\tau_{j}s_{j_{i}j}}{x_{j_{i}}+x_{j}}$ to $x_{j_{i}}^{d}$ for some $d\geq 1$, the only resulting term without $x_{j}$ is $\theta x_{j_{i}}^{d-1}$ in both cases.
-Hence, we only need to replace $\theta$ with $2\theta$ in Theorem˜6.1 to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$.
-∎
-Define the infinite dimensional matrix $\mathcal{M}^{D}\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$.
-Next, we consider the $\theta N\rightarrow c\in\mathbb{C}$ regime.
-Theorem 8.11.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
-Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-Then, statements (a) and (c) of Theorem˜8.21 are equivalent.
-Proof.
-This follows from Theorem˜8.10 and the proof of Theorem˜6.10 after $\Gamma$ and $\mathcal{M}^{A}$ are replaced with $\Gamma_{\text{even}}$ and $\mathcal{M}^{D}$, respectively. ∎
-Theorem 8.12.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$.
-Then, Theorem˜6.14 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $N\theta$ replaced by $2N\theta$ is true.
-Proof.
-Similarly to the proof of Theorem˜8.10, since the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$.
-Then, to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$, we only need to replace $N\theta$ with $2N\theta$ in Theorem˜6.14.
-∎
-Theorem 8.13.
-Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$.
-Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$.
-Then, in Theorem˜8.26, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
-Proof.
-This follows from Theorem˜8.12 and the proof of Theorem˜6.17 after $\Gamma$ and $c$ replaced with $\Gamma_{\text{even}}$ and $2c$, respectively. ∎
-Remark 8.14.
-Theorems 8.10, 8.11, 8.12, and 8.13 also follow from the analogous results for the type BC root system after setting $\theta_{1}=0$.
-Next, we consider the computation of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$, which is considerably more complicated.
-First, we exhibit a more direct computation in the case where $p_{\lambda},p_{\nu}=p_{(2)}^{k}$ and the multiplicity function is nonnegative by using Theorem˜8.1.
-Lemma 8.15.
-Suppose $k\geq 0$. Then, for all $\theta\in\mathbb{R}_{\geq 0}$,
-$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$
-Proof.
-Suppose $\theta\geq 0$. By Theorem˜8.1,
-$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})h^{2}wdx,$
-where $c_{N}$, $h$, and $w$ are defined with respect to the type $D$ root system with constant multiplicity $\theta$.
-By Lemma˜8.8, $e$ is harmonic. Then, using [dunkl_integralkernel]*Proposition 3.9 gives that
-$\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k}=(-1)^{k}k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)e,$
-where $\gamma=N(N-1)\theta$ and $L_{n}^{(\alpha)}$ denotes the generalized Laguerre polynomial. Hence,
-$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)\right)^{2}(eh)^{2}wdx.$
-For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$.
-Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B^{N}(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent.
-We similarly have that
-$\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}=(c^{\prime}_{N})^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{N}{2}+\gamma^{\prime}-1\right)}\left(\frac{p_{2}}{2}\right)\right)^{2}(h^{\prime})^{2}wdx$
-However, it is evident that $\gamma^{\prime}=\gamma+N$ and $(h^{\prime})^{2}=(eh)^{2}$, if we assume that the positive roots for the type B root system are $\{[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}:i\in[N]\}$ in addition to the positive roots for the type D root system.
-Then, we have that
-$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}$ $=\frac{c^{\prime}_{N}}{c_{N}}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}$
-$=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-1}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$
-We compute the leading order terms of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ in the following theorem.
-The leading order terms are obviously not polynomials, but can be expressed as a product involving gamma functions and polynomials.
-In particular, the product $\prod_{i=1}^{N-k}(1+2(i-1))\theta$ which appears in the following theorem can also be expressed as $(2\theta)^{N-k}\frac{\Gamma(\frac{1}{2\theta}+N-k)}{\Gamma(\frac{1}{2\theta})}$.
-Theorem 8.16.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=\prod_{i=1}^{N-k}(1+2(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(2N\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
-$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}+R(N,\theta)\bigg),$
-where $R\in\mathbb{Q}[x,y]$ satisfies:
-  1. (1)
-    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
-  2. (2)
-    The degree $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$.
-8.3. Proof of Theorem˜8.16
-We follow the framework of the proof of Theorem˜6.1. Observe that
-$ep_{\lambda},ep_{\nu}]=\mathcal{D}(p_{\lambda})\mathcal{D}(e)ep_{\nu}=\sum_{j_{1},\ldots,j_{\ell(\lambda)}=1}^{N}\prod_{i=1}^{\ell(\lambda)}(\mathcal{D}_{j_{i}}-\partial_{j_{i}}+\partial_{j_{i}})^{\lambda_{i}-1}\partial_{j_{i}}\mathcal{D}_{1}\cdots\mathcal{D}_{N}ep_{\nu}.$
-First, we consider case work on the operators in $\mathcal{D}(p_{\lambda})$.
-Assume that $\{j_{i}\}_{1\leq i\leq\ell(\lambda)}=[m_{1}]$.
-Furthermore, assume that the terms $\mathcal{D}_{j_{i}}-\partial_{j_{i}}$ have switches from $j_{i}$ to some $j\in[m_{2}]$ for some $m_{2}\geq m_{1}$, and that each element in $m_{2}]\backslash[m_{1}]$ appears in a switch.
-Moreover, we can replace the sets $m_{1}]$ and $m_{2}]\backslash[m_{1}]$ with any disjoint subsets of $[N]$ with size $m_{1}$ and $m_{2}-m_{1}$, respectively, and obtain the same result.
-Next, we consider the operators $\mathcal{D}_{j}$ in $\mathcal{D}_{1}\cdots\mathcal{D}_{N}$.
-First, observe that by symmetry, we assume that the switches in $\mathcal{D}_{j}$ are from $j$ to some element of $\{j+1,\ldots,N\}$ for $j\in[N]$.
-This is because prior to applying $\mathcal{D}_{j}$, we have applied $\mathcal{D}_{j+1},\ldots,\mathcal{D}_{N}$, so we are applying $\mathcal{D}_{j}$ to a polynomial which is invariant under switches between two distinct variables in $\{x_{1},\ldots,x_{j}\}$.
-To justify this, using Lemma˜2.2 gives that for $h\in H$,
-$h\mathcal{D}(x_{j+1}\cdots x_{N})ep_{\nu}=\mathcal{D}(h(x_{j+1}\cdots x_{N}))h(ep_{\nu})=\mathcal{D}(h(x_{j+1}\cdots x_{N}))ep_{\nu}.$
-If $h$ corresponds to a switch between two distinct variables in $\{x_{1},\ldots,x_{j}\}$, then $h(x_{j+1}\cdots\\ x_{N})=x_{j+1}\cdots x_{N}$, so we would have that $\mathcal{D}(x_{j+1}\cdots x_{N})ep_{\nu}$ is invariant under the action of $h$.
-Furthermore, each operator $\mathcal{D}_{j}$ either deletes the factor of $x_{j}$ in $e$ or applies $\partial_{j}$ to some $p_{\nu_{l}}$.
-Assume that there exists $j\in\{m_{2}+1,\ldots,N\}$ such that $\mathcal{D}_{j}$ applies $\partial_{j}$ to some $p_{\nu_{l}}$.
-Let $j_{\text{max}}$ be the largest such $j$.
-After applying $\mathcal{D}_{j_{\text{max}}}$, the exponent of $x_{j_{\text{max}}}$ will be positive.
-Then, we have that we will never be able to remove the factor of $x_{j_{\text{max}}}$.
-When applying $\mathcal{D}_{j}$ for some $j\in\{1,\ldots,j_{\text{max}}-1\}$, we will not be able to remove the factor with a switch since the exponent of $x_{j}$ is also positive.
-Furthermore, there are no switches to $j_{\text{max}}$ in the operators in $\mathcal{D}(p_{\lambda})$ by assumption.
-Then, for the contribution to be nonzero, we delete $x_{j}$ for each $j\in\{m_{2}+1,\ldots,N\}$, which results in multiplying by a factor of
-$\prod_{j=1}^{N-m_{2}}(1+2(j-1)\theta)=\prod_{j=N-k}^{N-m_{2}}(1+2(j-1)\theta)\prod_{j=1}^{N-k}(1+2(j-1)\theta);$
-note that $m_{2}\leq k$. The leading order term of this expression is
-(20) $(2\theta N)^{k-m_{2}}\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
-Afterwards, the expression we must evaluate is
-$\mathcal{D}(p_{\lambda})\mathcal{D}_{1}\cdots\mathcal{D}_{m_{2}}x_{1}\cdots x_{m_{2}}p_{\nu}.$
-Let $d_{1}$ be the number of derivatives in $\mathcal{D}(p_{\lambda})$, which does not include the $\ell(\lambda)$ derivatives at the starts of the blocks.
-Furthermore, let $d_{2}$ be the number of derivatives in $\mathcal{D}_{j}$ for $j\in[m_{1}]$ and let $d_{3}$ be the number of derivatives in $\mathcal{D}_{j}$ for $j\in[m_{2}]\backslash[m_{1}]$.
-The total number of derivatives is $d_{1}+d_{2}+d_{3}+\ell(\lambda)$; recall that this quantity must be at least $\ell(\nu)$ for the contribution to be nonzero.
-We have that the number of choices for the set $m_{2}]$ is $O(N^{m_{2}})$; afterwards, the number of ways to choose the operators in $\mathcal{D}(p_{\lambda})$ is $O(1)$.
-Note that if we choose a new set to replace $m_{2}]$, then we can reorder the operators in $\mathcal{D}_{1}\cdots\mathcal{D}_{N}$ accordingly using the commutativity property of the Dunkl operators.
-The total number of switches is $k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)$, which will multiply the contribution a factor of $(2\theta)^{k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)}$.
-Furthermore, the operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ will contribute an additional factor of $O(N^{m_{2}-d_{2}-d_{3}})$ for deciding the indices of their switches.
-Thus, using (20) gives that the total contribution is, without including $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$,
-$O(N^{2m_{2}-d_{2}-d_{3}}\theta^{k+m_{2}-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k-m_{2}})=O(N^{m_{2}-d_{2}-d_{3}}\theta^{k-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k}).$
-Additionally, observe that the number of switches in $\mathcal{D}(p_{\lambda})$ is $k-d_{1}-\ell(\lambda)$, so
-$m_{2}-m_{1}\leq k-d_{1}-\ell(\lambda).$
-Therefore, the total contribution is
-$O(N^{m_{1}}(\theta N)^{k-d_{1}-d_{2}-d_{3}-\ell(\lambda)}(\theta N)^{k}).$
-Since $m_{1}\leq\ell(\lambda)$ and $d_{1}+d_{2}+d_{3}+\ell(\lambda)\geq\ell(\nu)$, the total contribution is $O(N^{\ell(\lambda)}(\theta N)^{2k-\ell(\nu)})$.
-To achieve the leading order term, we must have that $m_{1}=\ell(\lambda)$, meaning that the $j_{i}$ are distinct, and that $d_{1}+d_{2}+d_{3}+\ell(\lambda)=\ell(\nu)$, meaning that the total number of derivatives is $\ell(\nu)$ and therefore that each derivative is assigned to a distinct $p_{\nu_{l}}$.
-In addition, we must have have that for $j\in[m_{2}]$, if $\mathcal{D}_{j}$ is not $\partial_{j}$, then it is equivalent to $2\theta Nd_{j}$; this is based on the factor of $O(N^{m_{2}-d_{2}-d_{3}})$ mentioned previously.
-That is, each of these $\mathcal{D}_{j}$ will delete the factor of $x_{j}$ in $e$, which will multiply the leading order term of the contribution by $2\theta N$.
-It is also possible that $\mathcal{D}_{j}$ is the switch from $j$ to $j^{\prime}$ such that $\mathcal{D}_{j^{\prime}}$ is a derivative which has already been assigned to some $p_{\nu_{l}}$ so that the exponent of $x_{j^{\prime}}$ is positive before applying $\mathcal{D}_{j}$.
-Then, $\mathcal{D}_{j}$ will not delete the factor of $x_{j}$, however the number of choices for $j^{\prime}$ is $O(1)$ so this case will not contribute to the leading order term.
-Furthermore, we must have that $m_{2}-m_{1}=k-d_{1}-\ell(\lambda)$, which means that the switches in $\mathcal{D}(p_{\lambda})$ are to distinct indices which are also not elements of $\{j_{i}\}_{1\leq i\leq\ell(\lambda)}$.
-Since $m_{1}=\ell(\lambda)$, this is equivalent to $m_{2}=k-d_{1}$.
-Furthermore, for the remainder term $R$, recall that to delete the variables $x_{j}$ for $j\in\{m_{2}+1,\ldots,N\}$, we multiply by a factor of
-$\prod_{j=N-k+1}^{N-m_{2}}(1+2(j-1)\theta)\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
-In particular, after we remove the term $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$, we obtain a polynomial with degree that is independent of $N$.
-Afterwards, we can deduce that the remainder term $R$ is a polynomial in $\mathbb{Q}[x,y]$ by using casework.
-Furthermore, for (1), we still have that the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$, by the argument that after selecting $j_{1},\ldots,j_{\ell(\lambda)}$ and deleting the variables $x_{j}$ for $j\in\{m_{2}+1,\ldots,N\}$, each factor of $N$ must be accompanied by a factor of $\theta$.
-For (2), we also have that in $R$, the degree of $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$, since we have removed the leading order term which is of order $N^{\ell(\lambda)}(\theta N)^{2k-\ell(\nu)}$.
-Let $\mathcal{S}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
-  1. (1)
-[…]
-  2. (2)
-    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-d_{1}(s)-\ell(\lambda)$, where $d_{1}(s)\leq\ell(\nu)-\ell(\lambda)$ is the number of derivatives in $s$ other than the ones that appear at $1+\lambda_{1}+\cdots+\lambda_{i-1}$ for some $i\in[\ell(\lambda)]$.
-The sequence $s\in\mathcal{S}$ records the locations of the derivatives, since the indices of the switches are fixed after the locations are selected. Then, we have that the leading order term is
-$\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{k}\sum_{s\in\mathcal{S}}N^{-d_{1}(s)}(2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{1}\circ\cdots\circ\mathcal{D}_{k-d_{1}(s)}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
-$=\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{k}\sum_{s\in\mathcal{S}}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-Note that we have reversed the order of $\mathcal{D}_{1},\ldots,\mathcal{D}_{k-d_{1}(s)}$ using the commutativity of the Dunkl operators.
-The purpose of this is to simplify the notation; the actual content of the proof does not change.
-We explain how we arrive at this formula.
-First, observe that $m_{2}=k-d_{1}(s)$.
-Then, recall that the term $(2\theta N)^{d_{1}(s)}$ is included in (20) as $(2\theta N)^{k-m_{2}}$ and that the term $N^{k-d_{1}(s)}$ is from the number of choices for $m_{2}]=[k-d_{1}(s)]$, which can be replaced by any ordered list of $k-d_{1}(s)$ distinct elements of $[N]$.
-It suffices to compute
-(21) $\sum_{s\in\mathcal{S}}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-For $s\in\mathcal{S}$, let $\mathcal{R}^{\prime}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
-  1. (1)
-    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or the switch from $i$ to an element of $\{k-d_{1}(s)+1,\ldots,N\}\cup[i-1]$.
-  2. (2)
-    The total number of derivatives in $r$ is $\ell(\nu)-\ell(\lambda)-d_{1}(s)$.
-Then, the leading order term of (21) is
-$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}^{\prime}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-Recall that for $r\in\mathcal{R}(s)$, if $r_{i}$ is a switch, then we may regard it as $2\theta Nd_{i}$ so that $s$ and $r$ contribute to the leading order term.
-Therefore, we modify the definition of $\mathcal{R}^{\prime}(s)$ as follows.
-For $s\in\mathcal{S}$, let $\mathcal{R}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
-  1. (1)
-    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or $2\theta d_{i}$.
-  2. (2)
-    The total number of derivatives in $r$ is $\ell(\nu)-\ell(\lambda)-d_{1}(s)$.
-Definition 8.17.
-For $r\in\mathcal{R}(s)$, let $d_{2}(r)$ denote the number of derivatives among the operators $r_{i}$ for $i\in[\ell(\lambda)]$ and $d_{3}(r)$ denote the number of derivatives among the operators $r_{i}$ for $\ell(\lambda)+1\leq i\leq k-d_{1}(s)$.
-Furthermore, let $d(r)\triangleq d_{2}(r)+d_{3}(r)$.
-We have that $d(r)+d_{1}(s)=\ell(\nu)-\ell(\lambda)$ for $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$. The leading order term of (21) is
-$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}N^{k-d_{1}(s)-d(r)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
-$=\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}N^{k-\ell(\nu)+\ell(\lambda)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-Note that the term $N^{k-d_{1}(s)-d(r)}$ accounts for the factors of $N$ arising from the switches among the operators of $r$, the number of which is $k-d_{1}(s)-d(r)$.
-Furthermore, we have that $k-d_{1}(s)-d(r)=k-\ell(\nu)+\ell(\lambda)$.
-Therefore, we compute
-(22) $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-For $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$, we let $\mathcal{P}(s,r)$ denote the pair $(S,R)$, where $S$ consists of the set of $i\in[\ell(\lambda)]$ such that $s_{i}$ is a derivative and $R$ consists of the set of $i\in[k-d_{1}(s)]$ such that $r_{i}$ is a derivative; note that $|\mathcal{P}(s,r)|=\ell(\nu)$.
-Let $\mathcal{H}(s,r)$ denote the set of bijections $\zeta:\mathcal{P}(s,r)\rightarrow[\ell(\nu)]$; for $\zeta\in\mathcal{H}(s,r)$, we let $\zeta_{S}$ and $\zeta_{R}$ denote $\zeta$ restricted to the first and second elements of $(S,R)$, respectively.
-For $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(s,r)$, we similarly define $c(s,r,\zeta)$.
-We start with $c=1$.
-We apply $r_{i}$ from $i=1$ to $k-d_{1}(s)$ and then $s_{i}$ from $i=1$ to $|\lambda|$.
-Suppose the current operator is in $r$.
-If the operator is $\partial_{i}$ and is assigned to $p_{\nu_{l}}$ by $\zeta_{R}$, then we multiply $c$ by $x_{i}\partial_{i}p_{\nu_{l}}$.
-Otherwise, if the operator is $2\theta d_{i}$, then we multiply $c$ by $2\theta$.
-Suppose the operator is in $s$.
-If the operator is $\partial_{i}$ and is assigned to $p_{\nu_{l}}$ by $\zeta_{S}$, then we multiply $c$ by $\partial_{i}p_{\nu_{l}}$.
-Otherwise, if the operator is the switch from $i$ to $j$, then we apply the operator to $c$.
-The final value of $c$ is $c(s,r,\zeta)$.
-Afterwards, (22) equals
-(23) $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s),\,\zeta\in\mathcal{H}(s,r)}(2\theta)^{d_{1}(s)}c(s,r,\zeta).$
-Suppose $i\in[\ell(\lambda)]$. Define $A_{i}(s)$ to be the set of $j$ such that there is a switch from $i$ to $j$ in $s$. Then,
-$\ell(\lambda)+1,\ldots,k-d_{1}(s)]=\bigsqcup_{i=1}^{\ell(\lambda)}A_{i}(s).$
-Furthermore, define $S_{i}(s,r,\zeta)$ to be the set of $l\in[\ell(\nu)]$ such that some $\partial_{i}$ is assigned to $p_{\nu_{l}}$ or $\partial_{j}$ is assigned to $p_{\nu_{l}}$ for some $j\in A_{i}(s)$; recall that we may have that $r_{i}=\partial_{i}$.
-Lemma 8.18.
-Suppose $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(r,s)$. If $c(s,r,\zeta)\not=0$, then
-$\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}=\lambda_{i}.$
-for all $i\in[\ell(\lambda)]$.
-Proof.
-Suppose $i\in[\ell(\lambda)]$.
-Note that the only way to remove a factor of $x_{i}$ or a factor of $x_{j}$ for $j\in A_{i}(s)$ is with the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$.
-Based on the algorithm to compute $c$, the total degree in the variables $x_{i}$ and $x_{j}$ for $j\in A_{i}(s)$ that we increase $c$ by is $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$.
-The total degree that the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ lower $c$ by is $\lambda_{i}$, so $\lambda_{i}\geq\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$.
-Conversely, the operators $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ will only lower the degrees of the variables $x_{i}$ and $x_{j}$ for $j\in A_{i}(s)$.
-Since the operators lower $c$ by a total degree of $\lambda_{i}$ and the total degree in the variables that we add to $c$ is $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}$, we must also have that $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}\geq\lambda_{i}$ for the contribution to be nonzero.
-∎
-Let $\mathcal{G}$ be the set of $(s,r,\zeta)$ for $s\in\mathcal{S}$, $r\in\mathcal{R}(s)$, and $\zeta\in\mathcal{H}(r,s)$ such that $\sum_{l\in S_{i}(s,r,\zeta)}\nu_{l}=\lambda_{i}$ for all $i\in[\ell(\lambda)]$.
-Suppose $(s,r,\zeta)\in\mathcal{G}$.
-Then, we define $C_{i}(s,r,\zeta)$ for $i\in[\ell(\lambda)]$ as the factor that $c$ is multiplied by when we apply
-$s_{\lambda_{1}+\cdots+\lambda_{i}}\circ\cdots\circ s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}\circ\prod_{j\in A_{i}(s)}r_{j}\circ r_{i},$
-where we apply the $r_{j}$s in $\prod_{j\in A_{i}(s)}r_{j}$ starting from lower values of $j$.
-Note that the blocks $s_{\lambda_{1}+\cdots+\lambda_{i}}\circ\cdots\circ s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}$ and $\prod_{j\in A_{i}(s)}r_{j}$ are consecutive in $s$ and $r$, respectively, although these two blocks and $r_{i}$ are not consecutive in (22).
-However, we can regard them as consecutive because the factors that operators associated with different $i\in[\ell(\lambda)]$ multiply $c$ by are independent.
-In particular, we have that
-$c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}C_{i}(s,r,\zeta).$
-For $i\in[\ell(\lambda)]$, define $d_{1}(s;\,i)$ to be the number of derivatives among $s_{\lambda_{1}+\cdots+\lambda_{i-1}+2},\ldots,\\ s_{\lambda_{1}+\cdots+\lambda_{i}}$; note that we do not include the derivative at $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1}$.
-Then, we have that
-(24) $(2\theta)^{d_{1}(s)}c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}(2\theta)^{d_{1}(s;\,i)}C_{i}(s,r,\zeta).$
-Similarly to Lemma˜6.5, we have that conditioned on the values of $S_{i}(s,r,\zeta)$, we can factorize the contributions.
-Lemma 8.19.
-For $\nu^{\prime}\in\Gamma_{\text{even}}$, let $c_{\nu^{\prime}}$ denote the quantity $\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s),\,\zeta\in\mathcal{H}(r,s)}(2\theta)^{d_{1}(s)}c(s,r,\\ \zeta)$ from (23) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$.
-Suppose $\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}$ such that $\sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}$ for all $i\in[\ell(\lambda)]$.
-Then,
-$\sum_{\begin{subarray}{c}(s,r,\zeta)\in\mathcal{G},\\ S_{i}(s,r,\zeta)=S_{i},\,1\leq i\leq\ell(\lambda)\end{subarray}}(2\theta)^{d_{1}(s)}c(s,r,\zeta)=\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:\,l\in S_{i}))}.$
-Proof.
-We describe the general idea of the proof, which is the same as that of Lemma˜6.5.
-Suppose that for $i\in[\ell(\lambda)]$, we are given a sequence $(s_{i},r_{i},\zeta_{i})$ which has the structure for $(s,r,\zeta)$ when $\lambda$ is set as $(\lambda_{i})$ and $\nu$ is set as $\gamma((\nu_{l}:l\in S_{i}))$.
-Then, we can combine the $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ for $i\in[\ell(\lambda)]$ to uniquely form an ordered tuple $(s,r,\zeta)\in\mathcal{G}$ such that $S_{i}(s,r,\zeta)=S_{i}$ for $i\in[\ell(\lambda)]$; we combine the $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ based on the conditions in the definition of $\mathcal{S}$.
-We can also uniquely decompose the ordered tuple $(s,r,\zeta)\in\mathcal{G}$ such that $S_{i}(s,r,\zeta)=S_{i}$ for $i\in[\ell(\lambda)]$ into ordered tuples $(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ for $i\in[\ell(\lambda)]$.
-In particular, $\mathbf{s}_{i}$ would correspond to $s_{\lambda_{1}+\cdots+\lambda_{i-1}+1},\ldots,s_{\lambda_{1}+\cdots+\lambda_{i}}$ and $\mathbf{r}_{i}$ would correspond to $r_{j}$ for $j\in A_{i}(s)$ and $r_{i}$.
-Afterwards, using (24) and the fact that $C_{i}(s,r,\zeta)=c(\mathbf{s}_{i},\mathbf{r}_{i},\zeta_{i})$ finishes the proof.
-∎
-Thus, (22) equals
-(25) $\begin{split}&\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}(2\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}\\ &=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\sum_{\begin{subarray}{c}(s,r,\zeta)\in\mathcal{G},\\ S_{i}(s,r,\zeta)=S_{i},1\leq i\leq\ell(\lambda)\end{subarray}}(2\theta)^{d_{1}(s)}c(s,r,\zeta)\\ &=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:\,l\in S_{i}))}.\end{split}$
-Lemma 8.20.
-For $\nu^{\prime}\in\Gamma_{\text{even}}$,
-$c_{\nu^{\prime}}=(2\theta)^{2|\nu^{\prime}|-\ell(\nu^{\prime})}\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC^{\prime}(|\nu^{\prime}|),\,b(\pi)=\nu^{\prime}}\mathbf{1}\{z(\pi)=0\}.$
-Proof.
-First, we compute the exponent of $2\theta$ in $(2\theta)^{d_{1}(s)}c(s,r,\zeta)$.
-The number of switches in $s$ is $|\nu^{\prime}|-d_{1}(s)-1$ and the number of switches in $r$ is $|\nu^{\prime}|-d_{1}(s)-d(r)=|\nu^{\prime}|-\ell(\nu^{\prime})+1$.
-The total number of switches is then $2|\nu^{\prime}|-\ell(\nu^{\prime})-d_{1}(s)$.
-Hence, the exponent of $2\theta$ is $2|\nu^{\prime}|-\ell(\nu^{\prime})$.
-Following this, the general idea is the same as that of Lemma˜6.7, however there are two differences.
-The first difference is due to the possibility that $r_{1}=\partial_{1}$.
-This case would correspond to the first block $B_{1}$ having length $1+\nu_{l}$; note that the length is $1+\nu_{l}$ rather than $\nu_{l}$ since there is an additional factor of $x_{1}$, so after applying $\partial_{1}$ to $p_{\nu_{l}}$, we would obtain the term $\nu_{l}x_{1}^{\nu_{l}}$ rather than $\nu_{l}x_{1}^{\nu_{l}-1}$.
-However, since $s_{1}=\partial_{1}$, the second block $B_{2}$ would start at $2$.
-On the other hand, if $r_{1}=2\theta d_{1}$, then that would correspond to the first block $B_{1}$ having length $1$; in this case, the remaining blocks would be a noncrossing partition of $\{2,\ldots,|\nu^{\prime}|+1\}$.
-The second difference is due to the operators $r_{j}$ for $j\in A_{1}(s)$ which are derivatives.
-If $j\in A_{1}(s)$ and $r_{j}=\partial_{j}$ is assigned to $p_{\nu_{l}}$, after applying $r_{j}$ to $p_{\nu_{l}}$ we would obtain a factor of $\nu_{l}x_{j}^{\nu_{l}}$, where the exponent is $\nu_{l}$ rather than $\nu_{l}-1$ since there is an additional factor of $x_{j}$.
-Afterwards, when we apply the switch from $1$ to $j$, we will need to delete the power of $x_{j}$.
-However, note that the exponent of $x_{1}$ must be zero when we apply this switch, otherwise we will not be able to delete the power of $x_{j}$ and the contribution will be 0.
-Furthermore, in order to delete the power of $x_{j}$ when we apply the switch, we will replace $x_{j}^{\nu_{l}}$ with $2\theta x_{1}^{\nu_{l}-1}$; this is straightforward to deduce because $\nu_{l}$ is even.
-The outcome is then equivalent to applying $2\theta\partial_{1}$ to $p_{\nu_{l}}$.
-We explain how this impacts the bijective correspondence.
-Suppose $\pi\in NC^{\prime}([|\nu^{\prime}|])$ and $i\in[|\nu^{\prime}|+1]$ such that $i\geq 3$ and $i$ is the minimal value of a block of $\pi$ of length $L$; recall that $i=1$ and $2$ correspond to $r_{1}$ and $s_{1}$, respectively.
-Then, we would have that $s_{i-1}$ is either a $\partial_{1}$ which is assigned to $p_{\nu_{l}}$ or a switch from $1$ to $j$ such that $r_{j}=\partial_{j}$ is assigned to $p_{\nu_{l}}$ for some $\nu_{l}$ such that $|\nu_{l}|=L$.
-However, note that for $c(s,r,\zeta)$ to be nonzero, we can only choose to set $r_{j}=\partial_{j}$ for $j\in A_{1}(s)$ such that the exponent of $x_{1}$ is zero when applying the switch from $1$ to $j$.
-That is, we must select a location $i\in\{3,\ldots,|\nu^{\prime}|+1\}$ for the switch in $\pi$ such that $i$ is the minimal element of the block that contains it and all earlier blocks do not contain elements greater than $i$; in other words, we must have that $b(i;\,\pi)=1$ and $d(i;\,\pi)$ equal to the size of the block that contains it.
-If such a location $i$ in $\pi$ exists, then for $s_{i-1}$, we must choose $\partial_{1}$ or the switch from $1$ to $j$ and set $r_{j}=\partial_{j}$, where $j$ is determined by the previous switches.
-However, the latter option is equivalent to applying $2\theta\partial_{1}$ rather than $\partial_{1}$.
-We have observed that after multiplying by $(2\theta)^{d_{1}(s)}$, the exponent of $2\theta$ is constant.
-Therefore, when summing the contributions from selecting $2\theta\partial_{1}$ and $\partial_{1}$, we will obtain zero.
-Hence, any $\pi\in NC^{\prime}(|\nu^{\prime}|)$ with $z(\pi)\geq 1$ will have a contribution of zero.
-On the other hand, if $\pi\in NC^{\prime}(|\nu^{\prime}|)$ satisfies $z(\pi)=0$, then the contribution corresponding to it is $(2\theta)^{2|\nu^{\prime}|-\ell(\nu^{\prime})}\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}$.
-∎
-Using (25) and Lemma˜8.20, it suffices to compute
-$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}(2\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
-$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))(\#\,\pi\in NC^{\prime}(\lambda_{i})\text{ such that }z(\pi)=0\text{ and }\gamma(b(\pi))=\gamma((\nu_{j}:j\in S_{i})))$
-$=\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}(2\theta)^{2k-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
-$\prod_{i=1}^{\ell(\lambda)}(\#\,\pi\in NC^{\prime}(\lambda_{i})\text{ such that }z(\pi)=0\text{ and }\gamma(b(\pi))=\gamma_{i})$
-$=(2\theta)^{2k-\ell(\nu)}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i}),\,z(\pi)=0}\prod_{B\in b(\pi)}x_{|B|}.$
-Suppose $k\geq 2$ and $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC^{\prime}(k)$, where the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $q\in[m-1]$.
-Then, we show that $z(\pi)=0$ if and only if $k+1\in B_{1}\cup B_{2}$.
-If $k+1\in B_{1}$, then clearly $z(\pi)=0$; if $k+1\in B_{2}$, then $B_{1}=\{1\}$ and it is also the case that $z(\pi)=0$.
-Next, assume that $k+1\notin B_{1}\cup B_{2}$.
-If $B_{1}\not=\{1\}$, let $M$ be the maximal element of $B_{1}$.
-Then, $M<k+1$ and it is evident that $b(M+1;\,\pi)=0$ and $d(M+1;\,\pi)$ equals the size of the block that contains $M+1$.
-Therefore, $z(\pi)\geq 1$.
-Next, assume that $B_{1}=\{1\}$.
-If we let $M$ be the maximal element of $B_{2}$, then we can similarly deduce that $b(M+1;\,\pi)=0$ and $d(M+1;\,\pi)$ equals the size of the block that contains $M+1$ to conclude that $z(\pi)\geq 1$.
-Afterwards, we observe that $b:\{\pi\in NC^{\prime}(k):z(\pi)=0\}\rightarrow NC(\{2,\ldots,k+1\})$ is a bijection.
-To show that $b$ is surjective, suppose $\pi^{\prime}=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(\{2,\ldots,k+1\})$, the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $q\in[m-1]$, and $k+1\in B_{i}$ for $i\in[m]$.
-If $i=1$, then let $\pi=\{1\}\sqcup\pi^{\prime}$.
-If $i>1$, then let $\pi$ be $\pi^{\prime}$ with $B_{i}$ replaced by $\{1\}\cup B_{i}$.
-In both cases, we have that $\pi\in NC^{\prime}(k)$, $z(\pi)=0$, and $b(\pi)=\pi^{\prime}$.
-Next, we show that $b$ is injective.
-We use the same notation for $\pi^{\prime}$.
-Suppose $\pi\in NC^{\prime}(k)$, $z(\pi)=0$, and $b(\pi)=\pi^{\prime}$.
-If $i=1$, then we must have that the block containing $1$ in $\pi$ is $\{1\}$ because the block containing $2$ also contains $k+1$.
-Hence, there is a unique choice for $\pi$.
-Next, assume that $i>1$.
-Then, we must have that the block that contains $1$ in $\pi$ also contains $k+1$.
-Therefore, the block that contains $1$ in $\pi$ is $B_{i}\cup\{1\}$ and we similarly have that there is a unique choice for $\pi$.
-Hence, the value of (25) is
-$(2\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}.$
-After multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)N^{2k-\ell(\nu)+\ell(\lambda)}$, we obtain the formula in the statement of the theorem.
-8.4. Proof of part (C) of Theorem˜1.1
-We prove part (C) of Theorem˜1.1, see Corollary˜8.22, as well as a generalization, see Theorem˜8.21.
-Recall that $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$ for $\lambda,\nu\in\Gamma_{\text{even}}$.
+8.2. Proof of part (C) of Theorem˜1.2
+We obtain the leading order terms of $ep_{\lambda},ep_{\nu}\\ ]_{D^{N}(\theta)}$ in the following lemma by using the leading order terms of the type BC Dunkl bilinear form.
+The leading order terms can be expressed as a product involving gamma functions and polynomials.
+For example, the product $\prod_{i=1}^{N}(1+2(i-1)\theta)$ can also be expressed as $(2\theta)^{N}\frac{\Gamma(\frac{1}{2\theta}+N)}{\Gamma(\frac{1}{2\theta})}$.
+Lemma 8.8 ([dunkl_singular_poly]*Corollary 4.5).
+Suppose $\lambda,\nu\in\Gamma_{\text{even}}$. Then,
+$ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=\prod_{i=1}^{N}(1+2(i-1)\theta)[p_{\lambda},p_{\nu}]_{BC^{N}(\theta,1)}.$
+Furthermore, for $\theta_{0},\theta_{1}\in\mathbb{C}$,
+$ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{N}(1+2(i-1)\theta_{0}+2\theta_{1})[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1}+1)}.$
+We prove part (C) of Theorem˜1.2, see Corollary˜8.10, as well as a generalization, see Theorem˜8.9.
+Recall that $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$ for $\lambda,\nu\in\Gamma_{\text{even}}$ and $\mathcal{M}^{D}=\mathcal{M}^{BC}(0)$.
 Theorem 8.9.
@@
 Proof.
-See Theorem˜8.11 for the equivalence of (a) and (c).
-For the equivalence of (b) and (d), we can use Theorem˜8.16 and follow the method of the proof of Theorem˜6.10, after dividing by the nonzero quantity $\prod_{j=1}^{N}(1+2(j-1)\theta)$.
-In particular, we have that $e,F_{N}]_{D^{N}(\theta)}=(d_{0}(N)+1)[e,e]_{D^{N}(\theta)}=(d_{0}(N)+1)\prod_{j=1}^{N}(1+2(j-1)\theta))$, so we can deduce that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}$ and
-$\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
-are equivalent. ∎
+The equivalence of (a) and (c) follows from Theorem˜7.4 with $c=0$.
+For the equivalence of (b) and (d), we note that $\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}(N)+1$ and for $\nu\in\Gamma_{\text{even}}$,
+$\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=\left[p_{\nu},\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)\right]_{BC^{N}(\theta,1)}$
+by Lemma˜8.8. Afterwards, the equivalence also follows from Theorem˜7.4 with $c=0$. ∎
 Corollary 8.10.
@@
 Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) and (d) are equivalent.
-8.5.
-Next, we consider the asymptotics of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ in the $\theta N\rightarrow c\in\mathbb{C}$ regime.
-Afterwards, we arrive at results similar to those of the previous subsection for this regime, see Theorem˜8.26 and Corollary˜8.27.
-For $\pi\in NC(k+1)$, define
-$W^{D;\,\text{odd}}(\pi)(x)\triangleq$ $(1+\mathbf{1}\{b(2;\,\pi)=0\}(d(2;\,\pi)-1))\times$
-$\prod_{i\in[k+1],\,i\geq 3,\,b(i;\,\pi)=0}(x+d(i;\,\pi))\left(\frac{1}{1+x}\right)^{z(\pi)+\mathbf{1}\{f(\pi)>1\}}.$
-Theorem 8.23.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=\prod_{i=1}^{N-k}(1+2(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(1+2N\theta)^{k}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\times$
-$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(2N\theta)\prod_{B\in b(\pi)}x_{|B|}$
-$+R(N,\theta)\bigg),$
-where $R\in\mathbb{Q}[x,y]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
-Proof.
-We follow the same framework as the proof of Theorem˜8.16. The only difference is that in this case, we view $\theta N$ as having the same weight as a constant to compute the leading order term.
-We use the same notation for $m_{1}$, $m_{2}$, $d_{1}$, $d_{2}$, and $d_{3}$.
-Define $d_{1}^{\prime}$, $d_{2}^{\prime}$, and $d_{3}^{\prime}$ to be the number of derivatives in $\mathcal{D}(p_{\lambda})$ other than the $\ell(\lambda)$ derivatives at the starts of the blocks, $\mathcal{D}_{j}$ for $j\in[m_{1}]$, and $\mathcal{D}_{j}$ for $j\in[m_{2}]\backslash[m_{1}]$, respectively, that we assign to some $p_{\nu_{l}}$.
-Then, we have that $d^{\prime}_{1}+d^{\prime}_{2}+d^{\prime}_{3}=\ell(\nu)-\ell(\lambda)$ and $d_{i}^{\prime}\leq d_{i}$ for $i\in[3]$.
-The factor of $\prod_{j=1}^{N-m_{2}}(1+2(j-1)\theta)$ remains the same. However, its leading order term is now
-(26) $(1+2\theta N)^{k-m_{2}}\prod_{j=1}^{N-k}(1+2(j-1)\theta).$
-The number of ways to select $m_{2}]$ is $O(N^{m_{2}})$ and the number of ways to choose the operators in $\mathcal{D}(p_{\lambda})$ is $O(1)$.
-Furthermore the switches in $\mathcal{D}(p_{\lambda})$ multiply the contribution by a factor of $(2\theta)^{k-d_{1}-\ell(\lambda)}$.
-The operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ which are not a derivative assigned to some $p_{\nu_{l}}$ will contribute a factor of $(1+2N\theta)^{m_{2}-d_{2}^{\prime}-d_{3}^{\prime}}$.
-Therefore, using (26) gives that the total contribution is, without including $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$,
-$O(N^{m_{2}}\theta^{k-d_{1}-\ell(\lambda)}(1+2N\theta)^{k-d_{2}^{\prime}-d_{3}^{\prime}}).$
-We still have that $m_{2}-m_{1}\leq k-d_{1}-\ell(\lambda)$, so the total contribution is
-$O(N^{m_{1}}(\theta N)^{k-d_{1}-\ell(\lambda)}(1+2N\theta)^{k-d_{2}^{\prime}-d_{3}^{\prime}}).$
-Furthermore, $m_{1}\leq\ell(\lambda)$, so multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$ gives the leading order term, which is $O(N^{\ell(\lambda)})\prod_{j=1}^{N-k}(1+2(j-1)\theta)$.
-To achieve the leading order term, we must have that $m_{2}-m_{1}=k-d_{1}-\ell(\lambda)$, which means that the switches in $\mathcal{D}(p_{\lambda})$ are to distinct indices not in $m_{1}]$, and that $m_{1}=\ell(\lambda)$, which means that the $j_{i}$ are distinct.
-Furthermore, each of the operators $\mathcal{D}_{j}$ for $j\in[m_{2}]$ which is not a derivative assigned to some $p_{\nu_{l}}$ should delete $x_{j}$ by either applying $\partial_{j}$ or $2N\theta d_{j}$.
-Let $\mathcal{S}^{\prime}$ denote the set of sequences of operators $s=\{s_{i}\}_{1\leq i\leq|\lambda|}$ such that:
-  1. (1)
-[…]
-  2. (2)
-    The $j$th switch is from some element of $\ell(\lambda)]$ (which is determined by condition (1)) to $\ell(\lambda)+j$ for $1\leq j\leq k-d_{1}(s)-\ell(\lambda)$, where $d_{1}(s)$ is the number of derivatives in $s$ other than the ones that appear at $1+\lambda_{1}+\cdots+\lambda_{i-1}$ for some $i\in[\ell(\lambda)]$.
-The sequence $s\in\mathcal{S}^{\prime}$ records the locations of the derivatives, since the indices of the switches are fixed after the locations are selected.
-Then, we have that the leading order term is
-$\prod_{j=1}^{N-k}(1+2(j-1)\theta)\sum_{s\in\mathcal{S}^{\prime}}N^{k-d_{1}(s)}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ$
-$\mathcal{D}_{1}\circ\cdots\circ\mathcal{D}_{k-d_{1}(s)}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}$
-$=\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}\sum_{s\in\mathcal{S}^{\prime}}N^{k-d_{1}(s)-\ell(\lambda)}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ$
-$\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-Recall that since $m_{2}=k-d_{1}(s)$, the term $(1+2\theta N)^{d_{1}(s)}$ is included in (20) as $(1+2\theta N)^{k-m_{2}}$ and the term $N^{k-d_{1}(s)}$ is from the number of choices for $m_{2}]=[k-d_{1}(s)]$, which can be replaced by any ordered list of $k-d_{1}(s)$ distinct elements of $[N]$.
-Define $\mathcal{S}$ to be the set of $s\in\mathcal{S}^{\prime}$ with each switch multiplied by a factor of $N$; this is equivalent to replacing $\theta$ with $\theta N$.
-Then, since the number of switches in $s\in\mathcal{S}^{\prime}$ is $k-d_{1}(s)-\ell(\lambda)$, the leading order term is
-$\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}\sum_{s\in\mathcal{S}}(1+2\theta N)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ\mathcal{D}_{k-d_{1}(s)}\circ\cdots\circ\mathcal{D}_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-For $s\in\mathcal{S}$, let $\mathfrak{R}^{\prime}(s)$ denote the set of sequences of operators $r=\{r_{i}\}_{1\leq i\leq k-d_{1}(s)}$ such that:
-  1. (1)
-    For $1\leq i\leq k-d_{1}(s)$, $r_{i}$ is either $\partial_{i}$ or $(1+2\theta N)d_{i}$.
-  2. (2)
-    If $1\leq i\leq k-d_{1}(s)$ and $r_{i}=\partial_{i}$, then $r_{i}$ must be assigned to some $p_{\nu_{l}}$.
-Note that we use a slight abuse of notation, since condition (2) should be a condition on the injective function $\zeta$ which assigns the derivatives of $s\in\mathcal{S}$ and $r\in\mathfrak{R}^{\prime}(s)$ to the $p_{\nu_{l}}$ rather than a condition on $r\in\mathfrak{R}^{\prime}(s)$.
-However, we avoid using the notation for $\zeta$ for simplicity.
-Then, the leading order term, without including $\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}$, is
-(27) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}^{\prime}(s)}(1+2N\theta)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-The reasoning behind conditions (1) and (2) is that if $r_{i}$ is not a derivative which is assigned to some $p_{\nu_{l}}$, then to contribute to the leading order term, it must delete the variable $x_{i}$ and multiply the contribution by a factor of $1+2N\theta$.
-In particular, we have that each derivative in $r\in\mathfrak{R}^{\prime}(s)$ must be assigned to some $p_{\nu_{l}}$, while this is not necessarily true for a derivative in $s\in\mathcal{S}$.
-For $s\in\mathcal{S}$, let $\mathfrak{R}(s)$ denote the set of $r\in\mathfrak{R}^{\prime}(s)$ with each of its operators multiplied by a factor of $\frac{1}{1+2\theta N}$.
-After using (27), the leading order term, without including $\prod_{j=1}^{N-k}(1+(j-1)2\theta)N^{\ell(\lambda)}$, is
-$(1+2N\theta)^{k}\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-Therefore, we compute
-(28) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}.$
-To finish the proof, we can use the framework as the proof of Theorem˜8.16.
-In particular, for $\nu^{\prime}\in\Gamma_{\text{even}}$, let $c_{\nu^{\prime}}$ denote the quantity (28) when $\lambda$ is set as $(|\nu^{\prime}|)$ and $\nu$ is set as $\nu^{\prime}$.
-We similarly have that
-(29) $\sum_{s\in\mathcal{S},\,r\in\mathfrak{R}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu}=\sum_{\begin{subarray}{c}[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)},\\ \sum_{l\in S_{i}}|\nu_{l}|=\lambda_{i}\,\forall i\in[\ell(\lambda)]\end{subarray}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{l}:l\in S_{i}))}.$
-Lemma 8.24.
-For $\nu^{\prime}\in\Gamma_{\text{even}}$,
-$c_{\nu^{\prime}}=\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC(|\nu^{\prime}|+1),\,\gamma(b(\pi))=\nu^{\prime}}W^{D;\,\text{odd}}(\pi)(2N\theta).$
-Proof.
-The proof follows the method of the proof of Lemma˜8.20.
-First, we state the bijective correspondence.
-Suppose $s\in\mathcal{S}$, $r\in\mathfrak{R}(s)$, and $(s,r,\zeta)$ corresponds to $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC(k+1)$, where the minimal element of $B_{q}$ is less than the minimal element of $B_{q+1}$ for $1\leq q\leq m-1$; we describe $\pi$ based on $(s,r,\zeta)$.
-We have not defined $\zeta$ yet; in this setting, $\zeta$ assigns each derivative in $r$ and some of the derivatives in $s$ to a $p_{\nu^{\prime}_{l}}$ so that each $p_{\nu^{\prime}_{l}}$ for $l\in[\ell(\nu^{\prime})]$ is assigned to once.
-The block $B_{1}$ corresponds to the operator $r_{1}$; if $r_{1}=\frac{1}{1+2N\theta}\partial_{1}$, then $r_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$ and $B_{1}$ has size $|\nu^{\prime}_{l}|+1$.
-Otherwise, $r_{1}=d_{1}$ and $B_{1}$ has size $1$.
-Afterwards, for $1\leq i\leq|\nu^{\prime}|$, when $s_{i}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$ or when $s_{i}$ is a switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to some $p_{\nu^{\prime}_{l}}$, then there is a block in $\pi$ with size $\nu^{\prime}_{l}$ and minimal element $i+1$.
-Note that in contrast with Lemma˜8.20, we use $NC(k+1)$ rather than $NC^{\prime}(k)$.
-The reason for this is that $s_{1}=\partial_{1}$ is not necessarily assigned to a $p_{\nu^{\prime}_{l}}$, so $B_{2}$ does not necessarily have minimal element $2$.
-Next, we justify the weight $W^{D;\,\text{odd}}(\pi)(2N\theta)$ for each $\pi\in NC(k+1)$.
-First, the reasoning for the coefficient of $\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}$ is the same.
-If $r_{1}=\frac{1}{1+2N\theta}\partial_{1}$, then we must multiply the contribution by a factor of $\frac{1}{1+2N\theta}$; this is the reasoning for the factor $(1+2N\theta)^{-\mathbf{1}\{f(\pi)>1\}}$.
-Otherwise, if $r_{1}=d_{1}$, then we do not need to multiply by an additional factor.
-Suppose $i\geq 3$ and $b(i;\,\pi)=0$.
-Then, $s_{i-1}$ is a switch or an unassigned derivative.
-If $s_{i-1}$ is the switch from $1$ to $j$, then because $b(i;\,\pi)=0$, $r_{j}=d_{j}$.
-In this case, we would multiply the contribution by a factor of $2N\theta+d(i;\,\pi)$, where the term $2N\theta$ is due to $s_{i-1}$.
-Suppose $b(2;\,\pi)=0$. Then, because $s_{1}=\partial_{1}$ is unassigned, we would multiply the contribution by a factor of $d(2;\,\pi)$.
-Next, suppose $i\geq 2$ and $b(i;\,\pi)=1$.
-Suppose $d(i;\,\pi)$ is greater than the size of the block that contains $i$.
-Then, we would have that $s_{i-1}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$, so we would multiply the contribution by a factor of $1$.
-Note that we could also have that $s_{i-1}$ is the switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to some $p_{\nu^{\prime}_{l}}$, but because $d(i;\,\pi)$ is greater than the size of the block that contains $i$, the contribution in this case would be zero.
-As mentioned earlier, this is due to the fact that the degree of $x_{1}$ before applying $s_{i-1}$ is positive, so we will not be able to eliminate the factor of $x_{j}$ arising from $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$.
-Next, suppose $d(i;\,\pi)$ equals the size of the block that contains it; recall that this is equivalent to all earlier blocks having no elements greater than $i$.
-Then, if $s_{i-1}=\partial_{1}$ is assigned to some $p_{\nu^{\prime}_{l}}$, we would similarly multiply the contribution by a factor of $1$.
-On the other hand, if $s_{i-1}$ is the switch from $1$ to $j$ and $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ is assigned to $p_{\nu^{\prime}_{l}}$, then we would multiply the contribution by a factor of $\frac{2N\theta}{1+2N\theta}$.
-Recall that in this case, $s_{i-1}$ must modify $x_{j}^{\nu^{\prime}_{l}}$ to a multiple of $x_{i}^{\nu^{\prime}_{l}-1}$; in order to do so, we will obtain $2N\theta x_{i}^{\nu^{\prime}_{l}-1}$ since $\nu^{\prime}_{l}$ is even.
-Since $r_{j}=\frac{1}{1+2N\theta}\partial_{j}$ has a factor of $\frac{1}{1+2N\theta}$, we obtain a contribution of $\frac{2N\theta}{1+2N\theta}$.
-Then, the total contribution in this case is $1-\frac{2N\theta}{1+2N\theta}=\frac{1}{1+2N\theta}$, which is the reasoning for the factor $(1+2N\theta)^{-z(\pi)}$.
-∎
-Using (29) and Lemma˜8.24, it suffices to compute
-$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}+1),\\ \gamma(b(\pi))=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W^{D;\,\text{odd}}(\pi)(2N\theta)$
-$=\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}+1),\\ \gamma(b(\pi))=\gamma_{i}\end{subarray}}W^{D;\,\text{odd}}(\pi)(2N\theta)$
-$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(2N\theta)\prod_{B\in b(\pi)}x_{|B|}.$
-After multiplying by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)(1+2N\theta)^{k}N^{\ell(\lambda)}$, we obtain the formula in the statement of the theorem. ∎
-Define the infinite dimensional matrix $\mathcal{W}^{D;\,\text{odd}}(x)\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}[x]$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$,
-$\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)\triangleq(1+x)^{|\nu|}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(x)\prod_{B\in b(\pi)}x_{|B|}$
-The following lemma is straightforward to deduce.
-Lemma 8.25.
-Suppose $\lambda,\nu\in\Gamma_{\text{even}}$ and $|\lambda|=|\nu|$.
-  1. (A)
-    If $\lambda=\nu$, then $\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)=(1+x)^{|\nu|}\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}W^{D;\,\text{odd}}(\{1\}\cup\{2,\ldots,\nu_{l}+1\})(x)$.
-  2. (B)
-    If $\ell(\lambda)\geq\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{D;\,\text{odd}}_{\lambda\nu}(x)=0$.
+8.3. Leading order terms of the type D Dunkl bilinear form in the high temperature regime
 Theorem 8.11.
@@
     and for all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
-    $=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
+    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}$
+    $=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(c,1)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
 Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
+Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer less than 1-1, then (d) implies (b).
 Proof.
-For the implications involving (a) and (c), see Theorem˜8.13.
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$.
-Using Theorem˜8.23 gives that $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ is divisible by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$.
-Then, after dividing by this nonzero quantity, we can use Theorem˜8.23 and the same analysis as in the proof of Theorem˜6.17 to deduce that (b) implies (d).
-Furthermore, we similarly have that if $2c$ is not a negative integer, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]$ is invertible for all $k\geq 1$, so (d) implies (b) if this is the case.
+The implications between (a) and (c) follow from Theorem˜7.8 with $c=0$, since $\mathcal{W}^{BC}(c,0)=\mathcal{W}^{A}(2c)$.
+For the implications between (b) and (d), we apply Lemma˜8.8 as we did to prove Theorem˜8.9 and then apply Theorem˜7.8 with $c=1$.
 ∎
@@
     and for all $\nu\in\Gamma_{\text{even}}$,
-    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
-    $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(|\nu|+1)}W^{D;\,\text{odd}}(\pi)(2c)\prod_{B\in b(\pi)}(1+2c)^{|B|}|B|d_{(|B|)},$
-    where $d_{(k)}=0$ for odd positive integers $k$.
+    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}$
+    $=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(c,1)\prod_{B\in\pi}|B|d_{(|B|)}.$
 Then, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
-Remark 8.28.
-As mentioned in Remark˜8.14, the equivalences between (a) and (c) of Theorems˜8.21 and 8.26 and Corollaries˜8.22 and 8.27 follow from the analogous results for the type BC root system after setting $\theta_{1}=0$.
+Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer less than 1-1, then (d) implies (b).
 9. Asymptotics of the coefficients of Bessel functions
-In this section, we determine the asymptotics of the coefficients of Bessel functions in each of the regimes mentioned in Subsection˜1.2.
+In this section, we determine the asymptotics of the coefficients of Bessel functions in each of the regimes mentioned in Subsection˜1.4.
 Note that we determine the asymptotics for the coefficients of the terms with a fixed degree and the analyses for different degrees are separate.
@@
     The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
-    $N^{-\ell(\nu)}(\mathcal{W}^{BC}(2c_{0},2c_{1})[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
+    $N^{-\ell(\nu)}(\mathcal{W}^{BC}(c_{0},c_{1})[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
   2. (b)
-    The diagonal entry $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda})^{-1}$.
+    The diagonal entry $\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\lambda})^{-1}$.
   3. (c)
-    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\nu}^{-1}=0$.
+    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\nu}^{-1}=0$.
 Proof.
@@
 If $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$, $c_{0}$ is not a negative integer, and $2c_{0}+2c_{1}$ is not a negative odd integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
-$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{BC}([r_{l}])(2c_{0},2c_{1})\right)^{-1}\times$
+$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{BC}([r_{l}])(c_{0},c_{1})\right)^{-1}\times$
 $(1+o_{N}(1))r(a)r(x)+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}\times$
-$(\mathcal{W}^{BC}(c)[k]^{-1}_{rs}+\mathcal{W}^{BC}(c)[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
+$(\mathcal{W}^{BC}(c_{0},c_{1})[k]^{-1}_{rs}+\mathcal{W}^{BC}(c_{0},c_{1})[k]^{-1}_{sr}+o_{N}(1))r(a)s(x).$
 Proof.
@@
 9.3. Coefficients for the $D^{N}$ root system
-The framework of this subsection is analogous to that of the $A^{N-1}$ root system.
-Lemma 9.9.
-Suppose $|\theta N|\rightarrow\infty$.
-Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$.
-Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
-  1. (a)
-    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
-    $(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
-  2. (b)
-    The diagonal entry $\mathcal{M}^{D}[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{M}^{D}[k]_{\lambda\lambda})^{-1}$.
-  3. (c)
-    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{D}[k]_{\lambda\nu}^{-1}=0$.
-Proof.
-See Theorem˜8.10. ∎
-Lemma 9.10.
-Suppose $|\theta N|\rightarrow\infty$.
-Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$.
-Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$.
-Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
-The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
-$(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)^{-1}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
-Proof.
-See Theorem˜8.16. ∎
-Corollary 9.11.
-If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then
-$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
-and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
-$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}(\theta N)^{-k}\theta^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}2^{r_{l}-1}r_{l}\right)^{-1}\times$
-$\left(1+o_{N}(1)+(1+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\right)r(a)r(x)$
-$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}(\theta N)^{-k}(\theta N)^{\min(\ell(r),\ell(s))}N^{-\max(\ell(r),\ell(s))}\times$
-$\bigg[\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1)$
-$+(\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
-Proof.
-See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10.
-Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$.
-Furthermore, we obtain the term $\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$ in $J_{a}^{D^{N}(\theta)}$ from $e,J_{a}^{D^{N}(\theta)}]_{D^{N}(\theta)}=e(a)$.
-∎
-Lemma 9.12.
-Suppose $\theta N\rightarrow c\in\mathbb{C}$ and $2c$ is not a negative integer.
-Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$.
-Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
-  1. (a)
-    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
-    $N^{-\ell(\nu)}(\mathcal{W}^{A}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
-  2. (b)
-    The diagonal entry $\mathcal{W}^{A}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{A}(2c)[k]_{\lambda\lambda})^{-1}$.
-  3. (c)
-    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{A}(2c)[k]_{\lambda\nu}^{-1}=0$.
-Proof.
-See Theorem˜8.12. ∎
-Lemma 9.13.
-Suppose $\theta N\rightarrow c\in\mathbb{C}$, $2c$ is not a negative integer, and assume that $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large.
-Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$.
-Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
-  1. (a)
-    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
-    $N^{-\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)^{-1}(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
-  2. (b)
-    The diagonal entry $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda})^{-1}$.
-  3. (c)
-    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\nu}^{-1}=0$.
-Proof.
-See Theorem˜8.23. ∎
-Corollary 9.14.
-If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then
-$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
-and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
-$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}\right)^{-1}\bigg[\prod_{l=1}^{\ell(r)}W^{A}([r_{l}])(2c)^{-1}+o_{N}(1)$
-$+\left(\prod_{l=1}^{\ell(r)}W^{D;\,\text{odd}}(\{1\}\cup\{2,\ldots,r_{l}+1\})(2c)^{-1}+o_{N}(1)\right)\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)r(x)$
-$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}\bigg[\mathcal{W}^{A}(2c)[k]^{-1}_{rs}+\mathcal{W}^{A}(2c)[k]^{-1}_{sr}+o_{N}(1)$
-$+(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{rs}+\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
-Proof.
+The framework of this subsection is analogous to that of the $A^{N-1}$ root system. By [demni_type_d], we have that
+(20) $J_{a}^{D^{N}(\theta)}(x)=J_{a}^{BC^{N}(\theta,0)}(x)+\frac{e(a)e(x)}{\prod_{i=1}^{N}(1+2(i-1)\theta)}J_{a}^{BC^{N}(\theta,1)}(x).$
+Therefore, it suffices to consider the coefficients of the Bessel function for the $BC^{N}$ root system.
 10. Applications
 In this section, we discuss applications of the main results of this paper.
-Some settings that we consider include the weak convergence of the measures in Conjecture˜1.6 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
+Some settings that we consider include the weak convergence of the measures in Conjecture˜1.5 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
 10.1. Weak convergence to the free convolution
-The results of this subsection are based on the assumption of Conjecture˜1.6.
+The results of this subsection are based on the assumption of Conjecture˜1.5.
 Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems.
@@
 Assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
-Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.6.
+Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.5.
 Then, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
 Proof.
-Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.4. Then, we have that
+Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.3. Then, we have that
 $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}(\epsilon_{i}^{1}+\epsilon_{i}^{2})}d\mu_{a_{1}}^{\text{sym}}(\epsilon^{1})d\mu_{a_{2}}^{\text{sym}}(\epsilon^{2})=\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)d\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}(a).$
@@
 $\mu^{\text{sym}}_{z}\left(B(z,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z_{i_{j}},\frac{r}{2})\right)=0,$
-which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.4. Thus, $u$ is in the support of $\epsilon$.
+which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.3. Thus, $u$ is in the support of $\epsilon$.
 However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$.
-Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.4, the result follows.
+Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.3, the result follows.
 ∎
-In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$.
+In the following corollary, we prove the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of power sums after assuming that $a$ and $b$ converge in terms of power sums as $N\rightarrow\infty$.
 Afterwards, it is straightforward to deduce Corollaries˜1.7 and 1.8.
@@
   • (A)
-    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{d}^{a}$ and $m_{d}^{b}$ are real numbers such that
-    $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}b(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{b}$
-    for $d\geq 1$. Let $a_{(k)}$ and $b_{(k)}$ for $k\geq 1$ be real numbers that solve
+    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{d}^{a}$ and $m_{d}^{b}$ are real numbers such that for all $d\geq 1$,
+    $\lim_{N\rightarrow\infty}\frac{p_{(d)}(a(N))}{(\theta N)^{d}N}=m_{d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{p_{(d)}(b(N))}{(\theta N)^{d}N^{1}}=m_{d}^{b}.$
+    Let $a_{(k)}$ and $b_{(k)}$ for $k\geq 1$ be real numbers that solve
     $m_{d}^{a}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|a_{(|B|)}\text{ and }m_{d}^{b}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}|B|b_{(|B|)}$
-    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
-    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{d}\right]}{(\theta N)^{d}N}=\sum_{\pi\in NC(d)}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
+    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma$,
+    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[p_{\lambda}(c)\right]}{(\theta N)^{|\lambda|}N^{\ell(\lambda)}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
   • (B)
@@
     $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
-    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
-    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta_{0}N)^{2d}N}$
-    $=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
+    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[p_{\lambda}(c)\right]}{(\theta_{0}N)^{|\lambda|}N^{\ell(\lambda)}}$
+    $=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\text{even}}(\lambda_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
   • (C)
@@
     $m_{2d}^{b}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|b_{(|B|)}$
-    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
-    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta N)^{2d}N}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
+    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma_{\text{even}}$,
+    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[p_{\lambda}(c)\right]}{(\theta N)^{|\lambda|}N^{\ell(\lambda)}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\text{even}}(\lambda_{i})}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
 Remark 10.3.
@@
 Then, we can express $J_{a(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}a_{\lambda}(N)p_{\lambda}\right)$ and $J_{b(N)}^{A^{N-1}(\theta)}=\exp\left(\sum_{\lambda\in\Gamma_{N}}b_{\lambda}(N)p_{\lambda}\right)$ over $\Omega_{N}$.
-For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.14.
-Thus, part (D) of Lemma˜10.15 implies that for all $\lambda\in\Gamma$,
+For $N\geq 2$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15.
+Thus, part (D) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
 (21) $\begin{split}\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]&=\left[p_{\lambda},J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}\right]_{A^{N-1}(\theta)}\\ &=\left[p_{\lambda},\exp\left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)\right]_{A^{N-1}(\theta)};\end{split}$
@@
 $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}[p_{\lambda}(c_{1},\ldots,c_{N})]}{N^{\ell(\lambda)}(\theta N)^{|\lambda|}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}(a_{(|B|)}+b_{(|B|)}).$
-(B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.22, respectively, and follow the same idea as the proof of (A). ∎
+(B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.10, respectively, and follow the same idea as the proof of (A). ∎
 With this result, we are prepared to prove Corollary˜1.7.
 Proof of Corollary˜1.7.
-(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$.
-Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition.
-Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$.
-It follows from [billingsley]*Theorem 30.2 that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in distribution to $\mu$.
+(A): By part (A) of Corollary˜10.2, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to $\mu$ as $N\rightarrow\infty$.
 (B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$.
 It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero.
-Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$.
-Afterwards, we can follow the steps of (A).
+Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 2}$ converges in terms of power sums to $\tilde{\mu}$.
 ∎
@@
 Proof of Corollary˜1.8.
-[…]
+We use part (B) of Corollary˜10.2.
 Suppose $d\geq 1$ and $\pi\in NC^{\text{even}}(2d)$.
@@
 the equation also holds after replacing $a_{(|B|)}$ with $b_{(|B|)}$.
-By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\\ \left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\mu$.
+By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto-\frac{x}{\theta_{0}N})$ converges in terms of power sums to $\mu$.
 ∎
 10.2. Upper bounds of the magnitudes of Bessel functions
-In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.4.
+In preparation for studying the uniform convergence of Bessel functions, we prove various upper bounds on them.
+Lemma 10.4.
+Suppose $\theta\in\Theta(\mathcal{R})$ and $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$. Then,
+$|J_{a}^{\mathcal{R}(\theta)}(x)|\leq\exp\left(\max_{h\in H(\mathcal{R})}\text{Re}(\left\langle ha,x\right\rangle)\right).$
+Proof.
+By [dunklbound]*Theorem 3.1,
+$\left(\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}E_{ha}^{\mathcal{R}(\theta)}(x)\right)^{2}\leq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}E_{ha}^{\mathcal{R}(\theta)}(x)^{2}\leq\exp\left(2\max_{h\in H(\mathcal{R})}\text{Re}(\left\langle ha,x\right\rangle)\right),$
+which finishes the proof. ∎
 Lemma 10.5.
@@
 Proof.
-First, observe that from Theorem˜1.4,
-$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
-for a measure $\mu_{a}$ over $\mathbb{R}^{N}$ which is supported on the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$. In particular,
-$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon).$
+First, observe that from Theorem˜1.3,
+$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)$
+for a measure $\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}$ over $\mathbb{R}^{N}$ which is supported over the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$.
+In particular,
+$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\right\rVert_{2}$
+$\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
 Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ and $\sum_{h\in H}c_{h}=1$. Then, for $h^{\prime}\in H$,
@@
 Proof.
-(A): We use Theorem˜1.4.
+(A): We use Theorem˜1.3.
 Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$.
 First, observe that
-$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\mu_{a}(\epsilon)\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon).$
-Next, note that since $\mu_{a}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
-$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
+$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)$
+$\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
+Next, note that since $\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
+$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
 Suppose $\epsilon=\sum_{h\in H}c_{h}ha$, where $c_{h}\in[0,1]$ for $h\in H$ satisfy $\sum_{h\in H}c_{h}=1$. Furthermore, suppose $\nu\in\Gamma$. By Muirhead’s inequality and the triangle inequality,
@@
 Hence,
-(31) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
+(22) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
 (B): The equation (22) becomes
-$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
+$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
 which completes the proof. ∎
@@
 ∎
-10.3. The uniform convergence of Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime
+10.3. The uniform convergence of Bessel functions in the high temperature regime
 First, we introduce the following well-known result; we include the proof, which is also well-known, for completeness. A similar argument is used in [rank_infinity].
@@
     Suppose that as $N\rightarrow\infty$, $\theta_{0}N\rightarrow c_{0}\in\mathbb{R}_{\geq 0}$ and $\theta_{1}\rightarrow c_{1}\in\mathbb{R}_{\geq 0}$. Also, suppose $c_{(2k)}$ for $k\geq 1$ solve
-    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}$
+    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{BC}(\pi)(c_{0},c_{1})\prod_{B\in\pi}|B|c_{(|B|)}$
     for $d\geq 1$.
@@
 Proof.
-(A): By Lemma˜10.5, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
+(A): By Lemma˜10.6, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
 Suppose $k\geq 1$ and $r\in\Gamma[k]$. We compute the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ as $N\rightarrow\infty$. First, note that for $s\in\Gamma$,
@@
 Thus, the $N\rightarrow\infty$ limit of the coefficient of $r(x)$ in $J_{a(N)}^{A^{N-1}(\theta)}$ is the coefficient of $r(x)$ in $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$.
-Afterwards, by part (B) of Lemma˜10.7, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
-(B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and Corollary˜9.14 for (C).
+Afterwards, by part (B) of Lemma˜10.8, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
+(B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and (20) and Corollary˜9.8 for (C).
 For (C), observe that the terms with all odd degrees are eliminated over $\Omega$.
@@
 In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system.
-The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.13.
-First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
+The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.14.
+First, we prove the analogues of Theorems˜6.10, 7.4 and 8.9.
 Theorem 10.10.
@@
 Proof.
-See Theorems˜8.10 and 8.16. ∎
-Theorem˜10.12 is an example of an application of the results of this paper.
+See Theorem˜7.1 with $c=0$ and Lemma˜8.8. ∎
+Theorem˜10.13 is an example of an application of the results of this paper.
 Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity].
 Part (C) has not appeared previously, although it is straightforward to deduce.
-See Remark˜10.13 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
+See Remark˜10.14 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
 Theorem 10.13.
@@
 Proof.
-(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.4 and Hölder’s inequality gives that
+(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.3 and Hölder’s inequality gives that
 $\left|J^{A^{N-1}(\theta)}_{a(N)}(x_{1},\ldots,x_{r},0,\ldots,0)\right|$ $=\left|\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{r}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)\right|$
@@
 $=\prod_{i=1}^{r}J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)^{\frac{1}{r}}.$
-Afterwards, using Lemma˜10.6 gives that
+Afterwards, using Lemma˜10.7 gives that
 $J^{A^{N-1}(\theta)}_{a(N)}(r\text{Re}(x_{i}),0,\ldots,0)$ $\leq 1+\sum_{k=1}^{\infty}\frac{|c_{k}(a)|}{(\theta N)^{k}}|r\text{Re}(x_{i})|^{k}$
@@
 Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity].
-We can then apply part (B) of Lemma˜10.7 to conclude the proof.
-(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.6 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.6 and Corollary˜9.11. ∎
+We can then apply part (B) of Lemma˜10.8 to conclude the proof.
+(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.7 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.7, (20), and Corollary˜9.6. ∎
 Remark 10.14.
 From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered.
-Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.12 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$.
+Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.13 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$.
 This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
 10.5. Bessel generating functions for exponentially decaying probability measures
-Rather than consider the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ for a single value of $a$, we can consider the average of $J_{a}^{\mathcal{R}(\theta)}(x)$ over a distribution of $a$.
-The resulting function is referred to as the Bessel generating function and has been studied previously in [gaussianfluctuation, matrix, limits_prob_measures, rectangularmatrix].
-This notion is also related to the Dunkl transform discussed in [dunklbound].
-Suppose $\mu$ is a Borel probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
-$G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)].$
-However, we must place restrictions on $\mu$ so that the Bessel generating function converges.
+Recall that the Bessel generating function is defined in Definition˜1.1.
+We must place restrictions on $\mu$ so that the expected value converges.
 In particular, we consider the class of exponentially decaying probability measures studied in [limits_prob_measures], which is a modification of the class of measures studied in [matrix].
@@
 Furthermore, we can show that the integral is holomorphic over $B_{R}(0)$ by following [matrix]*Proof of Lemma 2.9.
-Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
+Note that we have proved the nonsymmetric version of (A), which can be used to prove (A).
 The goal is to prove using induction that for any sequence $\{i_{j}\}_{1\leq j\leq m}$ of elements of $[N]$, we have that for $x\in B_{R}(0)$,
@@
 Remark 10.17.
-The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system.
+The papers [matrix, limits_prob_measures] only consider the $A^{N-1}$ root system.
 However, (A) and (D) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
@@
 However, we require the exponentially decaying condition to prove (C) and (D).
-Using Lemma˜10.15 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system.
+Using Lemma˜10.16 and part (A) of Theorem˜1.2, we are able to deduce the following corollary for the $A^{N-1}$ root system.
 It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems.
@@
 Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 2$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{N}\}$, define the coefficients $c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in part (B) of Lemma˜10.15.
+For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ as they are defined in part (B) of Lemma˜10.16.
 Then, the following are equivalent.
@@
     For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  2. (a’)
-    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   2. (b)
@@
     $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC(\nu_{i})}\prod_{B\in\pi}|B|c_{(|B|)}.$
-Condition (b) of Corollary˜10.17 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
+Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc].
 When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers.
@@
 Corollary 10.19.
-Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$.
-Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ have nonnegative real part for all $N\geq 1$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
-For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$, define the coefficients $c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.15.
+[…]
+For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.16.
 Then, the following are equivalent.
@@
     For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  2. (a’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   2. (b)
@@
     $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta_{0}N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
-For the type D analogue of Corollary˜10.17, we adjust the notation used to match the notation of part (C) of Theorem˜1.1.
-Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.15.
+For the type D analogue of Corollary˜10.18, we adjust the notation used to match the notation of part (C) of Theorem˜1.2.
+Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.16.
 Lemma 10.20.
@@
 There exist unique constants $d_{0}^{D^{N}(\theta)}(\mu)\\$ $\in\mathbb{C}$ and $c_{b}^{D^{N}(\theta)}(\mu),d_{b}^{D^{N}(\theta)}(\mu)\in\mathbb{C}$ for $b\in\mathcal{B}$ such that
-$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
+$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{D^{N}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{D^{N}(\theta)}(\mu)b\right)$
 over a neighborhood of the origin.
@@
 Assume that $\theta\in\mathbb{C}$ has nonnegative real part for all $N\geq 2$ and $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-Define the coefficient $d_{0}^{D^{N}(\theta)}(\mu)$, the coefficients $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, and the coefficients $c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N}),d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$ as they are defined in Lemma˜10.19.
+Define the coefficients $d_{0}^{D^{N}(\theta)}(\mu)$ and $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ as they are defined in Lemma˜10.20 for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$.
 Consider the following statements.
@@
     For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}\\ =0$ if $\ell(\lambda)\geq 2$.
-  2. (a’)
-    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   2. (b)
-    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
-  4. (b’)
-    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
+    It is the case that $\lim_{N\rightarrow\infty}d_{0}^{D^{N}(\theta)}(\mu_{N})=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
   3. (c)
@@
     $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[e(a)p_{\nu}\left(\frac{a}{\theta N}\right)\right]}{N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
+Then, (a) and (c) are equivalent and (b) and (d) are equivalent.
+Remark 10.22.
+It is straightforward to establish analogues of Corollaries˜10.18, 10.19 and 10.21 in the $\theta N\rightarrow c\in\mathbb{C}$ regime.
 Furthermore, we can generalize Corollary˜1.7 and Corollary˜1.8 to when $a(N)$ and $b(N)$ are randomly distributed. We first state the following generalization of Lemma˜10.1.
@@
 Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative.
-Suppose $\mu_{1}$, $\mu_{2}$, and $\mu$ are Borel probability measures over $\mathbb{R}^{N}$ such that $G_{\mu}^{\mathcal{R}(\theta)}=G_{\mu_{1}}^{\mathcal{R}(\theta)}G_{\mu_{2}}^{\mathcal{R}(\theta)}$ over an open subset of $i\mathbb{R}^{N}$.
+Suppose $\mu_{1}$, $\mu_{2}$, and $\mu$ are exponentially decaying Borel probability measures over $\mathbb{R}^{N}$ such that $G_{\mu}^{\mathcal{R}(\theta)}=G_{\mu_{1}}^{\mathcal{R}(\theta)}G_{\mu_{2}}^{\mathcal{R}(\theta)}$ over an open subset of $i\mathbb{R}^{N}$.
 Then, every element of $\text{supp}(\mu)$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$ for $i\in\{1,2\}$.
@@
 Then, for all $a_{i}\in\text{supp}(\mu_{i})$, $\Pr_{\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}}[\epsilon_{i}\in B(u,r)]=0$, which is a contradiction to (25).
-Recall from Theorem˜1.4 that for $a\in\mathbb{R}^{N}$, the support of $\mu_{a}^{\text{sym}}$ is a subset of the convex hull of $H(\mathcal{R})a$.
+Recall from Theorem˜1.3 that for $a\in\mathbb{R}^{N}$, the support of $\mu_{a}^{\text{sym}}$ is a subset of the convex hull of $H(\mathcal{R})a$.
 By considering when $r\rightarrow 0$, we have that $u$ must be in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$.
@@
     1. 1.
-      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     2. 2.
-      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     For $N\geq 2$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
-Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
+Then, the measure $\mu_{N}(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ with finite moments in terms of power sums.
   2. (B)
     Suppose $\theta\in\mathbb{C}$ is positive.
-Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true.
 For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin.
-Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
+Then, the measure $\mu_{N}(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in terms of power sums.
 Proof.
-(A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution.
-Instead of using Theorem˜1.1, we use Corollary˜10.17.
-(B): We use Lemma˜10.21 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying.
-Afterwards, we follow the proof of (A) for the remaining steps of the proof.
-∎
+(A): We can follow the proof of Corollary˜1.7. Instead of using Theorem˜1.2, we use Corollary˜10.18.
+(B): We use Lemma˜10.23 to deduce that for $N\geq 2$, $\mu^{N}$ is compactly supported and therefore exponentially decaying. Afterwards, we follow the proof of (A). ∎
 In the following two corollaries, we use notation defined in Corollary˜1.8.
@@
     1. 1.
-      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta_{0}N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     2. 2.
-      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta_{0}N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     For $N\geq 2$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
-    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
+    Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in terms of power sums.
   2. (B)
     Suppose $\theta_{0}$ and $\theta_{1}$ are nonnegative.
-Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true.
 For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
-    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
+    Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in terms of power sums.
 Remark 10.26.
@@
     1. 1.
-      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     2. 2.
-      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
+      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
     For $N\geq 2$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin.
-Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
+Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta N})$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in terms of power sums.
   2. (B)
     Suppose $\theta\in\mathbb{C}$ is positive.
-Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true.
+Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true.
 For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin.
-Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
+Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta N})$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in terms of power sums.
 Example 10.28.
 As an example, we consider the Dyson Brownian motion with a random initial condition.
-For $\theta>0$, the Dyson Brownian motion $DBM(\theta)$ over $\mathbb{R}^{N}$ with initial values $(a_{1}(0),\ldots,a_{N}(0))$ is the unique strong solution to
+For $\theta\geq\frac{1}{2}$, the Dyson Brownian motion $DBM(\theta)$ over $\mathbb{R}^{N}$ with initial values $(a_{1}(0),\ldots,a_{N}(0))$ is the unique strong solution to
 […]
 where $B_{i}$, $1\leq i\leq N$ are independent standard Brownian motions.
-For $t\geq 0$, let $\mu_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
+For $t\geq 0$, let $\mu^{\theta}_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
 (27) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
@@
 Suppose $\eta$ is a probability measure over $\mathbb{R}^{N}$.
-Let $\mu_{t}(\eta)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ if we assume that the initial values $(a_{1}(0),\ldots,a_{N}(0))$ are distributed according to $\eta$ independently of the Brownian motions.
+Let $\mu^{\theta}_{t}(\eta)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ if we assume that the initial values $(a_{1}(0),\ldots,a_{N}(0))$ are distributed according to $\eta$ independently of the Brownian motions.
 In particular, we define
-$\mu_{t}(\eta)\triangleq\operatorname{\mathbb{E}}_{a\sim\eta}[\mu_{t}(a)].$
+$\mu^{\theta}_{t}(\eta)\triangleq\operatorname{\mathbb{E}}_{a\sim\eta}[\mu^{\theta}_{t}(a)].$
 Lemma 10.29.
-Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$.
-Then, for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$ and
-$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
+Assume that $\theta\geq\frac{1}{2}$ and $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$.
+Then, for all $t\geq 0$, $\mu^{\theta}_{t}(\eta)$ exponentially decays at any rate less than $R$ and
+$G_{\mu^{\theta}_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
 over $B_{R}(0)$.
 Proof.
-First, we show that for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
+First, we show that for all $t\geq 0$, $\mu^{\theta}_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
 Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $t$ conditioned on the initial value $a$ is at most
@@
 Then, we have that for all $\epsilon>0$,
-$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
-$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
 and
-$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
-$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
-This implies that $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$.
-For the computation of the Bessel generating function $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}$, we have that
-$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}(\eta)(a).$
-for $x\in\mathbb{C}^{N}$. Because $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
-$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}(a)(a(t))\right)d\eta(a).$
-Afterwards, applying (36) and part (A) of Lemma˜10.15 implies that this expression equals
+$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
+$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
+This implies that $\mu_{t}^{\theta}(\eta)$ exponentially decays at any rate less than $R$.
+For the computation of the Bessel generating function $G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}$, we have that
+$G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}^{\theta}(\eta)(a).$
+for $x\in\mathbb{C}^{N}$. Because $\mu_{t}^{\theta}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
+$G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}^{\theta}(a)(a(t))\right)d\eta(a).$
+Afterwards, applying (27) and part (A) of Lemma˜10.16 implies that this expression equals
 $\int_{\mathbb{R}^{N}}J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)d\eta(a)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x).$
@@
 $G_{H_{\theta}}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{1}{2\theta}\sum_{i=1}^{N}x_{i}^{2}\right)$
-over $\mathbb{C}^{N}$.
+over $\mathbb{C}^{N}$; this formula also follows from Theorem˜8.1.
 For $t>0$, let $H_{\theta}^{t}$ be the pushforward of $H_{\theta}$ with respect to multiplication by $t$, so that $H_{\theta}^{t}$ is a random variable over $\mathbb{R}^{N}$ with density at $a$ proportional to $\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{2\theta}\prod_{i=1}^{N}e^{-\frac{\theta a_{i}^{2}}{2t^{2}}}$.
@@
 over $\mathbb{C}^{N}$.
-Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.27 implies that
-(37) $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
-over a neighborhood of the origin for exponentially decaying Borel probability measures $\eta$. This example satisfies Conjecture˜1.6 for randomly distributed $a_{1}$ and $a_{2}$.
+Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.29 implies that
+(28) $G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
+over a neighborhood of the origin for exponentially decaying Borel probability measures $\eta$. This example satisfies Conjecture˜1.5 for randomly distributed $a_{1}$ and $a_{2}$.
 Corollary 10.30.
-Suppose $\theta>0$ and that for $N\geq 1$, $\eta_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{R}^{N}$ such that as $N\rightarrow\infty$, $\operatorname{\mathbb{E}}_{a\sim\eta_{N}}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the distribution $\eta$ over $\mathbb{R}$.
-Then, for $\alpha\geq 0$, $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the free convolution of $\eta$ and the semicircular law multiplied by $\sqrt{\frac{\alpha}{\theta}}$ as $N\rightarrow\infty$.
-If $\eta$ is compactly supported, then $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges weakly to the free convolution as $N\rightarrow\infty$.
+Suppose $\theta\geq\frac{1}{2}$ and that for $N\geq 2$, $\eta_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{R}^{N}$ such that $\eta_{N}(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to the distribution $\eta$ over $\mathbb{R}$ with finite moments.
+Then, for $\alpha\geq 0$, $\mu_{\alpha N}^{\theta}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to the free convolution of $\eta$ and the semicircle law multiplied by $\sqrt{\frac{\alpha}{\theta}}$ as $N\rightarrow\infty$.
 Proof.
 Assume that $\alpha>0$.
-We can follow the method of the proof of Corollary˜10.22.
-First, since condition (b) of Corollary˜10.17 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well.
-The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$.
-This does not follow from (37) and Lemma˜10.21 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
-Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.27, $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$. Then, because
-$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$
-over $B_{R}(0)$ by Lemma˜10.27, applying Corollary˜10.17 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$.
-Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$.
-If $\eta$ is compactly supported, then the free convolution is compactly supported and therefore uniquely determined by its moments by Carleman’s condition, which implies that the sequence of probability measures also converges weakly to the free convolution after applying [billingsley]*Theorem 30.2.
-If $\alpha=0$, then the convergence in terms of moments is clear and the weak convergence follows from [billingsley]*Theorem 30.2. ∎
+We can follow the method of the proof of Corollary˜10.24.
+First, since condition (b) of Corollary˜10.18 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well.
+The next step is to deduce that $\mu_{\alpha N}^{\theta}(\eta_{N})$ is exponentially decaying for all $N\geq 2$.
+This does not follow from (28) and Lemma˜10.23 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
+Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.29, $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ exponentially decays at any rate less than $R$. Then, because
+$G_{\mu_{\alpha N}^{\theta}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$
+over $B_{R}(0)$ by Lemma˜10.29, applying Corollary˜10.18 again gives that the $k$th free cumulant of $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$.
+Hence, the probability measures $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\eta$ and the semicircle law multiplied by $\frac{\alpha}{\theta}$ in terms of power sums as $N\rightarrow\infty$.
+∎
+Example 10.31.
+Next, we study the law of large numbers for the $\beta$-Laguerre ensemble.
+First, we compute the type BC Bessel generating function of the Chiral ensemble, which has the same distribution as the square roots of the $\beta$-Laguerre ensemble, where $\beta=2\theta$.
+We discuss its construction for $\theta\in\{\frac{1}{2},1,2\}$ and its definition for general values of $\theta\geq 0$ from [forrester]*Section 3.1.
+Suppose $M\geq N$ such that $M-N+1-\frac{1}{2\theta}\geq 0$.
+Let $X$ denote a random $M\times N$ matrix whose independent entries are real, complex, or real quaternion numbers with densities:
+  • $\frac{1}{\sqrt{2\pi}}e^{-\frac{x^{2}}{2}}$ if $\theta=\frac{1}{2}$,
+  • $\frac{1}{\pi}e^{-|x|^{2}}$ if $\theta=1$,
+  • $\frac{4}{\pi^{2}}e^{-2|x|^{2}-2|y|^{2}}$ if $\theta=2$. Note that we parameterize a quaternion as $(x,y)\in\mathbb{C}^{2}$.
+The Chiral ensemble is defined as the positive eigenvalues of
+$H=\begin{bmatrix}0_{M\times M}&X\\ X^{H}&0_{N\times N}\end{bmatrix}$.
+This matrix almost surely has $N$ negative eigenvalues, MNM-N zero eigenvalues, and $N$ positive eigenvalues.
+Define the probability distribution $C_{\theta,M,t}^{N}$ over $\mathbb{R}_{\geq 0}^{N}$ to have the density of $a$ proportional to
+$\prod_{i=1}^{N}a_{i}^{2\theta(M-N+1)-1}e^{-\frac{a_{i}^{2}}{2t}}\prod_{1\leq i<j\leq N}|a_{i}^{2}-a_{j}^{2}|^{2\theta}.$
+Then, $C_{\theta,M,t}^{N}$ is the density function of the positive eigenvalues of $H$ after they are rescaled when $\theta\in\{\frac{1}{2},1,2\}$.
+The following lemma is a straightforward generalization of [rectangularmatrix]*Proposition 5.21.
+Lemma 10.32.
+Suppose $\theta\geq 0$, $M\geq N$ such that $M-N+1-\frac{1}{2\theta}\geq 0$, and $t>0$. Then,
+$G_{C^{N}_{\theta,M,t}}^{BC^{N}(\theta,\theta(M-N+1)-\frac{1}{2})}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right).$
+Next, we discuss the $\beta$-Laguerre ensemble, which has the same distribution as the squares of the eigenvalues of the Chiral ensemble.
+Suppose $M\geq N$ such that $M-N+1-\frac{1}{\theta}\geq 0$.
+Define the probability distribution $L_{\theta,M,t}^{N}$ over $\mathbb{R}^{N}_{\geq 0}$ to have the density of $a$ proportional to
+$\prod_{i=1}^{N}a_{i}^{\theta(M-N+1)-1}e^{-\frac{a_{i}}{2t}}\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{\theta}.$
+Although we cannot generalize the random matrix that we previously defined for $\theta\in\{\frac{1}{2},1,2\}$, [betaensembles] computes a tridiagonal matrix whose eigenvalue distribution is given by $L_{\theta,M,t}^{N}$ for all $\theta>0$.
+Moreover, the distribution is the image of $C_{\theta,M,t}^{N}$ under the pushforward map $x\mapsto x^{2}$.
+Despite this relation, it remains challenging to compute the Bessel generating function for the $\beta$-Laguerre ensemble.
+Afterwards, we obtain Corollary˜10.33, which has been previously proved in Theorem 1.1 of the paper [betalaguerre].
+The paper proves that the $\beta$-Laguerre ensemble weakly converges to the Marchenko-Pastur law.
+It is straightforward to verify that the moments appearing in the corollary are the same as those of the Marchenko-Pastur law.
+Corollary 10.33.
+Assume that $\lim_{N\rightarrow\infty}\theta N=\infty$, $M\geq N$, $M-N+1-\frac{1}{2\theta}\geq 0$, $\lim_{N\rightarrow\infty}\frac{M}{N}=c\geq 1$, and $\lim_{N\rightarrow\infty}t=\alpha\geq 0$.
+Then, for all $\nu\in\Gamma$,
+$\lim_{N\rightarrow\infty}\frac{\mathbb{E}_{a\sim L^{N}_{\theta,M,t}}[p_{\nu}(a)]}{(\theta N)^{|\nu|}N^{\ell(\nu)}}=\prod_{i=1}^{\ell(\nu)}\sum_{p\in\mathcal{D}_{\nu_{i}}}c^{e(p)}(2\alpha)^{\nu_{i}}.$
+Proof.
+Use Corollary˜10.19, Lemma˜10.32, and the fact that $L_{\theta,M,t}^{N}$ is the image of $C_{\theta,M,t}^{N}$ under the pushforward map $x\mapsto x^{2}$. ∎
 11. Computing the Dunkl bilinear form with combinatorics
@@
 Remark 11.2.
-The previous result is not applicable to when $k+\ell(\lambda)-\ell(\nu)\geq N$, but if this is the case, then we must consider when the number of distinct indices is $N$.
+The previous result is not applicable when $k+\ell(\lambda)-\ell(\nu)\geq N$, but if this is the case, then we must consider when the number of distinct indices is $N$.
 Then, to obtain a summation formula for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ in (A), we must set $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ to be one of the summands.
@@
 Afterwards, we cannot ensure that the order of the Dunkl bilinear form is $N^{k+\ell(\lambda)-\ell(\nu)}$, if we assume that $\lambda$ and $\nu$ are fixed.
-Assuming that $\theta$ is fixed, the order of the expressions in Theorem˜11.1 matches the order of the expressions in Theorems˜6.1, 7.1 and 8.10, which is $N^{k+\ell(\lambda)-\ell(\nu)}$.
+Assuming that $\theta$ is fixed, the order of the expressions in Theorem˜11.1 matches the order of the expressions in Theorems˜6.1 and 7.1, which is $N^{k+\ell(\lambda)-\ell(\nu)}$.
 However, the leading order coefficients are not apparent from the formulas.
-12. Leading order terms of the type A and BC Dunkl bilinear forms after multiplying by $x_{1}\cdots x_{N}$
-In this section, we prove analogues of Theorems˜8.16 and 8.23. First, we prove the analogues of Theorem˜8.16 for the $A^{N-1}$ and $BC^{N}$ root systems.
-Theorem 12.1.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{A^{N}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(N\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
-$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}+R(N,\theta)\bigg),$
-where $R\in\mathbb{Q}[x,y]$ satisfies:
-  1. (1)
-    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
-  2. (2)
-    The degree $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$.
-Proof.
-The same method as the proof of Theorem˜8.16 can be used. ∎
-Corollary 12.2.
-Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$. For (B), assume that $\lambda,\nu\in\Gamma_{\text{even}}$. Suppose $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
-$(A)\,\,\,[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+(i-1)\theta)[p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$
-$(B)\,\,\,[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+2(i-1)\theta)[p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$
-Proof.
-(A) follows from Theorems˜6.1 and 12.1 and (B) follows from Theorems˜8.10 and 8.16. ∎
-Theorem 12.3.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})\bigg(N^{2k+\ell(\lambda)-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}(2\theta_{0})^{k-\ell(\nu)}\times$
-$\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
-$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}+R(N,\theta_{0},\theta_{1})\bigg),$
-where $R\in\mathbb{Q}[x,y,z]$ satisfies:
-  1. (1)
-    In each summand, the degree of $x$ is at most $\ell(\lambda)$ greater than the degree of $y$.
-  2. (2)
-    The degree of $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the sum of the degrees of $y$ and $z$ in each summand is at most $2k-\ell(\nu)$.
-  3. (3)
-    In no summand is the degree of $x$ $\ell(\lambda)$ greater than the degree of $y$ while the degrees of $y$ and $z$ add to $2k-\ell(\nu)$.
-Proof.
-First, we have that $h\mathcal{D}(e)ep_{\nu}=\mathcal{D}(he)hep_{\nu}=\mathcal{D}(e)ep_{\nu}$ for all $h\in H(BC^{N}(\theta_{0},\theta_{1}))$.
-Hence, $\mathcal{D}(e)ep_{\nu}$ has all even degrees.
-It follows that we can replace $\mathcal{D}(ep_{\lambda})$ with
-$\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}\mathcal{D}_{j}^{\lambda_{i}-1}\partial_{j}\mathcal{D}(e),$
-as we have done previously.
-After noting that $\theta_{1}$ has order $\theta_{0}N$, the same method as the proof of Theorem˜8.16 can be used. In this case, we can view the switch from $i$ to $j$ as
-$\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right).$
-We replace $\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}$ with $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}$ to compute the leading order term. Thus, we can view the switch from $i$ to $j$ as
-(39) $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right);$
-afterwards, we follow the same method. The goal is to compute
-$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu},$
-where $\mathcal{S}$ is defined in the same way except that the switch from $i$ to $j$ is replaced with (39) and $\mathcal{R}(s)$ is defined in the same way except that $2\theta d_{i}$ is replaced by $(2\theta_{0}+\frac{2\theta_{1}}{N})d_{i}$.
-We get that we can similarly reduce to the case that $\ell(\lambda)=1$.
-Hence, let $c_{\nu^{\prime}}$ be the value of (22) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$ for $\nu^{\prime}\in\Gamma_{\text{even}}$.
-The only difference is in Lemma˜8.20.
-Assume that $\pi\in NC^{\prime}(|\nu^{\prime}|)$ and $b(\pi)=\nu^{\prime}$.
-For $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$ that are associated with $\pi$, we have that the number of switches in $s$ which are not associated with a derivative in $r$ is $|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1$ and the number of switches in $r$ is $|\nu^{\prime}|-\ell(\nu^{\prime})+1$; recall Definition˜8.17 for the definition of $d_{3}(r)$.
-The switches in $r$ contribute a factor of $\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}$.
-On the other hand, the switches in $s$ which are not associated with a derivative in $r$ contribute a factor of $(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}(1+\frac{\theta_{1}}{\theta_{0}N})^{o(\pi)}$.
-The switches in $s$ which are associated with a derivative in $r$ contribute a factor of $(-2\theta_{0})^{d_{3}(r)}$.
-Hence, the total contribution is
-$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}(-2\theta_{0})^{d_{3}(r)}$
-$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-d_{3}(r)}(-1)^{d_{3}(r)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)+d_{3}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)},$
-where we have used $d_{1}(s)+d_{2}(r)+d_{3}(r)=\ell(\nu^{\prime})-1$.
-Observe that the number of locations for a switch in $s$ that is associated with a derivative in $r$ is $z(\pi)$.
-For each of these locations, we can decide to place the derivative in $s$ or in $r$.
-After summing over these choices, we obtain
-$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-z(\pi)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}-2\theta_{0}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}$
-$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-d_{2}(r)}$
-$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
-Thus, we obtain that
-$c_{\nu^{\prime}}=$ $\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC^{\prime}(|\nu^{\prime}|),\,b(\pi)=\nu^{\prime}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\times$
-$\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
-It follows that
-$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{j}:j\in S_{i}))}$
-$=\prod_{l=1}^{\ell(\nu)}\nu_{l}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{\lambda_{i}-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\times$
-$(2\theta_{0})^{\lambda_{i}-\ell(b(\pi))}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
-$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma_{i}\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
-$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
-$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
-$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}.$
-After multiplying by $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})N^{2k+\ell(\lambda)-\ell(\nu)}$, we obtain the desired result. ∎
-Next, we prove the analogues of Theorem˜8.23 for the $A^{N-1}$ and $BC^{N}$ root systems.
-Theorem 12.4.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(1+N\theta)^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
-$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(N\theta)\prod_{B\in\pi}x_{|B|}$
-$+R(N,\theta)\bigg),$
-where $R\in\mathbb{Q}[x,y]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
-Proof.
-We can prove this in the same way as we prove Theorem˜8.23. ∎
-For $\pi\in NC(k+1)$, define
-$W^{BC;\,\text{odd}}(\pi)(x,y)\triangleq$ $(1+\mathbf{1}\{b(2;\,\pi)=0\}(d(2;\,\pi)-1))\times$
-$\prod_{i\in[k+1],i\geq 3,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi))\left(\frac{1+y}{1+x+y}\right)^{z(\pi)}\times$
-$\left(\frac{1}{1+x+y}\right)^{\mathbf{1}\{f(\pi)>1\}}.$
-Theorem 12.5.
-Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then,
-$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2\theta_{1}+2(i-1)\theta_{0})\times$
-$\bigg(N^{\ell(\lambda)}(1+2\theta_{1}+2N\theta_{0})^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
-$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{BC;\text{odd}}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in b(\pi)}x_{|B|}$
-$+R(N,\theta_{0},\theta_{1})\bigg),$
-where $R\in\mathbb{Q}[x,y,z]$ satisfies the condition that in each of its summands, the degree of $x$ is at most $\ell(\lambda)-1$ greater than the degree of $y$.
-Proof.
-We can prove this theorem in the same way as we prove Theorem˜8.23. The main difference is that in the last step of the proof of Lemma˜8.24, the total contribution is
-$1-\frac{2N\theta_{0}}{1+2\theta_{1}+2N\theta_{0}}=\frac{1+2\theta_{1}}{1+2\theta_{1}+2N\theta_{0}},$
-which would correspond to $\frac{1+y}{1+x+y}$ rather than $\frac{1}{1+x+y}$. ∎